A Bibliography of Publications in IEEE Computer: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

01 October 2024 Version 2.17

Title word cross-reference

$1 [1345]. $1-Million [1505]. 10, 000_{2} [1089, 1111]. $250 [1529]. 3 [1406, 2087]. ? =? [916]. ^{3} [976]. χ [322]. Φ [144].

1 [105, 743, 2104]. 100} [882]. 12 [792]. 14001 [1521]. 1987 [1541]. 1990s [32, 177, 184]. 1996 [1286]. 1997 [1671]. 1999 [2005]. 1B [1733, 2022].

2 [339, 2039]. 2000 [1324, 1462, 1482]. 2001 [2058, 2060]. 21st [1379, 1608, 1844]. 25 [474]. 2nd [259, 301].

3 [1966]. 300-mm [1662]. 300-nm [1454]. 31 [2084]. 32-bit [648]. 3D [513, 701, 702, 1178, 1710, 1993]. 3DP [209].

3DSP [2145]. 3rd [314].

4 [1634]. 400 [1771]. 4Developers [1837]. 4GL [2145].

5 [1366]. 5.0 [2145]. 52 [231].

6000 [105]. 64 [1712].

7 [1988]. 754 [1621].

802 [71].

91 [1637]. 95 [900]. ’96 [1235]. ’98 [1626]. 99 [2005].

abolished [898]. Absorb [1393].

Abstraction [278]. Abstractions} [239]. Academia [1707]. Accelerate [2130].

Accelerated [1698]. Acceptance [1080]. Access [29, 83, 207, 413, 523, 822, 1028, 1107, 1603, 1662, 1822, 1873, 2015, 2065, 2069]. Accessible [1527]. Accessing [30]. Accidents [474]. Accommodating [383]. Account [1636]. Accountants [1720]. Accreditation [1649]. Achievement [1384, 1492, 1514, 1745, 1909]. Achievers [2034]. Achieving [368, 369, 688]. Acid [1894]. ACM [2084]. Acquires [1505]. Across [602, 914]. ACS [2104]. ACS-1 [2104]. Act [1236]. Action [902, 1874]. Activating [1928]. Active [1735]. ActiveX [1791, 1837, 1937, 1962]. Activities [855, 876, 896, 930, 961, 985, 1015, 1035, 1067, 1098, 1117, 1139, 1162, 1183, 1205, 1228, 1250, 1270, 1309, 1355, 1385, 1440, 1467, 1491, 1513, 1570, 1592, 1593, 1616, 1649, 1650, 1718, 1766, 1788, 1809, 1828, 1859, 1860, 1909, 1957, 2008, 2034, 2113]. Activity [838, 1114, 1128]. Acumen [1962]. AD [1462]. Ada [23, 216, 570, 604, 605, 704, 1771, 1790, 1791]. Adaptability [1923]. Adapting [411, 1974]. Adaptive [352, 443, 752]. Added [189]. addendum [231]. Address [104, 289, 1260]. Addresses [450, 1419]. Addressing [1491]. Adds [2126]. Adopted [1679]. Adopters [1487]. Adoption [1869]. Adopts [1634]. ADSL [2015]. Advance [1858]. Advanced [370, 1323, 1947, 1973]. Advances [186, 1180, 1219, 1670, 1783, 1829, 1879, 2047]. Advancing [1958]. Advent [1756]. Advertiser [858, 935, 955, 979, 1012, 1061, 1210, 2037, 2115, 2142]. Advertiser/Product [858, 935, 955, 979, 1012, 1061, 1210, 2037, 2115, 2142]. Advice [1365]. affects [757]. Afraid [1579]. Africa [737]. After [1368, 1637, 1993]. Aftermath [1850]. Again [1190, 1272, 1299]. Age [883, 968, 982, 1008, 1091, 1546, 1708, 1927, 1957]. Agencies [1974]. Agency [1454]. Agent [1388, 1824, 1825]. Agent-Based [1388, 1824]. Agents [265, 473, 726, 820, 1110, 1177, 1388, 1507, 1508, 1510, 1923].

Agentsheets [785]. Agerwala [1745]. Ages [937]. Aggarwal [1384]. Agile [1764]. Ago [1370]. Agrawal [1745]. Agree [899, 1454, 1482]. Agreement [1299, 2072, 2096]. Ahead [1577]. Ahpah [1519]. AI [264, 279, 720, 755, 1285, 1970, 1994]. Aid [90, 1716]. Aided [160, 999, 1939]. aim [626]. Aims [2046]. Aizu [2005]. al [273]. al. [50]. Albert [2077]. Alerting [317, 318]. Alerts [1281]. alewife [1262]. Algebra [157]. Algorithm [74, 143, 367, 634, 718, 1130, 1764]. Algorithm-Agile [1764]. Algorithm-Based [718]. Algorithms [3, 8, 26, 33, 44, 88, 139, 155, 169, 223, 242, 243, 633, 653, 702, 1438]. Alice [1147]. Allegory [1449]. Alliance [1299, 1822]. Alloc [591]. Alone [1895]. Alpha [635, 1667, 1999]. Alternative [1454]. Always [706, 2146]. am [753]. Amazing [1505]. AMC [2087]. Amendment [1649, 1860, 2034]. amendments [479]. America [1974]. American [94]. Amid [1848]. Amoeba [32]. Amok [1790]. Analog [161]. Analogy [1791]. analyses [1166]. Analysis [55, 122, 286, 287, 330, 331, 416, 427, 552, 951, 1646, 1653, 1666, 1700, 1701, 1771, 1791, 1841, 1892, 2031]. Analyzer [570]. Anark [2087]. anatomical [998]. Anderson [229]. Andreessen [1384]. Andrew [1342]. Andrews [274]. animat [352]. Animate [220]. Animation [74, 367, 1178]. Annasoft [1791]. Annealing [263]. anniversary [1232]. announced [1541]. Announces [1898]. Annual [372, 989, 1539, 1827, 2138]. Anser [66]. Answer [1332, 2058]. Answers [1337, 1979, 2044]. Antipiracy [1634]. Antivirus [1582]. ANTS [1929]. Any [1933]. Anyway [1079]. Anywhere [2044]. AOL [1505]. Aonix [1724, 1962, 1988]. Aphorisms [1428]. API [626, 1554, 1812, 1837, 1898, 2128]. APIs [2003]. Apple [1101, 1319, 1454, 1849, 1961]. Appliance [1481, 1482, 1857, 2095].

Appliance-on-Chip [2095]. Appliances [382, 1566, 1898, 2045]. Applicants [1356, 1593]. Application [66, 86, 92, 122, 168, 337, 591, 851, 879, 925, 952, 966, 1003, 1035, 1049, 1193, 1261, 1262, 1321, 1498, 1699, 1724, 1937, 2021, 2087]. Application-Level [591]. Applications [12, 209, 284, 357, 592, 620, 719, 743, 761, 769, 853, 860, 1132, 1135, 1258, 1388, 1397, 1513, 1639, 1640, 1644, 1765, 1771, 1824, 2003, 2069, 2087, 2134]. Applied [264, 2014]. Applying [332, 683, 1377, 1437]. Appointments [1885]. Appreciation [2140]. Approach [16, 135, 160, 195, 203, 257, 348, 373, 501, 588, 605, 771, 1077, 1290, 1347, 1490, 2126]. Approaches [352, 353]. Approval [1634]. Approve [2084]. Approved [2072]. Approves [1554, 1649, 1721, 2009, 2034]. Apps [1519, 2014]. April [1071]. Apt [359]. Aquarius [48, 211, 212]. Arabic [296, 297]. Architect [1985]. Architectural [280, 281, 308, 459, 619, 653, 826, 1341, 1458]. Architecture [14, 50, 81, 98, 102, 115, 136, 176, 177, 209, 221, 226, 235, 244, 245, 255, 259, 263, 266, 267, 269, 270, 272, 279, 320, 321, 325, 336, 358, 718, 930, 1260, 1380, 1437, 1536, 1614, 1673, 1712, 1782, 1802, 1807, 2072]. Architecture-Independent [98, 221]. Architectures [7, 18, 38, 39, 45, 55, 242, 243, 261, 271, 459, 618, 668, 719, 1053, 1170, 1460, 1772, 1983, 2041]. Archive [1881]. Archives [1616]. Area [76, 101, 455, 472, 1708, 1951]. arena [217]. Aren’t [1409, 1797, 2093, 2146]. Ariane [1293, 1366]. Arithmetic [1325]. ARM [1835]. ARPA [432]. Array [110, 247, 248, 802, 1291]. Arrays [5, 251, 252, 524, 586]. Arrogance [2098]. Art [1456, 1501]. Artery [124]. Article [442, 843, 885, 908, 945, 969, 995, 1025, 1048, 1103, 1123, 1126, 1149, 1171, 1192, 1215, 1237, 1408, 1433, 1455, 1483, 1506, 1530, 1555, 1603, 1607, 1635, 1663, 1685, 1711, 1737, 1758, 1779, 1851, 1876, 1901, 1926, 1950, 1976, 1997, 2025, 2050, 2075, 2099, 2131]. Artificial [353, 488, 1051, 1053, 1054, 1223, 2144]. Artisoft [2014]. Artlandia [1791]. AS/400 [1771]. Asbestos [1942]. ASC [717]. ASCII [1347]. Ashes [1963]. Asimov [541, 556]. Ask [2093]. Asphalt [1752]. Aspray [340]. Assembly [2007]. Assessment [468, 898, 1001]. Assistance [159]. Assistants [1252]. Associated [1691]. Associative [716–721]. Associative-Computing [717]. Assurance [1638, 1639, 1647, 1859, 1940]. assured [240]. assuring [1240]. Astronautics [105]. Asynchronous [143, 1511, 1580]. Athena [75]. atlases [998]. ATM [601, 1294, 1661, 1764, 1790, 1838, 1994]. Attack [1058]. Attacks [1236, 1528, 1931]. Attaining [1254]. attitudes [265]. Attract [1909]. Auction [1662]. Audio [69, 602]. Auditor [645, 840, 1146, 1417]. Augmenting [67]. August [1167]. Authentication [210, 231, 232]. Author [366]. Authoring [1154, 1724]. Automata [345]. Automated [315, 316, 384, 1031, 1361, 2088, 2117]. Automatic [213, 789, 890, 1743, 1791, 1988, 2108]. Automatically [1131, 1489]. Automating [188, 850]. Automation [896, 2113]. Autonomous [129, 206, 265, 1176, 1512, 1982]. Autorouter [1724]. Availability [178, 254, 804, 1351]. Available [29, 62, 567, 936, 1628, 1642, 1791, 2084]. Avilon [1962]. Avionics [1838]. Avoiding [1238]. Award [1384, 1492, 1514, 1571, 1593, 1671, 1745, 1829, 1860, 1933, 1985, 2009, 2034, 2102, 2111, 2140]. Awarded [1985]. Awards [1032, 1287, 1384, 1492, 1514, 1745, 1909, 1944, 2084, 2140]. Aware [2001]. Away [1841]. Axiomatic [2091].

B [229, 257, 906]. Baan [1962]. Babel [1728].

Bach [1655]. Back [992, 1071, 1122, 1145, 1213, 1233, 1255, 1450, 1505, 1549]. Backbone [1723]. Backfiring [959]. backplane [692]. bad [694]. Bal [283]. Balancing [1962]. Ball [1770]. Ballot [2009]. Balloting [1441, 1541]. Ban [1454]. Bandwidth [986, 1256, 1480, 1715]. Banking [66, 2023]. Banyan [1836]. Bar [25, 277]. Baring [1466]. Barrier [1297]. BASAR [820]. Base [205]. Based [43, 51, 99, 119–121, 138, 157, 160, 195, 294, 295, 306, 307, 317, 318, 349, 350, 381, 382, 394, 444, 523, 527, 667, 718, 886, 890, 910, 923, 974, 1028, 1110, 1114, 1134, 1217, 1288, 1352, 1354, 1388, 1400, 1412, 1430, 1490, 1570, 1624, 1645, 1660, 1688, 1700, 1742, 1798, 1805, 1812, 1824, 1828, 1837, 1858, 1932, 1965, 1980, 1990, 2000, 2003, 2030, 2109, 2145]. Bases [91, 165]. Basic [1822]. Battle [1345, 1454, 1554, 1872, 1971, 1973]. Battles [1849]. Be [26, 819, 898, 1443, 1890, 1897, 1970, 2070, 2092]. beans [1190]. Bearing [1444]. Beautiful [1497]. Become [1307]. Becomes [1873]. begin [1541]. Beginners [1336]. Begins [992, 1554, 1973]. behavior [352]. Behavioral [2002]. Behind [1731]. Beholder [844]. Being [1679, 1915]. Belief [754]. Bell [385, 557, 756, 1005, 1266, 1286, 1541, 1569, 1829]. Benchmark [106, 1214, 1669]. benchmarking [938]. Benchmarks [100, 700]. Benefit [1981]. Benefits [1452, 1798]. Bertsekas [301]. Berzins [239]. Best [834, 1390, 2040]. Better [303, 977, 1430, 1586, 1822, 2096]. Between [132, 685]. Beware [1363, 1364]. Beyond [310, 1349, 1378, 1395, 1447, 1462, 1710, 1781, 1903, 1954, 1984]. big [823, 1046, 1947, 2110]. Billings [810]. Billion [1345, 1460, 1461, 1464]. Billion-Transistor [1460, 1464]. Bina [1384]. Binary [748, 764, 782, 799, 823, 841, 866, 943, 967, 993, 1022, 1046, 1072, 1101, 1124, 1147, 1169, 1170, 1190, 1235, 1295, 1316, 1332, 1345, 1364, 1393, 1449, 1476, 1501, 1524, 1550, 1579, 1602, 1626, 1658, 1680, 1706, 1730, 1731, 1753, 1774, 1797, 1817, 1843, 1868, 1894, 1919, 1942, 1968, 1992, 2018, 2043, 2068, 2092, 2123, 2149]. Binding [1791]. Bioelectric [516]. Biologic [344]. Biological [193]. Biology [193]. Biomedical [122]. Biometric [1898]. Biometric-API [1898]. Biometrics [1734]. Biomolecular [346–348]. Bit [113, 247, 248, 648, 2122]. Biting [208]. Bitmapped [9]. Black [1450, 1791]. blackboard [267]. Blacklists [1875]. Blame [1366]. Blinded [2119]. Block [289]. Blocks [344]. Blue [1345, 1484, 1724, 1937, 1947]. Bluetooth [1710]. Board [479, 897, 1468, 1541, 1593, 1649, 1721, 1819, 2009, 2034, 2059, 2087]. boards [1094]. Boca [1039]. Body [997]. Bomb [1041]. book [241]. Books [1312, 1359, 1389, 1423, 1445, 1472, 1497, 1520, 1546, 1573, 1595, 1619, 1652, 1673, 1696, 1723, 1747, 1770, 1790, 1811, 1833, 1862, 1888, 1913, 1936, 1961, 1987, 2013, 2038, 2063, 2088, 2117, 2144]. Boom [1112]. Boosting [1414]. Booth [1384, 1492]. booting [753]. BOP [754]. Borland [1519, 2039]. Both [1677]. Bourbakis [1745]. Box [2039]. Boxes [1606]. Bracing [1855]. Brain [353, 906, 1000, 1020, 1895]. Branch [1383]. Breaking [194, 1297]. Breakthrough [1425]. Brew [1849]. Brews [1454]. Bridging [132, 1473]. Brief [1850, 1899, 1924, 1948, 1974, 1995, 2023, 2048, 2073, 2097, 2129]. Briefs [1299, 1319, 1345, 1376, 1407, 1432, 1454, 1482, 1505, 1529, 1554, 1583, 1606, 1634, 1662, 1684, 1710, 1736, 1757, 1778, 1801, 1822, 1849, 1873, 1874, 1898, 1923, 1947, 1973, 1994, 2022, 2047, 2072, 2096, 2128]. Brighter [208]. Bring [1345, 1994]. Bringing [1898, 2058, 2060]. Brings [1933]. Broadband [499, 1662]. Broadband-ISDN [499]. Broadcasting [304, 305]. Broker [1791]. Brokers [916]. Brook [1818]. Brooks [1370]. Brother [2110]. Browser [1777, 1778, 1822]. Browsers [1868]. Browsing [317, 318, 1822]. BT [1907]. BTU [1669]. Buckley [1225]. Buffer [41]. Bug [1798]. Bug-Free [1798]. Buggy [1798]. Build [742, 1884]. Builder [1829, 2145]. Building [110, 140, 344, 1104, 1150, 1522, 1551, 1652, 1706, 1882, 1903]. Built [83, 1240, 1242]. Built-in [83, 1240]. Bullet [208]. bulletin [1094]. Bundling [68]. burned [1073]. Bus [48, 1801]. Buses} [258, 692]. Business [993, 1181, 1321, 1381, 1473, 1485, 1505, 1723, 1811, 1904, 1908, 1974]. Businesses [1389]. Bust [1112]. Busyware [1793]. buy [742]. Buzz [2000]. Buzzword [1447]. Bylaws [479, 1649, 1860, 2034]. Bytecode [1780].

C [273, 660, 878, 1011, 1185, 1369, 1554, 1609, 1681, 1765, 1834, 1909, 2064, 2118]. C/C [1834, 2064]. C2 [1775]. C31 [1062]. Cable [1606]. Cache [39, 40, 42, 43, 51, 320, 321, 700, 1415, 1723, 1724, 1983]. Cache-Only [320, 321, 1983]. Caches [1409]. Caching [587, 1512]. CACI [1771]. CAD [284, 290, 388, 1654, 2064, 2087]. CAD-UL [2064, 2087]. Café [1164, 1519]. CAI [1791]. Cain [1492]. Calendar [863, 903, 940, 962, 987, 1019, 1042, 1069, 1120, 1143, 1165, 1186, 1209, 1231, 1253, 1386, 1420, 1442, 1469, 1493, 1516, 1540, 1543, 1572, 1594, 1617, 1651, 1672, 1694, 1722, 1746, 1767, 1789, 1810, 1831, 1861, 1886, 1911, 1934, 1959, 1986, 2012, 2035, 2061, 2085, 2114, 2141]. Calendaring [1547]. Calibration [1668]. Call [863, 903, 940, 962, 987, 1019, 1042, 1069, 1120, 1135, 1143, 1165, 1186, 1209, 1231, 1253, 1308, 1386, 1420, 1442, 1469, 1493, 1516, 1543, 1572, 1578, 1593, 1594, 1617, 1651, 1672, 1694, 1722, 1746, 1767, 1789, 1810, 1831, 1861, 1886, 1910, 1911, 1934, 1959, 1986, 2005, 2006, 2012, 2035, 2061, 2085, 2110, 2114, 2141]. Can [1373, 1890, 1895, 1933, 1987, 2044, 2070]. Candid [1445]. Candidate [583, 584, 1768]. Candidates [479, 964, 2058]. Cap [2022].

Capability [517, 750]. Capacity [1454, 1554]. Captioned [890]. Capturing [377, 1029, 1643]. Card [858, 935, 955, 979, 1012, 1061, 1454, 1801]. Cardiac [1001]. Cards [1157, 1376, 1750, 2146]. Care [197, 2070]. Career [864, 939, 963, 988, 1018, 1068, 1211, 1495, 1517, 1544, 1618, 1717, 1769, 1832, 1887, 1935, 1960, 2011, 2036, 2062, 2086, 2116, 2143]. Careers [1912]. careful [819]. Carnot [205]. Cars [1524, 1567, 1681]. Cartoons [1088]. CASE [3, 284, 299, 427, 487, 538, 687, 700, 738, 861, 948, 986, 1031, 1245, 1280, 1392, 1434, 1713, 1736, 1771, 2013]. Case-Driven [2013]. Cases [1811]. Cash [1180]. Casualties [1777]. Cataloging [395]. Catalyst [1458]. Catch [1753]. Catch-Up [1753]. Causality [1029]. cause [111]. Causes [2071]. cautionary [904]. Cautious [386]. CD [842]. CD-ROM [842]. CE [1371, 1791, 1937]. Celebrating [1340]. celebration [992]. Celesstin [290]. Cell [8]. Cell-Selection [8]. Cellular [1998]. Center [1771, 1973, 2094]. Centered [1106]. Centers [981, 1995]. Central [1693]. Centric [1173, 2064]. Century [1379, 1608, 1844, 1854]. Ceremony [2140]. Certainty [1282]. Certified [1812]. Certifying [1690]. Chabot [889]. Chain [84, 615, 2074]. chairs [873]. Challenge [550, 813, 899, 1036, 1223, 1683, 2045]. Challenges [457, 600, 603, 615, 642, 661, 674, 694, 708, 724, 738, 742, 755, 757, 774, 793, 808, 809, 835, 856, 875, 895, 938, 959, 983, 1009, 1038, 1066, 1096, 1114, 1137, 1296, 1513, 1533, 1557–1564, 1642, 1688, 1804, 1848, 1853, 1981, 2105, 2107]. Chameleon [2033]. Chance [1675]. Change [964, 1038, 1224, 1458, 1459, 1510, 1751, 1781, 2082]. Changes [1184, 1333]. Changing [1536, 1604, 1644, 1776, 1842, 1905]. Channel [217, 240, 260, 275, 285, 303, 360, 361, 407, 422, 440, 454, 471, 482, 496, 509, 522, 534, 548, 564, 579, 597, 612, 629, 644, 664, 678, 697, 713, 730, 745, 763, 781, 798, 812, 839, 865, 877, 904, 941, 964, 990, 1020, 1043, 1070, 1099, 1121, 1144, 1166, 1187, 1212, 1232, 1254, 1272, 1341, 1373, 1428, 1451, 1479, 1526, 1578, 1601, 1625, 1656, 1677, 1704, 1729, 1733, 1795, 1815, 1842, 1893, 1941, 1967]. channels [260]. Chapter [1933]. Chapters [1933, 2084]. Character [296, 297, 1961]. Characterization [106]. Chart [1519, 1718]. Chasing [1345]. Chatterbots [1994]. Check [892, 2070]. Checker [1962]. Checkers [59, 220]. Checking [2118]. checkpoint [394]. Chemistry [1913]. Chief [392, 409, 423, 566, 581, 598, 613, 631, 646, 665, 680, 698, 714, 731, 747, 965, 1045, 1356, 1593, 2112]. chill [1046]. China [871]. Chinese [194]. Chip [968, 1148, 1299, 1432, 1461, 1465, 1505, 1662, 1801, 1822, 1898, 2047, 2095, 2096, 2128, 2145]. Chips [426, 1871, 1923, 1980]. Chips-First [426]. Choice [1893, 2146]. Choices [1881]. Choosing [1230]. Cipher [113]. Circuit [2014, 2100]. Circuits [58, 143, 1686]. Citation [1982]. Cite [1600]. Citing [1478]. claim [810]. Claims [1154, 1910, 1943]. Class [636, 878, 1631]. Classical [825]. Classification [146, 538]. Classifications [11]. Classifying [200]. Classroom [2058, 2060]. Client [623, 801, 804, 805, 892, 914, 1031, 1133, 1388, 1771]. Client-Server [914, 1031, 1133, 1388, 1771]. Client/Server [623, 801, 804, 805]. Clients [1981]. Clinic [999]. Clipper [1148]. Clippinger [1492]. Clock [82, 1558]. Clouds [196]. Cluster [490, 1047, 2008]. Clustering [1784, 2033]. CMM [1526]. co [854]. co-inventor [854]. COBOL [1299, 1330, 2145]. Cocke [2101]. Cocoon [1320]. Code [25, 97, 194, 292, 605, 685, 959, 1321, 1388, 1399, 1489, 1490, 1515, 1724, 1812, 2084, 2145]. Code-Breaking [194]. Code-Free [1321]. codes [277]. Codesign [386–388, 539, 540, 554, 770, 1682, 1742]. CodeWizard [1401]. Codie [1933]. Coding [57, 2006]. Coed [238]. coexist [333]. Cognitive [605]. Coherence [40, 43, 51]. coherent [46]. Coldfire [2039]. collaboration [817]. Collaborative [621, 652, 658, 1397, 1744, 2056]. Collaboratories [1174]. Collapse [722, 2130]. Collections [1105]. Collective [972]. Color [367]. Colorado [498]. Colossus [1179, 1829]. COM [915, 1837, 1913, 2016, 2144]. COM-CORBA [1913]. Combat [2018]. Combining [1533]. Comdex [1316, 1863]. Come [906, 915, 1444, 1708]. Comes [1813, 1957]. Coming [1918, 2120]. Command [1995]. Commander [1092]. Commandments [806, 1043]. Comments [37, 38, 50]. Commerce [1095, 1345, 1379, 1380, 1388, 1696, 1865, 1908, 1923, 2087]. Commercial [35, 54, 1724]. Committee [300, 356, 639, 896, 1250, 1593]. Common [100, 1869]. CommonPoint [791]. Communication [116, 170, 487, 540, 735, 972, 1354, 1646, 1669, 1700, 1701, 1825, 1964]. Communications [11, 553, 701, 1037, 1162, 1188, 2014]. Communicator [1039]. Community [19, 37, 956, 1040]. Company [1505]. Comparative [114, 1166]. Comparing [517, 1362]. comparison [105]. Compete [2047]. competence [1329]. Competition [1320, 1885, 2111]. Compilation [1686]. Compiler [42, 211, 212, 279, 911, 1265, 1536, 1812, 1955, 2014, 2039]. Compiler-Directed [42]. Compilers [1398, 1496, 1538]. Compiling [97]. complete [1254]. Complex [654, 849, 899, 1787, 2068]. Complexities [2124]. Complexity [348, 685, 1355, 1436, 1559, 1602, 1750, 1883]. compliance [892]. Compliant [1615]. complies [162]. Component [790, 1496, 1610, 1645, 1653, 1688, 1700, 1724, 1791, 1837, 1867, 1918, 1937, 1962, 1966, 1988, 2000, 2002, 2003, 2007, 2014, 2016, 2067, 2072, 2122, 2147]. Component-Based [1645, 1688, 2000, 2003].

Component-Savvy [1700]. Components [117, 1326, 1581, 1678, 1690, 1692, 1724, 1867, 1918, 2001, 2014, 2044, 2054, 2064]. Compose [278]. Composer [159]. Composing [161]. Composite [1325]. Composition [155, 160, 189]. Comprehension [869]. Comprehensive [93]. Compressing [1497]. Compression [314]. Compromise [1275, 1299]. CompuServe [1505]. Computation [98, 180, 345, 513]. Computational [254, 344, 348, 511, 514, 516, 1947]. Computations [22, 139, 655]. compute [490]. Computer [1, 7, 18, 75, 99, 102, 119, 149– 151, 154, 160, 176, 178, 182, 215, 218, 229, 259, 261, 374, 375, 414, 415, 500, 515, 572, 616, 617, 705, 706, 727, 743, 761, 777, 795, 810, 828, 834, 860, 870, 879, 882, 894, 897, 924, 929, 930, 958, 961, 975, 982, 990, 993, 999, 1017, 1041, 1065, 1094, 1098, 1112, 1127, 1129, 1158, 1161, 1162, 1181, 1203, 1226, 1247, 1267, 1273, 1288, 1359, 1404, 1468, 1482, 1491, 1601, 1637, 1756, 1802, 1815, 1844, 1852, 1858, 1909, 1939, 1985, 2052, 2057, 2058, 2084, 2087, 2111, 2140]. Computer [19, 37, 38, 118, 156, 174, 177, 181, 219, 224, 225, 228, 256, 271, 275, 311, 325, 338, 355, 371, 382, 447, 477, 478, 527, 854, 893, 897, 899, 926, 998, 1022, 1101, 1116, 1136, 1159, 1179, 1222, 1250, 1257, 1281, 1313, 1419, 1429, 1436, 1441, 1468, 1540, 1542, 1665, 1681, 1693, 1798, 1828, 1830, 1853, 1859, 1860, 1885, 1933, 1958, 1985, 2009, 2010, 2034, 2058, 2059, 2084, 2087, 2102, 2112, 2139, 2140]. Computer-Aided [160, 999, 1939]. Computer-Based [99, 119, 527, 1288, 1828]. Computer-controlled [975]. Computer-Generated [149, 156, 961]. Computer-Human [1127]. Computer-Integrated [414]. computer-related [743, 761]. Computer-Supported [375, 616, 617]. Computer-System [828]. Computer-Vision-Related [224, 225].

Computers [54, 57, 428, 436, 450, 477, 589, 852, 972, 996, 1088, 1148, 1217, 1310, 1373, 1527, 1588, 1650, 1790, 2063, 2070]. computic [990]. Computing [2, 53, 61, 62, 103, 164– 166, 184, 193, 198, 214, 272, 327, 340–343, 346– 350, 362, 393, 429, 457, 460, 503, 518, 526, 550, 552, 603, 648, 672, 717, 722, 767, 773, 801, 956, 965, 973, 976, 985, 1021, 1026, 1044, 1071, 1100, 1130, 1133, 1145, 1167, 1173, 1176, 1213, 1216, 1233, 1255, 1274, 1300, 1303, 1320, 1360, 1388, 1402, 1498, 1507, 1531, 1534, 1556, 1584– 1587, 1590, 1664, 1738, 1744, 1778– 1780, 1803, 1864, 1878, 1903, 1951, 1957, 1965, 1998, 2006, 2008, 2051, 2056, 2111]. Computoons [1088]. Concept [217, 1681]. Concepts [53, 130, 363, 364, 618, 721]. Conceptual [187, 382, 636, 1006]. Concern [1076, 1850]. Concerns [2048]. Concinity [1388]. Concurrency [31, 1790]. Concurrent [81, 274, 308, 359, 374– 377, 379, 381, 384, 387, 412, 827]. Conducting [1315]. Conductor [154]. Conducts [1778]. Conference [1064, 1985, 2005, 2048]. Conferences [391, 405, 438, 452, 469, 507, 532, 545, 562, 610, 627]. Conferencing [76, 192]. Configurable [787, 1498, 1534]. Configuration [571]. Configuring [247, 248]. Conflict [1575, 1613]. Conformational [345]. Confusion [648, 1505]. Congestion [2128]. Connection [222, 877, 1860, 1885, 1909, 1933, 1958, 1985, 2009, 2034, 2058, 2084, 2111, 2140]. Connectionist [1155]. Connectivity [1230, 2024]. conquer [1190]. Consensus [276, 1575]. Consequence [1641]. consideration [431]. Considerations [427]. Consistency [41, 1263, 1738]. Consoles [1729]. Consortium [1937]. Constrained [413]. Constrained-Latency [413]. Constraint [267, 381, 2030]. Constraint-Based [381, 2030]. Constraints [1723]. Construction [850]. Consultant [1043, 1855]. consultation [817]. Consumable [1083]. Consumer [1310, 2111]. Container [592]. Containment [1995]. Contender [2008]. Content [886, 887, 890, 1134, 1395, 1570]. Content-Based [886, 890, 1134, 1570]. Contention [1327]. Context} [273]. Contingency [2073]. Continues [1923]. Continuing [1681]. Continuous [190]. Continuously [1642]. Contract [332, 1063, 1293, 2001]. Contractors [750]. contrasting [517]. Contribution [2140]. Control [34, 57, 223, 262, 266, 306, 307, 523, 1028, 1692, 1751, 1791, 2031]. controlled [975]. Controlling [473, 1359]. Controls [1791, 2014, 2118]. Controversy [1357, 1405, 1778]. Conventional [275, 330, 331]. Convergence [734, 1093, 1505, 1714, 1945]. Converging [91, 1550]. Conversational [1152]. Conversations [1445]. Conversion [63, 290]. Converting [605, 959, 2145]. convex [1261]. COOL [667]. Cooperate [2047]. Cooperative [473, 616, 617, 619]. Coordinated [188, 306, 307]. coordinating [12, 623]. Coordination [801]. Cope [2070]. Copy [1771]. Copying [2070]. Copyright [1267]. Copyrighted [1094]. Copyrights [366, 1299]. CORBA [915, 1351, 1400, 1482, 1791, 1913, 1966, 2004, 2072, 2144]. CORBA-Based [1400]. Core [1239, 1571, 1980]. Coronary [124]. Corporate [195, 1136]. Corporation [1854, 1899]. Corps [1949]. Correcting [411]. Correction [51, 231, 232, 1870, 2069, 2150, 2151]. correctly [743, 761]. Correctness [89]. Cospecification [554]. Cost [773, 1318, 1681, 1689, 1812, 2093]. Cost-Effective [773]. costing [1114]. Costs [1452]. COTS [1242, 1688, 1689, 1691, 1692, 1818, 1866, 1895]. Could [956, 1299, 1345, 1443, 1482, 2129]. Council [1228]. Counterpoint [141, 323, 324, 1075, 1076]. Counting [2095]. Countries [1454]. Counts [2122]. Coupled [39]. Court [1016, 1606]. cousin [1160]. Cover [1739, 1782–1784, 1804]. Coverage [688, 1818]. Cowboy [1655]. Cram [258]. crash [1235]. Cray [1246, 1571, 2102, 2140]. Crazy [1839]. Create [1389]. Created [1973]. Creates [1972]. Creating [654, 785, 1131, 2014, 2088]. Crediting [1920]. creep [1137]. Crisis [362, 1792, 2098]. Criteria [1603, 1649, 1839]. Critic [748, 764, 782, 799, 823, 841, 866, 943, 967, 993, 1022, 1046, 1072, 1101, 1124, 1147, 1169, 1170, 1190, 1235, 1295, 1316, 1332, 1364, 1393, 1449, 1476, 1501, 1524, 1550, 1579, 1602, 1626, 1658, 1680, 1706, 1730, 1731, 1753, 1774, 1797, 1817, 1843, 1868, 1894, 1919, 1942, 1968, 1992, 2018, 2043, 2068, 2092, 2123, 2149]. Critical [1242, 1374, 1643, 1974]. Crossroads [1319]. CRT [323, 324]. Crucial [1299]. Crying [2049]. cryptochip [1037]. Cryptographic [1760]. Cryptography [113, 1140, 1275, 1915, 2040]. Cryptologist [1829]. Crystal [1770]. CS [390, 402, 419, 433, 448, 466, 492, 529, 544, 558, 575, 606, 625, 662, 673, 690, 851, 925, 952, 966, 1003, 1032, 1049, 1193, 1201, 1266, 1307, 1356, 1384, 1418, 1441, 1468, 1492, 1514, 1541, 1571, 1593, 1615, 1649, 1671, 1693, 1718– 1721, 1745, 1768, 1829, 1933, 2021]. CSCW [618, 622]. CSS [1728]. CSTB [832]. Cultural [1913]. Cures [1582]. Curious [2101]. Current [156, 236, 792, 2108]. Curricula [198]. Curriculum [973, 1133, 1823, 2058, 2060, 2070]. Custom [767]. Customize [1865]. Customized [654]. Customizing [693, 707, 1305]. Cuts [1710]. Cyberfarce [1656]. Cybermyths [1627]. Cyberspace [1299]. Cybersquare [1508, 1665, 1737, 1781, 1802, 1902, 2100]. Cyberwar [1148]. Cycle [233, 910, 960, 1452]. Cyclic [143]. Cytoskeleton [345].

D [215, 301, 302, 326, 358, 1406, 2087]. Dangerous [1681]. DAP [1443]. Dart [2014]. Darwinism [1170]. Dash [237]. Data [15, 116, 191, 197, 200, 301, 314, 401, 651, 654, 751, 891, 947, 1182, 1218, 1409, 1453, 1512, 1537, 1589, 1771, 1790, 1791, 1805, 1834, 1991, 2020, 2028, 2030, 2031, 2039, 2124, 2133]. Data-Driven [751]. Data-Intensive [1512]. Data-Modeling [1790]. Data-Transfer [191]. Database [88, 130, 171, 199, 200, 207, 336, 396, 444, 721, 889, 920, 921, 1604, 1862]. Databases [8, 91, 284, 751, 918–921, 1652, 1977, 2029]. Dataflow [143, 668]. Dates [1322, 1365]. Dave [1913]. David [2139]. Davidson [1384]. Dawn [1481]. dawning [968]. Days [1450]. DCI [1829]. DDM [320, 321]. Dead [2043]. Deadlocking [1335]. Deadlocks [1372]. Dealing [1322, 1365]. Debat [1681]. Debate [1079, 1732]. Debugging [463, 1697]. Debunking [1427, 1580, 1627]. Debut [1554]. Debuts [1429]. Decade [175]. Decades [174, 1224]. December [1255]. Decent [1749]. Decision [622, 686, 1054, 1903]. Decision-Support [622]. deck [873]. Declaration [1996]. Decompiler [1519]. Deep [1484, 1640, 2100]. Deep-Space [1640]. Deep-Submicron [2100]. Deering [1921]. defect [1096]. defect-removal [1096]. defense [834]. Defies [1317]. definable [786]. Defined [1673]. Defining [493, 568, 1916]. Definition [55]. Definitional [319]. Déjà [2124]. Delay [1634, 1662]. Delayed [736, 1554]. Delivery [1837]. Delphi [902]. Demand [1066, 1299]. Demonstrated [2047]. Demonstration [733]. Demonstrational [310]. Denied [736]. Density [426]. Depend [1987]. Dependability [80, 828]. Dependencies [204]. Dependency [570]. Description [162]. Descriptor [1490]. Descriptor-Based [1490]. Deserved [1580]. Deserves [1421]. Design [9, 17, 21, 59, 107, 135, 198, 238, 246, 330– 332, 377–379, 382, 414, 428, 431, 450, 525, 569, 588, 637, 770, 771, 776, 814, 826, 896, 953, 1027, 1031, 1080, 1127, 1291, 1293, 1345, 1350, 1438, 1446, 1459, 1498, 1525, 1557– 1564, 1623, 1630, 1647, 1654, 1666, 1696, 1724, 1740, 1753, 1756, 1760, 1804, 1877, 1885, 1907, 2014, 2022, 2063, 2091, 2100, 2105, 2113, 2117]. designed [1347]. Designer [431]. Designers [131, 1296, 1777]. Designing [280, 281, 501, 738, 1999]. Designs [1650, 2111, 2145]. Designworld [384]. Desktop [191, 626, 866, 1984]. destruction [240]. Details [1723]. Detecting [1528]. Detection [1743, 1988]. Determining [774]. Determinism [1625]. Developer [1197]. Developers [132, 726, 1292, 1497, 1658, 1869, 2121]. Developing [542, 936, 1329, 2054]. Development [208, 395, 425, 728, 1004, 1082, 1097, 1172, 1194, 1321, 1388, 1428, 1497, 1518, 1519, 1545, 1587, 1636, 1653, 1674, 1688, 1697– 1699, 1724, 1771, 1791, 1798, 1812, 1834– 1836, 1889, 1914, 1937, 1962, 1981, 1988, 2000, 2014, 2038, 2039, 2055, 2063, 2064, 2081, 2087, 2090, 2118, 2128, 2145]. Developments [28, 179, 1633]. Develops [2072]. Device [1973]. Devices [975, 1304, 1856]. Diab [1771, 1812, 2039]. Diagnose [124]. Diagnosing [125]. Diagnosis [1613]. Diagramming [1988]. Diagrams [298]. Dialog [144]. Dialogue [1942]. Diary [1855, 1875, 1900, 1925, 1949, 1975, 1996, 2024, 2049, 2074, 2098, 2130]. Dicaf [1614]. Dice [380]. Dickens [1970]. Dictate [1625]. Dictionary [1159]. Dies [854, 1225, 1820]. difference [148]. Different [715, 1347, 1915, 1970]. Differentiated [1989]. Differs [1707]. Difficulties [1974]. Difficulty [685]. Diffie [1696]. Digital [186, 234, 357, 530, 1104, 1106–1109, 1163, 1170, 1183, 1200, 1252, 1299, 1546, 1552, 1554, 1570, 1587, 1695, 1696, 1770, 1879, 1882– 1884, 1982, 2014]. Digitopolis [1476]. Digits [292]. Dignity [1625]. Dilemma [1979]. Dimension [1316]. Dimensional [209, 247, 248, 277]. Dimensions [328, 329]. dining [360]. dinner [360]. Direct [310, 397, 1780]. Directed [42]. Direction [1802]. Directions [2108]. Directories [2149]. Directory [43, 46, 47, 51, 1836]. Directory-Based [43, 51]. Directory-Enabled [1836]. Dirty [1925]. Disabled [1662]. Disappear [1843]. discipline [362]. disclosing [819]. Discovery [497, 498]. Discrete [143]. Discrete-Event [143]. discussed [26]. Discusses [1299]. discussion [1123]. Disk [585–587, 1554, 1670]. Disks [1454]. Display [309, 323, 324, 477, 947]. Displaying [702]. Dispute [1684]. Dissecting [1790]. Distance [2133]. Distributed [4, 14, 29, 32– 34, 46, 47, 50, 75, 82, 123, 136, 143, 164– 171, 181, 189, 196, 199, 203, 210, 216, 231, 232, 234, 262, 267, 276, 280, 281, 283, 285, 365, 393, 460, 462, 488, 499, 501, 623, 652, 801, 911, 914, 915, 946, 948, 1029, 1109, 1130, 1176, 1388, 1610, 1614, 1778, 1883, 1884, 1913, 2002, 2003, 2007, 2055, 2134, 2145]. Distributed-directory [46, 47]. distributed-memory [911]. Distributing [365]. Distribution [1399]. Distributive [1834]. Diverging [91]. Diverse [203, 1105]. diversity [389]. Divine [1170]. DLLs [2145]. DM [144]. DM-Dialog [144]. Do [3, 96, 111, 1373, 1774, 1955]. Document [142, 286, 287, 293–295]. Documentation [1494, 1798]. Documents [187, 298]. Does [1625, 1695]. Doesn’t [1910]. Doing [1174, 1181, 1381, 1723, 1811, 1900]. Domain [785, 787, 1319, 1357, 1405, 1421, 1454, 1554, 1633, 1736, 2071, 2096]. Domain-Name [1633, 1736, 2096]. Domain-Oriented [785]. Domain-Specific [787]. Don’t [1073, 1084, 1339, 1629, 1920, 2149]. Doorbells [1778]. doors [300]. Dot [702]. double [833]. double-edged [833]. Down [1369, 1505, 1925, 1975, 2063]. Draft [1649, 1898]. Draw [68]. drawbacks [1347]. Drawing [291]. Drawings [290]. Draws [1407]. Dreams [1312]. Drive [585, 853, 1505, 1822, 1869, 1981, 2047]. Driven [751, 826, 2013]. Driver [1681]. Drives [1821]. Driving [847]. DRS [294, 295]. DSP [425, 769, 1603, 1741, 1791]. DSPCard [1062]. DSPCard-C31 [1062]. DSPs [1497]. DT3155 [1011]. Duel [1898, 1923]. Dump [1898]. During [379, 869]. DVD [1505]. Dynamic [223, 1326, 1791, 2033]. Dynamical [257]. Dynamically [1156]. Dynamics [515].

E-Commerce

[1345, 1379, 1388, 1696, 1865, 1923, 2087]. e-mail [957, 1016, 1345, 1898]. E-pubs [1359]. E.-R [359]. Earlier [1580]. Early [427, 431, 1487, 1636, 1964]. Earn [1384, 1745]. Earnings [1317]. Earth [655]. Ease [2128]. Easier [1431]. Eastern [1693]. Easy [445]. Easy-to-Use [445]. EC [133]. ECBS99 [1858]. Eckert [854]. Eco [1380]. Ecology [1598]. Economic [1681]. Economics [285, 661, 1009, 1790]. Economy [1583, 1680]. EDA [1583]. Edge [1675, 1724]. edged [833]. Edison [1580]. Editing [2039]. Edition [978, 1519]. Editor [28, 39, 52, 61, 102, 119, 134, 142, 149, 164, 175, 199, 286, 341, 374, 392, 409, 423, 424, 441, 455, 456, 472, 484, 510, 511, 550, 566, 581, 583, 598, 613, 616, 631, 646, 647, 649, 665, 680, 682, 698, 714, 716, 731, 747, 766, 784, 844, 881, 886, 905, 942, 946, 965, 996, 1045, 1050, 1104, 1168, 1169, 1216, 1258, 1273, 1333–1339, 1356, 1394, 1403, 1436, 1593, 1631, 1844, 1937, 2105, 2112]. Editor-in [392, 714]. Editor-in-Chief [409, 423, 566, 581, 598, 613, 631, 646, 665, 680, 698, 731, 747, 965, 1045, 1356, 1593, 2112]. Editorial [1540]. Editors [218, 327, 1194, 1238, 1348, 1460, 1904, 2028]. edn [301, 314]. edn. [259]. Eds [284]. Educating [1510]. Education [2, 669, 792, 809, 997, 1085, 1110, 1118, 1249, 1328, 1568, 1773, 1943]. Educational [2052, 2054]. Edward [1384]. EEG [315, 316]. Effect [636, 1197, 1383, 1601].

Effective [227, 773]. Effectiveness [2083]. Effects [2106]. efficiency [1096]. Efficient [446, 1808]. Effort [1615, 1924]. Effortless [1799]. Efforts [12, 71, 1514, 1899, 2140]. EIC [536]. EiffelBase [1791]. Einstein [2077, 2135]. EJB [1988]. Elect [1541, 1819, 2058, 2059]. Election [504, 897, 980, 1032, 1266, 1308, 1468, 1593, 1721, 1745, 1768, 2057]. Elections [1158, 1201]. Electrical [1630]. Electrical-Engineering [1630]. Electromechanical [382]. Electronic [240, 317, 318, 424, 432, 1017, 1248, 1429, 1441, 1456, 1588, 1908, 2113]. Electronics [1310, 1766]. Element [513]. Elicitation [1154]. Elizabeth [1492]. Elvis [1656]. Embedded [59, 1309, 1358, 1446, 1545, 1698, 1740, 1753, 1771, 1812, 1835, 1891, 1962, 1980, 2039, 2064, 2087]. Embedded-Centric [2064]. Embedded-Core-Based [1980]. Embedded-Microprocessor [1740]. Embedded-System [1753]. Embrace [1335]. Embracing [2082]. Emerge [1482]. Emergence [1610, 1998]. Emerges [1981]. Emerging [401, 790, 1547, 1800]. Empirical [2077]. Employment [1112, 1553, 1848]. Emulation [2014]. Enabled [1567, 1836]. Enabling [1095]. Encalming [1476]. Encoding [277, 1347, 1749, 1798]. Encryption [26, 1299, 1345, 1454, 1505, 1554, 1710, 1755, 1764, 1778, 1898, 2022]. End [895, 1710, 1790, 1822]. End-user [895]. Endless [1520]. Endorses [2059]. Energy [461]. Engine [768, 1757, 1780]. Engineer [1706]. Engineered [1243]. Engineering [99, 126, 180, 183, 233, 239, 354, 373– 377, 381, 384, 387, 417, 493, 511, 527, 615, 689, 821, 899, 1065, 1083, 1084, 1098, 1207, 1219, 1228, 1288, 1313, 1360, 1446, 1449, 1498, 1510, 1515, 1568, 1598, 1630, 1637, 1649, 1654, 1702, 1745, 1750, 1775, 1793, 1828, 1838, 1858, 1891, 1933, 1936, 1939, 1952, 1958, 1991, 2041, 2077, 2078, 2084, 2091, 2102]. Engineers [241, 275, 880, 1083, 1479, 1849]. Engines [705, 1362]. Enhance [2004]. ENIAC [854]. Enough [623, 1409, 1447, 1796]. Ensure [1818]. ensuring [671]. Enterprise [994, 1402, 1471, 1519, 1611, 1724, 1834, 1937, 1965, 1988]. Enterprise-Wide [1834]. Enterprising [1619]. Enters [937, 1684]. Entity [15]. Entity-Relationship [15]. Entrants [2084]. Entrepreneur [1829, 2111]. Entrepreneurs [1920]. Entropy [636]. Entrus [1358]. Entrust [1988]. Entry [294, 295]. Environment [6, 12, 123, 192, 204, 228, 382, 430, 463, 486, 489, 568, 571, 733, 769, 787, 826, 829, 847– 849, 1698, 1766, 1771, 1787, 1835, 1937, 1962, 2118, 2145]. Environments [400, 634, 789, 846, 1035, 1644, 2137]. Envisioning [1520]. EPA [658, 861]. epilepsy [315, 316]. Epilog [315, 316]. Equal [1603]. Equations [2135]. Equipment [2107]. Era [1226, 1464, 1552, 1597, 1724, 1734]. ERMA [1514]. Error [57, 59, 1988, 2118]. Error-Control [57]. Error-Detection [1988]. Errors [2070]. Essays [1216]. Establishing [49]. Estimates [411]. Estimating [950, 1066]. Estimation [56, 1812]. Estonia [500]. Ethernet [1023, 1387, 1634, 1661, 1723]. Etherphone [192]. ethical [19, 37, 141]. Ethics [118, 1515, 2084]. ETrust [1319]. Euro [1662]. Europe [1693]. European [1299, 1994]. Europeans [1662]. Europe’s [1606]. Evaluate [670]. Evaluating [689, 2083]. Evaluation [97, 103, 114, 246, 379, 569, 868, 946, 970, 1129]. Event [143, 461, 1860]. Every [2122]. Everyone [1943]. Everything [706, 1184, 1900]. everywhere [899]. Evolution [128, 396, 650, 869, 1218, 1229, 1599, 1807, 1953, 2016]. Evolutionary [2080]. Evolving [176, 362, 692, 832, 1088, 1206, 1491, 1631]. examining [743, 761]. Example [444]. Example-Based [444]. Examples [363, 364, 475]. exchange [401, 958]. Excogitate [2067]. Executable [1395, 1410]. Execution [489, 1564, 1780]. Executive [1593]. exemplar [1261]. Exhaustive [194]. Existing [1432]. Expanding [531, 1873]. Expansion [1421]. expect [1161]. Expensive [2068]. Experiment [376, 1665]. Experimental [102, 144, 721, 1664]. Experiments [138, 1681]. Expert [16, 114, 125, 152, 325, 1933, 2042, 2070]. Expert-System [16]. Expertise [1331, 2034, 2140]. Explained [1359, 1696]. Explanations [188]. Exploit [2067]. Exploiting [226, 567]. Exploration [2133]. Explore [151, 2067]. Explorer [1343, 2022]. Exploring [1585]. Export [542, 1345, 1505]. Exports [1278]. Exposing [2026]. Expressive [150, 152]. Extend [1347, 1393, 2073]. Extended [15]. Extending [69, 388, 526, 1984, 1991]. Extensions [909, 1778, 2132]. Extra [1537]. Extranet [1699]. Extranets [1375]. Extreme [2082]. extremely [26]. Eye [120, 306, 307, 477, 478, 1477]. eye-gaze [478].

F [325, 1492]. Face [1776]. Faces [276, 1115, 1513, 1605, 1974]. Facilities [170, 1413]. Facility [591]. Facing [1922]. Factor [1402, 1873]. Factors [1446]. Factory [1580, 1962, 2145]. Failure [202, 793, 1238, 1346, 1677, 1710, 1775]. Failure-Resilient [202]. Failures [137, 1743, 2129]. Fair [1745, 1933]. Fall [1985]. Family [2087]. far [707]. fashioned [260]. Fast [1227, 1345, 1868, 2068, 2130]. Fast-Lane [1868]. fast-track [1227]. Faster [1276, 1801]. Fault [4, 5, 52– 56, 58, 81, 82, 280, 281, 551, 828, 985, 1348– 1350, 1352–1354, 1548, 1970, 2106]. Fault-Tolerant [5, 52, 53, 56, 82, 280, 281, 551, 985, 1350, 1354]. Faults [1333]. FCC [1345]. FDDI [1354]. FDDI-Based [1354]. Fear [846]. Fears [1299]. Feature [484, 485, 488, 1511, 1512, 1533– 1538, 1739, 1782–1787, 1804]. Feature-Interaction [484, 488]. Features [650, 1742, 2044]. February [1021, 1123]. Federal [807, 1974]. Federated [1880]. Federating [1105]. Feedback [129, 306, 307]. Feel [1396, 2088]. Fejer [215]. Fellow [1289]. Fellows [1007, 1307, 1593, 1885]. Fernbach [1593]. Few [1070, 2070]. Fiber [116, 1480, 2145]. Fiber-optic [2145]. Fibre [182, 877]. Fibre-Optic [182]. Field [516, 704, 1795]. fields [26]. fifth [1232]. Fifty [1216]. Fight [1924]. Fighting [1436]. File [29, 1724]. files [1837]. Filling [1359]. fills [437]. Film [1299]. Finalist [1933]. Finally [1813]. Financial [1719, 1932, 2010]. Financially [2010]. Finding [601, 741]. Finite [513]. Finite-Element [513]. Fire [1407]. Firewall [1073, 1812]. Firewire [1813]. firm [982]. Firms [1505, 1529]. First [426, 1200, 1300, 1391, 1778, 1833]. Flag [719]. Flag-Oriented [719]. Flap [1923]. Flash [1698]. Flat [308]. flavors [860]. Flaw [1299, 1343, 2022]. Fletcher [1225]. Flex [140]. Flexible [140, 198, 267, 309, 1727]. flight [671]. Floating [1345]. Floating-Point [1345]. Flow [244, 245, 464, 512]. Flowers [1829]. Flows [513]. Fluid [464, 512]. Fluke [2118]. Focus [617, 1553, 1995, 2044]. Folding [2117]. Follow [1156]. Follow-up [1156]. Follows [154]. Food [2130]. Force [1183, 2087]. Forecast [2097]. Forecasting [1697]. Forefront [1946]. Foreign [1482]. Forest [1411]. forestall [240]. Forget [1920]. Form [786, 1981]. Form/formula [786]. Formal [73, 485, 487, 771, 806, 874, 1074– 1082, 1085, 1301, 1333]. Formalize [1075]. Formally [412]. Formed [1299]. formula [150, 786]. Fortieth [1340]. Fortran [1340]. Fortune [1814]. Forum [838, 855, 876, 896, 930, 961, 985, 1015, 1035, 1067, 1098, 1117, 1139, 1162, 1183, 1205, 1228, 1250, 1270, 1309, 1355, 1385, 1440, 1467, 1491, 1513, 1570, 1592, 1616, 1650, 1718, 1766, 1788, 1809, 1828, 1859, 1957, 1973, 1994, 2008, 2113]. Forward [1895]. Foundations [215, 923]. Founder [1829]. Four [174, 1794]. Fourteen [657]. Fourth [1463]. Fourth-Generation [1463]. Fractals [1497]. frame [1011]. Framework [74, 395, 459, 473, 539, 820, 970, 1151, 1198, 1275, 1397, 1791, 1796, 1835, 2053]. Frameworks [594, 607, 1197, 1305, 1877]. Francis [1492]. Fray [1345, 1684, 1778]. Free [346, 347, 1321, 1587, 1591, 1671, 1791, 1798, 1846, 2039, 2118, 2124]. Free-Space [1587, 1591]. Freeware [1776, 1846, 1869]. frequency [26]. Fridays [1450]. Friendly [1869, 1994]. Front [1275, 1297, 1317, 1343, 1374, 1430, 1452, 1480, 1503, 1527]. Frontier [1520]. frontiers [1118]. Fronts [1849]. FSN [813]. FTC [1736]. Fugue [153]. Fukushima [2005]. Fulfillment [1939]. Fun [1659]. Function [123, 959, 2063, 2076]. Functional [89, 153, 272, 378]. Functional-link [272]. functionalities [272]. functionality [904]. Fundamental [53, 1070]. Fundamentals [25]. Funding [807, 854, 1924]. further [638]. Fusion [197]. Fuss [1989]. Futile [1393]. Future [182, 183, 208, 214, 235, 386, 436, 877, 1102, 1221, 1272, 1406, 1556, 1568, 1576, 1581, 1718, 1818, 1859, 1864, 1865, 2027, 2058, 2060]. FutureBusiness [943]. Fuzzy [124, 257, 269, 270, 1306, 1359, 1913].

G [274, 314]. Gadget [1566]. Gains [1457]. Galaxy [167]. Gallager [301]. Galorath [1812]. Game [715, 2042]. Games [1157]. Gang [967]. Ganging [1657]. Gaps [132, 809, 1359]. Garnet [93]. Gary [1978]. Gate [1291]. Gate-Array [1291]. GATT [795]. gaze [477, 478]. Gbit [1634]. Gelernter [302]. Gen [2039]. General [524, 1178, 1533]. General-Purpose [524, 1178, 1533]. Generated [149, 156, 961, 1156]. Generation [188, 213, 789, 1268, 1463, 1529, 1801, 1812, 1921, 2020, 2108, 2127, 2145]. Generators [113]. Genetic [633, 634, 718, 1438, 2013]. Genetic-Algorithm [634]. Genetics [2044]. Genitor [1834]. Geographic [1956]. Geographically [1108]. Geometric [222]. Geometry [656]. George [1829]. Get [478, 795, 1073, 2149]. Gets [727, 1554, 1994]. Getting [693, 1136, 1239, 1925, 2088]. Giants [1445]. Gigabit [78, 1023, 1387, 1661]. Girl [1898]. GIS [1260]. Give [1822]. Giveaway [1603]. Gives [1634]. Global [300, 1052, 1303, 1697, 1757, 1899, 1952, 1992, 2023]. Glory [722]. Glyphmaker [654]. Go [1141, 1899, 2105]. GoAhead [1937, 2039]. Goal [308]. goes [875, 1185]. Golden [1571]. gonna [795]. Good [694, 834, 1365, 1447, 1680, 1796, 1829, 2070, 2147, 2151]. Goode [1745, 2034]. Gorden [1541]. Gordian [1791]. Gordon [385, 557, 756, 1005, 1266, 1286, 1569]. Government [822, 1840]. Governmental [1798]. Governors [897, 1468, 1649, 1721]. Grab [1675]. grabber [1011]. Grace [1059, 1505]. Grand [550]. Grant [1820]. granularity [155]. Grape [769]. Grape-II [769]. Graph [22, 1937]. Graphical [15, 93, 444]. Graphics [68, 337, 476, 904, 1250, 1448, 1714, 1791, 1962]. Gray [1655]. Great [2119]. Green [1521, 1724, 1756, 1798, 1835]. greener [1269]. GreenTree [1791]. Grid [2133]. Ground [929, 1778, 2096]. Group [11, 620, 1519, 1634, 1791, 1898, 2064, 2128]. Groups [11, 1822, 1898, 1923, 1924]. Groupware [738, 884, 1128]. Grove [1342]. Grown [1299]. Grows [2022]. Growth [1505, 1719]. GTE [1906]. Guest [28, 39, 52, 61, 102, 119, 134, 149, 164, 175, 199, 218, 286, 327, 341, 374, 424, 456, 484, 511, 550, 583, 616, 649, 682, 716, 766, 784, 844, 886, 946, 996, 1050, 1104, 1194, 1216, 1238, 1258, 1348, 1436, 1460, 1904, 2028, 2105]. GUI [1497, 1723, 1835, 1947, 1988]. Guidance [1412]. guide [1324]. Gupta [284].

H [259, 283, 1733, 2022]. H-1B [1733, 2022]. Hackers [1505]. Haddad [357]. half [638]. half-truths [638]. Hamlet [1472]. Hamming [1615]. Hampered [1924]. Hand [306, 307]. Hand-Eye [306, 307]. handbook [368, 369]. Handed [845]. Handsets [1981]. Handwriting [1306]. Handwritten [292]. Hans [1492]. Hard [835, 1314]. Hardware [21, 55, 268, 357, 386– 388, 539, 555, 637, 1080, 1484, 1497, 1653, 1660, 1674, 1682, 1686, 1698, 1724, 1771, 1791, 1812, 1835, 1889, 1914, 2014, 2145]. Hardware- [55]. Hardware-Based [1660]. Hardware-Software [386–388, 555, 1484, 1682]. Hardware/Software [539]. Harlequin [1937]. HARP [159]. HARTS [136]. Hashing [222]. Hasn’t [1454]. Hazards [958]. Headaches [1895]. Headed [672, 805, 867, 919, 942]. Heads [303, 1505]. Health [1700, 1701]. Heat [1505]. Heats [1554]. Heavy [1821]. Hector [109]. Heights [846]. Held [314]. Help [1718, 1724, 1924, 1994]. Henry [1514]. Herd [2040]. Here [1344, 1500]. Heresy [1187]. Heterogeneity [200]. Heterogeneous [199, 201, 206, 456, 457, 459, 460, 462, 463, 489, 888, 976]. Hewlett [2014]. Hierarchical [18, 34, 38, 2033]. Hierarchically [109]. Hierarchies [636]. High [116, 178, 211, 212, 214, 223, 259, 263, 393, 424, 426, 458, 461, 586, 588, 589, 701, 705, 721, 924, 943, 1290, 1351, 1359, 1454, 1554, 1563, 1583, 1591, 1606, 1638, 1639, 1641, 1790, 1804, 1805, 1820, 1978, 2008, 2034, 2096]. High-Assurance [1638, 1639]. High-Availability [178]. High-Capacity [1454, 1554]. High-Consequence [1641]. High-Density [426]. High-End [1790]. High-Energy [461]. High-Level [223, 458, 1359]. High-Performance [211, 212, 214, 259, 424, 586, 589, 1591, 1804, 1978, 2008, 2096]. high-powered [943]. High-Pressure [705]. High-Reliability [586]. High-Speed [116, 393, 588, 701, 721, 1606, 1805]. High-Tech [1290, 1583, 1820]. Higher [122, 1608, 2022]. Higher-Level [1608]. Highlights [2048]. Highly [29, 93, 110, 115]. Highs [1751]. Hills [1724, 1835]. Hintz [358]. HiProf [1388]. hire [1016]. Hiring [1390]. History [60, 617, 1021, 1044, 1059, 1071, 1100, 1145, 1167, 1213, 1222, 1233, 1255, 1648, 2144]. Hit [1741]. Hodson [325]. Holberton [1492]. Hold [1943]. Holodeck [1472]. Holographic [1589]. Holy [1626]. Home [1757, 1813, 1896, 1970, 2120]. Honor [1944]. Honors [1933]. Hopes [1822]. Hopper [1059]. Horowitz [284]. Horribly [2068]. Host [588, 1669, 1805]. Host-Based [1805]. hostile [753]. Hot [352, 353, 370, 386–388, 398–400, 416, 436, 449, 464, 477, 478, 490, 503, 518, 530, 543, 560, 573, 582, 658, 675, 693, 707, 726, 741, 758, 775, 792, 808, 853, 877, 899, 1008, 1073, 2113]. house [1136]. HP [943, 1261, 1275, 1505, 1684, 2047]. HP/ [1261]. HPCC [658, 1007]. HPCCI [832]. HTML [892, 1331, 1549, 1728, 1822, 1962, 2039]. HTTP [1801]. Huffman [2139]. Human [154, 1054, 1127, 1341, 2013]. Humidifier [1449]. Hurdles [1605]. Hurt [1454, 1895]. Hurts [1604]. Hybrid [429, 1155]. Hybrid-WSI [429]. Hybridization [1050]. Hype [1731, 1841]. Hypercube [88, 443]. Hyperlink [1757]. Hypermedia [162, 172, 1284]. Hypertension [125]. Hypertext [203, 273, 1826, 1970]. HyTime [172].

I-CMM [1526]. I/O

[190, 583, 584, 589–592, 734, 1132, 1898, 2072]. I/O-Intensive [592]. IA [1712]. IA-64 [1712]. IBM [105, 1554, 1637, 1849, 1988, 2104]. IC [1291, 1791, 2106, 2145]. Ichikawa [1745]. ICL [1700]. ICPP [2005]. ICPP-99 [2005]. ICs [1239]. ICSA [1812]. ICSA-Certified [1812]. Idea [2147, 2151]. Identifying [117, 1489, 1532]. IEE [101]. IEEE [71, 174, 851, 897, 925, 952, 966, 1003, 1007, 1049, 1193, 1201, 1307, 1356, 1468, 1541, 1593, 1597, 1621, 1671, 1858, 1860, 1893, 1941, 2021, 2058, 2059, 2120]. IEEE-CS [851, 925, 952, 966, 1003, 1049, 1193]. IEEE-SA [1597]. IEEE-USA [1893, 1941]. IETF [1752, 2128]. If [1332]. Ignore [1403]. II [660, 769, 1056, 1111, 1136, 1179, 1249, 1829]. III [931, 1923]. Image [218, 219, 221, 226, 286, 287, 317, 318, 702, 767, 886, 887, 1088, 1900, 1913, 2013]. Image-Based [317, 318]. Imager [1432]. Images [702, 889, 890]. Imaging [1300, 1962]. immersive [847]. Immigration [1024]. Immunity [1924]. Impact [436, 531, 750, 1222, 1583, 1666, 1860]. Impacts [1879]. Impediments [1081]. Imperfection [1775]. Implantable [121]. Implementation [108, 207, 358, 555, 1508, 1687]. Implementations [1259, 1791]. Implementing [33, 571, 1488]. Implications [541, 556, 653, 825, 2065]. Implicitness [1726]. Importance [991, 1915]. important [1347]. Imprecise [139]. Improve [587, 1658]. Improved [1398]. Improvement [583, 584, 687, 1009, 2079, 2088]. Improving [486, 1027, 1261, 1609, 1681, 1715, 2090]. Imprudent [1426]. IMSL [1011]. Inaugural [2140]. Incentives [1339]. Inclusion [1603]. Increasing [1558, 1739]. Incremental [1781]. Indecency [1034]. Independent [98, 221, 458, 1482, 1720, 1794, 2083]. Index [372, 858, 935, 955, 979, 989, 1012, 1061, 1210, 1539, 1827, 2037, 2070, 2115, 2138, 2142]. Indexing [890, 1982]. Indicators [1128, 1578]. Indigo [2145]. Individualizing [998]. Inductive [1280]. Industrial [550, 552, 1079–1082]. Industrialized [830]. Industry [66, 415, 599, 614, 632, 648, 666, 681, 699, 715, 732, 749, 765, 783, 800, 824, 842, 884, 907, 944, 968, 994, 1024, 1047, 1073, 1102, 1125, 1148, 1188, 1191, 1214, 1236, 1256, 1276, 1298, 1318, 1344, 1375, 1411, 1431, 1432, 1453, 1481, 1504, 1507, 1528, 1552, 1581, 1604, 1632, 1660, 1678, 1682, 1708, 1734, 1776, 1799, 1812, 1820, 1847, 1871, 1896, 1921, 1945, 1971, 2019, 2022, 2045, 2071, 2094, 2126]. inference [269, 270, 1259]. Infinity [2070]. Influence [1301]. Influenced [1878]. Info [883, 982]. InfoAccess [2039]. Infobus [1884]. Information [25, 101, 185, 193, 195, 203, 235, 277, 449, 541, 556, 693, 707, 737, 822, 875, 1013, 1057, 1090, 1091, 1106, 1220, 1282, 1435, 1509, 1566, 1632, 1657, 1744, 1856, 1904–1907, 1931, 1956, 2006]. Informedia [1107]. Infrastructure [362, 449, 1013, 1282, 1947]. infringement [810, 894]. inheritance [1115]. Initiative [658, 1671, 1822]. Injection [828, 1353, 1548]. Inktomi [1724]. Innovation [736, 1072, 1385, 1556, 1878, 1919, 1920]. Innovations [2039]. Innovator [2009, 2111]. Innovators [1944]. Input [67]. Inquiry [1110]. Inquiry-Based [1110]. Inspection [129, 414]. Inspector [2118]. Installation [1937, 2145]. InstallShield [1937]. Instruction [128, 410, 459, 1538]. Instruction-Level [1538]. Instruction-Set [410, 459]. Instrument [120]. Instrumentation [77]. Integrate [1865]. Integrated [116, 373, 414, 430, 1313, 1360, 1446, 1498, 1598, 1638, 1654, 1702, 1750, 1793, 1828, 1838, 1891, 1939, 1991, 2041, 2091]. Integrating [203, 376, 820, 1654, 1750, 1823]. Integration [205, 242, 243, 249, 250, 921, 1724, 1798, 1838, 1866, 2013]. integrity [1240]. Intel [926, 1317, 1345, 1505, 1736, 1790, 1822, 2034, 2093]. Intellectual [706, 860, 879]. Intelligence [257, 268, 352, 353, 488, 1223, 2144]. Intelligent [158, 159, 261, 266, 352, 414, 1107, 1177, 1404, 1507, 1614, 1684]. intensifies [19, 37]. Intensive [197, 592, 1512]. Inter [191]. Inter-Desktop [191]. Interacting [1110]. Interaction [176, 484, 488, 1536]. Interactions [485]. Interactive [93, 132, 186, 512, 655, 656, 733, 813, 814, 818, 923, 1001, 1150, 1152, 1153, 1155, 1185, 1309, 1406, 1765, 1923, 2031, 2087, 2136]. InterBase [489]. Interboard [701]. Interceptors [2004]. interchange [172, 1182]. Interchip [701]. Intercomponent [1646]. Interconnect [426, 1724]. Interconnected [1590]. Interconnection [346, 347, 443]. Interconnection-Free [346, 347]. Interconnects [701, 1591]. Interdatabase [204]. Interdependence [1996]. interest [101]. Interests [91, 1671]. Interface [46, 86, 179, 525, 588, 626, 845, 853, 1806–1808]. Interfaces [93, 310, 477, 478, 619, 786, 1127, 1178, 1430, 1527, 1786, 1804, 1847, 1922, 2002]. Intermediate [223, 227]. Intermediate- [223]. intermediate-level [227]. Internal [120]. International [356, 829, 936, 1275, 1812, 1885, 1933, 2005, 2006, 2145]. Internationalization [1697]. Internationalizing [126]. Internet [30, 497, 498, 602, 833, 861, 873, 892, 937, 957, 958, 975, 1014, 1040, 1047, 1064, 1095, 1118, 1140, 1163, 1174, 1184, 1206, 1214, 1230, 1249, 1267, 1269, 1282, 1283, 1285, 1298, 1299, 1312, 1315, 1342– 1345, 1359, 1362, 1375, 1376, 1380, 1392, 1402, 1405, 1448, 1454, 1475, 1481, 1497, 1500, 1503– 1505, 1519, 1549, 1553, 1567, 1600, 1605, 1606, 1624, 1657, 1681, 1683, 1705, 1709, 1728, 1757, 1762, 1772, 1773, 1794, 1809, 1812, 1820, 1841, 1865, 1890, 1915, 1938, 1952, 1973, 1989, 2014, 2015, 2040, 2065, 2069, 2081, 2089, 2092, 2119, 2145, 2146]. Internet [758, 1276, 1294, 1331, 1577, 1679, 1752, 1814, 1965, 2132]. Internet-Based [2145]. Internet-Enabled [1567]. Internet/Intranet [1359].

InternetExpress [191]. Internetworked [1448]. Interoperability [1109, 1736]. Interplanetary [2046]. Interplay [1501]. Interpretation [288]. Interpreting [193, 1312]. Interrupt [827]. Interview [1621, 1627, 1689, 1704, 1725, 1933, 1954, 1978, 2101]. Intranet [1359, 1444, 1699, 1907]. Intranet/Extranet [1699]. Intranets [1285, 1763]. Intro [1359]. Introduction [22, 28, 39, 52, 61, 64, 73, 102, 119, 134, 149, 164, 175, 199, 218, 219, 243, 261, 286, 327, 341, 342, 374, 424, 456, 484, 511, 550, 583, 585, 616, 649, 682, 716, 766, 784, 844, 886, 946, 996, 1050, 1104, 1194, 1216, 1238, 1258, 1348, 1436, 1460, 1557, 1564, 1904, 2028, 2105]. Intusoft [1812]. Inventing [1747]. Invention [819, 1580]. inventor [854]. Invents [1898]. Invest [1529]. Investigating [1301]. Investigation [158, 474]. invisible [1134]. Invitation [1074]. Involvement [432]. IONA [1791, 1812, 1836, 1988]. Iowa [847]. IP [907, 1161, 1268, 1684, 1821, 1873, 1897, 1921, 1970, 1994]. IP-Friendly [1994]. IPOs [1116, 1136]. IPv4 [1818]. IPv6 [1800, 1818, 1869]. IRAM [1464]. Irish [1898]. ISDN [499, 1628]. ISE [1791]. ISEF [2034]. Isn’t [623, 1369, 2124]. ISO [207, 517, 1521, 1554, 1634]. ISPs [1277]. Israel [1247]. Issue [219, 243, 261, 342, 1024, 1893]. Issued [1299]. Issues [130, 169, 236, 396, 590, 734, 808, 814, 1299, 1670, 1687, 2097]. Iteration [1196]. Iterative [525]. IXM2 [720].

J [50, 230, 1733]. J. [854]. Jack [1829]. Jade [458]. Jalote [373]. James [1492]. Jan [231]. January [241]. Japan [1320, 2005]. Japanese [66, 94, 447]. Java [956, 1123, 1141, 1164, 1172, 1173, 1303, 1311, 1332, 1334, 1337, 1345, 1358, 1369, 1372, 1388, 1395–1398, 1400– 1402, 1407, 1454, 1482, 1505, 1519, 1554, 1626, 1684, 1697, 1699, 1700, 1723, 1778, 1780, 1803, 1812, 1837, 1849, 1913, 1962, 1971, 2013, 2039, 2041, 2064, 2088, 2117, 2118, 2128]. Java- [1400]. Java-Based [1700, 1837]. JavaBean [1519]. JavaBeans [1619, 1724, 1790, 1791]. JavaOne [2039]. Jazz [158]. JClass [1519]. Jeopardy [1299]. Jewels [1030]. Job [1315, 1970]. Jobs [1526]. Joe [1730]. John [340, 1689, 2101]. Joins [1778]. jolt [956]. Jon [1820]. Jour [1696]. Journals [287]. Joust [1930]. JSTOR [1881]. JTAG [1698]. JTC [1454]. Judge [1505]. Juggler [262]. July [1145]. Jump [1992]. Jump-Starting [1992]. Jumps [1345]. Junkies [1497]. Just [1949, 2070, 2101].

Kahan [1621]. Kahn [1492]. KAI [1697]. Kandel [325]. Karlsson [1492]. keep [692]. Keeping [1292, 1477, 1497, 1509]. Ken [1954]. Kenneth [1541]. Kernel [30, 1782]. Key [26, 101, 343, 1116, 1140, 1328, 1639, 1873, 1943, 1964, 2058]. Keyboard [90]. Keyword [1822]. Killer [1443]. Kiosk [833, 861, 892, 937, 957, 975, 1014, 1040, 1064, 1095, 1118, 1140, 1163, 1184, 1206, 1230, 1249, 1269]. Kit [2118]. KL [1519, 1791, 2064]. know [706]. Knowledge [16, 91, 205, 382, 1154, 1611, 1624, 1723]. Knowledge-Based [382]. Knuth [326]. Kosko [257].

L [1492]. Labor [1731]. Laboratory [1970]. lacking [944]. Laker [1541]. LAN [734]. Land [288]. Lane [1868, 2063]. Language [15, 140, 150, 153, 162, 334, 335, 458, 667, 1150, 1151, 1155, 1334, 1336, 1592, 1623, 1636, 1765, 1873, 2014]. Languages [96, 444, 741, 785, 788, 1347, 1359, 1659, 1798, 1825]. Lanphier [1725]. Large [43, 51, 205, 793, 813, 849, 850, 1086, 1104, 1199, 1951, 2029]. Large-Scale [43, 51, 813, 849, 850, 1104]. Larry [1730]. LaserJet [87]. Lashing [1975]. Last [1368, 1632, 1833]. Late [1101, 1549]. Latecomers [383]. Latency [413]. Latin [1974]. Lattice [523]. Lattice-Based [523]. Latvia [500]. Launches [1299, 1615]. Lauterbach [1978]. Law [361, 706, 727, 743, 761, 777, 795, 810, 834, 860, 879, 894, 929, 958, 982, 1017, 1041, 1065, 1094, 1116, 1136, 1161, 1181, 1203, 1226, 1247, 1267, 1798, 1818, 1854]. Lawmakers [1924]. Laws [541, 556, 1070]. Lawsuits [2022]. Layer [1762]. Layered [2088]. Layers [1869]. Laying [1752]. Layoffs [415]. LDAP [1443]. Lead [1467, 1978]. Leaders [1510, 1522]. Leadership [2125]. Leading [626, 1562, 1675]. Lean [772]. Learn [983, 1508]. Learned [604, 1882]. Learning [271, 718, 818, 1392, 1705, 1772, 1888, 1890, 2053]. leaves [807]. Lector [309]. Lecture [2034]. Lectures [1392]. Legacy [605, 910, 1246, 1489, 1702, 1724, 1817, 1937, 2044]. legacy- [910]. Legacy-to-Web [1937]. Legal [821]. Legalities [1747]. Lemming [1730]. Lemmings [1775]. Lengthening [1845]. Leonard [1492]. Less [830, 1482, 1726, 1786, 1895, 1967]. Less-Industrialized [830]. Lessons [604, 1080, 1293, 1487, 1522, 1869, 1882]. Let [899]. Letter [1333–1339, 1477, 1478, 1627–1630]. Letters [441, 510, 647, 881, 905, 942, 1169, 1365– 1372, 1394, 1403, 1427, 1450, 1502, 1525, 1580, 1603, 1659, 1681, 1707, 1754, 1775, 1798, 1818, 1846, 1869, 1895, 1920, 1943, 1970, 2044, 2070, 2093, 2124]. Level [138, 223, 227, 458, 591, 769, 776, 1357, 1359, 1538, 1608, 1806]. Leveraging [1504, 2055]. Libraries [8, 530, 776, 878, 1011, 1104, 1570, 1879, 1982]. Library [317, 318, 1108, 1109, 1183, 1697, 1791, 1812, 1882–1884]. licensing [1226]. Lies [1703]. Life [233, 910, 960, 1560, 1640, 1902, 2117]. Lifeblood [1795]. Lifecycle [1962]. Lift [1454]. Light [313, 1077, 1078, 1345, 2017]. Lightweight [1078]. Lilac [142]. Limit [2022]. Limitations [1818]. Limited [1812]. limits [1072]. Linda [10]. Line [296, 297, 822, 831, 1437]. lines [959]. Link [272, 615, 1847, 2032]. Linking [1310, 1326, 1426]. Links [1671, 2074, 2145].

Linus [1798]. Linux [872, 1947, 1962, 1970]. Lip [2096]. Lip-reading [2096]. Liquid [1930]. listen [2070]. Listeners [1430]. Listening [1913]. Literate [145, 326]. Literate-Programming [145]. Literature [1105, 1880]. Lithuania [500]. Litigation [1432]. Little [860, 2070, 2122]. Liu [1492]. Live [2043]. Liver [1002]. Living [883, 906, 1123, 1546]. Load [365, 1812, 1962]. Load-Testing [1812]. Lobster [1937]. Local [76]. Locally [365]. Location [289]. Lock [343]. Lock-Key [343]. Logic [110, 211, 212, 230, 269, 270, 1085, 1385, 2109, 2144]. Logical [166, 1029]. Long [315, 316, 1640, 2043]. Long-Life [1640]. long-term [315, 316]. Look [161, 478, 904]. Lookaside [41]. Looking [992, 1071, 1102, 1122, 1145, 1213, 1233, 1255, 1895]. Looking-back [1255]. Looking.back [1021, 1044, 1100, 1167, 1189]. looks [941]. Loops [567]. Loses [1778]. Lots [1961]. Lotto [194]. Lotus [1699]. loving [660]. Low [26, 227, 1812, 1924]. low- [227]. Low-Cost [1812]. Lowering [1318]. Lows [1751]. Luddite [1475]. Luqi [239, 1514]. Lurching [1728].

M [258]. Mac [299, 2064]. Mach [31]. Machinability [379]. Machine [194, 222, 226, 257, 352, 353, 458, 718, 1262, 1411, 1847, 1992, 2104]. Machine-Independent [458]. Machines [104, 1008, 1466]. Macromedia [2087]. Magazine [1631]. Magic [1092, 1519]. magnetic [26]. Mail [957, 1016, 1345, 1898]. Mainframes [2043]. Mainstream [1345, 1741]. Maintainability [670]. Maintaining [751, 1562]. Maintenance [282, 411, 685, 869, 929, 1318, 1323, 1716, 1826, 1970]. Major [1180, 1604, 1778, 1849]. Make [726, 1430, 1451, 1482, 1970, 1994, 2070]. Makers [1662, 2096]. Makes [884, 1432]. Making [686, 1054, 1431, 1453, 1485, 1497, 1527, 1696, 1786, 1908, 1913, 1917, 2001]. man [880]. man-month [880]. Manage [684, 1355, 1856, 2145]. Management [42, 86, 130, 168, 183, 202, 308, 336, 396, 462, 571, 686, 816, 891, 994, 1038, 1064, 1195, 1218, 1292, 1329, 1378, 1390, 1400, 1473, 1522, 1611, 1622, 1624, 1676, 1723, 1727, 1783, 1791, 1818, 1836, 1866, 1917, 1964, 1988, 2017, 2064, 2066, 2078, 2080, 2118, 2121, 2145]. Managing [771, 1137, 1194, 1196, 1198, 1200, 1563, 1613, 1727, 1837, 1883, 2017, 2051, 2066]. Manipulation [15, 129, 310]. Manners [941, 1862]. Mantras [1875]. Manuals [1153]. Manufacturing [253, 1727, 1905]. Many [276, 572, 860, 1115, 1681]. Map [291, 432, 1691]. Mapping [193, 224, 225, 1200, 1938, 2144]. Maps [288]. March [1044, 2006]. Market [601, 766, 884, 1073, 1604, 1872, 1974, 2045, 2047]. Marketplace [1505]. Marriage [920]. Mars [1432]. Masses [1548]. Massively [222, 429, 589, 802, 972, 998]. Matching [222, 1526]. Materials [1094, 1108]. Math [1345]. Mathematica [1791]. Mathematical [215, 1084]. Mathematics [1083, 1084]. Matrix [22]. Matter [1478]. mature [362, 692]. Maturing [1586]. Maturity [517, 750, 1817]. Maximizing [1265]. May [1100, 1857]. Mayo [999]. MBAs [1473]. MBone [602]. McDermid [1689]. McDowell [1384, 1745]. McKnight [273]. MCM [424]. McMathematics [241]. MCMs [430]. Me [1733]. Meaning [2126]. Meaningful [1453]. Means [943, 1967]. Measure [868, 1910]. Measurement [77, 437, 687, 949, 1834]. Measures [688]. Measuring [1245]. Mechanical [290]. Mechanism [190, 1475]. Media [190, 406, 421, 439, 453, 470, 481, 495, 508, 521, 533, 547, 563, 578, 596, 611, 628, 643, 663, 677, 696, 712, 729, 744, 762, 780, 797, 1014, 1652, 1794, 2049, 2087]. Media-Independent [1794]. Mediators [235]. Medical [119, 123, 817, 1643]. Medicine [516]. mediocrity [1014]. Medium [785].

MedNet [817]. Meet [2144]. Meeting [1177]. meetings [957]. Meets [10, 1476]. Melissa [1972]. Member [1893, 1933, 2034]. Members [1274, 1307, 1342, 1429, 1541, 1671, 1732, 1755, 1819, 1909, 1944, 2125]. Membership [851, 925, 952, 966, 1003, 1049, 1193, 1593, 1631, 1649, 1748, 2021, 2072]. Memberships [1571]. Memorial [1745]. Memories [83, 349, 350]. Memory [33, 44, 45, 48, 109, 115, 169, 246, 320, 321, 394, 802, 803, 911, 1026, 1258, 1263, 1327, 1414, 1535, 1687, 1698, 1715, 1738, 1898, 1963, 1983, 2063]. memory-based [394]. Memory-Consistency [1738]. MEMS [1008, 1654, 1981, 2076]. Mentat [445]. Merced [1505]. Meritorious [2140]. Merwin [1492]. Mesh [1438]. Meshes [22]. Message [333, 392, 408, 409, 423, 455, 472, 483, 535, 536, 549, 565, 566, 580, 581, 598, 613, 630, 631, 645, 646, 665, 679, 680, 698, 714, 731, 746, 747, 965, 991, 1045, 1168, 1234, 1273, 1551, 1631, 1771, 1791, 1808, 1844, 1845, 1969, 1988]. Message-Queueing [1988]. Messaging [1836, 2014]. Meta [1771]. Meta-CASE [1771]. MetaCreation [1962, 2118]. Metamorphosis [410]. Metaphor [411]. Metaphors [1427]. Method [22, 828, 1093, 1314]. Methodologies [330, 331]. Methodology [437, 475, 569]. Methods [73, 395, 656, 806, 1074– 1082, 1084, 1301, 1306, 1333, 1403, 1499, 1638]. Metric [1834]. Metrics [49, 437, 670, 671, 682– 684, 686, 694, 1629, 1816]. Metrowerks [1835, 1962, 2064]. Mexico [870]. Michael [1909]. Microarchitecture [974, 1117, 1462]. Microarchitectures [1463]. Microcomputer [120, 258]. Microcomputer-Based [120]. Microcontrollers [358]. Microdynamics [1599]. microelectronics [1008]. microprocessing [1232]. Microprocessor [121, 1056, 1059, 1217, 1667, 1668, 1740, 1790, 1999]. Microprocessor-Based [121, 1217].

Microprogramming [1117]. Microsoft [915, 1185, 1482, 1565, 1580, 1606, 1710, 1729, 1736, 1774]. Microsupercomputer [107]. Microsystems [2014]. Microtest [2145]. Middle [937]. Middleware [1791, 1836, 1946]. Mike [1492]. Mile [1537, 1632]. military [400]. Millennium [1523, 1855, 1869, 2056]. Million [1529]. Mills [1033]. Mind [844, 904, 1847]. Minding [1862]. Mining [1453, 1716, 2028–2030, 2032, 2067]. MIPS [2014, 2026]. Mirror [302]. Misconceptions [638, 1977]. Miss [941]. missing [694]. Mission [1974]. Mistakes [1888]. Misunderstood [1394]. Mitigating [1729]. Mixed [1791]. Mixed-Signal [1791]. mm [1662]. MMM [1014]. Mobile [503, 603, 726, 1302, 1304, 1397, 1605, 1684, 1872, 1898]. Mobility [1294]. Modal [1251]. Model [15, 89, 120, 222, 266, 306, 307, 401, 517, 750, 997, 1097, 1182, 1230, 1360, 1391, 1637, 1713, 1742, 1798, 1817, 1864, 1956, 2072, 2090]. Model-Based [306, 307, 1742, 1798]. Model-Integrated [1360]. Modeler [1812]. Modeling [1, 80, 131, 187, 328, 329, 337, 499, 585, 658, 685, 1087, 1410, 1790, 1824, 2033, 2117]. Modelling [2013]. Models [345, 487, 490, 523, 651, 770, 801, 920, 1028, 1050, 1054, 1127, 1263, 1434, 1664, 1668, 1738, 2106]. Modem [1606]. Modems [1345]. Modern [340]. Module [425, 426, 428]. Modules [1304]. Molecular [193, 341–343, 345, 1180]. Monarch [21]. Monash [2118]. Monetary [1299]. Money [823, 2123]. Monitor [14, 50, 120, 1403, 1700, 1701]. Monitoring [123, 315, 316, 412, 831]. Monopolistic [1603]. monopoly [993]. Monotonic [416, 552]. Monstor [1554]. month [880]. Moore [1854]. Mop [727]. Mortal [2018]. Mosaic [703]. Motion [2124]. Motivating [1292]. Motivation [1050]. Motorola [1454, 1698, 2014]. Mountain [1791, 1826]. Move [884, 1662, 1684, 1871]. Moves [1736, 1946]. Moving [1378, 1463, 1710, 1799]. Mozart [1937]. MPEG [1634]. MPEG-4 [1634]. MS [1710]. MS-DOS [1710]. Much [1659]. Mulder [1909]. Multi [200, 1014, 1791]. multi-media [1014]. Multi-ORB [1791]. Multicasting [1709]. Multichip [425, 426, 428]. Multicode [1347]. Multicomponent [430]. Multicomputer [115]. multicomputers [911]. Multidatabase [201, 204, 236]. Multidatabases [202]. Multidimensional [2030]. Multigroup [622]. Multilingual [9, 1153, 1347, 1812]. Multimedia [76, 185–189, 191, 192, 590, 733, 808, 815– 818, 860, 879, 888, 1356, 1459, 1533, 1822]. Multimesh [22]. Multimodel [220]. Multiple [48, 623, 1385]. Multiple-bus [48]. Multiple-Valued [1385]. Multiprocessing [971, 1535]. Multiprocessor [45, 51, 77, 88, 109, 237, 768, 1241, 1265, 1465]. Multiprocessors [39, 40, 42–44, 80, 475, 1258, 1327, 1414, 1738]. Multisensor [129]. Multithreaded [1697]. Multithreading [668]. Multiuser [619]. Multiwavelength [2088]. Music [149–151, 154, 156, 157, 162, 961, 1131, 1497, 1616, 2051]. Musical [151, 152, 155, 161]. Mustang [1771]. mutually [240]. My [1295]. Myth [1404, 1815, 2052]. Mythical [880, 2026]. Myths [1794].

Nader [1730]. Name [715, 1357, 1405, 1421, 1454, 1554, 1633, 1736, 2071, 2096]. Named [1059, 2112]. Names [1016, 1319]. Nanometer [2105–2107]. nanotechnology [1180]. NASA [1991]. National [852]. Nationalism [871]. Nationalize [1580]. Nations [542, 830]. Native [1488]. Natural [1150]. Navigator [1000]. Navy [1200, 1775]. Nay [854]. NC [1123]. NCSA [953]. NDRAM [1345]. Nears [1345]. necessary [354]. Need [437, 693, 1084, 1085]. Needed [1331]. Needs [96, 362, 1833]. Nerds [1656]. Net [272, 932, 1016, 1034, 1095, 1173, 1181, 1345, 1389, 1398, 1454, 1653, 1662, 1699, 1710, 1723, 1724, 1752, 1791, 1812, 1822, 1836, 1873, 1889, 1898, 1914, 1937, 1962, 1988, 1994, 2014, 2039, 2047, 2087, 2128]. Net-Centric [1173]. Nethead [967]. Netherlands [1628]. netiquette [941]. Netopia [1566]. Nets [157, 1040]. Netscape [1482, 1723, 1778]. Netstorms [1235]. Network [65, 78, 125, 143, 486, 499, 726, 920, 1022, 1052, 1054, 1058, 1148, 1220, 1346, 1400, 1612, 1700, 1701, 1723, 1724, 1759, 1771, 1778, 1786, 1791, 1804, 1806, 1896, 1929, 1970, 1973, 2046, 2089, 2117, 2118, 2128, 2144, 2145]. Networked [816, 1309, 1891]. Networking [1471, 1660, 1757, 1799, 1853, 1871, 1873, 1945, 1981, 2047]. Networks [76, 84, 116, 181, 182, 189, 257, 284, 301, 393, 397, 443, 500, 588, 808, 830, 1026, 1051, 1053, 1055, 1354, 1438, 1504, 1528, 1537, 1708, 1735, 1764, 1821, 1853, 1923, 1928, 2022, 2088]. Neumann [340]. Neural [125, 257, 1051–1055, 1306, 1847]. neuro [998]. neuro-anatomical [998]. Neurocomputing [1050]. Neuronet [123]. Neurons [2070]. Neurophysiologic [123]. Neurosurgical [845]. Neurswing [158]. Never [904]. News [256, 300, 311, 338, 355, 356, 371, 390, 402, 419, 433, 448, 466, 492, 529, 544, 558, 575, 606, 625, 662, 673, 690, 1299, 1319, 1345, 1376, 1407, 1432, 1454, 1482, 1505, 1529, 1553, 1554, 1569, 1582, 1583, 1605, 1606, 1633, 1634, 1661, 1662, 1670, 1683, 1684, 1709, 1710, 1735, 1736, 1744, 1748, 1756, 1757, 1777, 1778, 1800, 1801, 1821, 1822, 1848, 1849, 1857, 1872, 1873, 1897, 1898, 1922, 1923, 1946, 1947, 1972, 1973, 1993, 1994, 2020, 2022, 2046, 2047, 2072, 2095, 2096, 2127, 2128]. Next [175, 866, 1089, 1111, 1268, 1274, 1375, 1425, 1443, 1529, 1801, 1826, 1854, 1921, 2039, 2056, 2145]. Next-Gen [2039]. Next-Generation [1268, 1529, 1801, 2145]. Nielsen [216]. Nifty [1792]. Nineties [179]. nm [1454]. No [1331, 1336, 1655, 1798, 1961, 2149]. Nominations [1289]. Nominees [897, 1468].

Nominess [1308]. Nonconformance [1792]. Nondiscrete [348]. None [1622]. nonimmersive [399]. Noninvasive [14, 50, 1001]. Norton [1062]. Norwegian [1037]. Notation [17]. Notebook [428]. Notes [2112, 2139]. Nothing [657, 1968]. Notion [69]. Novel [143, 810]. Novell [810]. November [1233, 2140]. Novera [1699]. NP [1254]. NP-complete [1254]. NRC [832]. NRC/CSTB [832]. NRL [1647]. NSK [1037]. NT [866, 994, 1697, 1782–1784, 1818]. Nuclear [450, 1507]. Numbers [1797, 1839]. Numerical [1011]. NV [2006].

O [190, 583, 584, 589–591, 734, 1132, 1898, 2072]. O-Intensive [592]. OADD [1358]. OB [1696]. Object [16, 17, 130, 165, 234, 238, 284, 327–331, 333– 337, 363, 364, 396, 414, 445, 460, 637, 667, 909, 912, 913, 916, 918– 923, 1006, 1031, 1036, 1063, 1093, 1115, 1133, 1138, 1160, 1194, 1197, 1198, 1207, 1229, 1251, 1271, 1293, 1330, 1382, 1391, 1410, 1439, 1474, 1486, 1487, 1523, 1576, 1623, 1678, 1726, 1736, 1803, 1812, 1816, 1837, 1867, 1877, 1918, 1962, 1966, 2013, 2016, 2067, 2122, 2147]. Object-Based [667, 923]. Object-Bases [165]. Object-Oriented [16, 17, 130, 234, 238, 284, 327–331, 334– 336, 363, 364, 396, 414, 445, 637, 909, 912, 918, 922, 1133, 1194, 1197, 1198, 1382, 1816, 1877]. ObjecTime [1812]. Objects [189, 304, 305, 791, 914– 917, 1109, 1176, 1427, 1488, 1884, 1913]. ObjectSystems [1039]. Obstacles [1532, 1556]. Occur [2129]. October [1213, 2084]. Off [296, 297, 842, 1690, 1702, 1710]. Off-Line [296, 297]. Off-the-Shelf [1690]. offense [834]. offered [1180]. Offers [1454, 1505, 1582, 2111]. Office [293, 743, 757, 761, 897, 898, 901, 1039, 1468, 1768, 1981]. Officer [1885]. Officers [1541, 1819].

Offices [2009]. OK [1554]. OLAP [2014]. old [260]. old-fashioned [260]. Olderog [359]. Oldfield [1514]. OLE [2145]. Olympic [1157]. On-Line [822, 831]. One [303, 1461, 1477, 1706]. Online [1406, 1413, 1881, 1933, 1993, 2053]. Only [320, 321, 1983]. onto [224, 225, 1343, 1871]. OO [909, 1195, 1196, 1199, 1200, 1305, 1489, 1490, 1937, 2007, 2091, 2144]. OO-VHDL [909]. OOAD [931]. OODBSs [920]. Open [217, 240, 260, 275, 280, 281, 285, 303, 336, 360, 361, 407, 422, 440, 454, 471, 482, 486, 496, 509, 522, 534, 548, 564, 579, 597, 612, 629, 644, 664, 678, 697, 713, 726, 730, 745, 763, 781, 798, 812, 839, 865, 904, 941, 964, 990, 1020, 1043, 1070, 1093, 1099, 1121, 1144, 1166, 1187, 1212, 1232, 1254, 1272, 1341, 1373, 1428, 1451, 1479, 1482, 1526, 1578, 1583, 1601, 1625, 1656, 1677, 1704, 1729, 1733, 1795, 1798, 1815, 1842, 1893, 1894, 1941, 1942, 1967, 1988]. Opening [300]. OpenMP [1963]. Opens [1973]. Operating [28, 31, 32, 167, 196, 592, 872, 954, 1035, 1206, 2145]. operation [1179]. Operational [319]. Operations [2124]. Operators [88]. Opportunities [457, 864, 939, 963, 988, 1018, 1068, 1211, 1495, 1517, 1544, 1618, 1717, 1769, 1832, 1848, 1860, 1887, 1935, 1960, 2011, 2036, 2062, 2084, 2116, 2143]. Opportunity [1523]. Optic [116, 182, 2145]. Optical [271, 349, 350, 393, 701, 721, 1000, 1584, 1588, 2088]. Optically [1590]. Optics [1586, 1587]. optimism [386]. Optimization [254, 269, 270, 1052, 1962]. optimized [1132]. Optimizing [279, 950, 1398, 1424]. Optoelectronic [1591]. Oracle [1898]. ORB [1791]. Orbits [161]. ORBs [1700]. Orchestras [2051]. Order [122, 957, 1136, 1939, 1943, 1994]. Organization [1424]. organizational [1329]. Organizations [2079]. Orientation [337, 460, 922]. Oriented [16, 17, 130, 234, 238, 284, 327–331, 334– 336, 363, 364, 396, 414, 445, 637, 719, 785, 909, 912, 918, 922, 1133, 1194, 1197, 1198, 1382, 1816, 1877, 1962, 2118]. Origins [340]. OS/2 [339]. OSI [1470]. Ostriches [1363, 1364]. Ostroff [230]. OT [917]. other [490]. Our [1312, 1551, 1795, 1845, 1888]. Outlines [2039]. Outlook [1564, 1592, 1860]. output [87]. outsource [742]. Outstanding [1933, 2140]. overcomes [1347]. Overdog [1798]. Overlay [426]. Overload [1657, 2070]. Overview [62, 100, 131, 398]. Overviews [156, 254, 379]. ownership [1163]. Oxymoron [1681].

P [215, 229, 238, 373]. Pace [692, 1509]. Package [427]. Packageability [431]. Packaging [370, 424, 431, 432, 1440, 1962]. Packard [2014]. PACT [376]. Pager [1710]. Painting [2118]. Pajamas [1942]. Palmtop [1188]. PAMITC [984]. Panix [1236]. Paper [293]. paperless [901]. Papers [863, 903, 940, 962, 987, 1019, 1042, 1069, 1120, 1143, 1165, 1186, 1209, 1231, 1253, 2006]. Papers/Calendar [863, 903, 940, 962, 987, 1019, 1042, 1069, 1120, 1143, 1165, 1186, 1209, 1231, 1253]. Paradigm [115, 145, 343, 717, 786, 911, 936]. Paradigms [319, 341, 342, 1588]. Paradise [2039]. Paradox [1681]. Paradyn [949]. Parallel [7, 21, 34, 96, 98, 103, 104, 110, 122, 213, 218– 220, 222–225, 228, 244, 245, 264, 304, 305, 429, 445, 458, 475, 490, 513, 589, 653, 667, 669, 719– 722, 735, 773, 802, 803, 876, 916, 917, 946, 947, 949– 951, 972, 973, 998, 1027, 1053, 1132, 1259, 1264, 1272, 1554, 1590, 1787, 1829, 1988, 2005, 2134]. Parallel-Processing [213, 224, 225]. Parallelism [31, 226, 227, 567, 752, 916, 917, 1538, 1963]. Parallelizing [1260]. Parametric [1697, 1812]. ParaSoft [1988, 2118]. Parliament [1994]. Parrish [1860, 2059]. Parsec [1787]. Parsons [357]. Part [541, 556, 743, 878, 899, 931, 1089, 1111, 1116, 1252]. Partial [97]. Participate [1695].

Participation [1541, 2005]. Partition [1092]. Partitioning [1612]. Partnership [1947]. party [726]. Passion [1132]. Passive [1641]. Past [1221, 1985]. Patent [743, 761, 810, 834, 1154, 1943]. patent-infringement [810]. Patents [727, 881, 1247, 1927, 1970, 2044]. Path [352, 1532]. Paths [1805]. Pattern [213, 741, 892, 2108]. Patterns [793, 1474, 1497, 1623, 1696, 1877, 2117]. Paving [1632]. PAWS [103]. Paying [2024]. Payment [1456]. PC [1047, 1175, 1318, 1431, 1505, 1843]. PCI [1801, 1984]. PCs [878, 1358, 1430, 1873, 2127]. PDL [1087]. pedagogues [818]. Peering [1770]. Pegasus [201]. Pentatonic [160]. Pentium [1241, 1923]. People [1526]. Perception [1330]. Perfect [1391, 1970]. Perfecting [411]. Performance [77, 103, 105, 137, 152, 211, 212, 214, 259, 333, 396, 424, 586, 587, 589, 653, 803, 868, 946– 951, 953, 1087, 1175, 1261, 1262, 1265, 1398, 1414, 1457, 1591, 1609, 1666–1668, 1670, 1804, 1978, 1984, 2008, 2096, 2117, 2134]. Performance-Measurement [77]. Peripheral [1786]. perish [834]. Perl [932]. Permits [1345]. Persistence [322]. Personal [1252, 1300, 1345, 1377]. Personality [1479, 1578]. Personalized [1303]. personnel [983]. Perspective [183, 253, 319, 514, 1006, 1224]. Perspectives [1584]. Petri [157]. phenomena [1022, 1123]. Phenomenon [754]. philosophers [360]. Philosophies [99]. Phone [1454]. Phones [1524]. Photocopyright [894]. Physical [1457]. Physics [461]. Picking [1798]. PicoJava [1780]. Picture [914]. pieces [879]. PIM [802]. Pining [1970]. Pioneer [1514, 1571, 1820, 1829]. Pioneering [1118]. Pioneers [1492, 1693]. PIPE [108]. Pipeline [1735]. Pipes [1660]. piracy [356]. Piranha [752, 1351]. Pitches [2047].

Pitfalls [1361]. Place [1164, 2070]. Placing [1366, 1833]. Plan [1170, 1421, 1454, 1551, 1554, 1633, 1634, 1757, 1845]. Planetary [161]. Planetweb [2039]. Planning [414, 845, 999, 1002, 1270, 1359, 1798]. Plans [1407, 1703, 2047, 2073, 2128]. plant [450]. Platform [30, 726, 767, 1930, 1937]. Platforms [1610]. Platinum [1141]. Player [220]. Playing [2042]. Plays [1753]. Plea [241, 772]. Plug [1866]. poem [1187]. Point [141, 323, 324, 1075, 1076, 1345]. Point-Counterpoint [1075, 1076]. Point/Counterpoint [141, 323, 324]. points [959]. polaris [1264]. Policy [300, 356, 871]. Pop [2044]. popular [1066]. Portability [86, 92]. Portable [1135, 1750]. Porting [704]. Position [477, 1562]. Posix [60, 531, 1864]. Post [1681]. Post-2008 [1681]. Postal [289]. Postel [1820]. potential [226]. Potholes [1435]. Power [450, 906, 1676, 1726, 1967]. power-plant [450]. powered [943]. Powerful [299]. PowerPC [635, 1102, 2145]. Practical [2078, 2117]. Practice [274, 417, 1079–1082, 1314, 1391, 1592]. Practices [1133, 1507, 1531, 1585, 1664, 1679, 1738, 1779, 1780, 1803, 1903, 2051]. Practicing [1065]. Practitioners [1788]. Pragmatism [871]. Pragmatix [1771]. Pray [1866]. Precedents [1875]. Predicate [1512]. Predicated [1564]. Predict [1648]. Predicting [804, 1244]. Predictions [1317]. Predictive [90]. Prefetching [1409]. Premia [2039]. Preparation [1116, 1913]. Prepares [1778]. Preparing [133, 2020]. Presence [137]. Present [1221]. Presented [1909, 2034, 2140]. president [408, 535, 549, 565, 580, 630, 679, 746, 991, 1234, 1541, 1551, 1845, 1985, 2058, 2059]. president-elect [1541, 2058, 2059]. Presidents [1969]. Presper [854]. Press [2034]. Pressure [705, 1852]. Presto [1970]. Prevention [2118]. Preview [900, 1966, 1985]. Price [1697, 2024, 2093].

Prime [1661]. Primer [1790]. Principles [274, 1474, 1825]. Print [1600]. Printed [1131]. printer [87]. Privacy [1067, 1683, 1696, 1755, 1923, 2070]. Privasuite [1358]. Private [1504, 1736, 1763, 1778]. Privately [1338]. Prize [385, 557, 756, 1005, 1180, 1266, 1286, 1505, 1541, 1569, 1829, 1985]. Pro [1241]. probabilistic [1259]. Probe [1700, 1701]. Problem [484, 621, 1239, 1254, 1803, 1924]. Problems [111, 516, 589, 737, 1298, 1482, 1563, 2070, 2071]. Procedural [912]. Procedures [996]. Process [11, 60, 354, 414, 428, 687, 728, 874, 1009, 1199, 1227, 1377, 1378, 1599, 1648, 1676, 1962, 2079]. Process-Oriented [1962]. Processes [771, 960, 1004, 1554]. Processing [34, 122, 146, 170, 206, 213, 218, 219, 221, 224, 225, 244, 245, 264, 357, 445, 456, 669, 716, 721, 767, 802, 827, 876, 917, 1150, 1155, 1554, 1588, 1829, 1913, 1991, 2005, 2013]. Processor [5, 21, 108, 209, 394, 410, 720, 721, 1296, 1459, 1511, 1557–1564, 1710, 2124]. Processor- [394]. Processors [247, 248, 716, 827, 1272, 1463, 1464, 1529, 1533, 1552, 1580, 1741, 1835, 1978, 2014, 2047, 2087]. Produce [2039]. Product [13, 20, 27, 36, 72, 95, 112, 127, 147, 163, 173, 313, 339, 379, 401, 404, 418, 435, 451, 465, 480, 494, 506, 520, 528, 546, 561, 577, 593, 609, 624, 641, 659, 676, 695, 709, 725, 739, 759, 778, 794, 811, 837, 857–859, 878, 900, 933– 935, 954, 955, 977–979, 1010– 1012, 1039, 1061, 1062, 1092, 1119, 1142, 1164, 1185, 1208, 1210, 1252, 1291, 1437, 1818, 1910, 1981, 1988, 2037, 2115, 2142]. Product-Line [1437]. product-model [401]. Production [6, 1525, 1952, 2039]. Productivity [111, 757, 852, 1197, 1681, 1739, 2066]. Products [1141, 1358, 1388, 1401, 1411, 1422, 1444, 1471, 1496, 1518, 1519, 1545, 1724, 1771]. Profession [493, 1958]. Professional [978, 1519, 2084]. Professionalism [1941]. Professionals [2, 1112, 1491]. Professor [2112]. Profile [893, 1491]. Profiler [1388]. profiles [12]. Profiling [700]. Profitable [1389]. Program [49, 89, 138, 446, 537, 555, 869, 924, 1064, 1382, 1519, 1550, 1858, 1933, 2070]. Programmable [110, 1735, 1853]. Programmers [96, 912, 1299, 1655]. Programming [145, 150, 211, 212, 220, 274, 283, 304, 305, 319, 322, 326, 334, 335, 358, 458, 490, 634, 667, 757, 784, 786–789, 809, 895, 902, 912, 923, 1063, 1070, 1264, 1280, 1608, 1723, 1790, 2013, 2041, 2063, 2082, 2091]. Programs [359, 412, 475, 803, 950, 1686, 2084]. Progress [1928, 2048]. Progresses [1387]. Project [48, 71, 75, 156, 254, 379, 380, 1107, 1159, 1195, 1200, 1791, 1917, 1964, 2031, 2064, 2080, 2103]. Projects [684, 1196, 1199]. Prolog [211, 212, 308]. Promise [175, 791, 1778, 1923]. Promises [1822]. propagation [267]. Proper [2110]. property [706, 860, 879]. Proposal [1325, 1526, 2128]. Propose [1554]. Proposed [1070, 2022]. Proposes [1454, 1634]. Proprietary [1818]. Prospect [1470]. Prospects [253, 589, 737]. Protected [1808]. Protecting [1163, 1696]. Protection [706, 860, 879, 1161, 1771]. Protein [349, 350, 515]. Protein-Based [349, 350]. Protocol [47, 551, 553]. Protocols [443, 540, 1725, 1806]. Prototype [287]. Prototyping [223, 255, 620, 766, 768, 769, 1891]. ProtoVIew [1962, 1988]. provide [1161]. provides [602]. Providing [1678]. Pseudorandom [113]. Public [26, 300, 356, 486, 626, 1140, 1346]. Publicly [936]. Publishing [1184, 1257, 1331, 1794]. pubs [1359]. Pump [1647]. Purdue [2034]. Purpose [524, 881, 1178, 1419, 1533]. Pursuit [1682]. Push [1634]. Pushes [1662, 1822]. Pushing [1558]. Put [1549, 1868]. Putting [65, 879, 1723, 2144]. PX [70]. Pygmalion [1601].

QBIC [887]. QNX [954]. Qualifying [117]. Quality [133, 437, 468, 671, 688, 922, 1245, 1391, 1435, 1447, 1629, 1648, 1775]. Quench [1480]. Query [444, 887]. Question [155, 1332]. Questions [141, 1156, 1337, 1798, 2058]. Queue [1327]. Queueing [1988]. QuickTime [1634]. quiz [873].

R [258, 274, 284, 301, 325, 357, 359, 1953]. RAD [1488, 1917]. RADD [1519]. Rader [2112]. Ralph [1730]. Random [83, 702]. Random-Access [83]. Random-Dot [702]. Ranges [1947]. Rapid [255, 766, 768, 1001, 1497]. Rapidly [1644]. Rare [1030]. Rate [416, 552]. Rate-Monotonic [552]. Rating [1775]. Rational [1578, 1700, 1771, 1791, 1812]. Rationale [377, 378]. Raw [1466]. Raymond [1933]. RBOCs [1277]. RCS [266]. reach [975]. Reached [2022, 2072, 2096]. Reaching [1537]. React [1755]. Reactions [84]. Reactive [90, 267]. Read [3, 1451, 1857, 1933]. Reader [858, 935, 955, 979, 1012, 1061]. Readiness [2074]. Reading [292, 807, 1649, 1857, 2034, 2096]. Ready [1661, 1662, 1948]. Real [14, 23, 50, 123, 134–137, 140, 197, 216, 230, 262, 289, 325, 334, 335, 363, 364, 416, 501, 551– 553, 569, 604, 638, 755, 767, 825, 883, 906, 965, 1354, 1358, 1439, 1503, 1696, 1704, 1725, 1790, 1812, 1902, 1971, 1977, 2128]. Real-Life [1902]. Real-Time [14, 23, 50, 123, 134–136, 140, 197, 216, 230, 262, 325, 334, 335, 363, 364, 416, 501, 551– 553, 569, 604, 755, 767, 825, 1354, 1358, 1439, 1503, 1725, 1812, 1971, 1977, 2128]. Real-Time-Systems [137]. Real-World [1790]. Realities [1314, 1361, 1447, 1499, 1548, 1599, 1655, 1703, 1751, 1796, 1839, 1892, 1940, 1990, 2042, 2090, 2148]. Reality [185, 398– 400, 844, 951, 1047, 1160, 1330, 2070, 2148].

Really [354, 1970, 2040, 2147, 2148, 2151]. RealNetworks [1725]. Reasoning [888]. Rebol [2014]. Rebuts [1655]. Receive [1384]. Receives [1745, 1829, 1860, 1985, 2009, 2111]. Recipient [1384]. Recognition [64, 65, 291, 292, 294–297, 1306, 1518, 2096]. Recognize [1492]. Recognizes [1514, 1693, 1829, 2034, 2140]. Recognizing [1497]. Recombinant [151]. Recommendations [832]. Reconfigurable [224, 225, 1645, 1698]. Reconfiguration [5, 410]. Reconstruction [461]. Record [1317]. recovery [394]. Red [1073, 1929, 1962]. red-hot [1073]. Redefine [1974]. Redesign [591]. Redirector [1791]. Reducing [766, 1327]. Reduction [1383]. Reengineering [1434]. Refactoring [2088]. Referee [24]. Reference [266, 1182, 1359]. Referenced [1108]. References [1715]. Refining [1133]. Reflections [1257]. Reflexion [1434]. Reform [1328]. Reframing [1892]. Refuse [1658]. Register [288]. registration [133]. Rehearsal [997, 999]. Reining [1559]. Reinventing [1906]. Rejects [1454, 1505, 2128]. Rejoinder [1773]. Related [224, 225, 727, 743, 761, 2093]. Relational [88, 889, 920]. Relational-Database [88]. Relationship [15, 685]. Relationships [1916]. Relax [1943]. Release [1839]. Releases [1898, 2128]. Reliability [56, 131, 368, 369, 417, 586, 1238, 1243, 1244, 1348, 1797]. Reliable [368, 369, 2122]. Reminders [1281]. Remote [207, 753, 1185, 1937]. removal [1096]. removed [1160]. Renaissance [1285]. Replication [1352]. reply [37, 38, 50]. Report [85, 214, 351, 479, 645, 669, 704, 756, 832, 1017, 1146, 1274, 1316, 1342, 1416, 1417, 1429, 1720, 1732, 1755, 1819, 1873, 1924, 1928, 1944, 1974, 1995, 2010, 2023, 2048, 2125]. Reported [1684]. Reports [840]. Repositories [203]. Representation [17, 293, 378, 1055]. Request [916]. Requirements [11, 450, 568, 1137, 1323, 1643, 1892, 1990]. Requirements-Based [1990]. Rescuing [1616]. Research [102, 156, 362, 807, 829, 847, 870, 1082, 1511, 1512, 1565, 1592, 1707, 1742, 1785– 1787, 1802, 1809, 1822, 1852, 1854, 2027]. Researchers [1923]. Resection [1002]. Reshapes [1945]. Resilient [202]. Resistance [1393]. Resolution [488]. Resource [195, 205, 497, 498, 816, 1324, 1951]. Resource-Based [195]. Resources [30, 830]. Respect [2149]. Respond [96]. Responsibilities [366]. Responsive [848]. restored [1179]. Restrictions [1505]. Restructurable [249, 250]. Results [721, 825, 980, 1377, 1719, 1745, 2010]. Rethinking [1825, 2100]. Retrieval [886, 889–891, 1134, 1570]. Retrospect [1470]. reusability [1036]. Reusable [117, 1489]. Reuse [661, 910, 1198, 1363, 1364, 1446, 1485– 1488, 1490, 1943, 2066]. reuse-based [910]. Reveals [1299]. Reveiws [676]. Reversibility [1207]. Review [215, 216, 229, 230, 238, 239, 241, 257–259, 273, 274, 283, 284, 301, 302, 314, 325, 326, 340, 357– 359, 373]. Reviewer [2140]. Reviewers [1542, 1830]. Reviews [13, 20, 27, 36, 72, 95, 112, 127, 147, 163, 173, 313, 339, 404, 406, 418, 421, 435, 439, 451, 453, 465, 470, 480, 481, 494, 495, 506, 508, 520, 521, 528, 533, 546, 547, 561, 563, 577, 578, 593, 596, 609, 611, 624, 628, 641, 643, 659, 663, 677, 695, 696, 709, 712, 725, 729, 739, 744, 759, 762, 778, 780, 794, 797, 811, 837, 857, 878, 900, 933, 954, 977, 978, 1011, 1039, 1062, 1092, 1119, 1142, 1164, 1185, 1208, 1252, 1291]. Revised [450]. Revisionist [1636]. Revisited [231, 1775]. Revolution [1815, 1857, 2076]. Revolutionary [2044]. Rewarding [1338]. Rewards [1339]. Richard [1492, 1615]. Right [601, 1164]. RightPages [317, 318]. rights [1163]. Rigorous [1077, 2070]. RISC [217, 1944]. RISCs [635]. Rising [1850]. Risk [1378, 1990]. Risks [217, 1691]. risky [217]. Road [432, 648, 1090, 1314, 1435, 1895]. Roadblocks [1561]. Roadmap [1583]. Rob [1725]. Robert [957, 1492]. Robo [727]. Robo-Mop [727]. Robot [262]. Robotic [129, 306, 307]. Robotics [541, 556, 604]. Robust [291, 1130, 2109]. Rochester [220]. Rodgers [1733]. Rogue [1724]. ROI [1449]. Role [531, 1028, 1060, 1575, 1816, 1840, 1905, 2015]. Role-Based [1028]. rollback [394]. ROM [842]. room [1123]. Rose [2145]. Rounding [914]. Roundtable [867, 913, 1509, 1557–1564, 1953]. Routed [972]. Routers [1805]. Routing [397, 443]. RPM [768]. RS [105]. RS/6000 [105]. RSI [1771]. RST [1724]. RTOS [1724]. Rubin [1875, 1900, 1925, 1949, 1975, 1996, 2024, 2049, 2074, 2098, 2130]. Rules [751, 929, 957, 1066]. Run [1327, 1790]. Runs [1482]. Runtime [268, 412, 488, 1681, 1812]. Russian [1924].

S [259, 1384]. SA [1597]. Sabotage [1457]. Safe [1065, 1681]. Safety [118, 323, 324, 389, 450, 1242, 1641, 1643]. Safety-Critical [1242, 1643]. Sahni [1492]. Said [1370]. Sales [1454]. Salvation [1602]. same [715]. SAN [1988]. Sartaj [1492]. Satan [833]. Satellite [1302, 1454, 1710]. Saturn [1739]. Savvy [1700, 1962]. Sawyer [1937]. Say [657, 1531, 1726, 1948, 1949]. Says [1818]. Scala [1616]. Scalability [1457]. Scalable [29, 45, 46, 115, 526, 916, 1464]. Scale [43, 51, 242, 243, 246– 252, 254, 813, 849, 850, 1104, 1951]. Scales [160]. Scaling [475, 788]. Scan [2109]. Scan-Based [2109]. scenes [1123]. schedules [774, 1032]. Scheduling [23, 31, 139, 825, 971, 1547]. Schema [138, 1229]. Schematic [200]. scheme [46, 47]. Schemes [40, 555]. Scholar [2034].

School [924, 2034]. Schussel [1829]. Science [215, 362, 447, 511, 572, 655, 854, 870, 924, 990, 1098, 1174, 1359, 1501, 1601, 1681, 1745, 1933, 2028, 2102]. Scientific [96, 97, 180, 512, 657, 861, 1105, 1880, 1918]. Scientists [275, 1665]. Scope [1419]. Scores [157]. Scoresynth [157]. Scripting [1608]. Scripts [1723]. SDK [1962]. SDKs [1496]. Sea [1505]. Seagate [1700, 1701]. Search [754, 1256, 1315, 1362, 1503, 1880, 2019]. Searching [1757, 2019]. Secrets [1016, 1497, 1790]. Section [504]. Secure [29, 1248, 1283, 1399]. Security [1, 486, 937, 944, 1058, 1067, 1299, 1304, 1319, 1376, 1691, 1692, 1734, 1759, 1761– 1763, 1812, 1823, 1940, 2118, 2145]. See [1732]. Seeing [1585, 1696]. Seek [1345, 1822, 1924, 2084]. Seeking [1534]. Seeks [1710, 1994]. segment [639]. SEI [517, 750]. Selected [1853]. Selection [8, 853]. Self [83, 333, 754, 1240]. Self-Belief [754]. self-test [1240]. Self-Testing [83]. Selling [1873]. Semantic [1956, 2126]. Semantic-Model [1956]. Semiconductor [1299, 1583, 1854]. Semiotics [229]. Sense [477, 1869, 1908]. Sensing [1130]. September [1710, 2005]. Sequencing [128]. Sequential [359, 803]. Serendipity [2028]. Series [2034]. Serios [1076]. Seriously [1759]. Server [190, 623, 801, 804, 805, 914, 953, 1031, 1133, 1214, 1388, 1699– 1701, 1771, 1962, 1988, 2039, 2045, 2072]. Servers [815, 1444]. Service [174, 191, 858, 935, 955, 979, 1012, 1061, 1091, 1503, 1514, 1593, 1662, 1671, 1710, 1784, 1860, 1869, 1985, 2047, 2140]. Services [487, 497, 1106, 1929, 1989, 2016]. Servoing [1413]. SES [1771]. Set [410, 459, 1278, 1606, 2039]. Set-Top [1278, 1606, 2039]. Sets [124, 1317, 1662]. Settlement [1606]. Setup [2145]. Seven [1428, 1933]. Several [1347, 1849]. Seymour [1246, 1571, 2102]. SGI [1849, 1988, 2014]. SGML [1497]. Shackles [1702]. Shades [1732]. Shake [956, 1633]. Shape [891]. Share [1671]. Shared [33, 44, 45, 48, 109, 115, 169, 304, 305, 383, 1026, 1258, 1260, 1263, 1327, 1414, 1535, 1963]. Shared-Memory [44, 45, 48, 109, 115, 1963]. Shares [2034]. shareware [1267]. Sharing [285, 1951]. Sheep [2049]. Sheet [2119]. Shelf [1690]. Shell [16]. Shells [2117]. Shift [1173]. Ship [1059]. Shipping [592]. Shoot [1295]. Short [2070]. Shortage [1374, 1684, 1731]. Should [1531, 1580, 1665, 1718, 1738, 2073]. Show [2140]. shows [389, 873]. Shrinking [1452]. Shuttle [671]. Side [883, 906, 943, 1341]. Sidney [1593]. Siege [1058]. Signal [122, 357, 1552, 1791]. significant [12]. Silicon [1871]. Silver [208]. SIMD [227, 2132]. Similar [891]. Similar-Shape [891]. Simovici [215]. Simple [1738, 1822]. Simpler [2127]. Simplicity [1550, 1967]. Simplify [2128]. simplifying [12]. simply [990]. Simulated [263]. Simulating [1004, 1053]. Simulation [4, 143, 499, 568, 776, 1015, 1667, 1771, 1787, 1812, 2023, 2136, 2145]. Simulation-Based [1812]. Simulations [515]. Simulator [847, 1791]. Simulators [1439]. Single [247, 248, 702, 968, 1465, 1724, 2087]. Single-Bit [247, 248]. Single-Board [2087]. Single-Chip [968, 1465]. Single-Image [702]. Single-Source [1724]. Sixpack [1730]. skills [983]. Sky [1724]. Slate [479, 1721]. Slow [1798]. Small [382, 1008, 1919, 1974, 2078, 2079]. Smallest [1898, 2072]. Smalltalk [1696]. Smart [1157, 1454, 1497, 1622, 1750, 1801, 2123, 2146]. Smart-Card [1801]. Smarter [1715]. Smith [1985]. Snap [264]. SNMP [944]. Snyder [1492]. Snyder-Holberton [1492]. Soar [1345]. SOC [1979, 1981]. Social [1128, 1879]. societies [260]. Society [174, 256, 285, 311, 338, 355, 371, 897, 1158, 1419, 1429, 1441, 1468, 1509, 1693, 1798, 1860, 1885, 1909, 1933, 1958, 1985, 2009, 2010, 2034, 2057– 2059, 2063, 2084, 2111, 2112, 2125, 2139, 2140]. SOCs [1981]. Soft [1791]. SoftQuad [2014]. SoftQuest [2064]. Software [81, 117, 183, 373, 375, 381, 387, 395, 411, 489, 555, 571, 600, 615, 642, 661, 670, 674, 683, 688, 694, 708, 724, 742, 757, 774, 790, 793, 809, 835, 856, 875, 880, 895, 922, 938, 942, 959, 970, 983, 1004, 1009, 1038, 1066, 1076, 1086, 1096, 1110, 1114, 1137, 1175, 1194, 1226, 1228, 1245, 1311, 1314, 1326, 1352, 1358, 1361, 1377, 1388, 1437, 1447, 1454, 1479, 1484, 1499, 1501, 1510, 1515, 1525, 1599, 1649, 1655, 1682, 1697, 1703, 1716, 1727, 1751, 1775, 1796– 1798, 1892, 1930, 1943, 1952, 1970, 1990, 2002, 2042, 2066, 2081, 2090, 2148]. Software [6, 17, 55, 111, 126, 239, 282, 368, 369, 386, 388, 389, 417, 468, 493, 539, 542, 682, 684, 687, 689, 705, 727, 772, 805, 867, 869, 910, 929, 958, 1030, 1083, 1170, 1172, 1207, 1219, 1243, 1244, 1292, 1305, 1363, 1364, 1425, 1428, 1450, 1466, 1494, 1519, 1548, 1581, 1630, 1648, 1653, 1658, 1673, 1674, 1688, 1690, 1696, 1724, 1743, 1747, 1771, 1791, 1812, 1834, 1839, 1846, 1889, 1914, 1927, 1936, 1937, 1940, 1953, 1958, 1962, 1964, 1987, 1988, 2022, 2038, 2039, 2044, 2054, 2055, 2063, 2064, 2077, 2078, 2084, 2087, 2088, 2117, 2118, 2122, 2144, 2145]. software [228, 267, 354, 356, 437, 671, 821, 960, 1046, 1519, 1699, 1724, 1771, 1791, 1812, 1836, 1937, 1962, 2014, 2039, 2145]. Software-Based [1352]. Software-Fault-Tolerant [55]. Software-Process-Improvement [687]. software-quality [437]. software-related [727]. Software-Reliability-Engineered [1243]. SoftWired [1837]. Solaris [1988]. Sole [854]. solid [436]. solid-state [436]. Solution [1697, 1724, 1778, 2071]. Solutions [741, 1099, 1298, 1322, 1388, 1519, 1534]. Solving [621, 1803, 2135]. Something [1968]. Sonet [1821]. Sonic [1812]. Sony [2018]. Sorokin [2145]. Sought [1356]. Sound [153, 234, 367]. SoundWorks [234]. Source [138, 685, 854, 859, 934, 1010, 1724, 1798, 1894, 1942, 1988]. Source-Level [138].

SourceAgain [1519]. Sources [1346]. South [737]. Space [655, 671, 757, 1260, 1298, 1299, 1587, 1591, 1640]. Spacecraft [1991]. SPAP [279]. Spark [2000]. Spatial [262, 845]. Spatio [189]. Spatio-Temporal [189]. Speaker [64]. speaking [753]. Spec [700, 868, 1583, 1606]. Special [214, 219, 243, 261, 342, 669, 756, 832, 1514]. Specialization [1560]. Specific [787]. Specification [1, 826, 936, 1047, 1646, 2002, 2128]. Specification-Driven [826]. Specifications [485]. Specified [412]. Specifier [73]. Specifying [204]. Specs [1634, 1898, 1923]. specter [1014]. Spectral [122]. Spectrum [914]. Speech [63–68, 144, 1152, 1518, 1531, 2047, 2096]. Speech-to-Speech [144]. SpeechActs [1151]. Speed [116, 263, 393, 588, 701, 721, 1184, 1256, 1388, 1563, 1606, 1805, 1923, 2017]. Spell [2118]. Spell-Checking [2118]. Spert [1056]. Spert-II [1056]. Spin [1900]. Spiral [1713]. Splits [1405]. Spoken [1151, 1155]. Spoken-Language [1151, 1155]. Spreadsheets [1903]. Spurs [2022]. SQL [207]. Square [1710]. SRC [2145]. Stability [1719]. Stableware [1793]. Staff [1649]. Stage [1662, 2094]. Stake [2096]. Stakeholder [1916]. Standard [6, 26, 49, 162, 172, 207, 282, 293, 401, 437, 450, 568, 960, 1060, 1248, 1325, 1494, 1521, 1554, 1710, 1778, 1801, 1873, 1994, 2128]. standardization [101]. standardize [626]. Standardizing [1725]. Standards [6, 12, 19, 26, 35, 49, 60, 71, 79, 94, 101, 111, 118, 126, 141, 162, 172, 282, 323, 324, 354, 368, 369, 389, 401, 417, 437, 450, 468, 493, 502, 517, 531, 574, 595, 608, 626, 689, 692, 710, 728, 790, 862, 874, 893, 918, 927, 936, 986, 1013, 1037, 1060, 1097, 1135, 1159, 1182, 1204, 1222, 1227, 1248, 1268, 1290, 1310, 1328, 1357, 1387, 1407, 1443, 1454, 1470, 1494, 1521, 1547, 1554, 1575, 1597, 1615, 1621, 1675, 1695, 1725, 1749, 1772, 1792, 1813, 1822, 1840, 1864, 1898, 1916, 1923, 1947, 1963, 2119, 2120]. standards [37]. standards-development [1097]. Stanford [47, 237, 1884]. Stars [1656, 2046]. Start [660, 1933]. Starting [1992]. Startups [2044]. State [436, 977, 1456, 1943]. Statecharts [1410]. Statements [1768]. Status [669, 821, 876]. Stay [1482]. Staying [2040]. Steam [705]. Steganography [1585]. Stenosis [124]. Step [310, 401, 866, 1300, 1375, 2117]. Steps [221, 770]. Stereograms [702]. Steve [1921]. Still [944, 1347, 1993]. Stone [259]. stop [660]. Storage [413, 586, 815, 1589, 1708, 1783, 1803, 1818, 1923, 2094]. Stories [1936]. Story [1797]. strategic [1270]. Strategies [92, 195, 587, 1137, 1251, 1361]. Strategy [1195, 1454, 1647]. Stream [113, 591]. Stream-Cipher [113]. Streamed [1715]. Streaming [1725, 2132]. Streamlining [728]. Street [1778]. Stress [1475]. Stresses [1385]. Stretch [2103]. Stride [1845]. Stripping [1841]. Strives [1355]. Stroll [2063]. Strong [1345, 2010]. Stroustrup [1627, 1704]. Structure [1424, 2032]. Structured [16, 17, 109, 172, 912, 1129]. Structures [514, 1439, 1696]. Struggle [1452]. Student [1671, 2072, 2084, 2111]. Students [1343, 1467, 1933]. Studies [687]. Study [3, 60, 427, 487, 700, 738, 861, 948, 1245, 1392, 1434, 1508, 1713]. Stuff [1388]. Stumble [1343]. Style [151, 2119]. Styles [735, 1341]. Submicron [2100]. Submit [1778]. Subscribers [1505]. Subscriptions [1429]. Subsystem [583, 584]. Subsystems [586]. Success [793, 1150, 1254, 1371, 1551, 1677, 1707, 1775, 1936, 1964]. Successful [707, 2023]. Successfully [683]. Such [1798]. Sue [1943]. Suggestions [1367]. SUIF [1265]. Suitable [1705, 1890]. Suite [1724, 1771, 1962]. Suites [1496, 2039]. suits [873]. Summaries [442, 843, 885, 908, 945, 969, 995, 1025, 1048, 1103, 1126, 1149, 1171, 1192, 1215, 1237, 1408, 1433, 1455, 1483, 1506, 1530, 1555, 1607, 1635, 1663, 1685, 1711, 1737, 1758, 1779, 1851, 1876, 1901, 1926, 1950, 1976, 1997, 2025, 2050, 2075, 2099, 2131]. Summary [198]. Summer [1157]. Summit [1757]. Sun [1554, 1606, 1778, 1898, 1962, 2039]. superchip [1188]. Supercomputer [981]. Supercomputers [1849, 2135]. Supercomputing [214, 462, 715, 1224, 1513, 1541]. Superhighway [1090, 1632]. Superimposed [291]. Superlatives [1444]. supersafe [1037]. Superspeculative [1462]. supervised [271]. Support [31, 93, 268, 280, 281, 308, 337, 374, 592, 619, 622, 1153, 1284, 1421, 1671, 1738, 1903, 1947, 1956, 1970]. Supported [375, 616, 617]. supporters [626]. Supporting [70, 92, 459, 1294, 1486]. Supremacy [1973]. Surgery [996, 999, 1000]. Surgical [997, 1002]. Surrogate [1076]. Surveillance [1994]. Survey [7, 40, 85, 169, 351, 389, 397, 633, 734, 829, 1299, 1419, 1511, 1755, 1761]. surveyed [880]. Survival [1475]. Surviving [1612, 1931]. Svoboda [1571]. Swing [158]. Switch [601]. Switched [486, 1346]. Switching [1276, 1591]. sword [833]. Symantec [1519]. Symbolic [1055]. symmetrical [986]. Symposium [1858, 2006]. Sync [1397]. Synchronization [44, 82, 190]. Synchronous [620, 1765]. Synergistic [1313]. Synergy [1484]. Synthesis [153, 157, 430, 770, 1359, 1812, 2014]. System [16, 18, 31, 32, 38, 48, 67, 69, 74, 75, 81, 94, 114, 121, 123, 125, 131, 133, 144, 152, 154, 157, 159, 167, 196, 201, 208, 234, 272, 287, 288, 291, 292, 294, 295, 306, 307, 315, 316, 336, 462, 570, 587, 590, 592, 670, 769, 776, 817, 828, 872, 887, 921, 948, 954, 970, 1027, 1056, 1092, 1155, 1206, 1238, 1240, 1305, 1380, 1440, 1454, 1531, 1602, 1613, 1659, 1698, 1702, 1716, 1736, 1753, 1791, 1836, 1913, 1932, 1962, 1980, 1988, 2039, 2041, 2090, 2145]. System-Level [769, 776]. System-on-Chip [2145]. Systematic [501, 588, 1350, 2091]. Systems [14, 28, 52, 56, 82, 88, 99, 103, 119, 130, 132, 134, 135, 137, 140, 146, 164–166, 170, 171, 178, 181, 185, 186, 199, 200, 210, 216, 224, 235, 236, 257, 276, 278, 280, 286, 325, 364, 365, 376, 383, 396, 460, 489, 501, 527, 537, 555, 569, 621, 622, 735, 768, 825, 886, 916, 918, 946, 951, 971, 1029, 1035, 1129, 1242, 1288, 1309, 1313, 1350, 1358, 1359, 1436, 1439, 1446, 1449, 1456, 1512, 1545, 1590, 1598, 1638, 1639, 1641, 1642, 1645, 1653, 1700, 1701, 1787, 1812, 1828, 1852, 1858, 1891, 1903, 1970, 2007, 2027, 2076]. Systems} [50, 86, 118, 225, 230– 232, 261, 267, 271, 281, 283, 363, 416, 450, 551, 793, 814, 816, 818, 899, 976, 1697, 1724, 1771, 1812, 1834, 1956, 1962, 2064]. Systolic [22, 524]. Systolic-Type [22].

T [357, 1492]. T. [1733]. TAB [1745]. Tabak [358]. Table [2070]. Take [1467, 1759, 1896, 2063]. TakeFive [1724]. Takes [842, 1874, 2094]. Taking [1290, 1854, 2063]. Talarian [1988]. Tale [635, 904, 2121]. Taligent [791]. Talk [68]. Tamagotchi [1524]. Tangled [1426]. Tango [74]. Tanner [1724]. Tape [1929]. Target [1603]. Targets [2022]. Task [24, 244, 245, 971, 1183]. Task-Flow [244, 245]. Tasks [224, 225, 623]. Taulbee [85, 351]. taxes [810]. Taxing [1299]. Taxonomic [18, 38]. Taxonomy [5, 165, 236, 537, 976, 1115]. Taylor [1384, 1492]. TC [876, 930, 985, 1015, 1035, 1067, 1288, 1355, 1385, 1440, 1513]. TCP [907]. TCP/IP [907]. tea [807]. Teacher [2009]. Teaching [818, 1271, 1508]. Team [849, 1898, 1913]. Teams [377]. Teamwork [2144]. Tech [1290, 1583, 1820]. Technical [287, 298, 639, 838, 855, 876, 896, 930, 961, 985, 1015, 1035, 1067, 1098, 1117, 1139, 1162, 1183, 1205, 1228, 1250, 1270, 1309, 1355, 1384, 1385, 1440, 1467, 1491, 1492, 1513, 1514, 1570, 1592, 1616, 1650, 1718, 1745, 1766, 1788, 1809, 1828, 1858, 1859, 1909, 1941, 1957, 2008, 2113].

Technique [246]. Techniques [5, 56, 59, 213, 388, 397, 637, 1238, 1271, 1353, 1383, 1409, 1458]. Techno [1473]. Technological [1879]. Technologies [62, 91, 737, 792, 1099, 1697, 1723, 1791, 1801, 1812, 1836, 1896, 1988, 2019, 2109, 2145]. Technologists [2070]. Technology [63, 66, 176, 195, 424, 426, 429, 514, 538, 913, 921, 1006, 1036, 1057, 1063, 1093, 1115, 1138, 1160, 1207, 1229, 1251, 1279, 1293, 1302, 1330, 1391, 1411, 1432, 1474, 1476, 1480, 1486, 1487, 1523, 1553, 1554, 1564, 1576, 1582, 1586, 1605, 1623, 1633, 1653, 1661, 1670, 1678, 1683, 1698, 1700, 1709, 1710, 1724, 1726, 1732, 1735, 1756, 1777, 1790, 1791, 1800, 1801, 1816, 1821, 1833, 1837, 1848, 1867, 1872, 1897, 1904–1906, 1918, 1922, 1937, 1946, 1962, 1966, 1972, 1981, 1988, 1993, 2006, 2014, 2016, 2020, 2046, 2047, 2067, 2094, 2095, 2106, 2107, 2122, 2127, 2147]. Technology [101, 370, 436, 541, 556, 692, 835, 842, 861, 898, 1060, 1271, 1664, 1698, 1792, 1881]. Technosoft [2118]. Telecommunications [485, 1236]. Teleimmersive [2137]. Telemetry [121]. Telephone [65, 1346, 1454, 2065, 2069]. Telephony [1214, 1345, 1873, 1897, 2014]. Teleporting [1303]. telepresence [1134]. television [814]. Tell [1726]. Temporal [189, 230]. Temporaries [1609]. Ten [806, 1043, 1370]. Tentative [2096]. Terabyte [1882]. teraflop [926]. Teraflops [1297]. Terasys [802]. term [315, 316]. Terminal [2014]. Terminology [217]. terms [1541]. Test [213, 569, 892, 1238, 1240, 1359, 1724, 1778, 1812, 1894, 1902, 1962, 1970, 2105, 2107– 2109, 2145]. Test-Pattern [2108]. Testable [59]. Testbeds [78]. Tester [2145]. Testing [83, 531, 688, 922, 1239, 1243, 1251, 1361, 1796, 1812, 1980, 1990, 2088, 2106, 2117, 2145]. Tests [1716]. Tetherless [1778]. Text [63, 294, 295, 309, 1347, 1749, 1798]. Text-to-Speech [63]. Thanks [1367]. Thanksgiving [360]. Their [650]. Theme [1533–1538]. Theorems [2070]. Theoretical [721, 923]. Theory [23, 25, 229, 272, 357]. Therac [474]. Therac-25 [474]. Therapeutic [996]. There [1798]. Thermal [1791]. Thermodynamics [515]. Thesaurus [2118]. These [899, 1993]. Thin [1981]. Thing [1798]. Things [2117]. third [726, 1924]. third-party [726]. Thirst [1480]. Thompson [1954, 1970]. Those [1347, 1619]. Threat [1275, 1972]. Threaten [1299]. Three [209, 398–400, 1224, 1492, 2121]. Three-Dimensional [209]. Throughout [2129]. Throwing [1702]. thumb [1066]. Thuraisingham [1492]. TI [1188]. Tightly [39]. Time [14, 23, 50, 123, 134– 137, 140, 166, 197, 216, 230, 262, 289, 325, 334, 335, 363, 364, 416, 501, 551– 553, 569, 604, 638, 755, 766, 767, 825, 860, 883, 906, 1029, 1354, 1358, 1368, 1439, 1458, 1503, 1661, 1706, 1725, 1759, 1809, 1812, 1822, 1971, 1977, 2081, 2128, 2144]. Timed [553]. Timed-Token [553]. Times [1452]. Timetable [1308, 1593]. Timing [138, 1058]. TimingDesigner [978]. tiny [1190]. Titles [1649]. Today [275, 1014, 1344, 1863, 2119]. Together [879, 2144]. Token [553]. Tolerance [54, 58, 81, 1348, 1349, 1352]. Tolerant [5, 52, 53, 55, 56, 82, 280, 281, 551, 985, 1350, 1354]. tolls [1041]. Tom [1937]. Tommy [1829]. Tomorrow [1014, 1344, 1650, 1863]. Tonga [1454]. Too [1101, 1549, 1659]. Tool [12, 103, 138, 278, 570, 787, 949, 1153, 1351, 1415, 1697, 1698, 1700, 1724, 1771, 1785, 1791, 1812, 1834, 1837, 1933, 1937, 1962, 1988, 2014, 2039, 2064, 2087, 2118, 2145]. Toolkit [1937, 1988, 2064, 2087]. Toolkits [1988]. Tools [56, 114, 168, 279, 388, 618, 845, 931, 946, 1291, 1323, 1324, 1353, 1471, 1488, 1497, 1519, 1574, 1583, 1596, 1620, 1653, 1654, 1674, 1697– 1701, 1724, 1748, 1771, 1791, 1812, 1834– 1837, 1863, 1889, 1903, 1914, 1937, 1962, 1988, 2014, 2039, 2064, 2087, 2118, 2145, 2150]. Top [915, 1278, 1357, 1606, 2039, 2111]. Top-Level [1357]. Topics [352, 353, 370, 386–388, 398–400, 416, 436, 449, 464, 477, 478, 490, 503, 518, 530, 543, 560, 573, 582, 658, 675, 693, 707, 726, 741, 758, 775, 792, 808, 853, 877, 899, 1008, 2113]. tor [281]. Toshiba [1898]. Total [1723]. Tough [2105]. TouringMachines [265]. Towrad [1987]. Trace [1463]. Traceability [1199]. Tracing [104, 446, 1382]. Track [120, 1227]. Tracking [650]. trade [1016]. Trading [1932]. Traffic [1821, 1994]. Training [849, 1052, 1133]. Transaction [146, 170, 202, 206, 1248]. Transaction-Processing [170]. transducer [251, 252]. Transfer [191, 835, 861]. Transform [1709]. Transformation [2070]. Transforming [1904]. Transim [2145]. Transistor [1460, 1464]. Transistors [1461, 1681]. Transition [910, 1057, 1060]. Translating [1686]. Translation [41, 144, 1152]. Translation-Lookaside [41]. Transportation [1614]. Travel [2097]. Traveled [1895]. Traveling [2144]. TreadMarks [1026]. Treasurer [483, 645, 840, 1146, 1416]. Treating [125, 846]. Trends [370, 599, 614, 632, 648, 666, 681, 699, 715, 732, 749, 765, 783, 800, 824, 842, 884, 907, 922, 944, 968, 994, 1024, 1047, 1073, 1102, 1125, 1148, 1188, 1191, 1214, 1236, 1256, 1276, 1298, 1302, 1318, 1344, 1375, 1431, 1453, 1481, 1504, 1528, 1535, 1552, 1557– 1564, 1581, 1604, 1632, 1660, 1682, 1708, 1734, 1740, 1776, 1799, 1820, 1847, 1871, 1896, 1921, 1945, 1971, 2019, 2045, 2071, 2094, 2126]. Trial [1157]. Tribes [1524]. Tribute [1615, 1909]. Tripp [1492]. True [1970]. Truly [1347]. Trust [1478, 1499, 1955]. Trusted [1678]. Trusting [1304]. Truth [1732]. truths [638]. Tsai [50, 1492]. TTP [551]. Tuning [803, 1175]. Turing [1902]. Turnaround [1454, 1849]. turning [991]. Tutorial [815, 831, 1051, 1263, 1858].

Tutorials [1933]. TV

[813, 927, 1295, 1299, 1554]. Twenty [1232]. Twenty-fifth [1232]. twice [1160]. Two [142, 247, 248, 277, 303, 352, 353, 635, 845, 1451, 1909, 1944]. Two-Dimensional [247, 248, 277]. Two-Handed [845]. Two-View [142]. Type [22, 1972]. Types [899, 1479]. typesetting [148]. Typing [90].

U [2053]. U.S. [1580]. UbiNet [2092]. Ubiquitous [518, 1303, 2024, 2092]. Ubiquity [1500]. UL [2064, 2087]. Ultimate [1475]. Ultrafast [1849]. Ultrareliable [135]. UML [1141, 1812, 2013]. Uncertainty [1800]. Uncovering [2074]. Underestimating [1681]. Undergraduate [973, 1671]. Understanding [218, 219, 226, 298, 1348]. Unfolding [185]. Unicode [1347, 1394]. Union [1299]. Uniprocessor [1461]. Uniscape [1697, 1812]. Unit [197]. units [323, 324]. Universal [1091, 1334, 1873]. Universities [447]. University [498, 2005, 2118]. UNIX [10, 526, 878, 1299, 1697, 1822, 1944, 1954, 1970, 2117]. Unix-to-NT [1697]. Unpredictable [1282]. Unseen [1585]. Unveil [1505, 1710, 1849]. Unveils [1454, 1554]. Update [312, 403, 420, 434, 467, 491, 505, 519, 559, 576, 640, 691, 711, 723, 740, 760, 779, 796, 807, 836, 854, 880, 898, 907, 928, 956, 981, 1007, 1034, 1059, 1091, 1113, 1134, 1158, 1159, 1179, 1180, 1202, 1225, 1246, 1266, 1307, 1356, 1384, 1418, 1441, 1468, 1492, 1514, 1541, 1571, 1593, 1615, 1649, 1671, 1693, 1719–1721, 1745, 1829]. Upgrade [944, 1937]. Upgrading [1818]. Urban [2117]. Urgency [19, 37]. URLs [1478, 1600]. US$25 [2111]. USA [1893, 1941, 2006]. Usability [233, 1027]. Usage [1791]. USB [1279]. Use [227, 445, 1081, 1345, 1431, 1432, 1441, 1454, 1505, 1578, 1811, 1913]. used [450]. Useful [1085]. User [86, 93, 179, 525, 786, 895, 1178, 1304, 1508, 1549, 1806, 1833, 1836]. user-definable [786]. User-Interface [179, 525]. User-Level [1806]. Users [132, 1073]. Uses [120]. Using [16, 97, 110, 124, 129, 140, 143, 151, 205, 249, 250, 254, 267, 278, 279, 304, 305, 670, 684, 686, 892, 921, 998, 1064, 1087, 1109, 1176, 1305, 1438, 1609, 1648, 1688, 1691, 1713, 1884, 2004, 2033, 2144]. Utilities [1062, 1948].

V [239]. Validating [671, 828, 1664]. Validation [1241, 2083]. Value [189, 874]. Value-Added [189]. Valued [1385]. Values [1419]. VBX [2044]. VDSM [2109]. Vector [1056]. Vegas [2006]. Vehicle [1412]. Vending [1992]. Vendors [1345, 1452, 1454, 1604, 1710, 1872, 2095]. Verification [89, 359, 2083]. Verilog [1812, 1962, 2063]. Version [339, 978, 1345, 1515]. Versioning [2003]. Versus [319, 915, 1333, 1474, 1525, 2018]. Very [263, 2029]. VHDL [909]. Video [602, 887, 1107, 1882]. Videoconferencing [758]. VideoSoft [2118]. Videotape [1703]. Vietnam [1057]. View [94, 142]. Viewpoint [431]. Viewpoints [184]. views [398–400]. Vigilence [2110]. village [300]. Virtual [285, 398–400, 829, 844, 846, 848–850, 867, 878, 913, 951, 997, 1047, 1504, 1509, 1557–1564, 1687, 1729, 1763, 1807, 2013, 2053, 2088, 2117]. Virtual-Environment [829]. Virtual-Human [2013]. Virtual-U [2053]. Virtue [2134]. Virus [1972]. Visa [2022]. Visas [1482, 1733, 1820]. Visibility [2087]. Vision [218–220, 223– 228, 1342, 1411, 1412, 1522, 1714, 2119]. Vision-Based [1412]. Visual [306, 307, 323, 324, 784–789, 902, 947, 1185, 1311, 1413, 1519, 1700, 1724, 1824, 2039]. Visualizable [656]. Visualization [180, 463, 512, 537, 649, 652, 653, 655, 657, 658, 861, 948, 974, 1001, 1086, 1382, 1415, 1744, 1785, 2133, 2134, 2136, 2137]. Visualization-Based [974]. Visualizations [654]. Visualizing [464, 650]. VLSI [8, 58, 249, 250, 284, 431, 1788]. VMW [974]. Vocabulary [621]. Voice [61, 62, 70, 116, 1430, 1822]. Voice-Based [1430]. Voice-Data [116]. Volume [215, 476, 651, 1827, 2138]. Volumetric [651]. Volunteer [1860, 2140]. Voyager [1388]. VR [853]. VRML [1345, 1482, 1710]. vs [1277, 1630]. VSIA [1979]. Vu [2124]. Vulnerabilities [1760]. Vulnerability [1299]. VXML [1973]. Vying [1978].

W [340, 357]. W3C [1634, 1662]. Wafer [242, 243, 246–252, 254, 1662]. Wafer-Scale [242, 243, 246–252, 254]. Wafers [1454]. Wah [1745]. Waiting [1993]. Walk [1458]. Walk-Time [1458]. Wall [1778]. want [1073]. wanted [706]. Wants [1822, 1994]. WAP [2013]. War [1482, 1626, 1777, 1778]. warehouse [875]. Warfare [1931]. Warning [1299]. Wars [1505, 1822]. Was [1332, 1347]. wasteland [1147]. Watch [1294, 1315, 1331, 1362, 1392, 1402, 1448, 1475, 1500, 1549, 1577, 1600, 1624, 1657, 1679, 1705, 1728, 1752, 1773, 1794, 1814, 1841, 1865, 1890, 1915, 1938, 1965, 1989, 2015, 2040, 2065, 2089, 2119, 2146]. Watered [1369]. watermarking [1163]. Wave [1724]. Wavelets [1696]. Way [977, 1560, 1798, 1961]. Ways [657, 1856]. Weak [2074]. weakest [615]. Wearable [1300, 1856, 1957]. Weaves [1426]. Web [703, 820, 953, 1135, 1172, 1184, 1206, 1214, 1284, 1319, 1331, 1388, 1426, 1448, 1482, 1567, 1624, 1679, 1684, 1705, 1723, 1724, 1757, 1761, 1773, 1778, 1792, 1812, 1814, 1822, 1841, 1868, 1890, 1932, 1937, 1962, 1965, 1994, 2032, 2039, 2063, 2072, 2087, 2118, 2126]. Web-Based [1624, 1812, 1932, 1965]. Web-Oriented [2118]. Web-Savvy [1962]. Web-Searching [1757]. Web-to-Legacy [1724]. WebCompiler [1962]. Weighty [1359]. Welcome [1273, 2070]. Wheelchairs [1020]. Wheels [1567, 1868]. Where [171, 672, 805, 823, 867, 906, 919, 942, 2123, 2144]. White [1791]. Who [1579, 1671]. Whole [1797]. whom [753, 1041]. Wibu [1771]. Wide [703, 820, 953, 1284, 1834, 1951]. Wide-Area [1951]. Widens [1660]. Will [906, 915, 1161, 1190, 1345, 1457, 1459, 1480, 1505, 1709, 1710, 1820, 1843, 1869, 1897, 1970, 2047, 2092, 2130]. William [1621]. Window [67, 69, 214, 383, 2088]. Windows [313, 626, 866, 900, 905, 1062, 1371, 1782– 1784, 1791, 1937]. Windoze [1798]. Winners [385, 557, 756, 1005, 1266, 1286, 1541, 1569, 1671, 1745]. wins [1541]. Wintel [1579, 2018]. WinWin [1713]. WIPO [1299]. Wired [1147, 1381]. Wireless [1188, 1294, 1537, 1605, 1662, 1710, 1981, 2020, 2092, 2120]. Wiretap [2128]. wisdom [275]. Without [1392, 1929]. Witnessed [1902]. Women [1671]. wonderland [1147]. Work [65, 616, 617, 848, 1106, 1339, 1387, 1485, 1554, 1558, 1629, 1675, 1712, 1913, 1917, 1923]. Work-Centered [1106]. Workbench [158, 848, 974]. Worked [1368]. Worker [1684, 1732]. Workers [1374, 1482]. Workforce [1706]. Working [2117]. Workloads [1459]. Workplace [2063]. Workstation [9, 294, 295]. Workstation-Based [294, 295]. Workstations [70, 1026, 1554, 1858]. World [703, 753, 820, 953, 965, 1147, 1190, 1284, 1309, 1381, 1631, 1696, 1790, 1798, 1863, 1924, 1975, 2088]. World-Class [1631]. Worlds [302, 850]. Worm [2022]. Wormhole [397, 972]. Wormhole-Routed [972]. worrying [660]. Worth [1477]. worthwhile [1166]. Wrapping [1454]. write [35]. Writes [1451]. Writing [1153]. Wrong [1123, 1169]. wrung [727]. WSI [244, 253, 255, 429]. WST [245]. WW [1179, 1829]. WWW [892, 1014, 1221].

X [30]. X-kernel [30]. XML [1634, 1696, 1728, 1794, 1923, 1947, 2014, 2047]. Xynetix [1724].

Y2K [1365, 1367, 1553, 1615, 1778, 1785, 1849, 1850, 1855, 1869, 1873– 1875, 1899, 1900, 1913, 1924, 1925, 1948, 1949, 1974, 1975, 1995, 1996, 2023, 2024, 2048, 2049, 2073, 2074, 2097, 2098, 2129, 2130]. Y2K-Compliant [1615]. Y2KA [1850]. ya [795]. Yea [854]. Year [1295, 1322–1324, 1482, 1577, 1897, 2026]. Yearlong [992]. Years [1089, 1111, 1216, 1257, 1274, 1370, 1637, 1993]. YES [1949]. Yield [246]. yonder [313]. Yourdon [238]. Youth [1098]. you’ve [706]. Yuforic [1467].

ZIP [292]. ZS [105]. ZS-1 [105]. Zuse [1033].

References

McLean:1990:SMC

[1] John McLean. The specification and modeling of computer security. Computer, 23(1):9–16, January 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parnas:1990:ECP

[2] David Lorge Parnas. Education for computing professionals. Computer, 23 (1):17–22, January 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crosby:1990:HDW

[3] Martha E. Crosby and Jan Stelovsky. How do we read algorithms? A case study. Computer, 23(1):25–35, January 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Markas:1990:DFS

[4] Tassos Markas, Mark Royals, and Nick Kanopoulos. On distributed fault simulation. Computer, 23(1):40–52, January 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chean:1990:TRT

[5] Mengly Chean and Jose A. B. Fortes. A taxonomy of reconfiguration techniques for fault-tolerant processor arrays. Computer, 23(1):55–69, January 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buckley:1990:SSE

[6] F. J. Buckley. Standards: a standard environment for software production. Computer, 23(1):75–77, January 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Duncan:1990:SPC

[7] Ralph Duncan. A survey of parallel computer architectures. Computer, 23 (2):5–16, February 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foo:1990:DCA

[8] Simon Y. Foo and Yoshiyasu Takefuji. Databases and cell-selection algorithms for VLSI cell libraries. Computer, 23 (2):18–30, February 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walters:1990:DBM

[9] Richard F. Walters. Design of a bitmapped multilingual workstation. Computer, 23(2):33–41, February 1990.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leler:1990:LMU

[10] W. Leler. Linda meets Unix. Computer, 23(2):43–54, February 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liang:1990:PGG

[11] Luping Liang, Samuel T. Chanson, and Gerald W. Neufeld. Process groups and group communications: Classifications and requirements. Computer, 23(2): 56–66, February 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:AEP

[12] Anonymous. Applications environment profiles: A significant tool for simplifying and coordinating standards efforts. Computer, 23(2):69–??, February 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:PRa

[13] Anonymous. Product reviews. Computer, 23(2):78–??, February 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tsai:1990:NAM

[14] Jeffrey J. P. Tsai, Kwang-Ya Fang, and Horng-Yuan Chen. A noninvasive architecture to monitor real-time distributed systems. Computer, 23 (3):11–23, March 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Czejdo:1990:GDM

[15] Bogdan Czejdo, Ramez Elmasri, Marek Rusinkiewicz, and David W. Embley. A graphical data manipulation language for an extended entityrelationship model. Computer, 23(3): 26–36, March 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leung:1990:ESU

[16] K. S. Leung and M. H. Wong. An expert-system shell using structured knowledge: An object-oriented approach. Computer, 23(3):38–47, March 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wasserman:1990:OSD

[17] Anthony I. Wasserman, Peter A. Pircher, and Robert J. Muller. The object-oriented structured design notation for software design representation. Computer, 23(3):50–63, March 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dasgupta:1990:HTS

[18] Subrata Dasgupta. A hierarchical taxonomic system for computer architectures. Computer, 23(3):64–74, March 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McFarland:1990:SUE

[19] Michael C. McFarland. Standards: Urgency of ethical standards intensifies in computer community. Computer, 23 (3):77–81, March 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:PRb

[20] Anonymous. Product reviews. Computer, 23(3):87–??, March 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rettberg:1990:MPP

[21] Randall D. Rettberg, William R. Crowther, Philip P. Carvey, and Raymond S. Tomlinson. The Monarch Parallel Processor hardware design. Computer, 23(4):18–28, 30, April 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moreno:1990:MCS

[22] Jaime H. Moreno and Tomas Lang. Matrix computations on systolic-type meshes: An introduction to the multimesh graph method. Computer, 23 (4):32–51, April 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sha:1990:RTS

[23] Lui Sha and John B. Goodenough. Real-time scheduling theory and Ada. Computer, 23(4):53–62, April 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1990:TR

[24] Alan Jay Smith. The task of the referee. Computer, 23(4):65–71, April 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pavlidis:1990:FBC

[25] Theo Pavlidis, Jerome Swartz, and Ynjiun P. Wang. Fundamentals of bar code information theory. Computer, 23 (4):74–86, April 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:SEL

[26] Anonymous. A standard for extremely low frequency magnetic fields; standards for public key encryption algorithms to be discussed. Computer, 23 (4):95–??, April 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:PRc

[27] Anonymous. Product reviews. Computer, 23(4):105–??, April 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boykin:1990:GEI

[28] Joseph Boykin and Susan J. LoVerso. Guest editors’ introduction: Recent developments in operating systems. Computer, 23(5):5–8, May 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Satyanarayanan:1990:SSH

[29] Mahadev Satyanarayanan. Scalable, secure, and highly available distributed file access. Computer, 23(5):9–18, 20–21, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peterson:1990:XPA

[30] Larry L. Peterson, Norman C. Hutchinson, Sean W. O’Malley, and Herman C. Rao. The X-kernel: A platform for accessing Internet resources. Computer, 23(5):23–33, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Black:1990:SSC

[31] David L. Black. Scheduling support for concurrency and parallelism in the Mach operating system. Computer, 23 (5):35–43, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mullender:1990:ADO

[32] Sape J. Mullender, Guido van Rossum, Andrew S. Tanenbaum, Robbert van Renesse, and Hans van Staveren. Amoeba: A distributed operating system for the 1990s. Computer, 23 (5):44–53, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stumm:1990:AID

[33] Michael Stumm and Songnian Zhou. Algorithms implementing distributed shared memory. Computer, 23(5): 54–64, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feitelson:1990:DHC

[34] Dror G. Feitelson and Larry Rudolph. Distributed hierarchical control for parallel processing. Computer, 23(5): 65–77, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:HWC

[35] Anonymous. How not to write commercial standards. Computer, 23 (5):79–??, May 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:PRd

[36] Anonymous. Product reviews. Computer, 23(5):87–??, May 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Newman:1990:CRU

[37] E. Newman and M. C. McFarland. Comments, with reply, on ‘Urgency of ethical standards intensifies in computer community’ (standards). Computer, 23(6):6, June 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmidt:1990:CRH

[38] U. Schmidt and S. Dasgupta. Comments, with reply on ‘A hierarchical taxonomic system for computer architectures’. Computer, 23(6):6, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dubois:1990:GEI

[39] Michel Dubois and Shreekant Thakkar. Guest editors’ introduction: Cache architectures in tightly coupled multiprocessors. Computer, 23(6):9–11, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stenstrom:1990:SCC

[40] Per Stenstrom. A survey of cache coherence schemes for multiprocessors. Computer, 23(6):12–24, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Teller:1990:TBC

[41] Patricia J. Teller. Translationlookaside buffer consistency. Computer, 23(6):26–36, June 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheong:1990:CDC

[42] Hoichi Cheong and Alexander V. Veidenbaum. Compiler-directed cache management in multiprocessors. Computer, 23(6):39–47, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chaiken:1990:DCC

[43] David Chaiken, Craig Fields, Kiyoshi Kurihara, and Anant Agarwal. Directorybased cache coherence in large-scale multiprocessors. Computer, 23(6):49– 58, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See corrigendum [51].

Graunke:1990:SAS

[44] Gary Graunke and Shreekant Thakkar. Synchronization algorithms for sharedmemory multiprocessors. Computer, 23 (6):60–69, June 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thakkar:1990:NDS

[45] Shreekant Thakkar, Michel Dubois, Anthony T. Laundrie, and Gurindar S. Sohi. Scalable shared-memory multiprocessor architectures. Computer, 23 (6):71–74, June 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

James:1990:DSS

[46] David V. James, Anthony T. Laundrie, Stein Gjessing, and Gurindar S. Sohi. Distributed-directory scheme: Scalable coherent interface. Computer, 23(6):74–77, June 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thapar:1990:DSS

[47] Manu Thapar and Bruce Delagi. Distributed-directory scheme: Stanford distributed-directory protocol. Computer, 23(6):78–80, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlton:1990:MBS

[48] Michael Carlton and Alvin Despain. Multiple-bus shared-memory system: Aquarius project. Computer, 23(6): 80–83, June 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buckley:1990:SES

[49] F. J. Buckley. Standards: Establishing a standard metrics program. Computer, 23(6):85–86, June 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ford:1990:CRN

[50] R. Ford. Comments, with reply, on “A noninvasive architecture to monitor real-time distributed systems”, by J. Tsai et al. Computer, 23(7):12–13, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chaiken:1990:CDC

[51] D. Chaiken, C. Fields, K. Kurihara, and A. Agarwal. Correction to ‘Directory-based cache coherence in large-scale multiprocessor’. Computer, 23(7):13, July 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [43].

Singh:1990:GEI

[52] Adit D. Singh and Singaravel Murugesan. Guest editors’ introduction: Fault-tolerant systems. Computer, 23 (7):15–17, July 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nelson:1990:FCF

[53] Victor P. Nelson. Fault-tolerant computing: Fundamental concepts. Computer, 23(7):19–25, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siewiorek:1990:FTC

[54] Daniel P. Siewiorek. Fault tolerance in commercial computers. Computer, 23(7):26–37, July 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laprie:1990:DAH

[55] Jean-Claude Laprie, Jean Arlat, Christian Beounes, and Karama Kanoun. Definition and analysis of hardwareand software-fault-tolerant architectures. Computer, 23(7):39–51, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geist:1990:REF

[56] Robert Geist and Kishor S. Trivedi. Reliability estimation of fault-tolerant systems: Tools and techniques. Computer, 23(7):52–61, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fujiwara:1990:ECC

[57] Eiji Fujiwara and Dhiraj K. Pradhan. Error-control coding in computers. Computer, 23(7):63–72, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koren:1990:FTV

[58] Israel Koren and Adit D. Singh. Fault tolerance in VLSI circuits. Computer, 23(7):73–83, July 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCluskey:1990:DTT

[59] E. J. McCluskey. Design techniques for testable embedded error checkers. Computer, 23(7):84–88, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isaak:1990:SHP

[60] J. Isaak. Standards: the history of Posix: a study in the standards process. Computer, 23(7):89–92, July 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kamel:1990:GEI

[61] Ragui Kamel. Guest editor’s introduction: Voice in computing. Computer, 23(8):8–??, August 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Strathmeyer:1990:VCO

[62] Carl R. Strathmeyer. Voice in computing: An overview of available technologies. Computer, 23(8):10–15, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OMalley:1990:TCT

[63] Michael H. O’Malley. Text-to-speech conversion technology. Computer, 23 (8):17–23, August 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peacocke:1990:ISS

[64] Richard D. Peacocke and Daryl H. Graf. An introduction to speech and speaker recognition. Computer, 23(8): 26–33, August 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lennig:1990:PSR

[65] Matthew Lennig. Putting speech recognition to work in the telephone network. Computer, 23(8):35–41, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nakatsu:1990:AAS

[66] Ryohei Nakatsu. Anser: An application of speech technology to the Japanese banking industry. Computer, 23(8): 43–48, August 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmandt:1990:AWS

[67] Chris Schmandt, Mark S. Ackerman, and Debby Hindus. Augmenting a window system with speech input. Computer, 23(8):50–56, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Salisbury:1990:TDB

[68] Mark W. Salisbury, Joseph H. Hendrickson, Terence L. Lammers, Caroline Fu, and Scott A. Moody. Talk and draw: Bundling speech and graphics. Computer, 23(8):59–65, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ludwig:1990:ENW

[69] Lester F. Ludwig, Natalio Pincever, and Michael Cohen. Extending the notion of a window system to audio. Computer, 23(8):66–72, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kamel:1990:PSV

[70] Ragui Kamel, Kamyar Emami, and Robert Eckert. PX: Supporting voice in workstations. Computer, 23(8): 73–80, August 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:IPS

[71] Anonymous. IEEE Project 802 standards efforts. Computer, 23(8):84–??, August 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:PRe

[72] Anonymous. Product reviews. Computer, 23(8):103–??, August 1990. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wing:1990:SIF

[73] Jeanette M. Wing. A specifier’s introduction to formal methods. Computer, 23(9):8, 10–22, 24, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stasko:1990:TFS

[74] John T. Stasko. Tango: A framework and system for algorithm animation. Computer, 23(9):27–39, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Champine:1990:PAD

[75] George A. Champine, Daniel E. Geer, Jr., and William N. Ruh. Project Athena as a distributed computer system. Computer, 23(9):40–51, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ziegler:1990:MCL

[76] Chaim Ziegler and Gerald Weiss. Multimedia conferencing on local area networks. Computer, 23(9):52–61, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mink:1990:MPI

[77] Alan Mink, Robert J. Carpenter, George G. Nacht, and John W. Roberts. Multiprocessor performancemeasurement instrumentation. Computer, 23(9):63–75, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:GNT

[78] Anonymous. Gigabit network testbeds. Computer, 23(9):77–80, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:S

[79] Anonymous. Standards. Computer, 23 (9):82–??, September 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Das:1990:DMM

[80] Chita R. Das, Jeffrey T. Kreulen, Matthew J. Thazhuthaveetil, and Laxmi N. Bhuyan. Dependability modeling for multiprocessors. Computer, 23 (10):7–19, October 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ancona:1990:SAF

[81] Massimo Ancona, Gabriella Dodero, Vittoria Gianuzzi, Andrea Clematis, and Eduardo B. Fernandez. A system architecture for fault tolerance in concurrent software. Computer, 23(10): 23–32, October 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramanathan:1990:FCS

[82] Parameswaran Ramanathan, Kang G. Shin, and Ricky W. Butler. Faulttolerant clock synchronization in distributed systems. Computer, 23(10): 33–42, October 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Franklin:1990:BSR

[83] Manoj Franklin and Kewal K. Saluja. Built-in self-testing of random-access memories. Computer, 23(10):45–56, October 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Manber:1990:CRN

[84] Udi Manber. Chain reactions in networks. Computer, 23(10):57–63, October 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1990:TSR

[85] David Gries and Dorothy Marsh. The 1988–89 Taulbee Survey Report. Computer, 23(10):65–71, October 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:UIM

[86] Anonymous. User interface management systems and application portability. Computer, 23(10):73–??, October 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1990:POL

[87] T. L. (Frank) Pappas. PostScript output on a LaserJet printer. Computer, 23(10):82–88, October 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frieder:1990:MAR

[88] Ophir Frieder. Multiprocessor algorithms for relational-database operators on hypercube systems. Computer, 23(11):13–28, November 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zelkowitz:1990:FCM

[89] Marvin V. Zelkowitz. A functional correctness model of program verification. Computer, 23(11):30–39, November 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Darragh:1990:RKP

[90] John J. Darragh, Ian H. Witten, and Mark L. James. The reactive keyboard: A predictive typing aid. Computer, 23(11):41–49, November 1990.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Freundlich:1990:KBD

[91] Yehudah Freundlich. Knowledge bases and databases: Converging technologies, diverging interests. Computer, 23 (11):51–57, November 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mooney:1990:SSA

[92] James D. Mooney. Strategies for supporting application portability. Computer, 23(11):59–70, November 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1990:GCS

[93] Brad A. Myers, Dario A. Giuse, Roger B. Dannenberg, Brad Vander Zanden, David S. Kosbie, Edward Pervin, Andrew Mickish, and Philippe Marchal. Garnet: Comprehensive support for graphical, highly interactive user interfaces. Computer, 23(11):71– 85, November 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:AVJ

[94] Anonymous. An American’s view of the Japanese standards system. Computer, 23(11):87–??, November 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1990:PRf

[95] Anonymous. Product reviews. Computer, 23(11):94–??, November 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pancake:1990:DPL

[96] Cherri M. Pancake and Donna Bergmark. Do parallel languages respond to the needs of scientific programmers? Computer, 23(12):13– 23, December 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berlin:1990:CSC

[97] Andrew Berlin and Daniel Weise. Compiling scientific code using partial evaluation. Computer, 23(12):25–37, December 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skillicorn:1990:APC

[98] David B. Skillicorn. Architectureindependent parallel computation. Computer, 23(12):38–50, December 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawson:1990:PEC

[99] Harold W. Lawson. Philosophies for engineering computer-based systems. Computer, 23(12):52–63, December 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weicker:1990:OCB

[100] Reinhold P. Weicker. An overview of common benchmarks. Computer, 23 (12):65–75, December 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kemp:1990:SIT

[101] Alasdair Kemp. Standards: Information technology standardization is key area of interest to IEE. Computer, 23 (12):76–78, December 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patt:1991:GEI

[102] Yale N. Patt. Guest editor’s introduction: Experimental research in computer architecture. Computer, 24(1): 14–??, January 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pease:1991:PPE

[103] Daniel Pease, Arif Ghafoor, Ishfaq Ahmad, David L. Andrews, Kamal Foudil-Bey, Thomas E. Karpinski, Mohammad A. Mikki, and Mohamed Zerrouki. PAWS: A performance evaluation tool for parallel computing systems. Computer, 24(1):18–29, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stunkel:1991:ATP

[104] Craig B. Stunkel, Bob Janssens, and W. Kent Fuchs. Address tracing for parallel machines. Computer, 24(1): 31–38, January 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mangione-Smith:1991:PCI

[105] William Mangione-Smith, Santosh G. Abraham, and Edward S. Davidson. A performance comparison of the IBM RS/6000 and the Astronautics ZS-1. Computer, 24(1):39–46, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conte:1991:BC

[106] Thomas M. Conte and Wen mei W. Hwu. Benchmark characterization. Computer, 24(1):48–56, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mudge:1991:DM

[107] Trevor N. Mudge, Richard B. Brown, William P. Birmingham, Jeffrey A. Dykstra, Ayman I. Kayssi, Ronald J. Lomax, Oyekunle A. Olukotun, Karem A. Sakallah, and Raymond A. Milano. The design of a microsupercomputer. Computer, 24(1):57– 64, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Farrens:1991:IPP

[108] Matthew K. Farrens and Andrew R. Pleszkun. Implementation of the PIPE processor. Computer, 24(1): 65–70, January 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vranesic:1991:HHS

[109] Zvonko G. Vranesic, Michael Stumm, David M. Lewis, and Ron White. Hector: A hierarchically structured sharedmemory multiprocessor. Computer, 24 (1):72–79, January 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gokhale:1991:BUHb

[110] Maya Gokhale, William Holmes, Andrew Kopser, Sara Lucas, Ronald Minnich, Douglas Sweely, and Daniel Lopresti. Building and using a highly parallel programmable logic array. Computer, 24(1):81–89, January 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:DSC

[111] Anonymous. Do standards cause software productivity problems? Computer, 24(1):97–??, January 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PRa

[112] Anonymous. Product reviews. Computer, 24(1):107–??, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zeng:1991:PBG

[113] Kencheng Zeng, Chung-Huang Yang, Dah-Yea Wei, and T. R. N. Rao. Pseudorandom bit generators in streamcipher cryptography. Computer, 24(2): 8–17, February 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mettrey:1991:CEE

[114] William Mettrey. A comparative evaluation of expert system tools. Computer, 24(2):19–31, February 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheriton:1991:PHS

[115] David R. Cheriton, Hendrik A. Goosen, and Patrick D. Boyle. Paradigm: A highly scalable shared-memory multicomputer architecture. Computer, 24 (2):33–46, February 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mukherjee:1991:IVC

[116] Biswanath Mukherjee. Integrated voice-data communication over highspeed fiber optic networks. Computer, 24(2):49–58, February 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Caldiera:1991:IQR

[117] Gianluigi Caldiera and Victor R. Basili. Identifying and qualifying reusable software components. Computer, 24(2): 61–71, February 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McFarland:1991:SES

[118] Michael C. McFarland. Standards: Ethics and the safety of computer systems. Computer, 24(2):72–75, February 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kriewall:1991:GEI

[119] Timothy J. Kriewall and John M. Long. Guest editors’ introduction: Computer-based medical systems. Computer, 24(3):9–13, March 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1991:EMM

[120] Glenn A. Myers, Keith R. Sherman, and Lawrence Stark. Eye monitor: Microcomputer-based instrument uses an internal model to track the eye. Computer, 24(3):14–21, March 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fernald:1991:MIT

[121] Kenneth W. Fernald, Todd A. Cook, Thomas K. Miller, III, and John J. Paulos. A microprocessor-based implantable telemetry system. Computer, 24(3):23–30, March 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Manolakos:1991:PPB

[122] Elias S. Manolakos, Haris M. Stellakis, and Dana H. Brooks. Parallel processing for biomedical signal processing: Higher order spectral analysis — an application. Computer, 24 (3):33–43, March 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krieger:1991:NDR

[123] Don Krieger, Gerald Burk, and Robert J. Sclabassi. Neuronet: A distributed real-time system for monitoring neurophysiologic function in the medical environment. Computer, 24 (3):45–55, March 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cios:1991:UFS

[124] Krzysztof J. Cios, Inho Shin, and Lucy S. Goodenday. Using fuzzy sets to diagnose coronary artery stenosis. Computer, 24(3):57–63, March 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Poli:1991:NNE

[125] Riccardo Poli, Stefano Cagnoni, Riccardo Livi, Giuseppe Coppini, and Guido Valli. A neural network expert system for diagnosing and treating hypertension. Computer, 24(3):

64–71, March 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:ISE

[126] Anonymous. Internationalizing software engineering standards. Computer, 24(3):74–??, March 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PRb

[127] Anonymous. Product reviews. Computer, 24(3):86–??, March 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krick:1991:EIS

[128] Robert F. Krick and Apostolos Dollas. The evolution of instruction sequencing. Computer, 24(4):5–15, April 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abidi:1991:ARI

[129] Mongi A. Abidi, Richard O. Eason, and Rafael C. Gonzalez. Autonomous robotic inspection and manipulation using multisensor feedback. Computer, 24(4):17–31, April 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bertino:1991:ODM

[130] Elisa Bertino and Lorenzo Martino. Object-oriented database management systems: Concepts and issues. Computer, 24(4):33–47, April 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reibman:1991:RMO

[131] Andrew L. Reibman and Malathi Veeraraghavan. Reliability modeling: An overview for system designers. Computer, 24(4):49–57, April 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grudin:1991:ISB

[132] Jonathan Grudin. Interactive systems: Bridging the gaps between developers and users. Computer, 24 (4):59–69, April 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PEU

[133] Anonymous. Preparing for EC 1992 in the US through quality system registration. Computer, 24(4):70–??, April 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krishna:1991:GEI

[134] C. M. Krishna and Y. H. Lee. Guest editors’ introduction: Real-time systems. Computer, 24(5):10–11, May 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lala:1991:DAU

[135] Jaynarayan H. Lala, Richard E. Harper, and Linda S. Alger. A design approach for ultrareliable real-time systems. Computer, 24(5):12–22, May 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shin:1991:HDR

[136] Kang G. Shin. HARTS: A distributed real-time architecture. Computer, 24 (5):25–35, May 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Muppala:1991:RPP

[137] Jogesh K. Muppala, Steven P. Woolet, and Kishor S. Trivedi. Real-time-systems performance in the presence of failures. Computer, 24(5):37–47, May 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Park:1991:EPT

[138] Chang Yun Park and Alan C. Shaw. Experiments with a program timing tool based on source-level timing schema. Computer, 24(5):48–57, May 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:1991:ASI

[139] Jane W. S. Liu, Kwei-Jay Lin, Wei-Kuan Shih, Albert Chuang shi Yu, Jen-Yao Chung, and Wei Zhao. Algorithms for scheduling imprecise computations. Computer, 24(5):58–68, May 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kenny:1991:BFR

[140] Kevin B. Kenny and Kwei-Jay Lin. Building flexible real-time systems using the Flex language. Computer, 24 (5):70–78, May 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PCQ

[141] Anonymous. Point/counterpoint: Questions on ethical standards. Computer, 24(5):80–??, May 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brooks:1991:LTD

[142] Kenneth P. Brooks. Lilac: A twoview document editor. Computer, 24 (6):7–19, June 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBenedictis:1991:NAD

[143] Erik DeBenedictis, Sumit Ghosh, and Meng-Lin Yu. A novel algorithm for discrete-event simulation: Asynchronous distributed discrete-event simulation algorithm for cyclic circuits using a dataflow network. Computer, 24(6):21–33, June 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kitano:1991:DES

[144] Hiroaki Kitano. ΦDM-Dialog: An experimental speech-to-speech dialog translation system. Computer, 24 (6):36–50, June 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cordes:1991:LPP

[145] David Cordes and Marcus Brown. The literate-programming paradigm. Computer, 24(6):52–61, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leff:1991:CTP

[146] Avraham Leff and Calton Pu. A classification of transaction processing systems. Computer, 24(6):63–76, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PRc

[147] Anonymous. Product reviews. Computer, 24(6):88–??, June 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1991:TTD

[148] T. L. (Frank) Pappas. TEX typesetting with a difference. Computer, 24 (6):92–93, June 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baggi:1991:GEI

[149] Denis L. Baggi. Guest editor’s introduction: Computer-generated music. Computer, 24(7):6–??, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:1991:FPL

[150] David P. Anderson and Ron Kuivila. Formula: A programming language for expressive computer music. Computer, 24(7):12–21, July 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cope:1991:RMU

[151] David Cope. Recombinant music: Using the computer to explore musical style. Computer, 24(7):22–28, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1991:TES

[152] Margaret L. Johnson. Toward an expert system for expressive musical performance. Computer, 24(7):30–34, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dannenberg:1991:FFL

[153] Roger B. Dannenberg, Christopher Lee Fraley, and Peter Velikonja. Fugue: A functional language for sound synthesis. Computer, 24(7):36–42, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morita:1991:CMS

[154] Hideyuki Morita, Shuji Hashimoto, and Sadamu Ohteru. A computer music system that follows a human conductor. Computer, 24(7):44–53, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smoliar:1991:AMC

[155] Stephen W. Smoliar. Algorithms for musical composition: A question of granularity. Computer, 24(7):54–56, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smoliar:1991:POC

[156] Stephen W. Smoliar, Goffredo Haus, Alberto Sametti, Denis L. Baggi, Antonio Camurri, Corrado Canepa, Marcello Frixione, Renato Zaccaria, Yap Siong Chua, and Robert Keefe. Project overviews: Current research in computer-generated music. Computer, 24(7):54–75, July 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haus:1991:SSS

[157] Goffredo Haus and Alberto Sametti. Scoresynth: A system for the synthesis of music scores based on Petri nets and a music algebra. Computer, 24 (7):56–60, July 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baggi:1991:NIW

[158] Denis L. Baggi. Neurswing: An intelligent workbench for the investigation of swing in jazz. Computer, 24 (7):60–64, July 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Camurri:1991:HSI

[159] Antonio Camurri, Corrado Canepa, Marcello Frixione, and Renato Zaccaria. HARP: A system for intelligent composer’s assistance. Computer, 24 (7):64–67, July 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chua:1991:CBP

[160] Yap Siong Chua. Composition based on pentatonic scales: A computer-aided approach. Computer, 24(7):67–71, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keefe:1991:CMA

[161] Robert Keefe. Composing by musical analog: A look at planetary orbits. Computer, 24(7):72–75, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Newcomb:1991:SSM

[162] Steven R. Newcomb. Standards: Standard Music Description Language complies with hypermedia standard. Computer, 24(7):76–79, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PRd

[163] Anonymous. Product reviews. Computer, 24(7):81–??, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singhal:1991:GEI

[164] Mukesh Singhal and Thomas L. Casavant. Guest editor’s introduction: Distributed computing systems. Computer, 24(8):12–16, August 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:1991:OTD

[165] Bruce E. Martin, Claus H. Pedersen, and James Bedford-Roberts. An object-bases taxonomy for distributed computing systems. Computer, 24(8): 17–27, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fidge:1991:LTD

[166] Colin Fidge. Logical time in distributed computing systems. Computer, 24(8): 28–33, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sinha:1991:GDO

[167] Pradeep K. Sinha, Mamoru Maekawa, Kentaro Shimizu, Xiaohua Jia, Hyo Ashihara, Naoki Utsunomiya, Kyu S. Park, and Hirohiko Nakano. The Galaxy distributed operating system. Computer, 24(8):34–41, August 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marzullo:1991:TDA

[168] Keith Marzullo, Robert Cooper, Mark D. Wood, and Kenneth P. Birman. Tools for distributed application management. Computer, 24(8): 42–51, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nitzberg:1991:DSM

[169] Bill Nitzberg and Virginia Lo. Distributed shared memory: A survey of issues and algorithms. Computer, 24 (8):52–60, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mafla:1991:CFD

[170] Enrique Mafla and Bharat Bhargava. Communication facilities for distributed transaction-processing systems. Computer, 24(8):61–66, August 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ozsu:1991:DDS

[171] M. Tamer Özsu and Patrick Valduriez. Distributed database systems: Where are we now? Computer, 24(8): 68–78, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goldfarb:1991:SHS

[172] Charles F. Goldfarb. Standards: Hy-Time: a standard for structured hypermedia interchange. Computer, 24 (8):81–84, August 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:PRe

[173] Anonymous. Product reviews. Computer, 24(8):96–??, August 1991. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1991:ICS

[174] Merlin G. Smith. IEEE Computer Society: Four decades of service, 1951– 1991. Computer, 24(9):6–12, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aylor:1991:GEI

[175] James H. Aylor, Roy L. Russo, and Bruce D. Shriver. Guest editors’ introduction: The promise of the next decade. Computer, 24(9):15– 17, September 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hennessy:1991:CTA

[176] John L. Hennessy and Norman P. Jouppi. Computer technology and architecture: An evolving interaction. Computer, 24(9):18–29, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1991:CA

[177] Harold S. Stone and John Cocke. Computer architecture in the 1990s. Computer, 24(9):30–38, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gray:1991:HCS

[178] Jim Gray and Daniel P. Siewiorek. High-availability computer systems.

Computer, 24(9):39–48, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marcus:1991:UDN

[179] Aaron Marcus and Andries van Dam. User-interface developments for the nineties. Computer, 24(9):49–57, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nielson:1991:VSE

[180] Gregory M. Nielson. Visualization in scientific and engineering computation. Computer, 24(9):58–66, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wittie:1991:CND

[181] Larry D. Wittie. Computer networks and distributed systems. Computer, 24 (9):67–76, September 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Green:1991:FFC

[182] Paul E. Green. The future of fibre-optic computer networks. Computer, 24(9): 78–87, September 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Basili:1991:FGS

[183] Victor R. Basili and John D. Musa. The future engineering of software: A management perspective. Computer, 24(9): 90–96, September 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:1991:VC

[184] P. Treleaven, H. Aiso, M. A. Meth, A. K. Jones, R. H. Sprague, Jr., T. Van Duzer, G. E. Moore, A. C. Weaver, V. Cerf, M. R. Mercer, K. Boykin, D. Cheriton, T. G. Lewis, Luqi, and L. P. Deutsch. Viewpoints (computing in the 1990s). Computer, 24(9): 98–113, September 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:MIS

[185] Anonymous. Multimedia information systems: The unfolding of a reality. Computer, 24(10):6–??, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fox:1991:AID

[186] Edward A. Fox. Advances in interactive digital multimedia systems. Computer, 24(10):9–21, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meghini:1991:CMM

[187] Carlo Meghini, Fausto Rabitti, and Costantino Thanos. Conceptual modeling of multimedia documents. Computer, 24(10):23–30, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feiner:1991:AGC

[188] Steven K. Feiner and Kathleen R. McKeown. Automating the generation of coordinated multimedia explanations. Computer, 24(10):33–41, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Little:1991:SCD

[189] Thomas D. C. Little and Arif Ghafoor. Spatio-temporal composition of distributed multimedia objects for valueadded networks. Computer, 24(10): 42–50, October 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:1991:CMI

[190] David P. Anderson and George Homsy. A continuous media I/O server and its synchronization mechanism. Computer, 24(10):51–57, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Palaniappan:1991:IIM

[191] Murugappan Palaniappan and George Fitzmaurice. InternetExpress: An inter-desktop multimedia data-transfer service. Computer, 24(10):58–67, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vin:1991:MCE

[192] Herrick M. Vin, Polle T. Zellweger, Daniel C. Swinehart, and P. Venkat Rangan. Multimedia conferencing in the Etherphone environment. Computer, 24(10):69–79, October 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lander:1991:CMB

[193] Eric S. Lander, Robert Langridge, and Damian M. Saccocio. Computing in molecular biology: Mapping and interpreting biological information. Computer, 24(11):6–13, November 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Quisquater:1991:CLE

[194] Jean-Jacques Quisquater and Yvo G. Desmedt. Chinese Lotto as an exhaustive code-breaking machine. Computer, 24(11):14–22, November 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clemons:1991:CSI

[195] Eric K. Clemons. Corporate strategies for information technology: A resourcebased approach. Computer, 24(11): 23–32, November 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dasgupta:1991:CDO

[196] Partha Dasgupta, Richard J. LeBlanc, Jr., Mustaque Ahamad, and Umakishore Ramachandran. The Clouds Distributed Operating System. Computer, 24(11):34–44, November 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Factor:1991:RDF

[197] Michael Factor, David H. Gelernter, Craig E. Kolb, Perry L. Miller, and Dean F. Sittig. Real-time data fusion in the intensive care unit. Computer, 24(11):45–54, November 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tucker:1991:FDS

[198] Allen B. Tucker and Bruce H. Barnes. Flexible design: A summary of Computing Curricula 1991. Computer, 24 (11):56–66, November 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ram:1991:GEI

[199] Sudha Ram. Guest editor’s introduction: Heterogeneous distributed database systems. Computer, 24(12): 7–??, December 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:1991:CSD

[200] Won Kim and Jungyun Seo. Classifying schematic and data heterogeneity in multi database systems. Computer, 24(12):12–18, December 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:1991:PHM

[201] Rafi Ahmed, Philippe De Smedt, Weimin Du, William Kent, Mohammad A. Ketabchi, Witold A. Litwin, Abbas Rafii, and Ming-Chien Shan. The Pegasus Heterogeneous Multidatabase System. Computer, 24(12): 19–27, December 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Soparkar:1991:FTM

[202] Nandit Soparkar, Henry F. Korth, and Abraham Silberschatz. Failureresilient transaction management in multidatabases. Computer, 24(12):28– 36, December 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Noll:1991:IDI

[203] John Noll and Walt Scacchi. Integrating diverse information repositories: A distributed hypertext approach. Computer, 24(12):38–45, December 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rusinkiewicz:1991:SID

[204] Marek Rusinkiewicz, Amit Sheth, and George Karabatis. Specifying interdatabase dependencies in a multidatabase environment. Computer, 24 (12):46–53, December 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Collet:1991:RIU

[205] Christine Collet, Michael N. Huhns, and Wei-Min Shen. Resource integration using a large knowledge base in Carnot. Computer, 24(12):55– 62, December 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pu:1991:HAT

[206] Calton Pu, Avraham Leff, and Shu-Wie F. Chen. Heterogeneous and autonomous transaction processing. Computer, 24(12):64–72, December 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arnold:1991:SAI

[207] D. Arnold, P. Cannata, L. A. Glasson, G. Hallmark, B. McGuire, S. Newman, R. Odegard, and H. Sabharwal. SQL Access: an implementation of the ISO Remote Database Access Standard. Computer, 24(12):74– 78, December 1991. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harel:1992:BSB

[208] David Harel. Biting the silver bullet: Toward a brighter future for system development. Computer, 25(1):

8–20, January 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:1992:PAT

[209] Yulun Wang, Amante Mangaser, Partha Srinivasan, Steve Jordan, and Steven Butner. The 3DP: A processor architecture for three-dimensional applications. Computer, 25(1):25–36, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1992:ADS

[210] Thomas Y. C. Woo and Simon S. Lam. Authentication for distributed systems. Computer, 25(1):39–52, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See correction [232].

VanRoy:1992:HLP

[211] Peter Van Roy and Alvin M. Despain. High-performance logic programming with the Aquarius Prolog compiler. Computer, 25(1):54–68, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

VanRoy:1992:HPL

[212] Peter Van Roy and Alvin M. Despain. High-performance logic programming with the Aquarius Prolog compiler. Computer, 25(1):54–68, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klenke:1992:PTA

[213] Robert H. Klenke, Ronald D. Williams, and James H. Aylor. Parallelprocessing techniques for automatic test pattern generation. Computer, 25 (1):71–84, January 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1992:SSR

[214] Ware Myers. Supercomputing 91: Special report: High-performance computing as a window into the future. Computer, 25(1):87–90, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamming:1992:RMF

[215] R. W. Hamming. Review of Mathematical Foundations of Computer Science, Volume 1 (Fejer, P. A. and Simovici D. A; 1991). Computer, 25(1):134– 135, January 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RAD

[216] Janusz Zalewski. Review of Ada in Distributed Real-Time Systems (Nielsen, K; 1990). Computer, 25(1):135, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoevel:1992:OCT

[217] Lee Hoevel and Veljko Milutinovic. The Open Channel: Terminology risks with the RISC concept in the risky RISC arena. Computer, 25(1):136, January 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Choudhary:1992:GEI

[218] Alok Choudhary and Sanjay Ranka. Guest Editors’ introduction: Parallel processing for computer vision and image understanding. Computer, 25(2):

7–10, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Choudhary:1992:PPC

[219] Alok Choudhary and Sanjay Ranka. Parallel processing for computer vision and image understanding (special issue introduction). Computer, 25(2): 7–10, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marsh:1992:RCP

[220] Brian Marsh, Chris Brown, Thomas LeBlanc, Michael Scott, Tim Becker, Cesar Quiroz, Prakash Das, and Jonas Karlsson. The Rochester Checkers Player: Multimodel parallel programming for animate vision. Computer, 25 (2):12–19, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Webb:1992:STA

[221] Jon A. Webb. Steps toward architecture-independent image processing. Computer, 25(2):21–31, February 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rigoutsos:1992:MPM

[222] Isidore Rigoutsos and Robert Hummel. Massively parallel model matching: Geometric hashing on the Connection Machine. Computer, 25(2): 33–42, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wallace:1992:DCP

[223] A. M. Wallace, G. J. Michaelson, P. McAndrew, K. G. Waugh, and W. J. Austin. Dynamic control and prototyping of parallel algorithms for intermediate- and high-level vision. Computer, 25(2):43–53, February 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siegel:1992:MCT

[224] Howard Jay Siegel, James B. Armstrong, and Daniel W. Watson. Mapping computer-vision-related tasks onto reconfigurable parallel-processing systems. Computer, 25(2):54–63, February 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siegel:1992:MCV

[225] Howard Jay Siegel, James B. Armstrong, and Daniel W. Watson. Mapping computer-vision-related tasks onto reconfigurable parallel-processing systems. Computer, 25(2):54–63, February 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weems:1992:IUA

[226] Charles C. Weems, Edward M. Riseman, Allen R. Hanson, P. J. Narayanan, Ling Tony Chen, Larry S. Davis, Leah H. Jamieson, Edward J. Delp, Wang Chao-Chun, Juan Li, et al. Image understanding architecture: Exploiting potential parallelism in machine vision. Computer, 25(2): 65–68, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Narayanan:1992:EUS

[227] P. J. Narayanan, Ling Tony Chen, and Larry S. Davis. Effective use of SIMD parallelism in low- and intermediatelevel vision. Computer, 25(2):68– 73, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jamieson:1992:SEP

[228] Leah H. Jamieson, Edward J. Delp, Chao-Chun Wang, Juan Li, and Frank J. Weil. A software environment for parallel computer vision. Computer, 25(2):73–77, February 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Begg:1992:RTC

[229] Vivienne S. Begg. Review of A Theory of Computer Semiotics (Anderson, P. B; 1990). Computer, 25(2):118– 119, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RTL

[230] Janusz Zalewski. Review of Temporal Logic for Real-Time Systems (Ostroff, I J; 1990). Computer, 25(2): 119, February 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1992:ARC

[231] T. Y. C. Woo and S. S. Lam. ‘authentication’ revisited (correction and addendum to ‘Authentication’ for distributed systems, Jan. 92, 39–52). Computer, 25(3):10, March 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1992:CAD

[232] Thomas Y. C. Woo and Simon S. Lam. Correction: Authentication for distributed systems. Computer, 25 (3):10, March 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [210].

Nielsen:1992:UEL

[233] Jakob Nielsen. The usability engineering life cycle. Computer, 25(3): 12–22, March 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reichbach:1992:SOD

[234] Jonathan D. Reichbach and Richard A. Kemmerer. SoundWorks: An objectoriented distributed system for digital sound. Computer, 25(3):25–37, March 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wiederhold:1992:MAF

[235] Gio Wiederhold. Mediators in the architecture of future information systems. Computer, 25(3):38–49, March 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bright:1992:TCI

[236] M. W. Bright, A. R. Hurson, and Simin H. Pakzad. A taxonomy and current issues in multidatabase systems. Computer, 25(3):50–60, March 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lenoski:1992:SDM

[237] Daniel Lenoski, James Laudon, Kourosh Gharachorloo, Wolf-Dietrich Weber, Anoop Gupta, John Hennessy, Mark Horowitz, and Monica S. Lam. The Stanford Dash multiprocessor. Computer, 25(3):63–79, March 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Teruel:1992:ROO

[238] Alejandro Teruel. Review of Object-Oriented Design (Coed, P. and Yourdon, E.; 1991). Computer, 25(3): 125, March 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RSE

[239] Janusz Zalewski. Review of Software Engineering with Abstractions (Berzins, V, and Luqi). Computer, 25 (3):125–126, March 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stein:1992:OCH

[240] Richard Marion Stein. The Open Channel: How to forestall electronic mutually assured destruction. Computer, 25(3):128, March 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:JBR

[241] Anonymous. January book review: A plea for McMathematics for engineers. Computer, 25(4):4, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuchs:1992:WIA

[242] W. Kent Fuchs and Earl E. Swartzlander. Wafer-scale integration: Architectures and algorithms. Computer, 25(4):6–8, April 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuchs:1992:WSI

[243] W. Kent Fuchs and Earl E. Swartzlander, Jr. Wafer-scale integration: Architectures and algorithms (special issue introduction). Computer, 25 (4):6–8, April 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horst:1992:TAW

[244] Robert W. Horst. Task-flow architecture for WSI parallel processing. Computer, 25(4):10–18, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horst:1992:TFA

[245] Robert W. Horst. Task-flow architecture for WST parallel processing. Computer, 25(4):10–18, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yamashita:1992:DYE

[246] Koichi Yamashita and Shohei Ikehara. A design and yield evaluation technique for wafer-scale memory. Computer, 25 (4):19–27, April 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boubekeur:1992:CWS

[247] Ahmed Boubekeur, Jean-Luc Patry, Gabriele Saucier, and Jacques Trilhe. Configuring a wafer-scale twodimensional array of single-bit processors. Computer, 25(4):29–39, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boubekeur:1992:CWT

[248] Ahmed Boubekeur, Jean-Luc Patry, Gabriele Saucier, and Jacques Trilhe. Configuring a wafer-scale twodimensional array of single-bit processors. Computer, 25(4):29–39, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:1992:WIU

[249] Allan H. Anderson, Jack I. Raffel, and Peter W. Wyatt. Wafer-scale integration using restructurable VLSI. Computer, 25(4):41–47, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:1992:WSI

[250] Allan H. Anderson, Jack I. Raffel, and Peter W. Wyatt. Wafer-scale integration using restructurable VLSI. Computer, 25(4):41–47, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chapman:1992:WST

[251] Glenn H. Chapman, M. Parameswaran, and Marek Syrzycki. Wafer-scale transducer arrays. Computer, 25(4):50–56, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chapman:1992:WTA

[252] Glenn H. Chapman, M. Parameswaran, and Marek Syrzycki. Wafer-scale transducer arrays. Computer, 25(4):50–56, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maly:1992:PWM

[253] Wojciech Maly. Prospects for WSI: A manufacturing perspective. Computer, 25(4):58–65, April 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Landis:1992:POW

[254] David L. Landis, Nitin Nigam, Joseph W. Yoder, Vijay K. Jain, Hiroomi Hikawa, and D. C. Keezer. Project overviews: Wafer-scale optimization using computational availability. Computer, 25(4):66–71, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:1992:AWR

[255] Vijay K. Jain, Hiroomi Hikawa, and D. C. Keezer. An architecture for WSI rapid prototyping. Computer, 25 (4):71–75, April 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:CSNa

[256] Anonymous. Computer Society news. Computer, 25(4):76–78, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chou:1992:RNN

[257] Tsai-Chia Peter Chou. Review of Neural Networks and Fuzzy Systems: A Dynamical Systems Approach to Machine Intelligence (Kosko, B; 1992). Computer, 25(4):110, April 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RMB

[258] Janusz Zalewski. Review of Microcomputer Buses (Cram, R. M; 1991). Computer, 25(4):110–111, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:1992:RHP

[259] Rana Ejaz Ahmed. Review of High-Performance Computer Architecture Ahmed, 2nd edn. (Stone, H. S; 1990). Computer, 25(4):111, April 1992. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eggert:1992:OCN

[260] Paul Eggert. The Open Channel: New open channels, old-fashioned societies. Computer, 25(4):112, April 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Herath:1992:CAI

[261] Jayantha Herath. Computer architectures for intelligent systems (special issue introduction). Computer, 25(5):6–9, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rizzi:1992:DRC

[262] Alfred A. Rizzi, Louis L. Whitcomb, and Daniel E. Koditschek. Distributed real-time control of a spatial robot juggler. Computer, 25(5):12–24, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abramson:1992:VHS

[263] David Abramson. A very high speed architecture for simulated annealing. Computer, 25(5):27–36, May 1992.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moldovan:1992:SPP

[264] Dan Moldovan, Wing Lee, Changhwa Lin, and Minhwa Chung. Snap: Parallel processing applied to AI. Computer, 25(5):39–49, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferguson:1992:TAA

[265] Innes A. Ferguson. TouringMachines: Autonomous agents with attitudes. Computer, 25(5):51–55, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Albus:1992:RRM

[266] James S. Albus. RCS: A reference model architecture for intelligent control. Computer, 25(5):56–59, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

LePape:1992:UCP

[267] Claude Le Pape. Using constraint propagation in blackboard systems: A flexible software architecture for reactive and distributed systems. Computer, 25(5):60–62, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robinson:1992:HSR

[268] Ian N. Robinson. Hardware to support runtime intelligence. Computer, 25(5):63–66, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chineh:1992:OFL

[269] Tzi cker Chineh. Optimization of fuzzy logic inference architecture. Computer, 25(5):67–71, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chiueh:1992:OFL

[270] Tzi cker Chiueh. Optimization of fuzzy logic inference architecture. Computer, 25(5):67–71, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McAulay:1992:OCA

[271] Alastair D. McAulay. Optical computer architectures for supervised learning systems. Computer, 25(5):72–75, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pao:1992:FNC

[272] Yoh-Han Pao and Yoshiyasu Takefuji. Functional-link net computing: Theory, system architecture, and functionalities. Computer, 25(5):76–79, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Turano:1992:RHC

[273] August M. Turano. Review of Hypertext in Context (McKnight, C., et al; 1991). Computer, 25(5):110–111, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RCP

[274] Janusz Zalewski. Review of Concurrent Programming: Principles and Practice (Andrews, G. R; 1991). Computer, 25(5):111, May 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yan:1992:OCC

[275] Seung-Min Yan. The Open Channel: Conventional wisdom for today’s computer scientists and engineers. Computer, 25(5):112, May 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Turek:1992:MFC

[276] John Turek and Dennis Shasha. The many faces of consensus in distributed systems. Computer, 25(6):8–17, June 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pavlidis:1992:IET

[277] Theo Pavlidis, Jerome Swartz, and Ynjiun P. Wang. Information encoding with two-dimensional bar codes. Computer, 25(6):18–28, June 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garlan:1992:UTA

[278] David G. Garlan, Gail E. Kaiser, and David Notkin. Using tool abstraction to compose systems. Computer, 25 (6):30–38, June 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Azaria:1992:OCS

[279] Helnye Azaria and Adrian Dvir. An optimizing compiler for an SPAP architecture using AI tools. Computer, 25(6):39–48, June 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hariri:1992:ASD

[280] Salim Hariri, Alok Choudhary, and Behcet Sarikaya. Architectural support for designing fault-tolerant open distributed systems. Computer, 25 (6):50–62, June 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hariri:1992:AST

[281] Salim Hariri, Alok Choudhary, and Behcet Sarikaya. Architectural support tor designing fault-tolerant open distributed systems. Computer, 25 (6):50–62, June 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Edelstein:1992:SSS

[282] D. Vera Edelstein and Salvatore Mamone. Standards: A standard for software maintenance. Computer, 25 (6):82–83, June 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RPD

[283] Janusz Zalewski. Review of Programming Distributed Systems (Bal, H. E; 1990). Computer, 25(6):101–102, June 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Benedek:1992:ROO

[284] Z. Michael Benedek. Review of Object-Oriented Databases with Applications to CASE, Networks, and VLSI CAD (Gupta, R. and Horowitz, E., Eds; 1991). Computer, 25(6):102, June 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brakefield:1992:OCD

[285] James Brakefield. The open channel: Distributed virtual society and the economics of sharing. Computer, 25(6):104, June 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OGorman:1992:GEI

[286] Lawrence O’Gorman and Rangachar Kasturi. Guest editor’s introduction: Document image analysis systems. Computer, 25(7):5–8, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nagy:1992:PDI

[287] George Nagy, Sharad Seth, and Mahesh Viswanathan. A prototype document image analysis system for technical journals. Computer, 25(7):10–22, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boatto:1992:ISL

[288] Luca Boatto, Vincenzo Consorti, Monica Del Buono, Silvano Di Zenzo, Vincenzo Eramo, Alessandra Esposito, Francesco Melcarne, Marco Meucci, Andrea Morelli, Marco Mosciatti, Stefano Scarci, and Marco Tucci. An interpretation system for land register maps. Computer, 25(7):25–33, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Palumbo:1992:PAB

[289] Paul W. Palumbo, Sargur N. Srihari, Soh Jung, Ramalingam Sridhar, and Victor Demjanenko. Postal address block location in real time. Computer, 25(7):34–42, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaxiviere:1992:CCC

[290] Pascal Vaxivière and Karl Tombre. Celesstin: CAD conversion of mechanical drawings. Computer, 25(7):46–54, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shimotsuji:1992:RRS

[291] Shigeyoshi Shimotsuji, Osamu Hori, Mieko Asano, Kaoru Suzuki, Fumihiko Hoshino, and Toshiaki Ishii. A robust recognition system for a drawing superimposed on a map. Computer, 25 (7):56–59, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matan:1992:RHD

[292] Ofer Matan, Henry S. Baird, Jane Bromley, Christopher J. C. Burges, John S. Denker, Lawrence D. Jackel, Le Cun Yann, Edwin P. D. Pednault, William D. Satterfield, Charles E. Stenard, et al. Reading handwritten digits: A ZIP code recognition system. Computer, 25(7):59–63, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dengel:1992:POD

[293] Andreas Dengel, Rainer Bleisinger, Rainer Hoch, Frank Fein, and Frank Hones. From paper to office document standard representation. Computer, 25 (7):63–67, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amano:1992:DWB

[294] Tomio Amano, Akio Yamashita, Nobuyasu Itoh, Yoshinao Kobayashi, Shin Katoh, Kazuharu Toyokawa, and Hiroyasu Takahashi. DRS: A workstation-based document recognition system for text entry. Computer, 25(7):67–71, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amano:1992:DWD

[295] Tomio Amano, Akio Yamashita, Nobuyasu Itoh, Yoshinao Kobayashi, Shin Katoh, Kazuharu Toyokawa, and Hiroyasu Takahashi. DRS: A workstation-based document recognition system for text entry. Computer, 25(7):67–71, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goraine:1992:LAC

[296] Habib Goraine, Mike Usher, and Samir Al-Emami. Off-line Arabic character recognition. Computer, 25(7):71–74, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goraine:1992:OAC

[297] Habib Goraine, Mike Usher, and Samir Al-Emami. Off-line Arabic character recognition. Computer, 25(7):71–74, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Futrelle:1992:UDT

[298] Robert P. Futrelle, Ioannis A. Kakadiaris, Jeff Alexander, Catherine M. Carriero, Nikos Nikolakis, and Joseph M.

Futrelle. Understanding diagrams in technical documents. Computer, 25 (7):75–78, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:PCM

[299] Anonymous. Powerful CASE on the Mac. Computer, 25(7):84–87, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stapleton:1992:NCP

[300] Ross Alan Stapleton. News from the Committee on Public Policy — opening doors in the global village. Computer, 25(7):94–96, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chlamtac:1992:RDN

[301] Imrich Chlamtac. Review of Data Networks, 2nd edn. (Bertsekas, D., and Gallager, R; 1992). Computer, 25(7):110, July 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berkovec:1992:RMW

[302] James W. Berkovec. Review of Mirror Worlds (Gelernter, D; 1991). Computer, 25(7):111, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Estell:1992:OCT

[303] Robert G. Estell. The Open Channel: Two heads are better than one. Computer, 25(7):112, July 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanenbaum:1992:PPU

[304] Andrew S. Tanenbaum, M. Frans Kaashoek, and Henri E. Bal. Parallel programming using shared objects and broadcasting. Computer, 25(8): 10–19, August 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tannenbaum:1992:PPU

[305] Andrew S. Tannenbaum, M. Frans Kaashoek, and Henri E. Bal. Parallel programming using shared objects and broadcasting. Computer, 25(8): 10–19, August 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feddema:1992:MBV

[306] John T. F. Feddema, C. S. George Lee, and O. Robert Mitchell. Model-based visual feedback control for a hand-eye coordinated robotic system. Computer, 25(8):21–31, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feddema:1992:MVF

[307] John T. Feddema, C. S. George Lee, and O. Robert Mitchell. Model-based visual feedback control for a hand-eye coordinated robotic system. Computer, 25(8):21–31, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alkalaj:1992:ASG

[308] Leon Alkalaj, Tomas Lang, and Milos D. Ercegovac. Architectural support for goal management in Flat Concurrent Prolog. Computer, 25(8):

34–47, August 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raymond:1992:FTD

[309] Darrell R. Raymond. Flexible text display with Lector. Computer, 25(8): 49–60, August 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1992:DIS

[310] Brad A. Myers. Demonstrational interfaces: A step beyond direct manipulation. Computer, 25(8):61–73, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:CSNb

[311] Anonymous. Computer Society news. Computer, 25(8):74–76, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:U

[312] Anonymous. Update. Computer, 25 (8):77–79, August 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:PRW

[313] Anonymous. Product reviews: What light through yonder Windows 3.1? Computer, 25(8):81–86, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wegener:1992:RDC

[314] Al Wegener. Review of Data Compression, 3rd edn. (Held, G; 1991). Computer, 25(8):110–111, August 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Collura:1992:ESA

[315] Thomas F. Collura, Ernest C. Jacobs, Richard C. Burgess, and John P. Turnbull. The Epilog system: Automated long-term EEG monitoring for epilepsy. Computer, 25(9):5– 14, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Collura:1992:ESL

[316] Thomas F. Collura, Ernest C. Jacobs, Richard C. Burgess, and John P. Turnbull. The Epilog system — automated long-term EEG monitoring for epilepsy. Computer, 25(9):5– 14, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Story:1992:RIB

[317] Guy A. Story, Lawrence O’Gorman, David Fox, Louise Levy Schaper, and H. V. Jagadish. The RightPages imagebased electronic library for alerting and browsing. Computer, 25(9):17– 26, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Story:1992:RIE

[318] Guy A. Story, Lawrence O’Gorman, David Fox, Louise Levy Schaper, and H. V. Jagadish. The RightPages imagebased electronic library for alerting and browsing. Computer, 25(9):17– 26, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ambler:1992:OVD

[319] Allen L. Ambler, Margaret M. Burnett, and Betsy A. Zimmerman. Operational versus definitional: A perspective on programming paradigms. Computer, 25 (9):28–43, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hagersten:1992:DCM

[320] Erik Hagersten, Anders Landin, and Seif Haridi. DDM — A cache-only memory architecture. Computer, 25(9): 44–54, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hagersten:1992:DCO

[321] Erik H. Hagersten, Anders Landin, and Seif Haridi. DDM — A cache-only memory architecture. Computer, 25(9): 44–54, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sajeev:1992:PP

[322] A. S. M. Sajeev and A. J. Hurst. Programming persistence in χ. Computer, 25(9):57–66, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ashley:1992:SPC

[323] J. Robert Ashley and Fletcher Buckley. Standards: Point/counterpoint: Safety standards for CRT visual display units. Computer, 25(9):68–70, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ashley:1992:SSS

[324] J. R. Ashley and F. Buckley. Standards: Point/counterpoint: safety standards for CRT visual display units. Computer, 25(9):68–70, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RRT

[325] Janusz Zalewski. Review of Real-Time Expert Systems Computer Architecture (Hodson, R. F. and Kandel, A; 1991). Computer, 25(9):126, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birss:1992:RLP

[326] Robert C. Birss. Review of Literate Programming (Knuth, D. E; 1992). Computer, 25(9):126–127, September 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rine:1992:GEI

[327] David C. Rine and Bharat Bhargava. Guest Editors’ introduction: Objectoriented computing. Computer, 25(10): 6–10, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wegner:1992:DOM

[328] Peter Wegner. Dimensions of objectoriented modeling. Computer, 25(10): 12–20, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wegner:1992:DOO

[329] Peter Wegner. Dimensions of objectoriented modeling. Computer, 25(10): 12–20, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fichman:1992:OCA

[330] Robert G. Fichman and Chris F. Kemerer. Object-oriented and conventional analysis and design methodologies. Computer, 25(10):22–39, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fichman:1992:OOC

[331] Robert G. Fichman and Chris F. Kemerer. Object-oriented and conventional analysis and design methodologies. Computer, 25(10):22–39, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:1992:ADC

[332] Bertrand Meyer. Applying “design by contract”. Computer, 25(10):40– 51, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ungar:1992:OMP

[333] David Ungar, Randall B. Smith, Craig Chambers, and Urs Holzle. Object, message, and performance: how they coexist in Self. Computer, 25(10): 53–64, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ishikawa:1992:OOR

[334] Yutaka Ishikawa, Hideyuki Tokuda, and Clifford W. Mercer. An objectoriented real-time programming language. Computer, 25(10):66–73, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ishikawa:1992:ORP

[335] Yutaka Ishikawa, Hideyuki Tokuda, and Clifford W. Mercer. An objectoriented real-time programming language. Computer, 25(10):66–73, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wells:1992:AOO

[336] David L. Wells, José A. Blakeley, and Craig W. Thompson. Architecture of an open object-oriented database management system. Computer, 25(10): 74–82, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Egbert:1992:AGM

[337] Parris K. Egbert and William J. Kubitz. Application graphics modeling support through object orientation. Computer, 25(10):84–91, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:CSNc

[338] Anonymous. Computer Society news. Computer, 25(10):94–95, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:PRV

[339] Anonymous. Product reviews: OS/2 version 2.0. Computer, 25(10):97– 101, October 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamming:1992:RJN

[340] R. W. Hamming. Review of John von Neumann and the Origins of Modern Computing (Aspray, W; 1990). Computer, 25(10):126–127, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conrad:1992:GEI

[341] Michael Conrad. Guest editor’s introduction: Molecular computing paradigms. Computer, 25(11):6–9, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conrad:1992:MCP

[342] Michael Conrad. Molecular computing paradigms (special issue introduction). Computer, 25(11):6–9, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conrad:1992:MCL

[343] Michael Conrad. Molecular computing: The lock-key paradigm. Computer, 25(11):11–20, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Capstick:1992:BCB

[344] Myles H. C. Capstick, W. P. Liam Marnane, and Ronald Pethig. Biologic computational building blocks. Computer, 25(11):22–29, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hameroff:1992:MMC

[345] Stuart R. Hameroff, Judith E. Dayhoff, Rafael Lahoz-Beltra, Alexei V. Samsonovich, and Steen Rasmussen. Models for molecular computation: Conformational automata in the cytoskeleton. Computer, 25(11):30–39, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aoki:1992:IBC

[346] Takafumi Aoki, Michitaka Kameyama, and Tatsuo Higuchi. Interconnectionfree biomolecular computing. Computer, 25(11):41–51, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aoki:1992:IFB

[347] Takafumi Aoki, Michitaka Kameyama, and Tatsuo Higuchi. Interconnectionfree biomolecular computing. Computer, 25(11):41–51, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rambidi:1992:ACC

[348] Nikolay G. Rambidi. An approach to computational complexity: Nondiscrete biomolecular computing. Computer, 25(11):51–54, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birge:1992:PBO

[349] Robert R. Birge. Protein-based optical computing and memories. Computer, 25(11):56–67, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Birge:1992:POC

[350] Robert R. Birge. Protein-based optical computing and memories. Computer, 25(11):56–67, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1992:TSR

[351] David Gries and Dorothy Marsh. The 1990–91 Taulbee survey report. Computer, 25(11):69–75, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Todd:1992:HTT

[352] Peter M. Todd. Hot topics: Two approaches to machine intelligence — the animat path to intelligent adaptive behavior. Computer, 25(11):78, 81–??, November 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conrad:1992:HTT

[353] Michael Conrad. Hot topics: Two approaches to machine intelligence: Toward an artificial brain. Computer, 25 (11):79–80, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zelkowitz:1992:SSE

[354] Marvin V. Zelkowitz and Fletcher J. Buckley. Standards: Are software engineering process standards really necessary? Computer, 25(11):82– 84, November 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:CSNd

[355] Anonymous. Computer Society news. Computer, 25(11):86–90, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weisband:1992:NCP

[356] Suzanne P. Weisband and Seymour E. Goodman. News from the Committee on Public Policy — international software piracy. Computer, 25(11): 87–90, November 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Burleson:1992:RDS

[357] Wayne. Burleson. Review of Digital Signal Processing: Theory and Applications and Hardware (Haddad, R. A., and Parsons T. W; 1991). Computer, 25(11):126, November 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RMA

[358] Janusz Zalewski. Review of Microcontrollers: Architecture and Implementation and Programming (Hintz, K. and Tabak, D; 1992). Computer, 25(11): 126–127, November 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zalewski:1992:RVS

[359] Janusz Zalewski. Review of Verification of Sequential and Concurrent Programs (Apt, K. R. and Olderog, E.-R.; 1991). Computer, 25(11):127, October 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simmons:1992:OCD

[360] Chuck Simmons. The Open Channel: The dining philosophers at Thanksgiving dinner. Computer, 25(11): 128, November 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bliss:1992:OCL

[361] Robert D. Bliss. Open channel: It’s the law. Computer, 25(12):12, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adrion:1992:DCM

[362] W. Richards Adrion, Edward D. Lazowska, and Andries Van Dam. From discipline in crisis to mature science: Evolving needs for computing research infrastructure. Computer, 25(12):18– 24, December 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bihari:1992:OOR

[363] Thomas E. Bihari and Prabha Gopinath. Object-oriented real-time systems: Concepts and examples. Computer, 25(12):25–32, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bihari:1992:ORS

[364] Thomas E. Bihari and Prabha Gopinath. Object-oriented real-time systems: Concepts and examples. Computer, 25(12):25–32, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shivaratri:1992:LDL

[365] Niranjan G. Shivaratri, Phillip Krueger, and Mukesh Singhal. Load distributing for locally distributed systems. Computer, 25(12):33–44, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1992:CAR

[366] Harold S. Stone. Copyrights and author responsibilities. Computer, 25(12): 46–51, December 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). This paper discusses U.S. copyright law and its implications for the reuse of published text. Guidelines are given for the citation and quoting of such material.

Brown:1992:CSA

[367] Marc H. Brown and John Hershberger. Color and sound in algorithm animation. Computer, 25(12): 52–63, December 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siefert:1992:SRH

[368] D. M. Siefert and G. E. Stark. Standards: Software reliability handbook: Achieving reliable software. Computer, 25(12):64–66, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siefert:1992:SSR

[369] David M. Siefert and George E. Stark. Standards: Software reliability handbook: Achieving reliable software. Computer, 25(12):64–66, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balde:1992:HTT

[370] John W. Balde. Hot topics: Trends in advanced packaging technology. Computer, 25(12):67–68, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:CSNe

[371] Anonymous. Computer Society news. Computer, 25(12):71–72, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1992:AI

[372] Anonymous. 1992 annual index. Computer, 25(12):107–117, December 1992. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anneberg:1992:RIA

[373] Lisa Anneberg. Review of An Integrated Approach to Software Engineering (Jalote, P; 1991). Computer, 25 (12):118, December 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reddy:1993:GEI

[374] Y. V. Ramana Reddy, Kanakanahalli Srinivas, V. Jagannathan, and Raghu Karinthi. Guest editors’ introduction: Computer support for concurrent engineering. Computer, 26(1): 12–??, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dewan:1993:TCC

[375] Prasun Dewan and John Riedl. Toward computer-supported concurrent software engineering. Computer, 26(1): 17–27, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cutkosky:1993:PEI

[376] Mark R. Cutkosky, Robert S. Engelmore, Richard E. Fikes, Michael R. Genesereth, Thomas R. Gruber, W. S. Mark, J. M. Tenenbaum, and J. C. Weber. PACT: An experiment in integrating concurrent engineering systems. Computer, 26(1):28–37, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klein:1993:CDR

[377] Mark Klein. Capturing design rationale in concurrent engineering teams. Computer, 26(1):39–47, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chandrasekaran:1993:FRD

[378] B. Chandrasekaran, Ashok K. Goel, and Yumi Iwasaki. Functional representation as design rationale. Computer, 26(1):48–56, January 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:1993:POC

[379] Satyandra K. Gupta, Dana S. Nau, and Guangming Zhang. Project overviews: Concurrent evaluation of machinability during product design. Computer, 26 (1):61–63, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sriram:1993:DP

[380] Duvvuru Sriram and Robert Logcher. The MIT Dice Project. Computer, 26 (1):64–65, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowen:1993:CSC

[381] James Bowen and Dennis Bahler. Constraint-based software for concurrent engineering. Computer, 26(1): 66–68, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tong:1993:KCE

[382] Chris Tong and Andrew Gomory. A knowledge-based computer environment for the conceptual design of small electromechanical appliances. Computer, 26(1):69–71, January 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chung:1993:ALS

[383] Goopeel Chung, Kevin Jeffay, and Hussein Abdel-Wahab. Accommodating latecomers in shared window systems. Computer, 26(1):72–74, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huyn:1993:ACE

[384] Pierre Nam Huyn, Michael R. Genesereth, and Reed Letsinger. Automated concurrent engineering in Designworld. Computer, 26(1):74–76, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1993:GBP

[385] Alan H. Karp, Ken Miura, and Horst Simon. 1992 Gordon Bell Prize winners. Computer, 26(1):77–??, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Subrahmanyam:1993:HTH

[386] P. A. Subrahmanyam. Hot topics: Hardware-software codesign: Cautious optimism for the future. Computer, 26 (1):84–??, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buchenrieder:1993:HTH

[387] K. Buchenrieder. Hot topics: Hardware-software codesign: Codesign and concurrent engineering. Computer, 26(1):85–86, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Micheli:1993:HTH

[388] G. De Micheli. Hot topics: Hardwaresoftware codesign: Extending CAD tools and techniques. Computer, 26 (1):85–87, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhansali:1993:SSS

[389] P. V. Bhansali. Survey of software safety standards shows diversity. Computer, 26(1):88–89, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNa

[390] Anonymous. CS news. Computer, 26 (1):90–??, January 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ca

[391] Anonymous. Conferences. Computer, 26(1):112–??, January 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:ECM

[392] Anonymous. Editor-in chief’s message. Computer, 26(2):6–??, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:1993:DCH

[393] Ronald J. Vetter and David H. C. Du. Distributed computing with high-speed optical networks. Computer, 26(2): 8–18, February 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowen:1993:PMC

[394] Nicholas S. Bowen and Dhiraj K. Pradhan. Processor- and memorybased checkpoint and rollback recovery. Computer, 26(2):22–31, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karam:1993:CFS

[395] Gerald M. Karam and Ronald S. Casselman. A cataloging framework for software development methods. Computer, 26(2):34–44, 46, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurson:1993:ODM

[396] A. R. Hurson, Simin H. Pakzad, and Jia bing Cheng. Object-oriented database management systems: Evolution and performance issues. Computer, 26(2):48–58, 60, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ni:1993:SWR

[397] Lionel M. Ni and Philip K. McKinley. A survey of wormhole routing techniques in direct networks. Computer, 26(2):62–76, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pausch:1993:HTT

[398] R. Pausch. Hot topics: Three views of virtual reality: an overview. Computer, 26(2):79–80, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robertson:1993:HTT

[399] G. G. Robertson, S. K. Card, and J. D. Mackinlay. Hot topics: Three views of virtual reality: nonimmersive virtual reality. Computer, 26(2): 81, 83, February 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moshell:1993:HTT

[400] M. Moshell. Hot topics: Three views of virtual reality: virtual environments in the US military. Computer, 26(2): 81–82, February 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trapp:1993:SES

[401] G. Trapp. Standards: The emerging STEP standard for product-model data exchange. Computer, 26(2):85– 87, February 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNb

[402] Anonymous. CS news. Computer, 26 (2):88–??, February 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ua

[403] Anonymous. Update. Computer, 26 (2):90–??, February 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRa

[404] Anonymous. Product reviews. Computer, 26(2):100–??, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Cb

[405] Anonymous. Conferences. Computer, 26(2):117–??, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRa

[406] Anonymous. Media reviews. Computer, 26(2):126–??, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCa

[407] Anonymous. Open channel. Computer, 26(2):128–??, February 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PMa

[408] Anonymous. President’s message. Computer, 26(3):6–??, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:EMa

[409] Anonymous. Editor-in-chief’s message. Computer, 26(3):9–??, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athanas:1993:PRT

[410] Peter M. Athanas and Harvey F. Silverman. Processor reconfiguration through instruction-set metamorphosis. Computer, 26(3):11–18, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abdel-Hamid:1993:ACP

[411] Tarek K. Abdel-Hamid. Adapting, correcting, and perfecting software estimates: A maintenance metaphor.

Computer, 26(3):20–29, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sankar:1993:CRM

[412] Sriram Sankar and Manas Mandal. Concurrent runtime monitoring of formally specified programs. Computer, 26(3):32–41, March 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Staehli:1993:CSA

[413] Richard Staehli and Jonathan Walpole. Constrained-latency storage access. Computer, 26(3):44–53, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marefat:1993:OIC

[414] M. Marefat, Sandeep Malhotra, and R. L. Kashyap. Object-oriented intelligent computer-integrated design, process planning, and inspection. Computer, 26(3):54–65, March 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keaton:1993:LCI

[415] John Keaton. Layoffs in the computer industry. Computer, 26(3): 66–71, March 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Obenza:1993:HTR

[416] R. Obenza. Hot topics: Rate monotonic analysis for real-time systems. Computer, 26(3):73–74, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Everett:1993:SSR

[417] W. W. Everett and J. A. Musa. Standards: A software reliability engineering practice. Computer, 26(3): 77–79, March 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRb

[418] Anonymous. Product reviews. Computer, 26(3):81–??, March 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNc

[419] Anonymous. CS news. Computer, 26 (3):92–??, March 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ub

[420] Anonymous. Update. Computer, 26 (3):93–??, March 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRb

[421] Anonymous. Media reviews. Computer, 26(3):110–??, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCb

[422] Anonymous. Open channel. Computer, 26(3):112–??, March 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:EMb

[423] Anonymous. Editor-in-chief’s message. Computer, 26(4):6–??, April 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mukund:1993:GEI

[424] P. R. Mukund and J. F. McDonald. Guest Editor’s introduction: MCM — the high-performance electronic packaging technology. Computer, 26(4): 10–??, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Scannell:1993:DMM

[425] Robert K. Scannell and John K. Hagge. Development of a multichip module DSP. Computer, 26(4):13–21, April 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Daum:1993:OHI

[426] Wolfgang Daum, William E. Burdick, Jr., and Raymond A. Fillion. Overlay high-density interconnect: A chips-first multichip module technology. Computer, 26(4):23–29, April 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

LaPotin:1993:EPA

[427] David P. LaPotin, Toufie R. Mazzawy, and Marlin L. White. Early package analysis: Considerations and case study. Computer, 26(4):30–39, April 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Franzon:1993:MMD

[428] Paul D. Franzon and Robert J. Evans. A multichip module design process for notebook computers. Computer, 26 (4):41–49, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Habiger:1993:HMP

[429] Claus M. Habiger and R. Mike Lea. Hybrid-WSI: A massively parallel computing technology? Computer, 26 (4):50–61, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vemuri:1993:IMS

[430] Ranga Vemuri, Nand Kumar, Raghu Vutukuru, Prasad Subba Rao, Praveen Sinha, N. Ren, P. Mamtora, R. Mandayam, R. Vemuri, and J. Roy. An integrated multicomponent synthesis environment for MCMs. Computer, 26 (4):62–74, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dehkordi:1993:DPE

[431] Peyman H. Dehkordi and Donald W. Bouldin. Design for packageability: Early consideration of packaging from a VLSI designer’s viewpoint. Computer, 26(4):76–81, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Glasser:1993:RMA

[432] L. A. Glasser. A road map to ARPA involvement in electronic packaging. Computer, 26(4):82–86, April 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNd

[433] Anonymous. CS news. Computer, 26 (4):88–??, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Uc

[434] Anonymous. Update. Computer, 26 (4):88–??, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRc

[435] Anonymous. Product reviews. Computer, 26(4):97–??, April 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shur:1993:HTF

[436] M. Shur. Hot topics: Future impact of solid-state technology on computers. Computer, 26(4):103–104, April 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneidewind:1993:SNS

[437] N. F. Schneidewind. Standards: New software-quality metrics methodology standard fills measurement need. Computer, 26(4):105–106, April 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Cc

[438] Anonymous. Conferences. Computer, 26(4):107–??, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRc

[439] Anonymous. Media reviews. Computer, 26(4):118–??, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCc

[440] Anonymous. Open channel. Computer, 26(4):120–??, April 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:LEa

[441] Anonymous. Letters to the editor. Computer, 26(5):6–??, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:AS

[442] Anonymous. Article summaries. Computer, 26(5):9–??, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaughan:1993:ARP

[443] Patrick T. Gaughan and Sudhakar Yalamanchili. Adaptive routing protocols for hypercube interconnection networks. Computer, 26(5):12–23, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ozsoyoglu:1993:EGD

[444] Gultekin Ozsoyoglu and Huaqing Wang. Example-based graphical database query languages. Computer, 26(5):25–38, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grimshaw:1993:EOP

[445] Andrew S. Grimshaw. Easy-to-use object-oriented parallel processing with Mentat. Computer, 26(5):39–51, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larus:1993:EPT

[446] James R. Larus. Efficient program tracing. Computer, 26(5):52–61, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Notkin:1993:CSJ

[447] David Notkin and Richard D. Schlichting. Computer Science in Japanese universities. Computer, 26(5):62–70, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNe

[448] Anonymous. CS news. Computer, 26 (5):72–??, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:HTI

[449] Anonymous. Hot topics: Information infrastructure. Computer, 26 (5):75–??, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matras:1993:SRS

[450] J. R. Matras. Standards: Revised standard addresses design requirements for computers used in nuclear powerplant safety systems. Computer, 26 (5):76–79, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRd

[451] Anonymous. Product reviews. Computer, 26(5):81–??, May 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Cd

[452] Anonymous. Conferences. Computer, 26(5):95–??, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRd

[453] Anonymous. Media reviews. Computer, 26(5):110–??, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCd

[454] Anonymous. Open channel. Computer, 26(5):112–??, May 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:AEMa

[455] Anonymous. Area editor’s message. Computer, 26(6):6–??, June 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Freund:1993:GEI

[456] Richard F. Freund and Howard Jay Siegel. Guest editors’ introduction: Heterogeneous processing. Computer, 26(6):13–??, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khokhar:1993:HCC

[457] Ashfaq A. Khokhar, Viktor K. Prasanna, Muhammad E. Shaaban, and Cho-Li Wang. Heterogeneous computing: Challenges and opportunities. Computer, 26(6):18–27, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rinard:1993:JAH

[458] Martin C. Rinard, Daniel J. Scales, and Monica S. Lam. Jade: A highlevel machine-independent language for parallel programming. Computer, 26 (6):28–38, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Silberman:1993:AFS

[459] Gabriel M. Silberman and Kemal Ebcioglu. An architectural framework for supporting heterogeneous instruction-set architectures. Computer, 26(6):39–56, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nicol:1993:OOH

[460] John R. Nicol, C. Thomas Wilkes, and Frank A. Manola. Object orientation in heterogeneous distributed computing systems. Computer, 26(6):57–67, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rinaldo:1993:ERH

[461] Frank J. Rinaldo and Matthew R. Fausey. Event reconstruction in highenergy physics. Computer, 26(6): 68–77, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ghafoor:1993:DHS

[462] Arif Ghafoor and Jaehyung Yang. A distributed heterogeneous supercomputing management system. Computer, 26(6):78–86, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beguelin:1993:VDH

[463] Adam Beguelin, Jack Dongarra, Al Geist, and Vaidy Sunderam. Visualization and debugging in a heterogeneous environment. Computer, 26 (6):88–95, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosenblum:1993:HTV

[464] L. J. Rosenblum and F. H. Post. Hot topics: Visualizing fluid flow. Computer, 26(6):98–100, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRe

[465] Anonymous. Product reviews. Computer, 26(6):107–??, June 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNf

[466] Anonymous. CS news. Computer, 26 (6):115–??, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ud

[467] Anonymous. Update. Computer, 26 (6):116–??, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vollman:1993:SSQ

[468] T. E. Vollman. Standards: Software quality assessment and standards. Computer, 26(6):118–120, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ce

[469] Anonymous. Conferences. Computer, 26(6):120–??, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRe

[470] Anonymous. Media reviews. Computer, 26(6):134–??, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCe

[471] Anonymous. Open channel. Computer, 26(6):136–??, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:AEMb

[472] Anonymous. Area editor’s message. Computer, 26(7):6–??, July 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1993:FCC

[473] Kuo-Chu Lee, William H. Mansfield, Jr., and Amit P. Sheth. A framework for controlling cooperative agents. Computer, 26(7):8–16, July 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leveson:1993:ITA

[474] Nancy G. Leveson and Clark S. Turner. An investigation of the Therac-25 accidents. Computer, 26(7):18–41, July 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:1993:SPP

[475] Jaswinder Pal Singh, John L. Hennessy, and Anoop Gupta. Scaling parallel programs for multiprocessors: Methodology and examples. Computer, 26 (7):42–50, July 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaufman:1993:VG

[476] Arie Kaufman, Daniel Cohen, and Roni Yagel. Volume graphics. Computer, 26(7):51–64, July 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hutchinson:1993:HTG

[477] T. E. Hutchinson. Hot topics; eye gaze computer interfaces: computers that sense eye position on the display. Computer, 26(7):65, 67, July 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jacob:1993:HTC

[478] R. J. K. Jacob. Hot topics: eye-gaze computer interfaces: what you look at is what you get. Computer, 26 (7):65–66, July 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

CSNews:1993:CSB

[479] CS News. Candidates slate; Board report; bylaws amendments. Computer, 26(7):68–??, July 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRf

[480] Anonymous. Product reviews. Computer, 26(7):71–??, July 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRf

[481] Anonymous. Media reviews. Computer, 26(7):93–??, July 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCf

[482] Anonymous. Open channel. Computer, 26(7):96–??, July 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:TM

[483] Anonymous. Treasurer’s message. Computer, 26(8):8–??, August 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Griffeth:1993:GEI

[484] Nancy D. Griffeth and Yow-Jian Lin. Guest editors’ introduction: The feature-interaction problem. Computer, 26(8):14–18, August 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zave:1993:FIF

[485] Pamela Zave. Feature interactions and formal specifications in telecommunications. Computer, 26(8):20– 28, 30, August 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:1993:IPS

[486] Richard Kuhn, Patricia Edfors, Victoria Howard, Chuck Caputo, and Ted S. Phillips. Improving public switched network security in an open environment. Computer, 26(8):32–35, August 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fekete:1993:FMC

[487] Alan Fekete. Formal models of communication services: A case study. Computer, 26(8):37–47, August 1993.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Velthuijsen:1993:DAI

[488] Hugo Velthuijsen. Distributed artificial intelligence for runtime featureinteraction resolution. Computer, 26 (8):48–55, August 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bukhres:1993:IEE

[489] Omran A. Bukhres, Jiansan Chen, Weimin Du, Ahmed K. Elmagarmid, and Robert Pezzoli. InterBase: An execution environment for heterogeneous software systems. Computer, 26(8): 57–69, August 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilmarth:1993:HTC

[490] D. D. Wilmarth. Hot topics: The compute cluster and other parallel programming models. Computer, 26(8): 70–72, August 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ue

[491] Anonymous. Update. Computer, 26 (8):73–??, August 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNg

[492] Anonymous. CS news. Computer, 26 (8):74–??, August 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buckley:1993:SDS

[493] F. J. Buckley. Standards: defining software engineering as a profession. Computer, 26(8):76–78, August 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRg

[494] Anonymous. Product reviews. Computer, 26(8):84–??, August 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRg

[495] Anonymous. Media reviews. Computer, 26(8):101–??, August 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCg

[496] Anonymous. Open channel. Computer, 26(8):104–??, August 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Obraczka:1993:IRD

[497] Katia Obraczka, Peter B. Danzig, and Shih-Hao Li. Internet resource discovery services. Computer, 26(9):8– 22, September 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schwartz:1993:IRD

[498] Michael F. Schwartz. Internet resource discovery at the University of Colorado. Computer, 26(9):25– 35, September 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chai:1993:MDS

[499] Arthur Chai and Sumit Ghosh. Modeling and distributed simulation of a Broadband-ISDN network. Computer, 26(9):37–51, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pakstas:1993:CNE

[500] Algirdas Pakstas and Sonata Pakstiene. Computer networks in Estonia, Latvia, and Lithuania. Computer, 26 (9):53–64, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sha:1993:SAD

[501] Lui Sha and Shirish S. Sathaye. A systematic approach to designing distributed real-time systems. Computer, 26(9):68–78, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:S

[502] Anonymous. Standards. Computer, 26 (9):80–??, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Satyanarayanan:1993:HTM

[503] M. Satyanarayanan. Hot topics: Mobile computing. Computer, 26(9):81– 82, September 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:ES

[504] Anonymous. Election section. Computer, 26(9):83–??, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Uf

[505] Anonymous. Update. Computer, 26(9): 94–??, September 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRh

[506] Anonymous. Product reviews. Computer, 26(9):101–??, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Cf

[507] Anonymous. Conferences. Computer, 26(9):109–??, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRh

[508] Anonymous. Media reviews. Computer, 26(9):125–??, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCh

[509] Anonymous. Open channel. Computer, 26(9):128–??, September 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:LEb

[510] Anonymous. Letters to the editor. Computer, 26(10):4–??, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sameh:1993:GEI

[511] Ahmed Sameh and John Riganati. Guest editors’ introduction: Computational science and engineering. Computer, 26(10):8–??, October 1993. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woodward:1993:ISV

[512] Paul R. Woodward. Interactive scientific visualization of fluid flow. Computer, 26(10):13–25, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tezduyar:1993:PFC

[513] Tayfun Tezduyar, Shahrouz Aliabadi, Marek Behr, Andrew Johnson, and Sanjay Mittal. Parallel finite-element computation of 3D flows. Computer, 26(10):27–36, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Noor:1993:PCS

[514] Ahmed K. Noor and Samuel L. Venneri. A perspective on computational structures technology. Computer, 26 (10):38–46, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Case:1993:CSP

[515] David A. Case. Computer simulations of protein dynamics and thermodynamics. Computer, 26(10):47– 57, October 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1993:CMB

[516] Christopher R. Johnson, Robert S. MacLeod, and Michael A. Matheson. Computational medicine: Bioelectric field problems. Computer, 26(10): 59–67, October 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bamford:1993:SCI

[517] R. C. Bamford and W. J. Deibler, II. Standards: Comparing, contrasting ISO 9001 and the SEI capability maturity model. Computer, 26(10): 68–70, October 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiser:1993:HTU

[518] M. Weiser. Hot topics: ubiquitous computing. Computer, 26(10): 71–72, October 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ug

[519] Anonymous. Update. Computer, 26 (10):72–??, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRi

[520] Anonymous. Product reviews. Computer, 26(10):74–??, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRi

[521] Anonymous. Media reviews. Computer, 26(10):101–??, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCi

[522] Anonymous. Open channel. Computer, 26(10):104–??, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sandhu:1993:LAC

[523] Ravi S. Sandhu. Lattice-based access control models. Computer, 26(11): 9–19, November 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1993:GSA

[524] Kurtis T. Johnson, A. R. Hurson, and Behrooz Shirazi. General-purpose systolic arrays. Computer, 26(11):20– 31, November 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nielsen:1993:IUD

[525] Jakob Nielsen. Iterative user-interface design. Computer, 26(11):32–41, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBenedictis:1993:EUS

[526] Erik P. DeBenedictis and Stephen C. Johnson. Extending Unix for scalable computing. Computer, 26(11): 43–53, November 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

White:1993:SEC

[527] Stephanie White, Mack Alford, Julian Holtzman, Stephen Kuehl, Brian Mc-Cay, D. Oliver, D. Owens, C. Tully, and A. Willey. Systems engineering of computer-based systems. Computer, 26 (11):54–65, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRj

[528] Anonymous. Product reviews. Computer, 26(11):66–??, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNh

[529] Anonymous. CS news. Computer, 26 (11):76–??, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fox:1993:HTD

[530] E. A. Fox. Hot topics: Digital libraries. Computer, 26(11):79– 81, November 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leathrum:1993:SIR

[531] James F. Leathrum and Kathleen A. Liburdy. Standards: Impact of the role of testing in Posix is expanding. Computer, 26(11):81–82, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Cg

[532] Anonymous. Conferences. Computer, 26(11):83–??, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRj

[533] Anonymous. Media reviews. Computer, 26(11):101–??, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCj

[534] Anonymous. Open channel. Computer, 26(11):104–??, November 1993.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PMb

[535] Anonymous. President’s message. Computer, 26(12):4–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:EMc

[536] Anonymous. EIC’s message. Computer, 26(12):6–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roman:1993:TPV

[537] Gruia-Catalin Roman and Kenneth C. Cox. A taxonomy of program visualization systems. Computer, 26(12): 11–24, December 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuggetta:1993:CCT

[538] Alfonso Fuggetta. A classification of CASE technology. Computer, 26(12): 25–38, December 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:1993:FHS

[539] Sanjaya Kumar, James H. Aylor, Barry W. Johnson, and Wm. A. Wulf. A framework for hardware/ software codesign. Computer, 26(12): 39–45, December 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wenban:1993:CCP

[540] Alan S. Wenban, John W. O’Leary, and Geoffrey M. Brown. Codesign of communication protocols. Computer, 26 (12):46–52, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clarke:1993:ALR

[541] Roger Clarke. Asimov’s Laws of Robotics: Implications for Information Technology. Part 1. Computer, 26(12): 53–61, December 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Press:1993:SED

[542] Larry Press. Software export from developing nations. Computer, 26(12): 62–67, December 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:HT

[543] Anonymous. Hot topics. Computer, 26 (12):68–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:CNi

[544] Anonymous. CS news. Computer, 26 (12):69–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:Ch

[545] Anonymous. Conferences. Computer, 26(12):75–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:PRk

[546] Anonymous. Product reviews. Computer, 26(12):86–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MRk

[547] Anonymous. Media reviews. Computer, 26(12):99–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:OCk

[548] Anonymous. Open channel. Computer, 26(12):128–??, December 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PMa

[549] Anonymous. President’s message. Computer, 27(1):6–??, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sha:1994:GEI

[550] Lui Sha. Guest editor’s introduction: Industrial computing: A grand challenge. Computer, 27(1):12–13, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kopetz:1994:TPF

[551] Hermann Kopetz and Gunter Grunsteidl. TTP — a protocol for faulttolerant real-time systems. Computer, 27(1):14–23, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Klein:1994:RAR

[552] Mark H. Klein, John P. Lehoczky, and Ragunathan Rajkumar. Ratemonotonic analysis for real-time industrial computing. Computer, 27(1): 24–33, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Malcolm:1994:TPR

[553] Nicholas Malcolm and Wei Zhao. The timed-token protocol for real-time communications. Computer, 27(1):35– 41, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1994:CC

[554] Nam-Sung S. Woo, Alfred E. Dunlop, and Wayne Wolf. Codesign from cospecification. Computer, 27(1):42– 47, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:1994:PIS

[555] Rajesh K. Gupta, N. Coelho Claudionor, Jr., and Giovanni De Micheli. Program implementation schemes for hardware-software systems. Computer, 27(1):48–55, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clarke:1994:ALR

[556] Roger Clarke. Asimov’s Laws of Robotics: Implications for Information Technology. Part 2. Computer, 27(1): 57–66, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1994:GBP

[557] Alan H. Karp, Don Heller, and Horst Simon. 1993 Gordon Bell Prize winners. Computer, 27(1):69–??, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNa

[558] Anonymous. CS news. Computer, 27 (1):76–??, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Ua

[559] Anonymous. Update. Computer, 27 (1):79–??, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:HTa

[560] Anonymous. Hot topics. Computer, 27 (1):81–??, January 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRa

[561] Anonymous. Product reviews. Computer, 27(1):85–??, January 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Ca

[562] Anonymous. Conferences. Computer, 27(1):97–??, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRa

[563] Anonymous. Media reviews. Computer, 27(1):125–??, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCa

[564] Anonymous. Open channel. Computer, 27(1):128–??, January 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PMb

[565] Anonymous. President’s message. Computer, 27(2):5–??, February 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMa

[566] Anonymous. Editor-in-chief’s message. Computer, 27(2):8–??, February 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lilja:1994:EPA

[567] David J. Lilja. Exploiting the parallelism available in loops. Computer, 27 (2):13–26, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanir:1994:DRS

[568] Oryal Tanir and Suleyman Sevinc. Defining requirements for a standard simulation environment. Computer, 27 (2):28–34, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shere:1994:MDT

[569] Kenneth D. Shere and Rachelle A. Carlson. A methodology for design, test, and evaluation of real-time systems. Computer, 27(2):35–48, February 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baratta-Perez:1994:ASD

[570] Grace Baratta-Perez, Richard L. Conn, Charles A. Finnell, and Thomas J. Walsh. Ada System Dependency Analyzer tool. Computer, 27(2):49– 55, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buckley:1994:ISC

[571] Fletcher J. Buckley. Implementing a software configuration management environment. Computer, 27(2):56– 61, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Biermann:1994:CSM

[572] Alan W. Biermann. Computer science for the many. Computer, 27(2): 62–73, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:HTb

[573] Anonymous. Hot topics. Computer, 27 (2):76–??, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Sa

[574] Anonymous. Standards. Computer, 27 (2):80–??, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNb

[575] Anonymous. CS news. Computer, 27 (2):82–??, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Ub

[576] Anonymous. Update. Computer, 27 (2):84–??, February 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRb

[577] Anonymous. Product reviews. Computer, 27(2):92–??, February 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRb

[578] Anonymous. Media reviews. Computer, 27(2):117–??, February 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCb

[579] Anonymous. Open channel. Computer, 27(2):120–??, February 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PMc

[580] Anonymous. President’s message. Computer, 27(3):5–??, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMb

[581] Anonymous. Editor-in-chief’s message. Computer, 27(3):6–??, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:HTc

[582] Anonymous. Hot topics. Computer, 27 (3):10–??, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patt:1994:GEI

[583] Yale N. Patt. Guest editor’s introduction: The I/O subsystem — A candidate for improvement. Computer, 27 (3):14–??, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patt:1994:ISC

[584] Y. N. Patt. The I/O subsystem — a candidate for improvement. Computer, 27(3):15–16, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ruemmler:1994:IDD

[585] Chris Ruemmler and John Wilkes. An introduction to disk drive modeling. Computer, 27(3):17–28, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ganger:1994:DAH

[586] Gregory R. Ganger, Bruce L. Worthington, Robert Y. Hou, and Yale N. Patt. Disk arrays: High-performance, high-reliability storage subsystems. Computer, 27(3):30–36, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karedla:1994:CSI

[587] Ramakrishna Karedla, J. Spencer Love, and Bradley G. Wherry. Caching strategies to improve disk system performance. Computer, 27(3):38–46, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Steenkiste:1994:SAH

[588] Peter A. Steenkiste. A systematic approach to host interface design for high-speed networks. Computer, 27 (3):47–57, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

delRosario:1994:HPO

[589] Juan Miguel del Rosario and Alok N. Choudhary. High-performance I/O for massively parallel computers: problems and prospects. Computer, 27 (3):59–68, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reddy:1994:OIM

[590] A. L. Narasimha Reddy and James C. Wyllie. I/O issues in a multimedia system. Computer, 27(3):69–74, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krieger:1994:ASF

[591] Orran Krieger, Michael Stumm, and Ron Unrau. The Alloc Stream Facility: A redesign of application-level stream I/O. Computer, 27(3):75–82, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pasquale:1994:CSO

[592] Joseph Pasquale, Eric Anderson, and P. Keith Muller. Container shipping: Operating system support for I/Ointensive applications. Computer, 27 (3):84–93, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRc

[593] Anonymous. Product reviews. Computer, 27(3):97–??, March 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Fa

[594] Anonymous. Frameworks. Computer, 27(3):104–??, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Sb

[595] Anonymous. Standards. Computer, 27 (3):110–??, March 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRc

[596] Anonymous. Media reviews. Computer, 27(3):125–??, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCc

[597] Anonymous. Open channel. Computer, 27(3):128–??, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMc

[598] Anonymous. Editor-in-chief’s message. Computer, 27(4):6–??, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITa

[599] Anonymous. Industry trends. Computer, 27(4):8–??, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:SCa

[600] Anonymous. Software challenges. Computer, 27(4):12–??, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rooholamini:1994:FRA

[601] Reza Rooholamini, Vladimir Cherkassky, and Mark Garver. Finding the right ATM switch for the market. Computer, 27(4):16–28, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macedonia:1994:MPA

[602] Michael R. Macedonia and Donald P. Brutzman. MBone provides audio and video across the Internet. Computer, 27(4):30–36, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Forman:1994:CMC

[603] George H. Forman and John Zahorjan. The challenges of mobile computing. Computer, 27(4):38–47, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Steele:1994:ARR

[604] Robert D. Steele and Paul G. Backes. Ada and real-time robotics: Lessons learned. Computer, 27(4):49–54, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Scandura:1994:CLC

[605] Joseph M. Scandura. Converting legacy code into Ada: A cognitive approach. Computer, 27(4):55–61, April 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNc

[606] Anonymous. CS news. Computer, 27 (4):63–??, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Fb

[607] Anonymous. Frameworks. Computer, 27(4):65–??, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Sc

[608] Anonymous. Standards. Computer, 27 (4):68–??, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRd

[609] Anonymous. Product reviews. Computer, 27(4):73–??, April 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Cb

[610] Anonymous. Conferences. Computer, 27(4):96–??, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRd

[611] Anonymous. Media reviews. Computer, 27(4):101–??, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCd

[612] Anonymous. Open channel. Computer, 27(4):104–??, April 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMd

[613] Anonymous. Editor-in-chief’s message. Computer, 27(5):6–??, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITb

[614] Anonymous. Industry trends. Computer, 27(5):8–??, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1994:SCW

[615] C. Jones. Software challenges: the weakest link in the software engineering chain. Computer, 27(5):10–11, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Palmer:1994:GEI

[616] James D. Palmer and N. Ann Fields. Guest editors’ introduction: Computer-supported cooperative work. Computer, 27(5):15–17, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grudin:1994:CCW

[617] Jonathan Grudin. Computer-supported cooperative work: history and focus. Computer, 27(5):19–26, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reinhard:1994:CTC

[618] Walter Reinhard, Jean Schweitzer, Gerd Volksen, and Michael Weber. CSCW tools: Concepts and architectures. Computer, 27(5):28–36, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bentley:1994:ASC

[619] Richard Bentley, Tom Rodden, Pete Sawyer, and Ian Sommerville. Architectural support for cooperative multiuser interfaces. Computer, 27(5): 37–46, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tou:1994:PSG

[620] Ivan Tou, Steven Berson, Gerald Estrin, Yadran Eterovic, and Elsie Wu. Prototyping synchronous group applications. Computer, 27(5):48–56, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:1994:CSS

[621] Hsinchun Chen. Collaborative systems: Solving the vocabulary problem. Computer, 27(5):58–66, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Palmer:1994:MDS

[622] James D. Palmer, N. Ann Fields, and Peggy Lane Brouse. Multigroup decision-support systems in CSCW. Computer, 27(5):67–72, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liang:1994:WCS

[623] Ting-Peng Liang, Hsiangchu Lai, Nian-Shing Chen, Hungshiung Wei, and Meng Chang Chen. When client/ server isn’t enough: coordinating multiple distributed tasks. Computer, 27 (5):73–79, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRe

[624] Anonymous. Product reviews. Computer, 27(5):81–??, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNd

[625] Anonymous. CS news. Computer, 27 (5):88–??, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blus:1994:SPW

[626] G. Blus. Standards: Public Windows Interface supporters aim to standardize leading desktop API. Computer, 27(5):92–93, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Cc

[627] Anonymous. Conferences. Computer, 27(5):96–??, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRe

[628] Anonymous. Media reviews. Computer, 27(5):109–??, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCe

[629] Anonymous. Open channel. Computer, 27(5):112–??, May 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PMd

[630] Anonymous. President’s message. Computer, 27(6):5–??, June 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMe

[631] Anonymous. Editor-in-chief’s message. Computer, 27(6):6–??, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITc

[632] Anonymous. Industry trends. Computer, 27(6):10–??, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srinivas:1994:GAS

[633] M. Srinivas and Lalit M. Patnaik. Genetic algorithms: A survey. Computer, 27(6):17–26, June 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

RibeiroFilho:1994:GPE

[634] Jose L. Ribeiro Filho, Philip C. Treleaven, and Cesare Alippi. Geneticalgorithm programming environments. Computer, 27(6):28–43, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1994:PAT

[635] James E. Smith and Shlomo Weiss. PowerPC 601 and Alpha 21064: A tale of two RISCs. Computer, 27 (6):46–58, June 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dvorak:1994:CEE

[636] Joseph Dvorak. Conceptual entropy and its effect on class hierarchies. Computer, 27(6):59–63, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kumar:1994:OTH

[637] Sanjaya Kumar, James H. Aylor, Barry W. Johnson, and Wm. A. Wulf. Object-oriented techniques in hardware design. Computer, 27(6):64–70, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kurki-Suonio:1994:RTF

[638] Reino Kurki-Suonio. Real time: further misconceptions (or half-truths). Computer, 27(6):71–76, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

CsNews:1994:NTS

[639] CS News. New technical segment committee. Computer, 27(6):77–??, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Uc

[640] Anonymous. Update. Computer, 27 (6):78–??, June 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRf

[641] Anonymous. Product reviews. Computer, 27(6):81–??, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:SCc

[642] Anonymous. Software challenges. Computer, 27(6):90–??, June 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRf

[643] Anonymous. Media reviews. Computer, 27(6):102–??, June 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCf

[644] Anonymous. Open channel. Computer, 27(6):104–??, June 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:TMA

[645] Anonymous. Treasurer’s message and auditor’s report. Computer, 27(7):5–??, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMf

[646] Anonymous. Editor-in-chief’s message. Computer, 27(7):9–??, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:LE

[647] Anonymous. Letters to the editor. Computer, 27(7):12–??, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheek:1994:ITB

[648] M. Cheek. Industry trends: 32-bit computing: road to confusion? Computer, 27(7):13–14, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaufman:1994:GEI

[649] Arie E. Kaufman. Guest editor’s introduction: Visualization. Computer, 27(7):18–19, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Samtaney:1994:VFT

[650] Ravi Samtaney, Deborah Silver, Norman Zabusky, and Jim Cao. Visualizing features and tracking their evolution. Computer, 27(7):20–27, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ranjan:1994:VMV

[651] Vishwa Ranjan and Alain Fournier. Volume models for volumetric data. Computer, 27(7):28–36, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anupam:1994:DCV

[652] Vinod Anupam, Chandrajit Bajaj, Daniel Schikore, and Matthew Schikore. Distributed and collaborative visualization. Computer, 27 (7):37–43, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:1994:PVA

[653] Jaswinder Pal Singh, Anoop Gupta, and Marc Levoy. Parallel visualization algorithms: Performance and architectural implications. Computer, 27 (7):45–55, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ribarsky:1994:GCC

[654] William Ribarsky, Eric Ayers, John Eble, and Sougata Mukherjea. Glyphmaker: Creating customized visualizations of complex data. Computer, 27 (7):57–64, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hibbard:1994:IVE

[655] William L. Hibbard, Brian E. Paul, David A. Santek, Charles R. Dyer, Andre L. Battaiola, and Marie-Françoise Voidrot-Martinez. Interactive visualization of Earth and space science computations. Computer, 27(7):65–72, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hanson:1994:IMV

[656] Andrew J. Hanson, Tamara Munzner, and George Francis. Interactive methods for visualizable geometry. Computer, 27(7):73–83, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Globus:1994:FWS

[657] Al Globus and Eric Raible. Fourteen ways to say nothing with scientific visualization. Computer, 27(7):86–88, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rhyne:1994:HTC

[658] T. M. Rhyne. Hot topics: Collaborative modeling and visualization: an EPA HPCC initiative. Computer, 27 (7):92–93, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRg

[659] Anonymous. Product reviews. Computer, 27(7):97–??, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alger:1994:HSW

[660] J. Alger. How to stop worrying and start loving C++. II. Computer, 27 (7):104–106, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1994:SCE

[661] C. Jones. Software challenges: Economics of software reuse. Computer, 27 (7):106–107, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNe

[662] Anonymous. CS news. Computer, 27 (7):109–??, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRg

[663] Anonymous. Media reviews. Computer, 27(7):117–??, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCg

[664] Anonymous. Open channel. Computer, 27(7):120–??, July 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMg

[665] Anonymous. Editor-in-chief’s message. Computer, 27(8):5–??, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITe

[666] Anonymous. Industry trends. Computer, 27(8):8–??, August 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chandra:1994:COL

[667] Rohit Chandra, Anoop Gupta, and John L. Hennessy. COOL: An objectbased language for parallel programming. Computer, 27(8):13–26, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1994:DAM

[668] Ben Lee and A. R. Hurson. Dataflow architectures and multithreading. Computer, 27(8):27–39, August 1994.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1994:SRS

[669] Russ Miller. Special report: The status of parallel processing education. Computer, 27(8):40–??, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coleman:1994:UME

[670] Don Coleman, Dan Ash, Bruce Lowther, and Paul Oman. Using metrics to evaluate software system maintainability. Computer, 27(8):44–49, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneidewind:1994:VME

[671] Norman F. Schneidewind. Validating metrics for ensuring Space Shuttle flight software quality. Computer, 27 (8):50–57, August 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1994:WIC

[672] Ted G. Lewis. Where is computing headed? Computer, 27(8): 59–63, August 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNf

[673] Anonymous. CS news. Computer, 27 (8):64–??, August 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:SCe

[674] Anonymous. Software challenges. Computer, 27(8):66–??, August 1994.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:HTe

[675] Anonymous. Hot topics. Computer, 27 (8):67–??, August 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRh

[676] Anonymous. Product reveiws. Computer, 27(8):73–??, August 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRh

[677] Anonymous. Media reviews. Computer, 27(8):101–??, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCh

[678] Anonymous. Open channel. Computer, 27(8):104–??, August 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PMe

[679] Anonymous. President’s message. Computer, 27(9):5–??, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMh

[680] Anonymous. Editor-in-chief’s message. Computer, 27(9):6–??, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITf

[681] Anonymous. Industry trends. Computer, 27(9):8–??, September 1994.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khoshgoftaar:1994:GEI

[682] Taghi M. Khoshgoftaar and Paul Oman. Guest editors’ introduction: Software metrics. Computer, 27(9): 13–??, September 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grady:1994:SAS

[683] Robert B. Grady. Successfully applying software metrics. Computer, 27(9): 18–25, September 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weller:1994:UMM

[684] Edward F. Weller. Using metrics to manage software projects. Computer, 27(9):27–33, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lanning:1994:MRB

[685] David L. Lanning and Taghi M. Khoshgoftaar. Modeling the relationship between source code complexity and maintenance difficulty. Computer, 27 (9):35–40, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stark:1994:UMM

[686] George Stark, Robert C. Durst, and C. W. Vowell. Using metrics in management decision making. Computer, 27(9):42–48, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulish:1994:CSS

[687] Daniel J. Paulish and Anita D. Carleton. Case studies of software-process-improvement measurement. Computer, 27(9):50–57, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horgan:1994:ASQ

[688] Joseph R. Horgan, Saul London, and Michael R. Lyu. Achieving software quality with testing coverage measures. Computer, 27(9):60–69, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pfleeger:1994:ESE

[689] Shari Lawrence Pfleeger, Norman Fenton, and Stella Page. Evaluating software engineering standards. Computer, 27(9):71–79, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:CNg

[690] Anonymous. CS news. Computer, 27 (9):82–??, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Ud

[691] Anonymous. Update. Computer, 27(9): 92–??, September 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voigt:1994:SEB

[692] R. Voigt. Standards: Evolving buses: backplane buses mature to keep pace with technology. Computer, 27(9):94– 95, September 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berleant:1994:HTC

[693] D. Berleant and H. Berghel. Hot topics: Customizing information. 1. Getting what we need, when we need it. Computer, 27(9):96–98, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1994:SCS

[694] C. Jones. Software challenges: Software metrics: good, bad and missing. Computer, 27(9):98–100, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRi

[695] Anonymous. Product reviews. Computer, 27(9):102–??, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRi

[696] Anonymous. Media reviews. Computer, 27(9):125–??, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCi

[697] Anonymous. Open channel. Computer, 27(9):128–??, September 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMi

[698] Anonymous. Editor-in-chief’s message. Computer, 27(10):6–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITg

[699] Anonymous. Industry trends. Computer, 27(10):8–??, October 1994. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lebeck:1994:CPS

[700] Alvin R. Lebeck and David A. Wood. Cache profiling and the SPEC benchmarks: A case study. Computer, 27 (10):15–26, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Louri:1994:OIH

[701] Ahmed Louri and Hongki Sung. 3D optical interconnects for high-speed interchip and interboard communications. Computer, 27(10):27–37, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thimbleby:1994:DIA

[702] Harold W. Thimbleby, Stuart Inglis, and Ian H. Witten. Displaying 3D images: Algorithms for single-image random-dot stereograms. Computer, 27(10):38–48, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:1994:MWW

[703] Ronald J. Vetter, Chris Spell, and Charles Ward. Mosaic and the World Wide Web. Computer, 27(10):49– 57, October 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Skazinski:1994:PAR

[704] Joseph G. Skazinski. Porting Ada: A report from the field. Computer, 27 (10):58–64, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leveson:1994:HSE

[705] Nancy G. Leveson. High-pressure steam engines and computer software. Computer, 27(10):65–73, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1994:CLI

[706] I. H. Donner. Computer law: Intellectual property protection: everything you’ve always wanted to know. Computer, 27(10):74–75, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berleant:1994:CIH

[707] D. Berleant and H. Berghel. Hot topics: Customizing information. 2. How successful are we so far? Computer, 27(10):76–78, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:SCg

[708] Anonymous. Software challenges. Computer, 27(10):79–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRj

[709] Anonymous. Product reviews. Computer, 27(10):86–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Sf

[710] Anonymous. Standards. Computer, 27(10):92–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Ue

[711] Anonymous. Update. Computer, 27 (10):94–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRj

[712] Anonymous. Media reviews. Computer, 27(10):110–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCj

[713] Anonymous. Open channel. Computer, 27(10):112–??, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ECM

[714] Anonymous. Editor-in chief’s message. Computer, 27(11):5–??, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Snell:1994:ITS

[715] M. Snell. Industry trends: Supercomputing: same name, different game. Computer, 27(11):6–8, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krikelis:1994:GEI

[716] Anargyros Krikelis and Charles C. Weems. Guest editors’ introduction: Associative processing and processors. Computer, 27(11):12–17, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Potter:1994:AAP

[717] Jerry Potter, Johnnie Baker, Stephen Scott, Arvind Bansal, Chokchai Leangsuksun, and Chandra Asthagiri. ASC: An associative-computing paradigm. Computer, 27(11):19–25, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Twardowski:1994:AAG

[718] Kirk Twardowski. An associative architecture for genetic algorithm-based machine learning. Computer, 27(11): 27–38, November 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tavangarian:1994:FPA

[719] Djamshid Tavangarian. Flag-oriented parallel associative architectures and applications. Computer, 27(11):41– 52, November 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Higuchi:1994:IPA

[720] Tetsuya Higuchi, Kennichi Handa, Naoto Takahashi, Tatsumi Furuya, Hitoshi Iida, E. Sumita, O. Oi, and H. Kitano. The IXM2 parallel associative processor for AI. Computer, 27(11): 53–63, November 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Louri:1994:OAP

[721] Ahmed H. Louri and James A. Hatch, Jr. An optical associative parallel processor for high-speed database processing. Theoretical concepts and experimental results. Computer, 27(11):65– 72, November 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Furht:1994:PCG

[722] Borko Furht. Parallel computing: Glory and collapse. Computer, 27(11): 74–75, November 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Uf

[723] Anonymous. Update. Computer, 27 (11):76–??, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:SCh

[724] Anonymous. Software challenges. Computer, 27(11):79–??, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRk

[725] Anonymous. Product reviews. Computer, 27(11):81–??, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

White:1994:MAM

[726] J. E. White. Hot topics: Mobile agents make a network an open platform for third-party developers. Computer, 27 (11):89–90, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1994:CLR

[727] I. H. Donner. Computer law: Robo-Mop gets wrung out (software-related patents). Computer, 27(11):91–92, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:SSD

[728] Anonymous. Standards: Streamlining the standards development process. Computer, 27(11):93–94, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRk

[729] Anonymous. Media reviews. Computer, 27(11):109–??, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCk

[730] Anonymous. Open channel. Computer, 27(11):112–??, November 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:EMj

[731] Anonymous. Editor-in-chief’s message. Computer, 27(12):6–??, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:ITi

[732] Anonymous. Industry trends. Computer, 27(12):8–??, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rich:1994:DIM

[733] Charles Rich, Richard C. Waters, Carol Strohecker, Yves Schabes, William T. Freeman, Mark C. Torrance, Andrew R. Golding, and Michal Roth. Demonstration of an interactive multimedia environment. Computer, 27(12): 15–22, December 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sachs:1994:LIC

[734] Martin W. Sachs, Avraham Leff, and Denise Sevigny. LAN and I/O convergence: A survey of the issues. Computer, 27(12):24–32, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gross:1994:CSP

[735] Thomas Gross, Susan Hinrichs, David R. O’Hallaron, Thomas Stricker, and Atsushi Hasegawa. Communication styles for parallel systems. Computer, 27(12):34–44, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gore:1994:IDI

[736] Al Gore. Innovation delayed is innovation denied. Computer, 27(12): 45–47, December 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wood:1994:ITS

[737] Gay A. Wood, Seymour E. Goodman, and Jan Roos. Information technologies in South Africa: Problems and prospects. Computer, 27(12):48– 56, December 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davies:1994:CDG

[738] Mark Davies, Kris Pettersen, Shankar Srinivasen, and Katherine Kinsman. The challenges of designing groupware: A case study. Computer, 27(12):58– 62, December 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:PRl

[739] Anonymous. Product reviews. Computer, 27(12):65–??, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:Ug

[740] Anonymous. Update. Computer, 27 (12):73–??, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berczuk:1994:HTF

[741] S. Berczuk. Hot topics: Finding solutions through pattern languages. Computer, 27(12):75–76, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1994:SCB

[742] C. Jones. Software challenges: Build, buy, or outsource? Computer, 27(12): 77–78, December 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1994:POC

[743] I. H. Donner. Computer law: Is the Patent Office correctly examining computer-related patent applications? Part 1. Computer, 27(12): 78–79, December 1994. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:MRl

[744] Anonymous. Media reviews. Computer, 27(12):103–??, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1994:OCl

[745] Anonymous. Open channel. Computer, 27(12):120–??, December 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PM

[746] Anonymous. President’s message. Computer, 28(1):5–??, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:EM

[747] Anonymous. Editor-in-chief’s message. Computer, 28(1):6–??, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCa

[748] Anonymous. Binary critic. Computer, 28(1):8–??, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITa

[749] Anonymous. Industry trends. Computer, 28(1):9–??, January 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saiedian:1995:SCM

[750] Hossein Saiedian and Richard Kuzara. SEI capability maturity model’s impact on contractors. Computer, 28(1): 16–26, January 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gal:1995:MDR

[751] Avigdor Gal and Opher Etzion. Maintaining data-driven rules in databases. Computer, 28(1):28–38, January 1995.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carriero:1995:APP

[752] Nicholas Carriero, Eric Freeman, David Gelernter, and David Kaminsky. Adaptive parallelism and Piranha. Computer, 28(1):40–49, January 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lomas:1995:RBH

[753] Mark Lomas and Bruce Christianson. Remote booting in a hostile world: to whom am I speaking? Computer, 28 (1):50–54, January 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Glass:1995:SSB

[754] Robert L. Glass. In search of self-belief: The “BOP” phenomenon. Computer, 28(1):55–57, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Musliner:1995:CRA

[755] David J. Musliner, James A. Hendler, Ashok K. Agrawala, Edmund H. Durfee, Jay K. Strosnider, and C. J. Paul. The challenges of real-time AI. Computer, 28(1):58–66, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1995:SRG

[756] Alan H. Karp, Michael Heath, Don Heller, and Horst Simon. Special report: 1994 Gordon Bell Prize winners. Computer, 28(1):68–74, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCH

[757] Capers Jones. Software challenges: How office space affects programming productivity. Computer, 28(1):76– 77, January 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:1995:HTV

[758] Ronald J. Vetter. Hot topics: Videoconferencing on the Internet. Computer, 28(1):77–79, January 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRa

[759] Anonymous. Product reviews. Computer, 28(1):81–??, January 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:Ua

[760] Anonymous. Update. Computer, 28 (1):88–??, January 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLIa

[761] I. H. Donner. Computer law: Is the Patent Office correctly examining computer-related patent applications?

2. Computer, 28(1):93–94, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:MRa

[762] Anonymous. Media reviews. Computer, 28(1):117–??, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCa

[763] Anonymous. Open channel. Computer, 28(1):120–??, January 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCb

[764] Anonymous. Binary critic. Computer, 28(2):6–??, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITb

[765] Anonymous. Industry trends. Computer, 28(2):7–??, February 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dollas:1995:GEI

[766] Apostolos Dollas and Nick Kanopoulos. Guest editors’ introduction: Reducing the time to market through rapid prototyping. Computer, 28(2): 14–15, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athanas:1995:RIP

[767] Peter M. Athanas and A. Lynn Abbott. Real-time image processing on a custom computing platform. Computer, 28 (2):16–24, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barroso:1995:RRP

[768] Luiz Andre Barroso, Sasan Iman, Jaeheon Jeong, Koray Oner, Michel Dubois, and Krishnan Ramamurthy. RPM: A rapid prototyping engine for multiprocessor systems. Computer, 28 (2):26–34, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lauwereins:1995:GIS

[769] Rudy Lauwereins, Marc Engels, Marleen Ade, and J. A. Peperstraete. Grape-II: a system-level prototyping environment for DSP applications. Computer, 28(2):35–43, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ismail:1995:SSD

[770] Tarek Ben Ismail and Ahmed Amine Jerraya. Synthesis steps and design models for codesign. Computer, 28(2): 44–52, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baldwin:1995:FAM

[771] Reid A. Baldwin and Moon Jung Chung. A formal approach to managing design processes. Computer, 28(2): 54–63, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wirth:1995:PLS

[772] Niklaus Wirth. A plea for lean software. Computer, 28(2):64–68, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wood:1995:CPC

[773] David A. Wood and Mark D. Hill. Cost-effective parallel computing. Computer, 28(2):69–72, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCD

[774] Capers Jones. Software challenges: Determining software schedules. Computer, 28(2):73–75, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:HT

[775] Anonymous. Hot topics. Computer, 28 (2):75–??, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Richards:1995:SLS

[776] Mark A. Richards. Simulation libraries for system-level design. Computer, 28 (2):76–77, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CL

[777] Anonymous. Computer law. Computer, 28(2):78–??, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRb

[778] Anonymous. Product reviews. Computer, 28(2):81–??, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:Ub

[779] Anonymous. Update. Computer, 28 (2):87–??, February 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:MRb

[780] Anonymous. Media reviews. Computer, 28(2):108–??, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCb

[781] Anonymous. Open channel. Computer, 28(2):112–??, February 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCc

[782] Anonymous. Binary critic. Computer, 28(3):6–??, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITe

[783] Anonymous. Industry trends. Computer, 28(3):8–??, March 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Burnett:1995:GEI

[784] Margaret M. Burnett and David W. McIntyre. Guest editors’ introduction: Visual programming. Computer, 28 (3):14–16, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Repenning:1995:AMC

[785] Alexander Repenning and Tamara Sumner. Agentsheets: A medium for creating domain-oriented visual languages. Computer, 28(3):17–25, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kimura:1995:FFV

[786] Takayuki Dan Kimura, Ajay Apte, Samudra Sengupta, and Julie W. Chan. Form/formula: a visual programming paradigm for user-definable user interfaces. Computer, 28(3):27–35, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karsai:1995:CVP

[787] Gabor Karsai. A configurable visual programming environment: A tool for domain-specific programming. Computer, 28(3):36–44, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Burnett:1995:SVP

[788] Margaret M. Burnett, Marla J. Baker, Carisa Bohus, Paul Carlson, Sherry Yang, and Pieter van Zee. Scaling up visual programming languages. Computer, 28(3):45–54, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Costagliola:1995:AGV

[789] Gennaro Costagliola, Genoveffa Tortora, Sergio Orefice, and Andrea De Lucia. Automatic generation of visual programming environments. Computer, 28(3):56–66, March 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adler:1995:ESC

[790] Richard M. Adler. Emerging standards for component software. Computer, 28 (3):68–77, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1995:TCP

[791] Ware Myers. Taligent’s CommonPoint: The promise of objects. Computer, 28 (3):78–83, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bull:1995:HTC

[792] G. Bull, R. Geyer, D. Arkinn, and C. Appert. Hot topics: Current technologies in K–12 education. Computer, 28(3):84–85, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCP

[793] C. Jones. Software challenges: Patterns of large software systems: failure and success. Computer, 28(3):86–87, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRd

[794] Anonymous. Product reviews. Computer, 28(3):92–??, March 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLG

[795] I. H. Donner. Computer law: GATT is gonna get ya! Computer, 28(3): 98–100, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:Uc

[796] Anonymous. Update. Computer, 28 (3):99–??, March 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:MRc

[797] Anonymous. Media reviews. Computer, 28(3):117–??, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCc

[798] Anonymous. Open channel. Computer, 28(3):120–??, March 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCd

[799] Anonymous. Binary critic. Computer, 28(4):7–??, April 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITf

[800] Anonymous. Industry trends. Computer, 28(4):8–??, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adler:1995:DCM

[801] Richard M. Adler. Distributed coordination models for client/server computing. Computer, 28(4):14–22, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gokhale:1995:PMT

[802] Maya Gokhale, Bill Holmes, and Ken Iobst. Processing in memory: the Terasys massively parallel PIM array. Computer, 28(4):23–31, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martonosi:1995:TMP

[803] Margaret Martonosi, Anoop Gupta, and Thomas E. Anderson. Tuning memory performance of sequential and parallel programs. Computer, 28(4): 32–40, April 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wood:1995:PCS

[804] Alan Wood. Predicting client/server availability. Computer, 28(4):41–48, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1995:WCS

[805] Ted G. Lewis. Where is client/server software headed? Computer, 28 (4):49–55, April 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowen:1995:TCF

[806] Jonathan P. Bowen and Michael G. Hinchey. Ten commandments of formal methods. Computer, 28(4):56–63, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weingarten:1995:UFR

[807] Fred W. Weingarten. Update: Federal research funding: reading the tea leaves. Computer, 28(4):64–65, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liebeherr:1995:HTM

[808] Jorg Liebeherr. Hot topics: Multimedia networks: issues and challenges. Computer, 28(4):68–69, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCG

[809] Capers Jones. Software challenges: Gaps in programming education. Computer, 28(4):70–71, April 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLB

[810] Irah H. Donner. Computer law: Billings taxes Novell with a novel patent-infringement claim. Computer, 28(4):72–73, April 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRe

[811] Anonymous. Product reviews. Computer, 28(4):74–??, April 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCd

[812] Anonymous. Open channel. Computer, 28(4):96–??, April 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blank:1995:FCL

[813] C. Blank. The FSN challenge: largescale interactive TV. Computer, 28 (5):9–12, May 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Furht:1995:DII

[814] Borko Furht, Deven Kalra, Frederick L. Kitson, Arturo A. Rodriguez, and William E. Wall. Design issues for interactive television systems. Computer, 28(5):25–39, May 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gemmell:1995:MSS

[815] D. James Gemmell, Harrick M. Vin, Dilip D. Kandlur, P. Venkat Rangan, and Lawrence A. Rowe. Multimedia storage servers: a tutorial. Computer, 28(5):40–49, May 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nahrstedt:1995:RMN

[816] Klara Nahrstedt and Ralf Steinmetz. Resource management in networked multimedia systems. Computer, 28 (5):52–63, May 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simon:1995:MMM

[817] Robert Simon, Donald Krieger, Taieb Znati, Raymond Lofink, and Robert J. Sclabassi. Multimedia MedNet: a medical collaboration and consultation system. Computer, 28(5):65–73, May 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woolf:1995:MPI

[818] Beverly Park Woolf and Wendy Hall. Multimedia pedagogues: interactive systems for teaching and learning. Computer, 28(5):74–80, May 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:BCA

[819] I. H. Donner. Be careful about disclosing your invention. Computer, 28(5):83, May 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thomas:1995:BFI

[820] C. G. Thomas. BASAR: a framework for integrating agents in the World Wide Web. Computer, 28 (5):84–86, May 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:LSS

[821] C. Jones. Legal status of software engineering. Computer, 28(5):98–99, May 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weingarten:1995:OAG

[822] F. W. Weingarten. On-line access to government information. Computer, 28 (5):100–101, May 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1995:BCW

[823] T. Lewis. Binary critic: Where the big money is. Computer, 28(6):6–7, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITg

[824] Anonymous. Industry trends. Computer, 28(6):8–??, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stankovic:1995:ICS

[825] John A. Stankovic, Marco Spuri, Marco Di Natale, and Giorgio C. Buttazzo. Implications of classical scheduling results for real-time systems. Computer, 28(6):16–25, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tanir:1995:SAD

[826] Oryal Tanir, Vinod K. Agarwal, and P. C. P. Bhatt. A specificationdriven architectural design environment. Computer, 28(6):26–35, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walker:1995:IPC

[827] Wade Walker and Harvey G. Cragon. Interrupt processing in concurrent processors. Computer, 28(6):36–46, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clark:1995:FIM

[828] Jeffrey A. Clark and Dhiraj K. Pradhan. Fault injection: A method for validating computer-system dependability. Computer, 28(6):47–56, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boman:1995:ISV

[829] Duane K. Boman. International survey: Virtual-environment research. Computer, 28(6):57–65, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Press:1995:RNL

[830] Larry Press. Resources for networks in less-industrialized nations. Computer, 28(6):66–71, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schroeder:1995:OMT

[831] Beth A. Schroeder. On-line monitoring: A tutorial. Computer, 28 (6):72–78, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1995:SRE

[832] S. Hamilton. Special report: Evolving the HPCCI. recommendations from the NRC/CSTB 1995 Report. Computer, 28(6):79–81, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ram:1995:IKS

[833] P. Ram and D. K. Rand. Internet kiosk: Satan: double-edged sword. Computer, 28(6):82–83, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLP

[834] I. H. Donner. Computer law: Patent or perish: the best defense is a good offense. Computer, 28(6):84–85, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCWa

[835] C. Jones. Software challenges: Why is technology transfer so hard? Computer, 28(6):86–87, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:Ue

[836] Anonymous. Update. Computer, 28 (6):88–??, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRf

[837] Anonymous. Product reviews. Computer, 28(6):92–??, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAFa

[838] Anonymous. Technical activity forum. Computer, 28(6):100–??, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCe

[839] Anonymous. Open channel. Computer, 28(6):120–??, June 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAR

[840] Anonymous. Treasurer’s and auditor’s reports. Computer, 28(7):5–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCf

[841] Anonymous. Binary critic. Computer, 28(7):10–??, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bournellis:1995:ITC

[842] C. Bournellis. Industry trends: CD-ROM technology takes off. Computer, 28(7):11–12, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ASa

[843] Anonymous. Article summaries. Computer, 28(7):15–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pratt:1995:GEI

[844] David R. Pratt, Michael Zyda, and Kristen Kelleher. Guest editors’ introduction: Virtual reality — in the mind of the beholder. Computer, 28 (7):17–19, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goble:1995:TSI

[845] John G. Goble, Ken Hinckley, Randy Pausch, John W. Snell, and Neal F. Kassell. Two-handed spatial interface tools for neurosurgical planning. Computer, 28(7):20–26, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodges:1995:VET

[846] Larry F. Hodges, Rob Kooper, Thomas C. Meyer, Barbara O. Rothbaum, Dan Opdyke, Johannes J. de Graaff, James S. Williford, and Max M. North. Virtual environments for treating the fear of heights. Computer, 28 (7):27–34, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhl:1995:IDS

[847] Jon Kuhl, Douglas Evans, Yiannis Papelis, Richard Romano, and Ginger Watson. The Iowa Driving Simulator: an immersive research environment. Computer, 28(7):35–41, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kruger:1995:RWV

[848] Wolfgang Kruger, Christian-A. Bohn, Bernd Frohlich, Heinrich Schuth, Wolfgang Strauss, and Gerold Wesche. The responsive workbench: A virtual work environment. Computer, 28 (7):42–48, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mastaglio:1995:LCV

[849] Thomas W. Mastaglio and Robert Callahan. A large-scale complex virtual environment for team training. Computer, 28(7):49–56, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Polis:1995:ACL

[850] Michael F. Polis, Stephen J. Gifford, and David M. McKeown, Jr. Automating the construction of largescale virtual worlds. Computer, 28 (7):57–65, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:IMAa

[851] Anonymous. IEEE-CS membership application. Computer, 28(7):64A–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gomory:1995:NPC

[852] Ralph E. Gomory. National productivity and computers. Computer, 28 (7):66–72, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bolas:1995:HTA

[853] M. T. Bolas, I. McDowall, and R. Mead. Hot topics: Applications drive VR interface selection. Computer, 28(7):72–75, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weingarten:1995:USS

[854] Fred W. Weingarten. Update: Sole source for science funding: Yea or nay? ENIAC computer co-inventor J. Presper Eckert dies. Computer, 28 (7):76–78, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAFb

[855] Anonymous. Technical activities forum. Computer, 28(7):80–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:SC

[856] Anonymous. Software challenges. Computer, 28(7):81–??, July 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRg

[857] Anonymous. Product reviews. Computer, 28(7):84–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:APIa

[858] Anonymous. Advertiser/product index and reader service card. Computer, 28(7):88–??, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PSa

[859] Anonymous. Product source. Computer, 28(7):91–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLIb

[860] I. H. Donner. Computer law: Intellectual property protection for multimedia applications. 1. So many flavors, so little time. Computer, 28(7):92–93, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rhyne:1995:IKS

[861] T.-M. Rhyne. Internet kiosk: Scientific visualization and technology transfer: an EPA case study. Computer, 28(7):94–96, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:S

[862] Anonymous. Standards. Computer, 28(7):97–??, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CPCa

[863] Anonymous. Call for papers/calendar. Computer, 28(7):98–??, July 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:COa

[864] Anonymous. Career opportunities. Computer, 28(7):102–??, July 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCf

[865] Anonymous. Open channel. Computer, 28(7):104–??, July 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCW

[866] Anonymous. Binary critic — Windows 95: Next step to desktop NT? Computer, 28(8):8–??, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1995:WIS

[867] Ted Lewis, Dave Power, Bertrand Meyer, Jack Grimes, Mike Potel, Ron Vetter, Phil Laplante, Wolfgang Pree, Gustav Pomberger, Mark Hill, James Larus, David Wood, Hesham El-Rewini, and Bruce W. Weide. Where is software headed? A virtual roundtable. Computer, 28(8):20–32, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Giladi:1995:SPE

[868] Ran Giladi and Niv Ahituv. SPEC as a performance evaluation measure. Computer, 28(8):33–42, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vonMayrhauser:1995:PCD

[869] Anneliese von Mayrhauser and A. Marie Vans. Program comprehension during software maintenance and evolution. Computer, 28(8):44–55, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Estivill-Castro:1995:CSR

[870] Vladimir Estivill-Castro. Computer science research in Mexico. Computer, 28(8):56–62, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kraemer:1995:NPI

[871] Kenneth L. Kraemer and Jason Dedrick. From nationalism to pragmatism: IT policy in China. Computer, 28 (8):64–73, August 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bokhari:1995:LOS

[872] Shahid H. Bokhari. The Linux operating system. Computer, 28(8):

74–79, August 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murray:1995:DCS

[873] J. Murray. Of deck chairs, suits, and quiz shows (Internet). Computer, 28 (8):80–81, August 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:1995:SVF

[874] C. Severance. Standards: The value of the formal standards process. Computer, 28(8):82–83, August 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCWb

[875] Capers Jones. Software challenges: What goes into an information warehouse? Computer, 28(8):84–85, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAFc

[876] Anonymous. Technical activities forum — parallel processing TC: Status and activities. Computer, 28(8): 86–??, August 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jurgens:1995:HTF

[877] Clint Jurgens. Hot topics: Fibre Channel: a connection to the future. Computer, 28(8):88–90, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRC

[878] Anonymous. Product reviews — C++ class libraries, part 2; virtual Unix for PCs. Computer, 28(8):91–??, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLIc

[879] I. H. Donner. Computer law: Intellectual property protection for multimedia application. 2. Putting the pieces together. Computer, 28(8): 99–100, August 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:USE

[880] Anonymous. Update — software engineers surveyed; mythical man-month. Computer, 28(8):100–??, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:LEP

[881] Irah H. Donner. Letters to the editor: The purpose of patents. Computer, 28 (9):6, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Potes:1995:C

[882] Marilyn Potes. The computer 100 . Computer, 28(9):6–7, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1995:LRT

[883] T. Lewis. Living in real time, side A (what is the Info Age?). Computer, 28(9):8–10, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

English:1995:ITG

[884] Erin English. Industry trends: Groupware makes its market move. Computer, 28(9):11–14, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ASb

[885] Anonymous. Article summaries. Computer, 28(9):16–17, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gudivada:1995:GEI

[886] Venkat N. Gudivada and Vijay V. Raghavan. Guest editors’ introduction: Content-based image retrieval systems. Computer, 28(9):18–22, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flickner:1995:QIV

[887] Myron Flickner, Harpreet Sawhney, Wayne Niblack, Jonathan Ashley, Qian Huang, Byron Dom, Monika Gorkani, Jim Hafner, Denis Lee, Dragutin Petkovic, David Steele, and Peter Yanker. Query by image and video content: The QBIC system. Computer, 28 (9):23–32, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brink:1995:HMR

[888] Anne Brink, Sherry Marcus, and V. S. Subrahmanian. Heterogeneous multimedia reasoning. Computer, 28(9):33– 39, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ogle:1995:CRR

[889] Virginia E. Ogle and Michael Stonebraker. Chabot: Retrieval from a relational database of images. Computer, 28(9):40–48, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srihari:1995:AIC

[890] Rohini K. Srihari. Automatic indexing and content-based retrieval of captioned images. Computer, 28(9):49– 56, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mehrotra:1995:SSR

[891] Rajiv Mehrotra and James E. Gary. Similar-shape retrieval in shape data management. Computer, 28(9):57– 62, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:1995:IKU

[892] H. Berghel. Internet kiosk: Using the WWW Test Pattern to check HTML client compliance. Computer, 28(9): 63–65, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:1995:SWI

[893] C. Severance. Standards: What is a profile? (computer standards). Computer, 28(9):66, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donnar:1995:CLP

[894] I. H. Donnar. Computer law: Photocopyright infringement. Computer, 28 (9):67–68, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCE

[895] C. Jones. Software challenges: Enduser programming. Computer, 28(9): 68–70, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Israel:1995:TAF

[896] Michael Israel. Technical activities forum: Design automation technical committee. Computer, 28(9):71– 72, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ICS

[897] Anonymous. IEEE Computer Society election: Nominees for Computer Society office and Board of Governors. Computer, 28(9):73–82, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weingarten:1995:UOT

[898] Fred W. Weingarten. Update: Office of Technology Assessment to be abolished. Computer, 28(9):82–84, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stoyenko:1995:HTE

[899] Alex Stoyenko. Hot topics: Engineering complex computer systems: A challenge for computer types everywhere. Part I: Let’s agree on what these systems are. Computer, 28(9): 85–86, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eckhouse:1995:PRW

[900] Richard Eckhouse. Product reviews: Windows 95: A preview. Computer, 28(9):89–91, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sastry:1995:PO

[901] David C. Sastry and Cristian P. Masgras. A paperless office? Computer, 28(9):92–93, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cratty:1995:DVP

[902] John Cratty. Delphi: Visual programming in action. Computer, 28(9):93– 95, September 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CPCb

[903] Anonymous. Call for papers/calendar. Computer, 28(9):100–107, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carpenter:1995:OCN

[904] Paul Carpenter. The open channel: Never mind the functionality, look at the graphics! A cautionary tale. Computer, 28(9):112, September 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:LEWa

[905] Anonymous. Letters to the editor: Windows 95. Computer, 28(10):6– ??, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:LRT

[906] Anonymous. Living in real time, side B (where will the brain power come from?). Computer, 28(10):8– 10, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITT

[907] Anonymous. Industry trends: TCP/ IP update. Computer, 28(10):12– ??, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ASc

[908] Anonymous. Article summaries. Computer, 28(10):16–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swamy:1995:OOE

[909] Sowmitri Swamy, Arthur Molin, and Burt Covnot. OO-VHDL. Objectoriented extensions to VHDL. Computer, 28(10):18–26, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahrens:1995:TLR

[910] Judith D. Ahrens and Noah S. Prywes. Transition to a legacy- and reuse-based software life cycle. Computer, 28(10): 27–36, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Banerjee:1995:PCD

[911] Prithviraj Banerjee, John A. Chandy, Manish Gupta, Eugene W. Hodges IV, John G. Holm, Antonio Lain, Daniel J. Palermo, Shankar Ramaswamy, and Ernesto Su. The Paradigm compiler for distributed-memory multicomputers. Computer, 28(10):37–47, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khan:1995:OPS

[912] Emdad H. Khan, Mansoor Al-A’ali, and Moheb R. Girgis. Object-oriented programming for structured procedural programmers. Computer, 28(10): 48–57, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

El-Rewini:1995:OTV

[913] Hesham El-Rewini and Scott Hamilton. Object technology: A virtual roundtable. Computer, 28(10): 58–59, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shan:1995:DOR

[914] Yen-Ping Shan, Ralph Earle, and Skip McGaughey. Distributed objects: Rounding out the picture: Objects across the client-server spectrum. Computer, 28(10):60, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:1995:DOC

[915] Abdelsalam Helal and Ran Badrachalam. Distributed objects: COM versus CORBA: Will Microsoft come out on top? Computer, 28(10):61–62, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chien:1995:POS

[916] Andrew A. Chien. Parallelism and objects: Scalable parallel systems ? =? scalable object request brokers. Computer, 28(10):62–63, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grimshaw:1995:POP

[917] Andrew Grimshaw. Parallelism and objects: Parallel processing and OT. Computer, 28(10):63–64, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1995:ODO

[918] Byung S. Lee. Object databases: Object-oriented databases: Systems and standards. Computer, 28(10): 64–65, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wade:1995:ODW

[919] Andrew E. Wade. Object databases: Where are object databases headed? Computer, 28(10):65–66, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morse:1995:ODO

[920] Dave Morse. Object databases: OODBSs and the marriage of network and relational database models. Computer, 28(10):66–67, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elmagarmid:1995:ODU

[921] Ahmed Elmagarmid and Evaggelia Pitoura. Object databases: Using object technology for database system integration. Computer, 28(10):67–68, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Binder:1995:SQO

[922] Robert Binder. Software quality and object orientation: Trends in testing object-oriented software. Computer, 28 (10):68–69, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wegner:1995:TFI

[923] Peter Wegner. Theoretical foundations: Interactive foundations of object-based programming. Computer, 28(10):70– 72, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gal-Ezer:1995:HSP

[924] Judith Gal-Ezer, Catriel Beeri, David Harel, and Amiram Yehudai. A high school program in computer science. Computer, 28(10):73–80, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:IMAb

[925] Anonymous. IEEE-CS membership application. Computer, 28(10):80A– ??, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITC

[926] Anonymous. Intel’s teraflop computer. Computer, 28(10):81–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TS

[927] Anonymous. TV standards. Computer, 28(10):81–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:Uf

[928] Anonymous. Update. Computer, 28 (10):81–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:GRS

[929] I. H. Donner. Computer law: Ground rules for software maintenance. Computer, 28(10):84–85, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAFe

[930] Anonymous. Technical activities forum: Computer architecture TC. Computer, 28(10):85–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OTP

[931] Anonymous. OOAD tools, part III. Computer, 28(10):87–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PN

[932] Anonymous. Perl on the net. Computer, 28(10):87–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRh

[933] Anonymous. Product reviews. Computer, 28(10):87–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PSb

[934] Anonymous. Product source. Computer, 28(10):95–??, October 1995.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:APIb

[935] Anonymous. Advertiser/product index and reader service card. Computer, 28(10):96–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:1995:SPA

[936] J. L. Hill. Standards: Publicly available specification. A new paradigm for developing international standards. Computer, 28(10):97–98, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:1995:ISE

[937] R. Oppliger. Internet kiosk: Internet security enters the Middle Ages. Computer, 28(10):100–101, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCS

[938] C. Jones. Software challenges: Software benchmarking. Computer, 28(10): 102–103, October 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:COb

[939] Anonymous. Career opportunities. Computer, 28(10):104–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CPCc

[940] Anonymous. Call for papers/calendar. Computer, 28(10):111–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCM

[941] Anonymous. Open channel: Miss Manners looks at netiquette. Computer, 28 (10):120–??, October 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:LEWb

[942] Anonymous. Letters to the editor: Where is software headed? Computer, 28(11):5–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1995:BCH

[943] T. Lewis. Binary critic: HP means high-powered: FutureBusiness, side A. Computer, 28(11):6–8, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITS

[944] Anonymous. Industry trends: Security still lacking in SNMP upgrade. Computer, 28(11):10–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ASd

[945] Anonymous. Article summaries. Computer, 28(11):14–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pancake:1995:GEI

[946] Cherri M. Pancake, Margaret L. Simmons, and Jerry C. Yan. Guest editors’ introduction: Performance evaluation tools for parallel and distributed systems. Computer, 28(11):16–19, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heath:1995:VDP

[947] Michael T. Heath, Allen D. Malony, and Diane T. Rover. The visual display of parallel performance data. Computer, 28(11):21–28, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jakiela:1995:PVD

[948] Hank Jakiela. Performance visualization of a distributed system: A case study. Computer, 28(11):30– 36, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1995:PPP

[949] Barton P. Miller, Mark D. Callaghan, Jonathan M. Cargille, Jeffrey K. Hollingsworth, R. Bruce Irvin, Karen L. Karavanic, Krishna Kunchithapadam, and Tia Newhall. The Paradyn parallel performance measurement tool. Computer, 28(11):37–46, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fahringer:1995:EOP

[950] Thomas Fahringer. Estimating and optimizing performance for parallel programs. Computer, 28(11):47–56, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reed:1995:VRP

[951] Daniel A. Reed, Keith A. Shields, Will H. Scullin, Luis F. Tavera, and Christopher L. Elford. Virtual reality and parallel systems performance analysis. Computer, 28(11):57–67, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:IMAc

[952] Anonymous. IEEE-CS membership application. Computer, 28(11):64A– ??, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kwan:1995:NWW

[953] Thomas T. Kwan, Robert E. Mc-Grath, and Daniel A. Reed. NCSA’s World Wide Web server: Design and performance. Computer, 28(11):68– 74, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:PRQ

[954] Anonymous. Product reviews: QNX operating system. Computer, 28(11): 75–??, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:APIc

[955] Anonymous. Advertiser/product index and reader service card. Computer, 28(11):80–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlson:1995:UJJ

[956] Bob Carlson. Update: A jolt of Java could shake up the computing community. Computer, 28(11):81– 82, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berleant:1995:RRO

[957] D. Berleant and B. Liu. Internet kiosk: Robert’s rules of order for email meetings. Computer, 28(11):84– 85, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CLH

[958] Anonymous. Computer law: Hazards in Internet software exchange. Computer, 28(11):86–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1995:SCB

[959] C. Jones. Software challenges: Backfiring: converting lines of code to function points. Computer, 28(11):87– 88, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:1995:SLC

[960] R. Singh. The Software Life Cycle Processes standard. Computer, 28(11): 89–90, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAFf

[961] Anonymous. Technical activities forum: Computer-generated music. Computer, 28(11):91–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CPCd

[962] Anonymous. Call for papers/calendar. Computer, 28(11):93–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:COc

[963] Anonymous. Career opportunities. Computer, 28(11):100–??, November 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCCa

[964] Anonymous. Open channel: Candidates for change? Computer, 28(11): 112–??, November 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:EMR

[965] Anonymous. Editor-in-chief’s message: The real world of computing. Computer, 28(12):5–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:IMAd

[966] Anonymous. IEEE-CS membership application. Computer, 28(12):6–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:BCN

[967] Anonymous. Binary critic: The nethead gang. Computer, 28(12):8– ??, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ITD

[968] Anonymous. Industry trends: The dawning of the single-chip age. Computer, 28(12):10–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ASe

[969] Anonymous. Article summaries. Computer, 28(12):14–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boloix:1995:SSE

[970] Germinal Boloix and Pierre N. Robillard. A software system evaluation framework. Computer, 28(12): 17–26, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

El-Rewini:1995:TSM

[971] Hesham El-Rewini, Hesham H. Ali, and Ted Lewis. Task scheduling in multiprocessing systems. Computer, 28(12): 27–37, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKinley:1995:CCW

[972] Philip K. McKinley, Yih jia Tsai, and David F. Robinson. Collective communication in wormhole-routed massively parallel computers. Computer, 28(12): 39–50, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nevison:1995:PCU

[973] Christopher H. Nevison. Parallel computing in the undergraduate curriculum. Computer, 28(12):51–56, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Diep:1995:VVM

[974] Trung A. Diep and John Paul Shen. VMW: A visualization-based microarchitecture workbench. Computer, 28 (12):57–64, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:IKC

[975] Anonymous. Internet kiosk: Computercontrolled devices reach the Internet. Computer, 28(12):66–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ekmecic:1995:ETH

[976] Ilija Ekmecić, Igor Tartalja, and Veljko Milutinović. EM^{3}: a taxonomy of heterogeneous computing systems. Computer, 28(12):68–70, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ali:1995:PRB

[977] Mehbub Ali. Product reviews: A better way to state it. Computer, 28(12): 72–74, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Masgras:1995:PRT

[978] Cristian P. Masgras. Product reviews: TimingDesigner Version 3.0, Professional Edition. Computer, 28(12):74– 76, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:APId

[979] Anonymous. Advertiser/product index and reader service card. Computer, 28(12):80–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:ER

[980] Anonymous. Election results. Computer, 28(12):81–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:USC

[981] Anonymous. Update: Supercomputer centers. Computer, 28(12):81– ??, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1995:CLId

[982] I. H. Donner. Computer law: The Info Age law firm. Computer, 28(12): 86–87, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:SCH

[983] Anonymous. Software challenges: How software personnel learn new skills. Computer, 28(12):88–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:P

[984] Anonymous. PAMITC. Computer, 28 (12):90–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:TAFg

[985] Anonymous. Technical activities forum: Fault-tolerant computing TC. Computer, 28(12):90–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wright:1995:SAC

[986] T. Wright. Standards: A case for symmetrical bandwidth. Computer, 28(12): 93–94, December 1995. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:CPCe

[987] Anonymous. Call for papers/calendar. Computer, 28(12):95–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:COd

[988] Anonymous. Career opportunities. Computer, 28(12):106–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:AI

[989] Anonymous. 1995 annual index. Computer, 28(12):124–135, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OCCb

[990] Anonymous. Open channel: Computer science or simply ‘computics’? Computer, 28(12):136–??, December 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PMI

[991] Anonymous. President’s message: On the importance of turning 50. Computer, 29(1):6–??, January 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LBY

[992] Anonymous. Looking back: Yearlong celebration begins. Computer, 29(1):

7–??, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:BCC

[993] Ted Lewis. Binary critic: Computer business or monopoly. Computer, 29 (1):10–13, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITE

[994] Anonymous. Industry trends: Enterprise management on NT. Computer, 29(1):14–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASa

[995] Anonymous. Article summaries. Computer, 29(1):17–??, January 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rhodes:1996:GEI

[996] Michael L. Rhodes and Douglas D. Robertson. Guest editors’ introduction: Computers in surgery and therapeutic procedures. Computer, 29(1): 20–23, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hohne:1996:VBM

[997] K. H. Höhne, Bernhard Pflesser, Andreas Pommert, Martin Riemer, Thomas Schiemann, Rainer Schubert, and Ulf Tiede. A ‘virtual body’ model for surgical education and rehearsal. Computer, 29(1):25–31, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Christensen:1996:INA

[998] Gary E. Christensen, Michael I. Miller, Michael W. Vannier, and Ulf Grenander. Individualizing neuro-anatomical atlases using a massively parallel computer. Computer, 29(1):32–38, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robb:1996:CSP

[999] Richard A. Robb, Dennis P. Hanson, and Jon J. Camp. Computeraided surgery planning and rehearsal at Mayo Clinic. Computer, 29(1): 39–47, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adams:1996:ONB

[1000] Ludwig Adams, Achim Knepper, Dietrich Meyer-Ebrecht, Rainer Ruger, and Willem van der Brug. An optical navigator for brain surgery. Computer, 29 (1):48–54, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Solaiyappan:1996:IVR

[1001] Meiyappan Solaiyappan, Tim Poston, Pheng Ann Heng, Elliot R. McVeigh, Michael A. Guttman, and Elias A. Zerhouni. Interactive visualization for rapid noninvasive cardiac assessment. Computer, 29(1):55–62, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fishman:1996:SPL

[1002] Elliot K. Fishman, Brian S. Kuszyk, David G. Heath, Luomin Gao, and Brian Cabral. Surgical planning for liver resection. Computer, 29(1):64– 72, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IMAa

[1003] Anonymous. IEEE-CS membership application. Computer, 29(1):72–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hansen:1996:SSD

[1004] Gregory A. Hansen. Simulating software development processes. Computer, 29(1):73–77, January 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1996:GBP

[1005] Alan H. Karp, Michael Heath, and Al Geist. 1995 Gordon Bell Prize winners. Computer, 29(1):79–85, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:1996:OTC

[1006] Bertrand Meyer. Object technology: The conceptual perspective. Computer, 29(1):86–88, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UIF

[1007] Anonymous. Update: IEEE fellows; HPCC. Computer, 29(1):88–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koester:1996:HTM

[1008] David A. Koester, Karen W. Markus, and Mark D. Walters. Hot topics:

MEMS: small machines for the microelectronics age. Computer, 29(1): 93–94, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:SCE

[1009] Capers Jones. Software challenges:

The economics of process improvement. Computer, 29(1):95–97, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PS

[1010] Anonymous. Product source. Computer, 29(1):99–??, January 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRI

[1011] Anonymous. Product reviews: IMSL C numerical libraries; DT3155 frame grabber. Computer, 29(1):100–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:APIa

[1012] Anonymous. Advertiser/product index and reader service card. Computer, 29(1):104–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isaak:1996:SSI

[1013] J. Isaak. Standards: Standards for the information infrastructure. Computer, 29(1):105–106, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ciolek:1996:IKT

[1014] T. M. Ciolek. Internet kiosk: Today’s WWW — tomorrow’s MMM? the specter of multi-media mediocrity. Computer, 29(1):106–108, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFa

[1015] Anonymous. Technical activities forum: Simulation TC. Computer, 29(1): 109–??, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Galkin:1996:NNE

[1016] W. Galkin. Names on the Net, e-mail in court, and trade secrets for hire. Computer, 29(1):110–111, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CLE

[1017] Anonymous. Computer law: Electronic computer law report. Computer, 29(1): 110–??, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:COa

[1018] Anonymous. Career opportunities. Computer, 29(1):112–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCa

[1019] Anonymous. Call for papers/calendar. Computer, 29(1):129–??, January 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCW

[1020] Anonymous. Open channel: Wheelchairs for the brain. Computer, 29(1): 136–??, January 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LFC

[1021] Anonymous. looking.back: February in computing history. Computer, 29 (2):6–??, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:BCNa

[1022] Anonymous. Binary critic: The network computer phenomena. Computer, 29(2):8–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:GE

[1023] Anonymous. Gigabit Ethernet. Computer, 29(2):10–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITIa

[1024] Anonymous. Industry trends: Immigration issue. Computer, 29(2): 10–??, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASb

[1025] Anonymous. Article summaries. Computer, 29(2):14–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amza:1996:TSM

[1026] Cristiana Amza, Alan L. Cox, Sandhya Dwarkadas, Pete Keleher, Honghui Lu, Ramakrishnan Rajamony, Weimin Yu, and Willy Zwaenepoel. Tread-Marks: shared memory computing on networks of workstations. Computer, 29(2):18–28, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nielsen:1996:ISU

[1027] Jakob Nielsen and Jan Maurits Faber. Improving system usability through parallel design. Computer, 29(2):29– 35, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sandhu:1996:RAC

[1028] Ravi S. Sandhu, Edward J. Coyne, Hal L. Feinstein, and Charles E. Youman. Role-based access control models. Computer, 29(2):38–47, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raynal:1996:LTC

[1029] Michel Raynal and Mukesh Singhal. Logical time: Capturing causality in distributed systems. Computer, 29(2): 49–56, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parnas:1996:WSJ

[1030] David Lorge Parnas. Why software jewels are rare. Computer, 29(2): 57–60, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Desfray:1996:AOD

[1031] Philippe Desfray. Automated object design: The client-server case. Computer, 29(2):62–66, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CES

[1032] Anonymous. CS election schedules and awards. Computer, 29(2):68– ??, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:MZ

[1033] Anonymous. Mills, Zuse die. Computer, 29(2):68–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UIN

[1034] Anonymous. Update: Indecency on the Net. Computer, 29(2):68–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFb

[1035] Anonymous. Technical activities forum: TC on operating systems and application environments. Computer, 29 (2):74–??, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:1996:OTRa

[1036] Bertrand Meyer. Object technology:

The reusability challenge. Computer, 29(2):76–78, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pakstas:1996:SNN

[1037] Algirdas Pakstas and Sonata Pakstiene. Standards: NSK: A Norwegian cryptochip for supersafe communications. Computer, 29(2):78– 79, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:SCSa

[1038] Capers Jones. Software challenges: Software change management. Computer, 29(2):80–82, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRO

[1039] Anonymous. Product reviews: Object-Systems 1.1.0; Boca Office Communicator. Computer, 29(2):85–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IKCa

[1040] Anonymous. Internet kiosk: Community nets. Computer, 29(2):92– ??, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CLW

[1041] Anonymous. Computer law: For whom the bomb tolls. Computer, 29(2): 94–??, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCb

[1042] Anonymous. Call for papers/calendar. Computer, 29(2):97–??, February 1996.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCCa

[1043] Anonymous. Open channel: A consultant’s Ten Commandments. Computer, 29(2):120–??, February 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LMC

[1044] Anonymous. looking.back: March in computing history. Computer, 29 (3):6–??, March 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:EMa

[1045] Anonymous. Editor-in-chief’s message. Computer, 29(3):8–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:BCB

[1046] T. Lewis. Binary critic: The big software chill. Computer, 29(3):12–14, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITV

[1047] Anonymous. Industry trends: Virtual reality and the Internet; PC cluster specification. Computer, 29(3): 16–??, March 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASc

[1048] Anonymous. Article summaries. Computer, 29(3):20–??, March 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IMAb

[1049] Anonymous. IEEE-CS membership application. Computer, 29(3):22–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pal:1996:GEI

[1050] Sankar K. Pal and Pradip K. Srimani. Guest editors’ introduction: Neurocomputing: Motivation, models, and hybridization. Computer, 29(3): 24–28, March 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:1996:ANN

[1051] Anil K. Jain, Jianchang Mao, and K. M. Mohiuddin. Artificial neural networks: A tutorial. Computer, 29 (3):31–44, March 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shang:1996:GON

[1052] Yi Shang and Benjamin W. Wah. Global optimization for neural network training. Computer, 29(3):45–54, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serbedzija:1996:SAN

[1053] Nikola B. Serbedzija. Simulating artificial neural networks on parallel architectures. Computer, 29(3):56–63, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tan:1996:ANN

[1054] Chew Lim Tan, Tong Seng Quah, and Hoon Heng Teh. An artificial neural network that models human decision making. Computer, 29(3):64–70, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Setiono:1996:SRN

[1055] Rudy Setiono and Huan Liu. Symbolic representation of neural networks. Computer, 29(3):71–77, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wawrzynek:1996:SVM

[1056] John Wawrzynek, Krste Asanovic, Brian Kingsbury, David Johnson, James Beck, and Nelson Morgan. Spert-II: A vector microprocessor system. Computer, 29(3):79–86, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Do:1996:VIT

[1057] James Do, Phan Dinh Dieu, and Seymour E. Goodman. Vietnam: Information technology for the transition. Computer, 29(3):88–94, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

English:1996:NSU

[1058] Erin English and Scott Hamilton. Network security under siege: The timing attack. Computer, 29(3):95–97, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:USN

[1059] Anonymous. Update: Ship named for Grace Hopper; microprocessor history.

Computer, 29(3):99–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:SSR

[1060] Anonymous. Standards: Standards’ role in technology transition. Computer, 29(3):102–??, March 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:APIb

[1061] Anonymous. Advertiser/product index and reader service card. Computer, 29 (3):104–??, March 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRD

[1062] Anonymous. Product reviews: DSPCard-C31; Norton Utilities for Windows 95. Computer, 29(3):105–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKim:1996:OTP

[1063] James C. McKim, Jr. Object technology: Programming by contract. Computer, 29(3):109–111, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nicol:1996:IKC

[1064] D. M. Nicol. Internet kiosk: Conference program management using the Internet. Computer, 29(3):112–113, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowyer:1996:CLP

[1065] K. W. Bowyer and I. H. Donner. Computer law: Practicing “safe” engineering. Computer, 29(3):114–115, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:SCP

[1066] Capers Jones. Software challenges:

By popular demand: Software estimating rules of thumb. Computer, 29(3): 116–118, March 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFc

[1067] Anonymous. Technical activities forum: Security and privacy TC. Computer, 29(3):118–??, March 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:COb

[1068] Anonymous. Career opportunities.

Computer, 29(3):121–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCc

[1069] Anonymous. Call for papers/calendar.

Computer, 29(3):126–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCF

[1070] Anonymous. Open channel: A few (proposed) fundamental laws of programming. Computer, 29(3):136–??, March 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LBA

[1071] Anonymous. Looking back: April in computing history. Computer, 29 (4):6–??, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:BCL

[1072] Ted Lewis. Binary critic: The limits of innovation. Computer, 29(4):7–9, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

English:1996:ITU

[1073] Erin English. Industry trends: Users don’t want to get burned in redhot firewall market. Computer, 29 (4):10–12, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saiedian:1996:IFM

[1074] Hossein Saiedian. An invitation to formal methods. Computer, 29(4): 16–17, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:1996:FMP

[1075] Michael G. Hinchey and Jonathan P. Bowen. Formal methods: Pointcounterpoint: To formalize or not to formalize? Computer, 29(4): 18–19, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Glass:1996:FMP

[1076] Robert L. Glass. Formal methods: Point-counterpoint: Formal methods are a surrogate for a more serios software concern. Computer, 29(4):19, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:FML

[1077] Cliff B. Jones. Formal methods light:

A rigorous approach to formal methods. Computer, 29(4):20–21, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jackson:1996:FML

[1078] Daniel Jackson and Jeannette Wing.

Formal methods light: Lightweight formal methods. Computer, 29(4): 21–22, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hall:1996:IPW

[1079] Anthony Hall. Industrial practice:

What is the formal methods debate anyway? Computer, 29(4):22–23, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dill:1996:IPA

[1080] David L. Dill and John Rushby. Industrial practice: Acceptance of formal methods: Lessons from hardware design. Computer, 29(4):23–24, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holloway:1996:IPI

[1081] C. Michael Holloway and Ricky W.

Butler. Industrial practice: Impediments to industrial use of formal methods. Computer, 29(4):25–26, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zave:1996:IPF

[1082] Pamela Zave. Industrial practice: Formal methods are research, not development. Computer, 29(4):26–27, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:1996:EMC

[1083] Michael J. Lutz. Engineering mathematics: Consumable mathematics for software engineers. Computer, 29(4): 27–28, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parnas:1996:EMM

[1084] David L. Parnas. Engineering mathematics: Mathematical methods: What we need and don’t need. Computer, 29(4):28–29, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gries:1996:ENE

[1085] David Gries. Education: The need for education in useful formal logic. Computer, 29(4):29–30, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ball:1996:SVL

[1086] Thomas Ball and Stephen G. Eick. Software visualization in the large. Computer, 29(4):33–43, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vemuri:1996:PMU

[1087] Ranga Vemuri, Ram Mandayam, and Vijay Meduri. Performance modeling using PDL. Computer, 29(4): 44–53, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grupe:1996:CEI

[1088] Fritz H. Grupe. Computoons: The evolving image of computers in cartoons. Computer, 29(4):55–62, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:NYPa

[1089] Ted Lewis. The next 10, 000_{2} years:

Part I. Computer, 29(4):64–70, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1996:RIS

[1090] Ware Myers. On the road to the information superhighway. Computer, 29 (4):71–??, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UUS

[1091] Anonymous. Update: Universal service in the information age. Computer, 29(4):74–??, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRS

[1092] Anonymous. Product reviews: System commander and partition magic. Computer, 29(4):84–??, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sellers:1996:OTO

[1093] Brian Henderson-Sellers and Ian Graham. Object technology: OPEN: toward method convergence? Computer, 29(4):86–89, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CLCa

[1094] Anonymous. Computer law: Copyrighted materials on bulletin boards. Computer, 29(4):90–??, April 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neumann:1996:IKE

[1095] C. Neumann. Internet kiosk: Enabling commerce on the Net. Computer, 29 (4):91–92, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:SCSb

[1096] C. Jones. Software challenges: Software defect-removal efficiency. Computer, 29 (4):94–95, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:SNS

[1097] Anonymous.  Standards:  New

standards-development model. Computer, 29(4):96–??, April 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFd

[1098] Anonymous. Technical activities forum: Youth Forum in Computer Science and Engineering. Computer, 29 (4):97–??, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCN

[1099] Anonymous. Open channel: New solutions for new technologies. Computer, 29(4):112–??, April 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LBM

[1100] Anonymous. Looking.back: May in computing history. Computer, 29 (5):6–??, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:BCI

[1101] Anonymous. Binary critic: Is it too late for Apple Computer? Computer, 29 (5):8, 10–12, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITL

[1102] Anonymous. Industry trends: Looking into the PowerPC’s future. Computer, 29(5):14, 16–17, May 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASd

[1103] Anonymous. Article summaries. Computer, 29(5):18–??, May 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schatz:1996:GEI

[1104] Bruce Schatz and Hsinchun Chen.

Guest editors’ introduction: Building large-scale digital libraries. Computer, 29(5):22–26, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schatz:1996:FDC

[1105] Bruce Schatz, William H. Mischo, Timothy W. Cole, Joseph B. Hardin, Ann P. Bishop, and Hsinchun Chen. Federating diverse collections of scientific literature. Computer, 29(5):

28–36, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilensky:1996:TWD

[1106] Robert Wilensky. Toward workcentered digital information services. Computer, 29(5):37–44, May 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wactlar:1996:IAD

[1107] Howard D. Wactlar, Takeo Kanade, Michael A. Smith, and Scott M. Stevens. Intelligent access to digital video: Informedia project. Computer, 29(5):46–52, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1996:DLG

[1108] Terence R. Smith, Dan Andresen, Larry Carver, Ron Dolin, Christoph Frisher, James Frew, Michael Goodchild, Oscar Ibarra, Randall B. Kemp, Ravi Kothuri, Mary Larsgaard, Bangalore S. Manjunath, Doug Nebert, Jason Simpson, Alex Wells, Tao Yang, and Qi Zheng. A digital library for geographically referenced materials. Computer, 29(5):54–60, May 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). All but first author were missing from the original article; see erratum in July 1996 issue, p. 14.

Paepcke:1996:UDO

[1109] Andreas Paepcke, Steve B. Cousins, Hector Garcia-Molina, Scott W. Hassan, Steven P. Ketchpel, Martin Roscheisen, and Terry Winograd. Using distributed objects for digital library interoperability. Computer, 29 (5):61–68, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkins:1996:TIE

[1110] Daniel E. Atkins, William P. Birmingham, Edmund H. Durfee, Eric J. Glover, Tracy Mullen, Elke A. Rundensteiner, Elliot Soloway, Jose M. Vidal, Raven Wallace, and Michael P. Wellman. Toward inquiry-based education through interacting software agents. Computer, 29(5):69–76, May 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:NYPb

[1111] Ted Lewis. The next 10, 000_{2} years:

Part II. Computer, 29(5):78–86, May 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keaton:1996:EBB

[1112] John Keaton and Scott Hamilton. Employment 2005: Boom or bust for computer professionals? Computer, 29 (5):87–98, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:Ua

[1113] Anonymous. Update. Computer, 29 (5):99–??, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:SCA

[1114] Capers Jones. Software challenges: Activity based software costing. Computer, 29(5):103–104, May 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:1996:MFI

[1115] Bertrand Meyer. Object technology: The many faces of inheritance: a taxonomy of taxonomy. Computer, 29 (5):105–108, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tepper:1996:CLIa

[1116] Steven G. Tepper and Irah H. Donner. Computer law: IPOs, part I: preparation is the key. Computer, 29 (5):108–110, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferrens:1996:TAF

[1117] Matthew Ferrens. Technical activities forum: Microprogramming and microarchitecture. Computer, 29(5): 111–??, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krile:1996:IKP

[1118] Terry Krile, Paul Juell, and Ron Vetter. Internet kiosk: Pioneering on the new frontiers of education. Computer, 29(5):112–114, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRa

[1119] Anonymous. Product reviews. Computer, 29(5):117–??, May 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCd

[1120] Anonymous. Call for papers/calendar. Computer, 29(5):128–??, May 1996.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCb

[1121] Anonymous. Open channel. Computer, 29(5):136–??, May 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LB

[1122] Anonymous. Looking back. Computer, 29(6):6–??, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roberts:1996:WWJ

[1123] Ken Roberts. What’s wrong with Java? (discussion of the February 1996 article, the NC phenomena: scenes from your living room). Computer, 29(6):8–9, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:BC

[1124] Anonymous. Binary critic. Computer, 29(6):12–??, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITa

[1125] Anonymous. Industry trends. Computer, 29(6):18–??, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASe

[1126] Anonymous. Article summaries. Computer, 29(6):26–??, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gentner:1996:DMC

[1127] Donald R. Gentner and Jonathan Grudin. Design models for computerhuman interfaces. Computer, 29(6): 28–35, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ackerman:1996:SAI

[1128] Mark S. Ackerman and Brian Starr. Social activity indicators for groupware. Computer, 29(6):37–42, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bockle:1996:SEC

[1129] Günter Böckle, Hermann Hellwagner, Roland Lepold, Gerd Sandweg, Burghardt Schallenberger, Raimar Thudt, and Stefan Wallstab. Structured evaluation of computer systems. Computer, 29(6):45–51, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brooks:1996:RDC

[1130] Richard R. Brooks and S. Sitharama Iyengar. Robust distributed computing and sensing algorithm. Computer, 29(6):53–60, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rader:1996:CPM

[1131] Gary M. Rader. Creating printed music automatically. Computer, 29 (6):61–68, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See erratum in July 1996 issue, p. 14: Figures labelled 2, 3, and 4 should have been 3, 4, and 2.

Thakur:1996:POI

[1132] Rajeev Thakur, Alok Choudhary, Rajesh Bordawekar, Sachin More, and Sivaramakrishna Kuditipudi. Passion: optimized I/O for parallel applications. Computer, 29(6):70–78, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See erratum in July 1996 issue, p. 14, with corrected captions for Figure 9 and Table 2.

Douglas:1996:CPR

[1133] Patricia J. Douglas, George M. Alliger, and Robert Goldberg. Computing practices: Refining the curriculum: Client-server and object-oriented training. Computer, 29(6):80–84, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:1996:UCB

[1134] R. Jain. Update: From content-based retrieval to invisible telepresence. Computer, 29(6):85–86, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isaak:1996:CSP

[1135] Jim Isaak. A call for standards for portable Web applications. Computer, 29(6):89–90, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Recommends standardization of Java, JavScript, VB-Script, and Java’s byte codes.

Tepper:1996:CLIb

[1136] S. G. Tepper and I. H. Donner. Computer law: IPOs. II. getting the corporate house in order. Computer, 29 (6):90–91, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1996:SCSc

[1137] C. Jones. Software challenges: Strategies for managing requirements creep. Computer, 29(6):92–94, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OT

[1138] Anonymous. Object technology. Computer, 29(6):95–??, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFe

[1139] Anonymous. Technical activities forum. Computer, 29(6):99–??, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garfinkel:1996:IKP

[1140] S. L. Garfinkel. Internet kiosk: Public key cryptography. Computer, 29 (6):101–104, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:NPU

[1141] Anonymous. New products: UML and Java go Platinum. Computer, 29(6): 104–106, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRb

[1142] Anonymous. Product reviews. Computer, 29(6):107–??, June 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCe

[1143] Anonymous. Call for papers/calendar.

Computer, 29(6):113–??, June 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCc

[1144] Anonymous. Open channel. Computer, 29(6):120–??, June 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LBJ

[1145] Anonymous. Looking back: July in computing history. Computer, 29 (7):6–??, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAR

[1146] Anonymous. Treasurer’s and Auditor’s report. Computer, 29(7):8–??, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:BCA

[1147] T. Lewis. Binary critic: Alice in Wired World: wonderland or wasteland? Computer, 29(7):17–19, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaneshige:1996:ITC

[1148] Thomas Kaneshige. Industry trends:

Cyberwar; Clipper chip; network computers. Computer, 29(7):20–23, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Contains news brief: Lots of Java brewing at conference.

Anonymous:1996:ASf

[1149] Anonymous. Article summaries. Computer, 29(7):24–??, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Manaris:1996:INL

[1150] Bill Z. Manaris and Brian M. Slator. Interactive natural language processing: Building on success. Computer, 29(7):28–32, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:1996:SSF

[1151] Paul Martin, Frederick Crabbe, Stuart Adams, Eric Baatz, and Nicole Yankelovich. SpeechActs: A spokenlanguage framework. Computer, 29 (7):33–40, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Waibel:1996:ITC

[1152] Alex Waibel. Interactive translation of conversational speech. Computer, 29(7):41–48, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paris:1996:IST

[1153] Cecile Paris and Keith Vander Linden. An interactive support tool for writing multilingual manuals. Computer, 29(7):49–56, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sheremetyeva:1996:KEA

[1154] Svetlana Sheremetyeva and Sergei Nirenburg. Knowledge elicitation for authoring patent claims. Computer, 29(7):57–63, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wermter:1996:ISP

[1155] Stefan Wermter and Volker Weber. Interactive spoken-language processing in a hybrid connectionist system. Computer, 29(7):65–74, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moore:1996:DGF

[1156] Johanna D. Moore and Vibhu O. Mittal. Dynamically generated follow-up questions. Computer, 29(7):75–86, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1996:TAS

[1157] Ware Myers. On trial at the Summer Olympic Games: Smart cards. Computer, 29(7):88–91, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UCS

[1158] Anonymous. Update: Computer society elections. Computer, 29(7):92–??, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geraci:1996:SCD

[1159] A. Geraci. Standards: The Computer Dictionary Project: an update. Computer, 29(7):95, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:1996:OTRb

[1160] B. Meyer. Object technology: Reality: a cousin twice removed. Computer, 29(7):96–97, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1996:CLW

[1161] I. H. Donner. Computer law: Will your IP protection provide the protection you expect. Computer, 29(7):98, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Onvural:1996:TAF

[1162] R. A. Onvural and V. Srinivasan. Technical activities forum: Computer communications. Computer, 29(7): 99–100, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:1996:IKP

[1163] H. Berghel and L. O’Gorman. Internet kiosk: Protecting ownership rights through digital watermarking. Computer, 29(7):101–103, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krishna:1996:PRC

[1164] Dileep Krishna. Product reviews: Café: the right place for Java. Computer, 29 (7):104–108, July 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCf

[1165] Anonymous. Call for papers/calendar. Computer, 29(7):113–??, July 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCCb

[1166] Anonymous. Open channel: Are comparative analyses worthwhile? Computer, 29(7):120–??, July 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LAC

[1167] Anonymous. looking.back: August in computing history. Computer, 29 (8):6–??, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:EMb

[1168] Anonymous. Editor’s message. Computer, 29(8):8–??, August 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cleaveland:1996:LEW

[1169] Craig Cleaveland and Ted Lewis. Letters to the editor: What’s wrong with the binary critic? Computer, 29(8): 10–11, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Debate on the merits of the Java programming language.

Lewis:1996:BCS

[1170] Ted Lewis. Binary critic: Software architectures: Divine plan of digital Darwinism? Computer, 29(8): 13–15, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASg

[1171] Anonymous. Article summaries. Computer, 29(8):22–??, August 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yourdon:1996:JWS

[1172] Edward Yourdon. Java, the Web, and software development. Computer, 29 (8):25–30, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1996:JSN

[1173] Marc A. Hamilton. Java and the shift to net-centric computing. Computer, 29(8):31–39, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kouzes:1996:CDS

[1174] Richard T. Kouzes, James D. Myers, and William A. Wulf. Collaboratories: Doing science on the Internet. Computer, 29(8):40–46, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkins:1996:PSP

[1175] Mark Atkins and Ramesh Subramaniam. PC software performance tuning. Computer, 29(8):47–54, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bic:1996:DCU

[1176] Lubomir F. Bic, Munehiro Fukuda, and Michael B. Dillencourt. Distributed computing using autonomous objects. Computer, 29(8):55–61, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:1996:TIM

[1177] Hsinchun Chen, Andrea Houston, Jay Nunamaker, and Jerome Yen. Toward intelligent meeting agents. Computer, 29(8):62–69, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balaguer:1996:UIG

[1178] Jean-Francis Balaguer and Enrico Gobbetti. 3D user interfaces for generalpurpose 3D animation. Computer, 29 (8):71–78, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UWI

[1179] Anonymous. Update: WW II Colossus computer is restored to operation. Computer, 29(8):79, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UCP

[1180] Anonymous. Update: Cash prize offered for major advances in molecular nanotechnology. Computer, 29(8): 79–80, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Donner:1996:CLD

[1181] Irah H. Donner. Computer law: Doing business on the Net. Computer, 29 (8):85–86, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spring:1996:SRM

[1182] Michael B. Spring. Standards: Reference model for data interchange standards. Computer, 29(8):87–88, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adam:1996:TAF

[1183] Nabil R. Adam, Richard Holowczak, Milton Halem, Nand Lal, and Yelena Yesha. Technical activities forum: Digital Library Task Force. Computer, 29 (8):89–91, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hitchcock:1996:IKW

[1184] Steve Hitchcock. Internet kiosk: Web publishing: Speed changes everything. Computer, 29(8):91–93, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pappas:1996:PRM

[1185] T. L. Pappas. Product reviews: Microsoft Visual C++ goes interactive and remote. Computer, 29(8): 94–98, August 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCg

[1186] Anonymous. Call for papers/calendar. Computer, 29(8):102–??, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCH

[1187] Anonymous. Open channel: Heresy: A poem. Computer, 29(8):120–??, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITP

[1188] Anonymous. Industry trends: Palmtop wireless communications; TI superchip. Computer, 29(8):163–??, August 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:L

[1189] Anonymous. Looking.back. Computer, 29(9):6–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1996:BCW

[1190] Ted Lewis. Binary critic: Will tiny beans conquer the world again? Computer, 29(9):13–14, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses Java and its component technology, Java Beans.

Anonymous:1996:ITb

[1191] Anonymous. Industry trends. Computer, 29(9):16–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASh

[1192] Anonymous. Article summaries. Computer, 29(9):22–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IMAc

[1193] Anonymous. IEEE-CS membership application. Computer, 29(9):24–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fayad:1996:GEI

[1194] Mohamed E. Fayad and Marshall Cline. Guest Editors’ introduction: Managing object-oriented software development. Computer, 29(9):26–31, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sadr:1996:OPM

[1195] Babak Sadr and Patricia J. Dousette.

An OO project management strategy. Computer, 29(9):33–38, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:1996:MIO

[1196] John D. Williams. Managing iteration in OO projects. Computer, 29(9): 39–43, September 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moser:1996:EOO

[1197] Simon Moser and Oscar Nierstrasz. The effect of object-oriented frameworks on developer productivity. Computer, 29(9):45–51, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sparks:1996:MOO

[1198] Steve Sparks, Kevin Benner, and Chris Faris. Managing object-oriented framework reuse. Computer, 29(9):52– 61, September 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Corriveau:1996:TPL

[1199] Jean-Pierre Corriveau. Traceability process for large OO projects. Computer, 29(9):63–68, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shaw:1996:MUN

[1200] Kevin Shaw, Maria Cobb, Miyi Chung, and David Arctur. Managing the US Navy’s first OO digital mapping project. Computer, 29(9):69– 74, September 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ICE

[1201] Anonymous. IEEE and CS elections. Computer, 29(9):75–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:Ud

[1202] Anonymous. Update. Computer, 29(9):

84–??, September 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CL

[1203] Anonymous. Computer law. Computer, 29(9):85–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:S

[1204] Anonymous. Standards. Computer, 29 (9):86–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFf

[1205] Anonymous. Technical activities forum. Computer, 29(9):88–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reynolds:1996:IKE

[1206] Franklin D. Reynolds. Internet kiosk:

Evolving an operating system for the Web. Computer, 29(9):90–92, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OTR

[1207] Anonymous. Object technology: Reversibility in software engineering. Computer, 29(9):93–95, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRc

[1208] Anonymous. Product reviews. Computer, 29(9):96–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCh

[1209] Anonymous. Call for papers/calendar. Computer, 29(9):101–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:APIc

[1210] Anonymous. Advertiser/product index. Computer, 29(9):104–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:COc

[1211] Anonymous. Career opportunities. Computer, 29(9):107–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCd

[1212] Anonymous. Open channel. Computer, 29(9):112–??, September 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LBO

[1213] Anonymous. Looking back: October in computing history. Computer, 29 (10):6–??, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITIb

[1214] Anonymous. Industry trends: Internet telephony; Web server benchmark. Computer, 29(10):13–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASi

[1215] Anonymous. Article summaries. Computer, 29(10):20–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yau:1996:GEI

[1216] Stephen S. Yau. Guest Editor’s introduction: Essays on fifty years of computing. Computer, 29(10):24–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tredennick:1996:MBC

[1217] Nick Tredennick. Microprocessorbased computers. Computer, 29(10): 27–37, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gray:1996:EDM

[1218] Jim Gray. Evolution of data management. Computer, 29(10):38–46, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramamoorthy:1996:ASE

[1219] C. V. Ramamoorthy and Wei tek Tsai. Advances in software engineering. Computer, 29(10):47–58, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meleis:1996:TIN

[1220] Hanafy Meleis. Toward the information network. Computer, 29(10): 59–67, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berners-Lee:1996:WPP

[1221] Tim Berners-Lee. WWW: Past, present, and future. Computer, 29(10): 69–77, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robinson:1996:HIC

[1222] Gary S. Robinson and Carl Cargill. History and impact of computer standards. Computer, 29(10):79–85, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reddy:1996:CAI

[1223] Raj Reddy. The challenge of artificial intelligence. Computer, 29(10): 86–96, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woodward:1996:PST

[1224] Paul R. Woodward. Perspective on supercomputing: Three decades of change. Computer, 29(10):99–111, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UFB

[1225] Anonymous. Update: Fletcher Buckley dies. Computer, 29(10):112–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CLS

[1226] Anonymous. Computer law: Software licensing: A new era. Computer, 29 (10):114–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:SFS

[1227] Anonymous. Standards: A fast-track standards process. Computer, 29(10): 115–??, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFg

[1228] Anonymous. Technical activities forum: Technical Council on Software Engineering. Computer, 29(10):117– ??, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OTS

[1229] Anonymous. Object technology:

Schema evolution. Computer, 29(10): 119–??, October 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IKCb

[1230] Anonymous. Internet kiosk: Choosing a connectivity model. Computer, 29(10):122–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCi

[1231] Anonymous. Call for papers/calendar.

Computer, 29(10):130–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCT

[1232] Anonymous. Open channel: Twentyfifth anniversary of microprocessing. Computer, 29(10):152–??, October 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LBN

[1233] Anonymous. Looking back: November in computing history. Computer, 29(11):6–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PM

[1234] Anonymous. President’s message. Computer, 29(11):8–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:BCNb

[1235] Anonymous. Binary critic: Netstorms: The crash of ’96. Computer, 29(11): 12–??, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITT

[1236] Anonymous. Industry trends: Telecommunications Act of 1996; panix attacks. Computer, 29(11):16–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ASj

[1237] Anonymous. Article summaries. Computer, 29(11):23–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kapur:1996:GEI

[1238] Rohit Kapur, Thomas W. Williams, and Edward F. Miller. Guest Editors’ introduction: System test and reliability techniques for avoiding failure. Computer, 29(11):28–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murray:1996:TIG

[1239] Brian T. Murray and John P. Hayes.

Testing ICs: Getting to the core of the problem. Computer, 29(11):32– 38, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Konemann:1996:BSA

[1240] Bernd Könemann, Ben Bennetts, Najmi Jarwala, and Benoit Nadeau-Dostie. Built-in self-test: assuring system integrity. Computer, 29(11):39– 45, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marr:1996:MVP

[1241] Deborah T. Marr, Subramanian Natarajan, Shreekant Thakkar, and Richard Zucker. Multiprocessor validation of the Pentium Pro. Computer, 29(11):47–53, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Profeta:1996:SSB

[1242] Joseph A. Profeta III, Nikos P. Andrianos, Bing Yu, Barry W. Johnson, Todd A. DeLong, David Guaspari, and Damir Jamsek. Safety-critical systems built with COTS. Computer, 29(11):

54–60, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Musa:1996:ST

[1243] John D. Musa. Software-reliability-engineered testing. Computer, 29(11): 61–68, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wood:1996:PSR

[1244] Alan Wood. Predicting software reliability. Computer, 29(11):69–77, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drake:1996:MSQ

[1245] Thomas Drake. Measuring software quality: A case study. Computer, 29 (11):78–87, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:USC

[1246] Anonymous. Update: Seymour Cray legacy. Computer, 29(11):88– ??, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CLP

[1247] Anonymous. Computer law: Patents in Israel. Computer, 29(11):90–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:SSE

[1248] Anonymous. Standards: Secure Electronic Transaction standard. Computer, 29(11):92–??, November 1996.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IKEb

[1249] Anonymous. Internet kiosk: Education and Internet II. Computer, 29(11): 93–??, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFh

[1250] Anonymous. Technical activities forum: Computer Graphics Technical Committee. Computer, 29(11):95– ??, November 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OTM

[1251] Anonymous. Object technology:

Modal testing strategies. Computer, 29 (11):97–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:PRP

[1252] Anonymous. Product reviews: Personal Digital Assistants: Part 2. Computer, 29(11):100–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CPCj

[1253] Anonymous. Call for papers/calendar.

Computer, 29(11):106–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCA

[1254] Anonymous. Open channel: Attaining success is an NP-complete problem. Computer, 29(11):136–??, November 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:LDC

[1255] Anonymous. Looking-back — December in computing history. Computer, 29(12):8–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:ITS

[1256] Anonymous. Industry trends — in search of speed and bandwidth. Computer, 29(12):12–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Seaborn:1996:RYP

[1257] True Seaborn. Reflections on 23 years of publishing computer. Computer, 29 (12):24–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stenstrom:1996:GEI

[1258] Per Stenstrom and Frederik Dahlgren. Guest editors’ introduction: Applications for shared memory multiprocessors. Computer, 29(12):29–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kozlov:1996:PIP

[1259] Alexander V. Kozlov and Jaswinder Pal Singh. Parallel implementations of probabilistic inference. Computer, 29 (12):33–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shekhar:1996:PGS

[1260] Shashi Shekhar, Sivakumar Ravada, Vipin Kumar, Douglas Chubb, and Greg Turner. Parallelizing a GIS on a shared address space architecture. Computer, 29(12):42–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sterling:1996:IAP

[1261] Thomas Sterling, Phillip Merkey, and Daniel Savarese. Improving application performance on the HP/ convex exemplar. Computer, 29(12):50– ??, December 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chong:1996:APA

[1262] Frederic T. Chong, Beng-Hong Lim, Ricardo Bianchini, John Kubiatowicz, and Anant Agarwal. Application performance on the MIT alewife machine. Computer, 29(12):57–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adve:1996:SMC

[1263] Sarita V. Adve and Kourosh Gharachorloo. Shared memory consistency models: A tutorial. Computer, 29(12): 66–??, December 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blume:1996:PPP

[1264] William Blume, Ramon Doallo, Rudolf Eigenmann, John Grout, Jay Hoeflinger, Thomas Lawrence, Jaejin Lee, David Padua, Yunheung Paek, Bill Pottenger, Lawrence Rauchwerger, and Peng Tu. Parallel programming with polaris. Computer, 29(12):78–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hall:1996:MMP

[1265] Mary W. Hall, Jennifer M. Anderson, Saman p. Amarasinghe, Brian R. Murphy, Shih-Wei Liao, Edouard Bugnion, and Monica S. Lam. Maximizing multiprocessor performance with the SUIF compiler. Computer, 29(12): 84–??, December 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:UCE

[1266] Anonymous. Update — CS election; Gordon Bell prize winners. Computer, 29(12):102–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:CLCb

[1267] Anonymous. Computer law — copyright in shareware on the Internet. Computer, 29(12):110–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:SIN

[1268] Anonymous. Standards: IP nextgeneration. Computer, 29(12):112– ??, December 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:IKG

[1269] Anonymous. Internet kiosk — for a greener Internet. Computer, 29(12):

113–??, December 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:TAFi

[1270] Anonymous. Technical activities forum — technical activities strategic planning. Computer, 29(12):115–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OTT

[1271] Anonymous. Object technology — techniques for teaching object technology. Computer, 29(12):117–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1996:OCP

[1272] Anonymous. Open channel — parallel processors: The future again? Computer, 29(12):152–??, December 1996. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parrish:1997:EMW

[1273] Edward A. Parrish. Editor’s message:

Welcome to the new computer. Computer, 30(1):8, 10–12, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:RMN

[1274] Anonymous. Report to members: The next 50 years of computing. Computer, 30(1):16–21, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hedberg:1997:FIC

[1275] Sara Reese Hedberg. Up front: HP’s international cryptography framework: compromise or threat? Computer, 30 (1):28–30, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaughan-Nichols:1997:ITS

[1276] Steven J. Vaughan-Nichols. Industry trends: Switching to a faster Internet. Computer, 30(1):31–32, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:RVI

[1277] Anonymous. RBOCs vs. ISPs. Computer, 30(1):33–??, January 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:SE

[1278] Anonymous. Set-top exports. Computer, 30(1):33–??, January 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:UT

[1279] Anonymous. USB technology. Computer, 30(1):33–??, January 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Partridge:1997:CIP

[1280] Derek Partridge. The case for inductive programming. Computer, 30(1): 36–41, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Silverman:1997:CRA

[1281] Barry G. Silverman. Computer reminders and alerts. Computer, 30(1): 42–49, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blumenthal:1997:UCI

[1282] Marjory S. Blumenthal. Unpredictable certainty: The Internet and the information infrastructure. Computer, 30 (1):50–56, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkinson:1997:TMS

[1283] Randall J. Atkinson. Toward a more secure Internet. Computer, 30(1): 57–61, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bieber:1997:TSH

[1284] Michael Bieber and Fabio Vitali. Toward support for hypermedia on the World Wide Web. Computer, 30(1): 62–70, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

OLeary:1997:IIA

[1285] Daniel E. O’Leary. The Internet, intranets, and the AI renaissance. Computer, 30(1):71–78, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karp:1997:GBP

[1286] Alan H. Karp, Al Geist, and David Bailey. 1996 Gordon Bell Prize winners. Computer, 30(1):80–85, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:A

[1287] Anonymous. Awards. Computer, 30 (1):86, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:ECS

[1288] Anonymous. Engineering of computerbased systems TC. Computer, 30(1): 86–??, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:FN

[1289] Anonymous. Fellow nominations. Computer, 30(1):87, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:1997:STH

[1290] Charles Severance. Standards: Taking a high-tech approach to standards. Computer, 30(1):114, 116, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gopakumaran:1997:PRG

[1291] Bala Gopakumaran. Product reviews: Gate-array IC design tools. Computer, 30(1):117–119, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Whitaker:1997:MMK

[1292] Ken Whitaker. Management: Motivating and keeping software developers. Computer, 30(1):126–128, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jezequel:1997:OTD

[1293] Jean-Marc Jézéquel and Bertrand Meyer. Object technology: Design by contract: The lessons of Ariane. Computer, 30(1):129–130, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also letters [1366].

Varshney:1997:IWS

[1294] Upkar Varshney. Internet watch: Supporting mobility with wireless ATM. Computer, 30(1):131–133, January 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1997:BCY

[1295] Ted Lewis. Binary critic: The year I shoot my TV. Computer, 30(1):136, 134–135, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilkes:1997:CPD

[1296] M. V. Wilkes. Challenges for processor designers. Computer, 30(1): 10–??, January 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clark:1997:FBT

[1297] David Clark. Up front: Breaking the teraflops barrier. Computer, 30(2): 12–14, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goldberg:1997:ITI

[1298] Lee Goldberg. Industry trends: The Internet in space: Problems and solutions. Computer, 30(2):15–16, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1997:NBC

[1299] Lee Garber. News briefs: Crucial compromise launches digital TV. US encryption agreement in jeopardy. warning issued about UNIX flaw. WIPO discusses cyberspace copyrights. IT issues could threaten European Monetary Union. COBOL programmers in demand again. chip alliance formed. semiconductor film grown in space. survey reveals security fears and vulnerability. taxing the Internet. Computer, 30(2):18, 19, 22, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mann:1997:WCF

[1300] Steve Mann. Wearable computing: A first step toward personal imaging. Computer, 30(2):25–32, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pfleeger:1997:IIF

[1301] Shari Lawrence Pfleeger and Les Hatton. Investigating the influence of formal methods. Computer, 30(2): 33–43, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Comparetto:1997:TMS

[1302] Gary Comparetto and Rafols Ramirez. Trends in mobile satellite technology. Computer, 30(2):44–52, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wood:1997:GTJ

[1303] Kenneth R. Wood, Tristan Richardson, Frazer Bennett, Andy Harter, and Andy Hopper. Global teleporting with Java: Toward ubiquitous personalized computing. Computer, 30(2): 53–59, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pfitzmann:1997:TMU

[1304] Andreas Pfitzmann, Birgit Pfitzmann, and Michael Waidner. Trusting mobile user devices and security modules. Computer, 30(2):61–68, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Islam:1997:CSS

[1305] Nayeem Islam. Customizing system software using OO frameworks. Computer, 30(2):69–78, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gader:1997:NFM

[1306] Paul D. Gader, James M. Keller, Raghu Krishnapuram, Jung-Hsien Chiang, and Magdi Mohamed. Neural and fuzzy methods in handwriting recognition. Computer, 30(2):79–86, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:CUCa

[1307] Anonymous. CS update: CS members become IEEE fellows. Computer, 30(2): 88–91, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:ET

[1308] Anonymous. 1997 election timetable, call for nominess. Computer, 30(2): 89–90, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Subrahmanyam:1997:TAF

[1309] P. A. Subrahmanyam. Technical activities forum: Embedded networked systems for an interactive world. Computer, 30(2):92–93, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Severance:1997:SLC

[1310] Charles Severance. Standards: Linking computers and consumer electronics. Computer, 30(2):119, 121, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:SVJ

[1311] Anonymous. Software: Visual Java. Computer, 30(2):123, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:1997:NBI

[1312] Michael J. Lutz. New books: Interpreting our Internet dreams. Computer, 30 (2):125, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rozenblit:1997:IET

[1313] Jerzy W. Rozenblit and Sanjaya Kumar. Integrated engineering: Toward synergistic engineering of computer systems. Computer, 30(2):126– 127, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bach:1997:SRH

[1314] James Bach. Software realities: The hard road from method to practice. Computer, 30(2):129–130, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maltais:1997:IWC

[1315] Dwaine Maltais. Internet watch: Conducting a job search on the Internet. Computer, 30(2):131–133, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:1997:BCR

[1316] Sorel Reisman. Binary critic: Report from the Comdex dimension. Computer, 30(2):136, 134, 135, February 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Melford:1997:FID

[1317] Robert Melford. Up front: Intel defies predictions, sets earnings record. Computer, 30(3):12–14, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Korzeniowski:1997:ITL

[1318] Paul Korzeniowski. Industry trends — lowering the cost of PC maintenance. Computer, 30(3):15–16, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:NBE

[1319] Anonymous. News briefs — ETrust Web security: Apple at the crossroads:

New domain names. Computer, 30 (3):17–20, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

ParkerSmith:1997:CJC

[1320] Norris Parker Smith. Computing in Japan: From cocoon to competition. Computer, 30(3):26–33, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wegscheider:1997:TCB

[1321] Eric Wegscheider. Toward codefree business application development. Computer, 30(3):35–43, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:1997:DDS

[1322] Robert A. Martin. Dealing with dates: Solutions for the year 2000. Computer, 30(3):44–51, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Newcomb:1997:RAY

[1323] Philip H. Newcomb and Melvin Scott. Requirements for advanced year 2000 maintenance tools. Computer, 30(3): 52–57, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zvegintzov:1997:RGY

[1324] Nicholas Zvegintzov. A resource guide to Year 2000 tools. Computer, 30 (3):58–63, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:1997:CAP

[1325] W. Neville Holmes. Composite arithmetic: Proposal for a new standard. Computer, 30(3):65–73, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

ftp://ftp.comp.utas.edu.au/pub/
nholmes/ca/dsdf.ps;   ftp://ftp.
comp.utas.edu.au/pub/nholmes/ca/
dsrf.ps;   ftp://ftp.comp.utas.
edu.au/pub/nholmes/ca/dssf.ps.

Franz:1997:DLS

[1326] Michael Franz. Dynamic linking of software components. Computer, 30 (3):74–81, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dandamudi:1997:RRQ

[1327] Sivarama P. Dandamudi. Reducing run queue contention in shared memory multiprocessors. Computer, 30 (3):82–89, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoening:1997:SSK

[1328] James Schoening and Thomas Wheeler.

Standards: Standards — the key to education reform. Computer, 30(3): 116–117, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Curtis:1997:MDO

[1329] Bill Curtis, William E. Hefley, Sally Miller, and Michael Konrad. Management: Developing organizational competence. Computer, 30(3):122–124, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arranga:1997:OTC

[1330] Edmund C. Arranga and Frank P. Coyle. Object technology: Cobol: Perception and reality. Computer, 30(3): 126–128, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:1997:IWW

[1331] Ronald J. Vetter. Internet watch — Web publishing: No HTML expertise needed. Computer, 30(3): 129–131, March 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1997:BCI

[1332] Ted Lewis. Binary critic: If Java is the answer, what was the question? Computer, 30(3):136, 133–135, March 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See responses [1369–1372, 1403].

Zvegintzov:1997:LEC

[1333] Nicholas Zvegintzov, Shari Lawrence Pfleeger, and Les Hatton. Letter to the editor: Changes versus faults in formal methods. Computer, 30(4):6, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Breuel:1997:LEJ

[1334] Thomas M. Breuel and Ted Lewis. Letter to the editor: Java as universal language. Computer, 30(4):6–7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:1997:LED

[1335] Jihong Kim and Ted Lewis. Letter to the editor: Deadlocking embrace. Computer, 30(4):7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCormick:1997:LEN

[1336] John W. M. McCormick. Letter to the editor: No language for beginners. Computer, 30(4):7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Decries adoption of Java as a dominant language in beginning Computer Science courses.

Woytaszek:1997:LEQ

[1337] Tom Woytaszek and Ted Lewis. Letter to the editor: Questions Java answers. Computer, 30(4):7, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ozcan:1997:LER

[1338] Mehmet Bulent Ozcan and Ken Whitaker. Letter to the editor: Rewarding privately. Computer, 30(4):7– 8, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bast:1997:LER

[1339] John Bast. Letter to the editor: Rewards work, incentives don’t. Computer, 30(4):8, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:1997:CFF

[1340] J. A. N. Lee. Celebrating Fortran’s fortieth. Computer, 30(4):8, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:1997:OCH

[1341] Dirk Riehle. Open channel: The human side of architectural styles. Computer, 30(4):10, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1997:RMA

[1342] Scott Hamilton and Janet Wilson. Report to members: Andrew Grove’s vision for the Internet. Computer, 30 (4):14–15, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1997:FSS

[1343] Lee Garber. Up front: Students stumble onto Internet Explorer flaw. Computer, 30(4):18–20, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1997:ITI

[1344] Ben Smith. Industry trends: Tomorrow’s Internet is here today. Computer, 30(4):22–23, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1997:NBB

[1345] Lee Garber. News briefs: Binary version could bring VRML into the mainstream. FCC jumps into Internet fray. Java and floating-point math. Intel to design NDRAM. battle over net telephony. vendors seek fast modems. US permits export of strong encryption. E-commerce nears $1 billion. chasing the blue light. personal E-mail use will soar. Computer, 30(4):25–27, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:1997:SFP

[1346] D. Richard Kuhn. Sources of failure in the public switched telephone network. Computer, 30(4):31–36, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mudawwar:1997:MTM

[1347] Muhammed F. Mudawwar. Multicode:

A truly multilingual approach to text encoding: Unicode was designed to extend ASCII for encoding text in different languages, but it still have several important drawbacks. multicode overcomes those drawbacks. Computer, 30 (4):37–43, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also response and rebuttal [1394] and letter [1798].

Somani:1997:GEI

[1348] Arun K. Somani and Nitin H. Vaidya.

Guest Editors’ introduction: Understanding fault tolerance and reliability. Computer, 30(4):45–46, 50, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chou:1997:BFT

[1349] Timothy C. K. Chou. Beyond fault tolerance. Computer, 30(4):47–49, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Avizienis:1997:TSD

[1350] Algirdas Avizienis. Toward systematic design of fault-tolerant systems. Computer, 30(4):51–58, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maffeis:1997:PCT

[1351] Silvano Maffeis. Piranha: A CORBA tool for high availability. Computer, 30(4):59–66, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Guerraoui:1997:SBR

[1352] Rachid Guerraoui and André Schiper. Software-based replication for fault tolerance. Computer, 30(4):68–74, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tsai:1997:FIT

[1353] Mei-Chen Hsueh, Timothy K. Tsai, and Ravishankar K. Iyer. Fault injection techniques and tools. Computer, 30(4):75–82, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:1997:FTR

[1354] Biao Chen, Sanjay Kamat, and Wei Zhao. Fault-tolerant, real-time communication in FDDI-based networks. Computer, 30(4):83–90, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:1997:TAF

[1355] Mike Hinchey. Technical activities forum: TC strives to manage complexity. Computer, 30(4):91–92, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:1997:CUA

[1356] Carl Chang. CS update: Applicants sought for Editor-in-Chief of IEEE Multimedia. Computer, 30(4):92, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wiggins:1997:STL

[1357] Richard Wiggins. Standards: The top-level domain name controversy. Computer, 30(4):105–106, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1997:NPE

[1358] Scott Hamilton. New products: Embedded systems software: Real-time Java. embedded OADD. entrus for PCs. Privasuite. Computer, 30(4):107, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:1997:NBW

[1359] Michael J. Lutz. New books: Weighty computer science reference. high-level test synthesis. filling in the gaps. Epubs languages explained. fuzzy intro. Internet/intranet planning. controlling systems. Computer, 30(4):108, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sztipanovits:1997:IEM

[1360] Janos Sztipanovits and Gabor Karsai. Integrated engineering: Modelintegrated computing. Computer, 30 (4):110–111, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaner:1997:SRP

[1361] Cem Kaner. Software realities: Pitfalls and strategies in automated testing. Computer, 30(4):114–116, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kingoff:1997:IWC

[1362] Andrew Kingoff. Internet watch: Comparing Internet search engines. Computer, 30(4):117–118, April 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1997:SRO

[1363] W. Myers. Software reuse: ostriches beware. Computer, 30(4):119–120, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1997:BCS

[1364] Ware Myers. Binary critic: Software reuse: Ostriches beware. Computer, 30 (4):120, 119, April 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hemdal:1997:LGA

[1365] J. Erik Hemdal. Letters: Good advice for dealing with Y2K dates. Computer, 30(5):6, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holager:1997:LPB

[1366] Per Holager, Jean-Marc Jézéquel, and Bertrand Meyer. Letters: Placing the blame for Ariane 5. Computer, 30 (5):6, 8, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See [1293].

Spadaros:1997:LTY

[1367] Sally Spadaros. Letters: Thanks for the Y2K suggestions. Computer, 30(5):6, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Young:1997:LAA

[1368] David C. Young, Jean-Marc Jézéquel, and Bertrand Meyer. Letters: After all, it worked the last time. Computer, 30(5):8–9, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Olsen:1997:LWJ

[1369] Kurt Olsen. Letters: Why Java isn’t watered down C++. Computer, 30(5):9, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response to [1332]. See also [1403].

McReynolds:1997:LBS

[1370] David McReynolds. Letters: Brooks said it ten years ago. Computer, 30 (5):9–10, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response to [1332].

Gozum:1997:LWC

[1371] Marvine E. Gozum and Sorel Reisman.

Letters: Windows CE success. Computer, 30(5):10, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response to [1332].

Struble:1997:LDJ

[1372] Craig A. Struble and Ted Lewis. Letters: Deadlocks, and Java. Computer, 30(5):10, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response to [1332].

Holmes:1997:OCW

[1373] Neville Holmes. Open channel: What can computers do? Computer, 30 (5):11, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barlas:1997:FHC

[1374] Stephen Barlas. Up front: How critical is the shortage of IT workers? Computer, 30(5):14–16, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lawton:1997:ITE

[1375] George Lawton. Industry trends: Extranets: Next step for the Internet. Computer, 30(5):17, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1997:NBI

[1376] Lee Garber. News briefs: Internet security: Is it in the cards? Computer, 30(5):18–20, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferguson:1997:RAP

[1377] Pat Ferguson, Watts S. Humphrey, Soheil Khajenoori, Susan Macke, and Annette Matvya. Results of applying the personal software process. Computer, 30(5):24–31, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gemmer:1997:RMM

[1378] Art Gemmer. Risk management: Moving beyond process. Computer, 30 (5):33–43, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1997:CC

[1379] Scott Hamilton. E-commerce for the 21st century. Computer, 30(5):44–47, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tenenbaum:1997:ESI

[1380] Jay M. Tenenbaum, Tripatinder S.

Chowdhry, and Kevin Hughes. Eco system: An Internet commerce architecture. Computer, 30(5):48–55, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kambil:1997:DBW

[1381] Ajit Kambil. Doing business in the wired world. Computer, 30(5):56–61, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lange:1997:OOP

[1382] Danny B. Lange and Yuichi Nakamura. Object-oriented program tracing and visualization. Computer, 30 (5):63–70, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Uht:1997:BER

[1383] Augustus K. Uht, Vijay Sindagi, and Sajee Somanathan. Branch effect reduction techniques. Computer, 30 (5):71–81, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlson:1997:CUA

[1384] Bob Carlson. CS update: Andreessen, Bina receive 1997 McDowell award. Edward S. Davidson is 1996 Taylor Booth recipient. Aggarwal and Das earn technical achievement awards. Computer, 30(5):82–83, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kameyama:1997:TAF

[1385] Michitake Kameyama. Technical activities forum: Multiple-valued logic TC stresses innovation. Computer, 30 (5):83–85, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:CCa

[1386] Anonymous. Call and calendar. Computer, 30(5):85–89, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thompson:1997:SWP

[1387] Geoffrey O. Thompson. Standards: Work progresses on gigabit Ethernet. Computer, 30(5):95–96, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hamilton:1997:NPA

[1388] Scott Hamilton. New products: Agentbased distributed computing in Java: Voyager agents. E-commerce solutions for the Web: Concinity and Web Speed. software development: HiProf code profiler, distributed applications, client-server to the Web. stuff. Computer, 30(5):97–98, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:1997:NBHa

[1389] Michael J. Lutz. New books: How to create profitable businesses on the Net. Computer, 30(5):99, May 1997.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haas:1997:MHB

[1390] I. Mark Haas. Management: Hiring the best. Computer, 30(5):100–101, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meyer:1997:OTP

[1391] Bertrand Meyer. Object technology:

Practice to perfect: The quality first model. Computer, 30(5):102, 103, 105, 106, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maki:1997:IWL

[1392] William S. Maki and Ruth H. Maki.

Internet watch: Learning without lectures: A case study. Computer, 30 (5):107–108, May 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lewis:1997:BCA

[1393] Ted Lewis. Binary critic: Absorb and extend: Resistance is futile! Computer, 30(5):112, 109–111, May 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Do:1997:LEU

[1394] James Do and Muhammed Mudawwar.

Letters to the editor: Unicode misunderstood. Computer, 30(6):6, 9, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response and rebuttal to [1347].

Srinivas:1997:JBE

[1395] Kankanahalli Srinivas, V. “Juggy” Jagannathan, Y. V. Ramana Reddy, and Raghu Karinthi. Java and beyond: Executable content. Computer, 30(6):52, June 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gosling:1997:FJ

[1396] James Gosling. The feel of Java. Computer, 30(6):53–57, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Munson:1997:SJF

[1397] Jonathan P. Munson and Prasun Dewan. Sync: A Java framework for mobile collaborative applications. Computer, 30(6):59–66, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsieh:1997:ONC

[1398] Cheng-Hsueh A. Hsieh, Marie T. Conte, Teresa L. Johnson, John C. Gyllenhaal, and Wen mei W. Hwu. Optimizing NET compilers for improved Java performance. Computer, 30(6):67–75, June 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:1997:SCD

[1399] X. Nick Zhang. Secure code distribution. Computer, 30(6):76–79, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses distribution security problem for Java, Safe Tcl, and Safe Perl.

Leppinen:1997:JCB

[1400] Mika Leppinen, Pekka Pulkkinen, and Aapo Rautiainen. Java- and CORBAbased network management. Computer, 30(6):83–87, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:NPC

[1401] Anonymous. New products: CodeWizard for Java. Computer, 30(6):106, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ram:1997:IWE

[1402] Prabhu Ram and Robert Abarbanel.

Internet watch: Enterprise computing: The Java factor. Computer, 30(6): 115–119, June 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Welch:1997:LEI

[1403] Peter Welch. Letters to the editor:

Ignore the monitor methods! Computer, 30(7):7, 9, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Comments on [1332], and raises concerns about the inadequacy of the Java threads model, and points to http://www.cs.

bris.ac.uk/~alan/javappp.html as

a source for Java classes implementing C. A. R. Hoare’s Communicating Sequential Processes. See also [1369].

Holmes:1997:MIC

[1404] Neville Holmes. The myth of the intelligent computer. Computer, 30(7):8–??, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1997/pdf/r7008.pdf;
http://www.computer.org/computer/
co1997/r7008abs.htm.

Garber:1997:ISD

[1405] Lee Garber and Paul Korzeniowski. Internet splits over domain name controversy. Computer, 30(7):12–14, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7012.pdf;
http://www.computer.org/computer/
co1997/r7012abs.htm.

Flohr:1997:IOF

[1406] Udo Flohr. Interactive online 3D’s future. Computer, 30(7):15–16, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7015.pdf;
http://www.computer.org/computer/
co1997/r7015abs.htm.

Garber:1997:NBP

[1407] Lee Garber. News briefs: Plans for Java standards draws fire. Computer, 30(7):17, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Discusses the pending request by Sun to the Joint Technical Committee of the International Organization for Standardization and the International Electrotechnical Commission to standardize Java, and the opposition that has been raised by Sun’s competitors.

Anonymous:1997:AS

[1408] Anonymous. Article summaries. Computer, 30(7):20–21, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7020.pdf;

http://www.computer.org/computer/
co1997/r7020abs.htm.

VanderWiel:1997:WCA

[1409] Steven P. VanderWiel and David J.

Lilja. When caches aren’t enough: Data prefetching techniques. Computer, 30(7):23–30, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co1997/
pdf/r7023.pdf;   http://www.computer.
org/computer/co1997/r7023abs.htm.

Harel:1997:EOM

[1410] David Harel and Eran Gery. Executable object modeling with statecharts. Computer, 30(7):31–42, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7031.pdf;
http://www.computer.org/computer/
co1997/r7031abs.htm.

Conners:1997:MVT

[1411] Richard W. Conners, D. Earl Kline, Philip A. Araman, and Thomas H. Drayer. Machine vision technology for the forest products industry. Computer, 30(7):43–48, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7043.pdf;
http://www.computer.org/computer/
co1997/r7043abs.htm.

Bertozzi:1997:VBV

[1412] Massimo Bertozzi and Alberto Broggi.

Vision-based vehicle guidance. Computer, 30(7):49–55, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co1997/
pdf/r7049.pdf;   http://www.computer.
org/computer/co1997/r7049abs.htm.

Parvin:1997:VSO

[1413] Bahram Parvin, John R. Taylor, Daniel E. Callahan, William E. Johnston, and Ulrich Dahmen. Visual servoing for online facilities. Computer, 30(7):56–62, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7056.pdf;
http://www.computer.org/computer/
co1997/r7056abs.htm.

Stenstrom:1997:BPS

[1414] Per Stenström, Mats Brorsson, Fredrik Dahlgren, Håkan Grahn, and Michel Dubois. Boosting the performance of shared memory multiprocessors. Computer, 30(7):63–70, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7063.pdf;
http://www.computer.org/computer/
co1997/r7063abs.htm.

Deijl:1997:CVT

[1415] Eric van der Deijl, Gerco Kanbier, Olivier Temam, and Elena D. Granston. A cache visualization tool. Computer, 30(7):71–78, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7071.pdf;
http://www.computer.org/computer/
co1997/r7071abs.htm.

Anonymous:1997:TR

[1416] Anonymous. Treasurer’s report.

Computer, 30(7):80–81, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7080.pdf;
http://www.computer.org/computer/
co1997/r7080abs.htm.

Anonymous:1997:AR

[1417] Anonymous. Auditor’s report.

Computer, 30(7):81–85, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7081.pdf;
http://www.computer.org/computer/
co1997/r7081abs.htm.

Anonymous:1997:CU

[1418] Anonymous. CS update. Computer, 30(7):86–88, July 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1997/pdf/
r7086.pdf;   http://www.computer.
org/computer/co1997/r7086abs.htm.

Tripp:1997:SAC

[1419] Leonard L. Tripp. Survey addresses Computer Society’s scope, purpose, values. Computer, 30(7):89–90, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7089.pdf;
http://www.computer.org/computer/
co1997/r7089abs.htm.

Anonymous:1997:CC

[1420] Anonymous. Call and calendar.

Computer, 30(7):91–95, 97–99, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7091.pdf;
http://www.computer.org/computer/
co1997/r7091abs.htm.

Crocker:1997:DNE

[1421] Dave Crocker. Domain name expansion plan deserves support. Computer, 30(7):104–105, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7104.pdf;
http://www.computer.org/computer/
co1997/r7104abs.htm.

Anonymous:1997:NP

[1422] Anonymous. New products. Computer, 30(7):106–107, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7106.pdf;
http://www.computer.org/computer/
co1997/r7106abs.htm.

Anonymous:1997:NBb

[1423] Anonymous. New books. Computer, 30(7):109–??, July 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1997/pdf/
r7109.pdf;   http://www.computer.
org/computer/co1997/r7109abs.htm.

Oliver:1997:OOS

[1424] David W. Oliver. Optimizing the organization’s structure. Computer, 30(7):110–112, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7110.pdf;
http://www.computer.org/computer/
co1997/r7110abs.htm.

Meyer:1997:NSB

[1425] Bertrand Meyer. The next software breakthrough. Computer, 30(7): 113–114, July 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1997/pdf/
r7113.pdf;   http://www.computer.
org/computer/co1997/r7113abs.htm.

Lynch:1997:ILW

[1426] Patrick J. Lynch and Sarah Horton.

Imprudent linking weaves a tangled Web. Computer, 30(7):115–117, July 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1997/pdf/r7115.pdf;
http://www.computer.org/computer/
co1997/r7115abs.htm.

Anonymous:1997:La

[1427] Anonymous. Letters: Debunking metaphors and objects. Computer, 30(8):8, August 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8008.pdf.

Pescio:1997:OCS

[1428] Carlo Pescio. Open channel: Seven aphorisms of software development. Computer, 30(8):9, August 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r8009.pdf.

Anonymous:1997:RMC

[1429] Anonymous. Report to members: Computer Society debuts electronic subscriptions. Computer, 30(8):10– 11, August 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //neumann.computer.org/co/books/
                                        co1997/pdf/r8010.pdf.

Edwards:1997:FVB

[1430] John Edwards. Up front: Voice-based interfaces make PCs better listeners. Computer, 30(8):14–16, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://neumann.computer.org/co/
                                        books/co1997/pdf/r8014.pdf.

Vaughan-Nichols:1997:ITM

[1431] Steven J. Vaughan-Nichols. Industry trends: Making the PC easier to use. Computer, 30(8):17, 25, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://neumann.computer.org/co/
                                        books/co1997/pdf/r8017.pdf.

Garber:1997:NBM

[1432] Lee Garber. News briefs: Mars Imager makes new use of existing technology; chip industry litigation. Computer, 30 (8):20–22, 24, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //neumann.computer.org/co/books/
                                        co1997/pdf/r8020.pdf.

Anonymous:1997:ASa

[1433] Anonymous. Article summaries. Computer, 30(8):26–27, August 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r8026.pdf.

Murphy:1997:RRM

[1434] Gail C. Murphy and David Notkin.

Reengineering with reflexion models: A case study. Computer, 30 (8):29–36, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8029.pdf;   http://
www.computer.org/computer/co1997/
r8029abs.htm.

Strong:1997:PRI

[1435] Diane M. Strong, Richard Y. Wang, and Yang W. Lee. 10 potholes in the road to information quality. Computer, 30(8):38–46, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8038.pdf;   http://
www.computer.org/computer/co1997/
r8038abs.htm.

Stoyen:1997:GEI

[1436] Alex Stoyen. Guest editor’s introduction: Fighting complexity in computer systems. Computer, 30 (8):47–48, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8047.pdf;   http://
www.computer.org/computer/co1997/
r8047abs.htm.

Dikel:1997:ASP

[1437] David Dikel, David Kane, Steve Ornburn, William Loftus, and Jim Wilson. Applying software productline architecture. Computer, 30(8): 49–55, August 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8049.pdf;   http://
www.computer.org/computer/co1997/
r8049abs.htm.

Ko:1997:UGA

[1438] King-Tim Ko, Kit-Sang Tang, Cheung-Yau Chan, Kim-Fung Man, and Sam Kwong. Using genetic algorithms to design mesh networks. Computer, 30(8):56–61, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8056.pdf;   http://
www.computer.org/computer/co1997/
r8056abs.htm.

Kim:1997:OSR

[1439] K. H. (Kane) Kim. Object structures for real-time systems and simulators. Computer, 30(8):62– 70, August 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8062.pdf;   http://
www.computer.org/computer/co1997/
r8062abs.htm.

Anonymous:1997:TAFa

[1440] Anonymous. Technical activities forum: TC on system packaging. Computer, 30(8):71, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8071.pdf.

Anonymous:1997:CUCb

[1441] Anonymous. CS update: Computer Society to use electronic balloting. Computer, 30(8):72, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8072.pdf.

Anonymous:1997:CCb

[1442] Anonymous. Call and calendar. Computer, 30(8):75–79, 81–82, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic).  URL http://neumann.
computer.org/co/books/co1997/pdf/
r8075.pdf.

Severance:1997:SCL

[1443] Charles Severance. Standards: Could LDAP be the next killer DAP? Computer, 30(8):88–89, August 1997. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r8088.pdf.

Anonymous:1997:NPI

[1444] Anonymous. New products: Intranet servers come bearing superlatives. Computer, 30(8):90–91, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r8090.pdf.

Lutz:1997:NBC

[1445] Michael J. Lutz. New books: Candid conversations with giants. Computer, 30(8):92, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8092.pdf.

Mrva:1997:IER

[1446] Michael Mrva. Integrated engineering: Reuse factors in embedded systems design. Computer, 30(8):93– 95, August 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8093.pdf.

Bach:1997:SRG

[1447] James Bach. Software realities: Good enough quality: Beyond the buzzword. Computer, 30(8):96–98, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r8096.pdf.

Rhyne:1997:IWI

[1448] Theresa-Marie Rhyne, Don Brutzman, and Michael Macedonia. Internet watch: Internetworked graphics and the Web. Computer, 30 (8):99–101, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8099.pdf.

Sampson:1997:BCA

[1449] Mark E. Sampson. Binary critic: The allegory of the humidifier: ROI for systems engineering. Computer, 30(8): 104, 102, 103, August 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r8104.pdf.

Kayton:1997:LBD

[1450] Myron Kayton and Peter Wolstenholme. Letters: Back in the days of software “black fridays”. Computer, 30(9):6, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9006.pdf.

Duvanenko:1997:OCT

[1451] Victor J. Duvanenko. Open channel:

Two writes make a read. Computer, 30(9):8, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9008.pdf.

Simms:1997:FVS

[1452] David Simms. Up front: Vendors struggle with costs, benefits of shrinking cycle times. Computer, 30(9): 12–14, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9012.pdf.

Mullich:1997:ITD

[1453] Joe Mullich. Industry trends: Data mining: Making data meaningful. Computer, 30(9):18, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r9018.pdf.

Garber:1997:NBAa

[1454] Lee Garber. News briefs: Agency: Net use hasn’t hurt US phone system; Apple unveils turnaround strategy; JTC rejects Java standards plan; vendors plan 300-nm wafers; battle brews over smart card encryption; countries lift Internet telephone ban; Motorola proposes satellite system; vendors agree on high-capacity disks; “wrapping” software sales; Tonga offers domain name alternative. Computer, 30(9): 19–22, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9019.pdf.

Anonymous:1997:ASb

[1455] Anonymous. Article summaries. Computer, 30(9):26–27, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r9026.pdf.

Asokan:1997:SAE

[1456] N. Asokan, Phillipe A. Janson, Michael Steiner, and Michael Waidner. The state of the art in electronic payment systems. Computer, 30(9): 28–35, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9028.pdf;   http://
www.computer.org/computer/co1997/
r9028abs.htm.

Matzke:1997:WPS

[1457] Doug Matzke. Will physical scalability sabotage performance gains? Computer, 30(9):37–39, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://neumann.computer.org/co/
books/co1997/pdf/r9037.pdf;   http:

//www.computer.org/computer/co1997/
r9037abs.htm.

Fisher:1997:WTT

[1458] Joseph A. Fisher. Walk-time techniques: Catalyst for architectural change. Computer, 30(9):40–42, September 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9040.pdf;   http://
www.computer.org/computer/co1997/
r9040abs.htm.

Diefendorff:1997:HMW

[1459] Keith Diefendorff and Pradeep K.

Dubey. How multimedia workloads will change processor design. Computer, 30 (9):43–45, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9043.pdf;   http://
www.computer.org/computer/co1997/
r9043abs.htm.

Burger:1997:GEI

[1460] Doug Burger and James R. Goodman.

Guest editors introduction: Billiontransistor architectures. Computer, 30 (9):46–49, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9046.pdf.

Patt:1997:OBT

[1461] Yale N. Patt, Sanjay J. Patel, Marius Evers, Daniel H. Friendly, and Jared Stark. One billion transistors, one uniprocessor, one chip. Computer, 30 (9):51–57, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9051.pdf;   http://
www.computer.org/computer/co1997/
r9051abs.htm.

Lipasti:1997:SMB

[1462] Mikko H. Lipasti and John Paul Shen. Superspeculative microarchitecture for beyond AD 2000. Computer, 30(9): 59–66, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9059.pdf;   http://
www.computer.org/computer/co1997/
r9059abs.htm.

Smith:1997:TPM

[1463] James E. Smith and Sriram Vajapeyam. Trace processors: Moving to fourth-generation microarchitectures. Computer, 30(9):68– 74, September 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9068.pdf;   http://
www.computer.org/computer/co1997/
r9068abs.htm.

Kozyrakis:1997:SPB

[1464] Christoforos E. Kozyrakis, Stylianos Perissakis, David Patterson, Thomas Anderson, Krste Asanović, Neal Cardwell, Richard Fromm, Jason Golbus, Benjamin Gribstad, Kimberly Keeton, Randi Thomas, Noah Treuhaft, and Katherine Yelick. Scalable processors in the billion-transistor era: IRAM. Computer, 30(9):75– 78, September 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9075.pdf;   http://
www.computer.org/computer/co1997/
r9075abs.htm.

Hammond:1997:SCM

[1465] Lance Hammond, Basem A. Nayfeh, and Kunle Olukotun. A single-chip multiprocessor. Computer, 30(9): 79–85, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9079.pdf;   http://
www.computer.org/computer/co1997/
r9079abs.htm.

Waingold:1997:BIA

[1466] Elliot Waingold, Michael Taylor, Devabhaktuni Srikrishna, Vivek Sarkar, Walter Lee, Victor Lee, Jang Kim, Matthew Frank, Peter Finch, Rajeev Barua, Jonathan Babb, Saman Amarasinghe, and Anant Agarwal. Baring it all to software: Raw machines. Computer, 30(9):86–93, September 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9086.pdf;   http://
www.computer.org/computer/co1997/
r9086abs.htm.

Okazaki:1997:TAF

[1467] Chie Okazaki, Ivan Gimenez, Juan Noguera-Rodriguez, and Tadao Ichikawa. Technical activities forum: Students take lead in Yuforic. Computer, 30 (9):94–95, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9094.pdf.

Anonymous:1997:CUb

[1468] Anonymous. CS update: IEEE Computer Society Election: Nominees for Computer Society office and Board of Governors. Computer, 30 (9):97–104, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //neumann.computer.org/co/books/
                                        co1997/pdf/r9097.pdf.

Anonymous:1997:CCc

[1469] Anonymous. Call and calendar. Computer, 30(9):105–113, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://neumann.computer.org/co/
                                        books/co1997/pdf/r9105.pdf.

Foley:1997:SOR

[1470] Jerry Foley and Charles Severance. Standards: OSI retrospect and prospect. Computer, 30(9): 123–124, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //neumann.computer.org/co/books/
                                        co1997/pdf/r9123.pdf.

Garber:1997:NPN

[1471] Lee Garber. New products: Networking tools for the enterprise. Computer, 30(9):125–126, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://neumann.computer.org/co/
                                        books/co1997/pdf/r9125.pdf.

Lutz:1997:NBHb

[1472] Michael J. Lutz. New books: Hamlet on the holodeck. Computer, 30 (9):127, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9127.pdf.

Gurbaxani:1997:MBI

[1473] Vijay Gurbaxani and Janet Wilson.

Management: Bridging IT and business: Techno MBAs. Computer, 30 (9):128–129, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9128.pdf.

Pescio:1997:OTP

[1474] Carlo Pescio. Object technology: Principles versus patterns. Computer, 30 (9):130–131, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9130.pdf.

Mason:1997:IWS

[1475] L. John Mason. Internet watch:

Stress and the ultimate Luddite survival mechanism. Computer, 30(9): 132–133, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//neumann.computer.org/co/books/
co1997/pdf/r9132.pdf.

Lewis:1997:BCD

[1476] Ted Lewis. Binary critic: Digitopolis meets encalming technology. Computer, 30(9):136, 134–135, September 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gomez:1997:LWK

[1477] Dan Gomez. Letter: Worth keeping an eye one. Computer, 30(10):6, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx006.pdf.

Wooller:1997:LCU

[1478] Tim Wooller. Letter: Citing URLs: A matter of trust? Computer, 30 (10):6, October 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://pdf.

                                        computer.org/co/books/co1997/pdf/
                                        rx006.pdf.

Hardiman:1997:OCP

[1479] Lawrence T. Hardiman. Open channel: Personality types and software engineers. Computer, 30(10):10, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx010.pdf.

Clark:1997:FWN

[1480] David Clark. Up front: Will new fiber technology quench the thirst for bandwidth? Computer, 30(10):14–15, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx014.pdf.

Lawton:1997:ITD

[1481] George Lawton. Industry trends: Dawn of the Internet appliance. Computer, 30(10):16, 18, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx016.pdf.

Garber:1997:NBAb

[1482] Lee Garber. News briefs: Appliance war could make Web less open; Microsoft, Netscape agree of VRML; CORBA to stay independent of Java; US runs out of visas for foreign computer workers; new Year 2000 problems emerge. Computer, 30(10):20, 23–25, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx020.pdf.

Anonymous:1997:ASc

[1483] Anonymous. Article summaries. Computer, 30(10):26–27, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx026.pdf.

Hamilton:1997:DBH

[1484] Scott Hamilton and Lee Garber.

Deep Blue’s hardware-software synergy. Computer, 30(10):29–35, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx029.pdf;   http:
//www.computer.org/computer/co1997/
rx029abs.htm.

Jacobson:1997:MRB

[1485] Ivar Jacobson, Martin Griss, and Patrik Jonsson. Making the reuse business work. Computer, 30(10):36–42, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.

                                        org/co/books/co1997/pdf/rx036.pdf;
                                         http://www.computer.org/computer/
                                        co1997/rx036abs.htm.

Rine:1997:SRO

[1486] David C. Rine. Supporting reuse with object technology. Computer, 30(10):43–45, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx043.pdf;   http:
                                        //www.computer.org/computer/co1997/
                                        rx043abs.htm.

Fichman:1997:OTR

[1487] Robert G. Fichman and Chris F. Kemerer. Object technology and reuse: Lessons from early adopters. Computer, 30(10):47–59, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx047.pdf;   http:
                                        //www.computer.org/computer/co1997/
                                        rx047abs.htm.

Zubeck:1997:IRR

[1488] John C. Zubeck. Implementing reuse with RAD Tools’ native objects. Computer, 30(10):60–65, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx060.pdf;   http:
                                        //www.computer.org/computer/co1997/
                                        rx060abs.htm.

Etzkorn:1997:AIR

[1489] Letha H. Etzkorn and Carl G. Davis. Automatically identifying reusable OO legacy code. Computer, 30(10):66–71, October 1997.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx066.pdf;   http:
//www.computer.org/computer/co1997/
rx066abs.htm.

Damiani:1997:DBA

[1490] Ernesto Damiani, Maria Grazia Fugini, and Enrico Fusaschi. A descriptorbased approach to OO code reuse. Computer, 30(10):73–80, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx073.pdf;   http:
//www.computer.org/computer/co1997/
rx073abs.htm.

Baggi:1997:TAF

[1491] Dennis Baggi. Technical Activities Forum: Addressing the evolving profile of computer professionals. Computer, 30(10):84–85, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx084.pdf.

Anonymous:1997:CUc

[1492] Anonymous. CS update: Awards recognize three pioneers: Richard F. Clippinger, Francis Elizabeth Snyder-Holberton, Robert E. Kahn; James T. Cain: Richard E. Merwin award; Mike Liu: Richard E. Merwin award; Sartaj K. Sahni: Taylor Booth award; Thuraisingham and Tsai: Technical achievement awards; Leonard L. Tripp: Hans Karlsson award. Computer, 30(10):85–87, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx085.pdf.

Anonymous:1997:CCd

[1493] Anonymous. Call and calendar. Computer, 30(10):93–95, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx093.pdf.

Proha:1997:SSS

[1494] Vir Proha. Standards: A standard for software documentation. Computer, 30(10):97–98, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx097.pdf.

Anonymous:1997:COa

[1495] Anonymous. Career opportunities. Computer, 30(10):100–110, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx100.pdf.

Kroeker:1997:NPSa

[1496] Kirk L. Kroeker. New products: SDKs, compilers, and component suites. Computer, 30(10):111–113, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rx111.pdf.

Lutz:1997:NBM

[1497] Michael J. Lutz. New books: Making beautiful music with DSPs; for hardware junkies; smart tools; recognizing patterns; keeping Internet secrets; SGML for developers; compressing with fractals; rapid GUI development. Computer, 30(10):114, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx114.pdf.

Mangione-Smith:1997:IEA

[1498] William H. Mangione-Smith. Integrated engineering: Application design for configurable computing. Computer, 30(10):115–117, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx115.pdf.

Hohmann:1997:SRM

[1499] Luke Hohmann. Software realities:

In methods we trust? Computer, 30(10):119–121, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx119.pdf.

Hundt:1997:IWI

[1500] Reed Hundt. Internet watch: The Internet: From here to ubiquity. Computer, 30(10):122–124, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rx122.pdf.

Bollinger:1997:BCI

[1501] Terry Bollinger. Binary critic: The interplay of art and science in software. Computer, 30(10):128, 125–127, October 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1997/pdf/rx128.pdf.

Anonymous:1997:Lb

[1502] Anonymous.  Letters.  Com-

puter, 30(11):6, 11, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/ry006.pdf.

Lawton:1997:FSR

[1503] George Lawton. Up front: In search of real-time Internet service. Computer, 30(11):14–16, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/ry014.pdf.

Ortiz:1997:ITV

[1504] Sixto Ortiz Jr. Industry trends: Virtual private networks: Leveraging the Internet. Computer, 30(11):18– 20, November 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1997:NBJ

[1505] Lee Garber and David Clark. News briefs: Judge rejects US restrictions on export of encryption; Intel, HP unveil Merced chip; Java wars heat up; AOL acquires CompuServe subscribers; confusion in the DVD marketplace; ‘amazing grace’ heads to sea; PC firms back down on convergence; business use will drive Internet growth; company offers $1-million prize for hackers. Computer, 30(11):22–25, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/ry022.pdf.

Anonymous:1997:ASd

[1506] Anonymous. Article summaries. Computer, 30(11):26–27, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry026.pdf.

Wang:1997:CPI

[1507] Huaiqing Wang and Chen Wang. Computing practices: Intelligent agents in the nuclear industry. Computer, 30 (11):28–34, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maulsby:1997:CTA

[1508] David Maulsby and Ian H. Witten. Cybersquare: Teaching agents to learn: From user study to implementation. Computer, 30(11):36–44, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

El-Rewini:1997:KPI

[1509] Hesham El-Rewini, Michael C. Mulder, Peter A. Freeman, Gordon E. Stokes, Innes Jelly, Lillian Cassel, Doris K. Lidtke, Stefano Russo, Bernd J. Krämer, Jimmie E. Haines, and Joe Turner. Keeping pace with an information society: A virtual roundtable. Computer, 30(11):46–57, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry046.pdf;   http:
//www.computer.org/computer/co1997/
ry046abs.htm.

Garlan:1997:ACE

[1510] David Garlan, David P. Gluch, and James E. Tomayko. Agents of change:

Educating software engineering leaders. Computer, 30(11):59–65, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1997/pdf/ry059.pdf;
 http://www.computer.org/computer/
co1997/ry059abs.htm.

Werner:1997:RFA

[1511] Tony Werner and Venkatesh Akella. Research feature: Asynchronous processor survey. Computer, 30(11):67– 76, November 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://pdf.

computer.org/co/books/co1997/pdf/
ry059.pdf;   http://www.computer.
org/computer/co1997/ry059abs.htm.

Kamel:1997:RFP

[1512] Nabil N. Kamel. Research feature: Predicate caching for dataintensive autonomous systems. Computer, 30(11):77–73, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry077.pdf;   http:
//www.computer.org/computer/co1997/
ry077abs.htm.

Anonymous:1997:TAFc

[1513] Anonymous. Technical activities forum: Supercomputing Applications TC faces new challenges. Computer, 30(11):84, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry084.pdf.

Anonymous:1997:CUA

[1514] Anonymous. CS update: Award recognizes ERMA pioneer; Henry Oldfield; Technical Achievement Award, Luqi; service awards recognizes special efforts. Computer, 30(11):85–87, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1997/pdf/ry088.pdf;
http://www.computer.org/computer/
co1997/ry085.pdf.

Gotterbarn:1997:SEC

[1515] Don Gotterbarn, Keith Miller, and Simon Rogerson. Software engineering code of ethics, version 3. Computer, 30 (11):88–92, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:CCe

[1516] Anonymous. Call and calendar. Computer, 30(11):93–95, 97–99, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry093.pdf.

Anonymous:1997:COb

[1517] Anonymous. Career opportunities.

Computer, 30(11):100–124, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry100.pdf.

Kroeker:1997:NPSb

[1518] Kirk L. Kroeker. New products:

Speech recognition development. Computer, 30(11):125–129, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry125.pdf.

Kroeker:1997:NPJ

[1519] Kirk L. Kroeker. New products: Java and Internet development tools: Borland’s Internet Solutions Program, Magic Software Enterprise’s RADD for Enterprise and Internet apps, Symantec’s Visual Café for Java 2.0 professional development edition, Ahpah Software’s SourceAgain Java decompiler, KL Group’s JClass Chart 2.0 JavaBean. Computer, 30(11): 128–129, November 1997. CO-DEN CMPWAB. ISSN 0010-4841.

URL http://pdf.computer.org/co/
books/co1997/pdf/ry125.pdf.

Anonymous:1997:NBd

[1520] Anonymous. New books: Envisioning the endless frontier. Computer, 30(11):130, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry130.pdf.

Howe:1997:SIG

[1521] Roger Howe. Standards: ISO 14001: The green standard. Computer, 30(11):133–134, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry133.pdf.

Haas:1997:MBV

[1522] Mark Haas. Management: Building the vision: Lessons for all leaders. Computer, 30(11):135–136, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry133.pdf.

Creel:1997:OTO

[1523] Christopher Creel, Bertrand Meyer, and Philippe Stephan. Object technology: The opportunity of a millennium. Computer, 30(11):137–138, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/ry137.pdf.

Lewis:1997:BCC

[1524] Ted Lewis. Binary critic: Cars, phones, and Tamagotchi tribes. Computer, 30(11):144, 142–143, November 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:LOC

[1525] Anonymous. Letters: Software:

Design versus production. Computer, 30(12):6, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz006.pdf.

Prasad:1997:OCM

[1526] K. V. K. K. Prasad. Open channel: Matching people and jobs: The I-CMM proposal. Computer, 30(12):10, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz006.pdf.

Edwards:1997:FNI

[1527] John Edwards. Up front: New interfaces: Making computers more accessible. Computer, 30(12):12–14, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1997/pdf/rz012.pdf.

Herringshaw:1997:ITD

[1528] Chris Herringshaw. Industry trends: Detecting attacks on networks. Computer, 30(12):16–17, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz016.pdf.

Gerber:1997:NBF

[1529] Lee Gerber and David Clark. News briefs: Firms to invest $250 million in next-generation processors. Computer, 30(12):18–20, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz018.pdf.

Anonymous:1997:ASe

[1530] Anonymous. Article summaries. Computer, 30(12):22–23, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz022.pdf.

Bernsen:1997:CPW

[1531] Niels Ole Bernsen, Hans Dybkjær, and Laila Dybkjær. Computing practices: What should your speech system say? Computer, 30(12):25–31, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz025.pdf;   http:
//www.computer.org/computer/co1997/
rz025abs.htm.

Patt:1997:IOP

[1532] Yale N. Patt. Identifying obstacles in the path to more. Computer, 30 (12):32, December 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conte:1997:TFC

[1533] Thomas M. Conte, Pradeep K. Dubey, Matthew D. Jennings, Ruby B. Lee, Alex Peleg, Salliah Rathnam, Mike Schlansker, Peter Song, and Andrew Wolfe. Theme feature: Challenges to combining general-purpose and multimedia processors. Computer, 30(12):33–37, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz033.pdf;   http:
//www.computer.org/computer/co1997/
rz033abs.htm.

Mangione-Smith:1997:TFS

[1534] William H. Mangione-Smith, Brad Hutchings, David Andrews, André De-Hon, Carl Ebeling, Reiner Hartenstein, Oskar Mencer, John Morris, Krishna Palem, Viktor K. Prasanna, and Henk A. E. Spaanenburg. Theme feature: Seeking solutions in configurable computing. Computer, 30(12):38–43, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz038.pdf;   http:
//www.computer.org/computer/co1997/
rz038abs.htm.

Stenstrom:1997:TFT

[1535] Per Stenström, Erik Hagersten, David J. Lilja, Margaret Martonosi, and Madan Venugopal. Theme feature: Trends in shared memory multiprocessing. Computer, 30(12):44– 50, December 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adve:1997:TFC

[1536] Sarita V. Adve, Doug Burger, Rudolf Eigenmann, Alasdair Rawsthorne, Michael D. Smith, Catherine H. Gebotys, Mahmut T. Kandemir, David J. Lilja, Alok N. Choudhary, Jesse Z. Fang, and Pen-Chung Yew. Theme feature: Changing interaction of compiler and architecture. Computer, 30(12):51–58, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz051.pdf;   http:
//www.computer.org/computer/co1997/
rz051abs.htm.

Alkhatib:1997:TFW

[1537] Hasan S. Alkhatib, Chase Bailey, Mario Gerla, and James McCrae. Theme feature: Wireless data networks: Reaching the extra mile. Computer, 30(12):59–62, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz059.pdf;   http:
//www.computer.org/computer/co1997/
rz059abs.htm.

Schlansker:1997:TFC

[1538] MIchael Schlansker, Thomas M. Conte, James Dehnert, Kemar Ebcioğlu, Jesse Z. Fang, and Carol L. Thompson. Theme feature: Compilers for instruction-level parallelism. Computer, 30(12):63–69, December 1997.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz070.pdf.

Anonymous:1997:AI

[1539] Anonymous. Annual index. Computer, 30(12):70–79, 82–84, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz070.pdf.

Anonymous:1997:CEC

[1540] Anonymous. Computer 1998 editorial calendar. Computer, 30(12):84, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1997/pdf/rz084.pdf.

Anonymous:1997:CUM

[1541] Anonymous. CS update: Members elect officers and board members to begin terms in 1998; balloting and participation; Kenneth Laker wins in balloting for IEEE 1998 presidentelect; 1987 Gorden Bell Prize winners announced at supercomputing. Computer, 30(12):85–86, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz085.pdf.

Anonymous:1997:CR

[1542] Anonymous. Computer reviewers.

Computer, 30(12):87–89, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1997:CCf

[1543] Anonymous. Call and calendar. Computer, 30(12):90–93, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rz090.pdf.

Anonymous:1997:COc

[1544] Anonymous. Career opportunities. Computer, 30(12):94–121, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rz090.pdf.

Kroeker:1997:NPE

[1545] Kirk L. Kroeker. New products: Embedded systems development. Computer, 30(12):122–124, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rz122.pdf.

Lutz:1997:NBL

[1546] Michael J. Lutz. New books: Living in the digital age. Computer, 30(12):125, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rz125.pdf.

Dawson:1997:SEC

[1547] Frank Dawson. Standards: Emerging calendaring and scheduling standards. Computer, 30(12):126–128, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1997/pdf/rz126.pdf.

Voas:1997:SRF

[1548] Jeffrey Voas. Software realities: Fault injection for the masses. Computer, 30(12):129–130, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz126.pdf.

Magel:1997:IWI

[1549] Ken Magel. Internet watch: Is it too late to put the user back into HTML? Computer, 30(12):131–132, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz131.pdf.

Holmes:1997:BCC

[1550] Neville Holmes. Binary critic: Converging on program simplicity. Computer, 30(12):136, 134–135, December 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1997/pdf/rz136.pdf.

Carver:1998:PMB

[1551] Doris Carver. President’s message:

Building on success: Our 1998 plan. Computer, 31(1):4, 6, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1004.pdf.

Clark:1998:ITN

[1552] David Clark. Industry trends: New era for digital signal processors. Computer, 31(1):10–11, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1010.pdf.

Garber:1998:TNE

[1553] Lee Garber. Technology news: Employment in 1998: Focus on Y2K and the Internet. Computer, 31(1):14–18, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1014.pdf.

Garber:1998:NBS

[1554] Lee Garber. News briefs: Sun gets OK to propose Java standards; domain name plan delayed; encryption battle heats up; monstor processes to debut in 1998; work begins on API for digital TV; ISO approves C++ Standard; new standard for parallel processing workstations; IBM unveils high-capacity disk technology. Computer, 31(1):21–24, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1021.pdf.

Anonymous:1998:ASa

[1555] Anonymous. Article summaries. Computer, 31(1):26–27, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1026.pdf.

Clark:1998:IOF

[1556] David D. Clark, Edward A. Feigenbaum, Don Greenberg, Juris Hartmanis, Robert W. Lucky, Robert M. Metcalfe, Raj Reddy, Mary Shaw, and William Wulf. Innovation and obstacles: The future of computing. Computer, 31(1):29–38, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1029.pdf;   http:
//www.computer.org/computer/co1998/
r1029abs.htm.

Wilson:1998:VRC

[1557] Janet Wilson. Virtual roundtable:

Challenges and trends in processor design: Roundtable introduction. Computer, 31(1):39, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1039.pdf.

Tremblay:1998:VRC

[1558] Marc Tremblay. Virtual roundtable:

Challenges and trends in processor design: Increasing work, pushing the clock. Computer, 31(1):40–41, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1040.pdf.

Grohoski:1998:VRC

[1559] Greg Grohoski. Virtual roundtable:

Challenges and trends in processor design: Reining in complexity. Computer, 31(1):41–42, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1041.pdf.

Burgess:1998:VRC

[1560] Brad Burgess. Virtual roundtable:

Challenges and trends in processor design: Specialization: A way of life. Computer, 31(1):43, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1043.pdf.

Killian:1998:VRC

[1561] Earl Killian. Virtual roundtable: Challenges and trends in processor design: Challenges, not roadblocks. Computer, 31(1):44–45, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1044.pdf.

Colwell:1998:VRC

[1562] Robert Colwell. Virtual roundtable: Challenges and trends in processor design: Maintaining a leading position. Computer, 31(1):45–47, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1045.pdf.

Rubinfeld:1998:VRC

[1563] Paul I. Rubinfeld. Virtual roundtable: Challenges and trends in processor design: Managing problems at high speed. Computer, 31(1):47–48, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1047.pdf.

Hwu:1998:VRC

[1564] Wen mei Hwu. Virtual roundtable: Challenges and trends in processor design: Technology outlook: Introduction to predicated execution. Computer, 31(1):49–50, January 1998.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1049.pdf.

Hamilton:1998:IMR

[1565] Scott Hamilton. Inside Microsoft Research. Computer, 31(1):51–58, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1998/pdf/r1051.pdf;
 http://www.computer.org/computer/
co1998/r1051abs.htm.

Lewis:1998:IAG

[1566] Ted Lewis. Information appliances: Gadget netopia. Computer, 31(1):59–68, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1059.pdf;   http:
//www.computer.org/computer/co1998/
r1059abs.htm.

Jameel:1998:WWT

[1567] Akhtar Jameel, Matthias Stuempfle, Daniel Jiang, and Axel Fuchs. Web on wheels:toward Internet-enabled cars. Computer, 31(1):69–76, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1069.pdf;   http:
//www.computer.org/computer/co1998/
r1069abs.htm.

Lee:1998:EEF

[1568] Edward A. Lee and David G.

Messerschmitt. Engineering an education for the future. Computer, 31(1):77–85, January 1998.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1077.pdf;   http:
                                        //www.computer.org/computer/co1998/
                                        r1077abs.htm.

Karp:1998:NGB

[1569] Alan H. Karp, Ewing Lusk, and David H. Bailey. News: 1997 Gordon Bell Prize winners. Computer, 31(1):86–92, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1086.pdf.

Adam:1998:TAF

[1570] Nabil R. Adam and Aryya Gangopadhyay. Technical Activities Forum: Content-based retrieval in digital libraries. Computer, 31(1):93–94, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://pdf.computer.
                                        org/co/books/co1998/pdf/r1093.pdf.

Anonymous:1998:CUS

[1571] Anonymous. CS update: Seymour Cray Award; Pioneer Award for Svoboda; Golden Core memberships. Computer, 31(1):95–96, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r1095.pdf.

Anonymous:1998:CCa

[1572] Anonymous. Call and calendar. Computer, 31(1):97–101, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1998:NBa

[1573] Anonymous. New books. Computer, 31(1):128–129, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1128.pdf.

Anonymous:1998:NTa

[1574] Anonymous. New tools. Computer, 31(1):130–135, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1130.pdf.

Severance:1998:SCC

[1575] Charles Severance. Standards conflict and consensus: The role of standards. Computer, 31(1):138–139, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1138.pdf.

Meyer:1998:OTF

[1576] Bertrand Meyer. Object technology the future of object technology. Computer, 31(1):140–141, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1140.pdf.

Vetter:1998:IWI

[1577] Ron Vetter and Kirk L. Kroeker. Internet watch: The Internet in the year ahead. Computer, 31(1):143–144, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1998/pdf/r1143.pdf.

Kerth:1998:OCC

[1578] Norman L. Kerth, Jim Coplien, and Jerry Weinberg. Open channel: Call for the rational use of personality indicators. Computer, 31(1):146–147, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1146.pdf.

Lewis:1998:BCWa

[1579] Ted Lewis. Binary critic: Who’s afraid of Wintel? Computer, 31(1):152, 149–151, January 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r1152.pdf.

Weeks:1998:LSU

[1580] Bob Weeks, Brian Hackworth, Roland N. Ibbett, and Paul E. Black. Letters: Should the U.S. nationalize Microsoft? deserved debunking; earlier asynchronous processors; Edison’s invention factory. Computer, 31(2):4–6, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2004.pdf.

Kiely:1998:ITC

[1581] Don Kiely. Industry trends: Are components the future of software? Computer, 31(2):10–11, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2010.pdf.

Garber:1998:TNA

[1582] Lee Garber and Richard Raucci.

Technology news: Antivirus technology offers new cures. Computer, 31(2):12–14, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2012.pdf.

Anonymous:1998:NBO

[1583] Anonymous. News briefs: Open spec for EDA tools; semiconductor roadmap; high-tech impact on US economy. Computer, 31(2):16–18, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2016.pdf.

Caulfield:1998:POC

[1584] H. John Caulfield. Perspectives in optical computing. Computer, 31(2): 22–25, February 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://pdf.

computer.org/co/books/co1998/pdf/
r2022.pdf;   http://www.computer.
org/computer/co1998/r2022abs.htm.

Johnson:1998:CPE

[1585] Neil F. Johnson and Sushil Jajodia. Computing practices: Exploring steganography: Seeing the unseen. Computer, 31(2):26–34, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2026.pdf;   http:
//www.computer.org/computer/co1998/
r2026abs.htm.

Irakliotis:1998:OMT

[1586] Leo J. Irakliotis and Pericles A. Mitkas. Optics: A maturing technology for better computing. Computer, 31(2):36–37, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2036.pdf;   http:
//www.computer.org/computer/co1998/
r2036abs.htm.

Gourlay:1998:DFS

[1587] James Gourlay, Tsung-Yi Yang, Julian A. B. Dines, John F. Snowdon, and Andrew C. Walker. Development of free-space digital optics in computing. Computer, 31(2):38–44, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2038.pdf;   http:
//www.computer.org/computer/co1998/
r2038abs.htm.

Mitkas:1998:OPP

[1588] Pericles A. Mitkas, George A. Betzos, and Leo J. Irakliotis. Optical processing paradigms for electronic computers. Computer, 31(2):45–51, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2045.pdf;   http:
//www.computer.org/computer/co1998/
r2045abs.htm.

Psaltis:1998:HDS

[1589] Demetri Psaltis and Geoffrey W. Burr. Holographic data storage. Computer, 31(2):52–60, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2052.pdf;   http:
//www.computer.org/computer/co1998/
r2052abs.htm.

Ishikawa:1998:OIP

[1590] Masatoshi Ishikawa and Neil McArdle. Optically interconnected parallel computing systems. Computer, 31(2):61–68, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2061.pdf;   http:
//www.computer.org/computer/co1998/
r2061abs.htm.

Guilfoyle:1998:FSI

[1591] Peter S. Guilfoyle, John M. Hessenbruch, and Richard V. Stone. Free-space interconnects for highperformance optoelectronic switching. Computer, 31(2):69–75, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2069.pdf;   http:
//www.computer.org/computer/co1998/
r2069abs.htm.

Cooke:1998:TAF

[1592] Daniel E. Cooke. Technical activities forum: Outlook on language research and practice. Computer, 31(2):76–77, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2076.pdf.

Anonymous:1998:CUI

[1593] Anonymous. CS update: 1998 IEEE Fellows election timetable’ 1998 Executive Committee; call for Editor-in-Chief applicants; 1997 Sidney Fernbach Award; Membership Activities Board service award. Computer, 31(2):79–83, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r2079.pdf.

Anonymous:1998:CCb

[1594] Anonymous. Call and calendar. Computer, 31(2):84–87, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r2084.pdf.

Lutz:1998:NB

[1595] Michael J. Lutz. New books. Computer, 31(2):100, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r2100.pdf.

Kroeker:1998:NT

[1596] Kirk L. Kroeker. New tools. Computer, 31(2):101–105, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r2101.pdf.

Loughry:1998:SIS

[1597] Donald C. Loughry. Standards: The IEEE-SA: A new era for standards. Computer, 31(2):106, 110, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r2106.pdf.

Sage:1998:IET

[1598] Andrew P. Sage. Integrated engineering: Toward systems ecology. Computer, 31(2):107–110, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2107.pdf.

Bach:1998:SRM

[1599] James Bach. Software realities: Microdynamics of process evolution. Computer, 31(2):111–113, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2111.pdf.

Anonymous:1998:IWU

[1600] Anonymous. Internet watch: URLs in print: To cite or not to cite? Computer, 31(2):114–115, 117, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2114.pdf.

Carreira:1998:OCC

[1601] Joao Carreira and Joao Gabriel Silva.

Open channel: Computer science and the Pygmalion effect. Computer, 31(2):116–117, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2116.pdf.

Lawson:1998:BCS

[1602] Harold W. Lawson. Binary critic: Salvation from system complexity. Computer, 31(2):120, 118–119, February 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r2120.pdf.

McCabe:1998:LWG

[1603] Rich McCabe, Ted Lewis, Patrick Zazcyk, Brian L. Evans, Mihai Nadin, and Angela Burgess. Letters: When is a giveaway monopolistic? equal access to the OS; DSP article on target; criteria for inclusion. Computer, 31(3):4, 6, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3004.pdf.

Edwards:1998:ITCa

[1604] John Edwards. Industry trends: Changing database market hurts major vendors. Computer, 31(3):10–11, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1998/pdf/r3010.pdf.

Parker:1998:TNM

[1605] Tammy Parker. Technology news: Mobile wireless Internet technology faces hurdles. Computer, 31(3):12–14, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3012.pdf.

Anonymous:1998:NBM

[1606] Anonymous. News briefs: Microsoft, Sun, and set-top boxes; Europe’s high-speed Internet; Microsoft’s court settlement; new cable modem spec. Computer, 31(3):16–19, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3016.pdf.

Anonymous:1998:ASb

[1607] Anonymous. Article summaries.

Computer, 31(3):20–21, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3020.pdf.

Ousterhout:1998:SHL

[1608] John K. Ousterhout. Scripting:

Higher-level programming for the 21st century. Computer, 31(3):23–30, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3023.pdf;   http:
//www.computer.org/computer/co1998/
r3023abs.htm.

Dingle:1998:ICP

[1609] Adair Dingle and Thomas H. Hildebrandt. Improving C++ performance using temporaries. Computer, 31(3):31–41, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3031.pdf;   http:
//www.computer.org/computer/co1998/
r3031abs.htm.

Krieger:1998:EDC

[1610] David Krieger and Richard M.

Adler. The emergence of distributed component platforms. Computer, 31(3):43–53, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3043.pdf;   http:

                                        //www.computer.org/computer/co1998/
                                        r3043abs.htm.

OLeary:1998:EKM

[1611] Daniel E. O’Leary. Enterprise knowledge management. Computer, 31(3):54–61, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r3054.pdf;   http:
                                        //www.computer.org/computer/co1998/
                                        r3054abs.htm.

Melliar-Smith:1998:SNP

[1612] Peter Michael Melliar-Smith and Louise Elizabeth Moser. Surviving network partitioning. Computer, 31(3):62–68, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://pdf.computer.org/co/
                                        books/co1998/pdf/r3062.pdf;   http:
                                        //www.computer.org/computer/co1998/
                                        r3062abs.htm.

Sheppard:1998:MCS

[1613] John W. Sheppard and William R. Simpson. Managing conflict in system diagnosis. Computer, 31(3):69–76, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://pdf.computer.
                                        org/co/books/co1998/pdf/r3069.pdf;
                                         http://www.computer.org/computer/
                                        co1998/r3069abs.htm.

Utamaphethai:1998:DDA

[1614] Noppanunt Utamaphethai and Sumit Ghosh. Dicaf: A distributed architecture for intelligent transportation. Computer, 31(3):78–84, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3078.pdf;   http:
//www.computer.org/computer/co1998/
r3078abs.htm.

Anonymous:1998:CUCa

[1615] Anonymous. CS update: CS launches Y2K-compliant standards effort; tribute to Richard Hamming. Computer, 31(3):85–87, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3085.pdf.

Haus:1998:TAF

[1616] Goffredo Haus. Technical Activities Forum: Rescuing La Scala’s music archives. Computer, 31(3):88–89, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1998/pdf/r3088.pdf.

Anonymous:1998:CCc

[1617] Anonymous. Call and calendar. Computer, 31(3):89–90, 92, 94–95, 98–99, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://pdf.computer.
org/co/books/co1998/pdf/r3089.pdf.

Anonymous:1998:COa

[1618] Anonymous. Career opportunities.

Computer, 31(3):100–105, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3100.pdf.

Anonymous:1998:NBb

[1619] Anonymous. New books: All about those enterprising JavaBeans.

Computer, 31(3):106, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3106.pdf.

Anonymous:1998:NTb

[1620] Anonymous. New tools. Computer, 31(3):108–113, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3108.pdf.

Severance:1998:SII

[1621] Charles Severance. Standards: IEEE 754: An interview with William Kahan. Computer, 31(3):114–115, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3114.pdf.

Bennis:1998:MNU

[1622] Warren Bennis and Patricia Ward Biederman. Management: None of us is as smart as all of us. Computer, 31(3):116–117, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3116.pdf.

Gil:1998:OTD

[1623] Joseph (Yossi) Gil and David H. Lorenz. Object technology: Design patterns and language design. Computer, 31(3):118–120, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3118.pdf.

Maurer:1998:IWW

[1624] Hermann Maurer. Internet watch:

Web-based knowledge management. Computer, 31(3):122–123, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3122.pdf.

Stein:1998:OCD

[1625] Richard Marlon Stein. Open channel: Does determinism dictate dignity? Computer, 31(3):124, March 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://pdf.computer.org/co/
books/co1998/pdf/r3124.pdf.

Lewis:1998:BCJa

[1626] Ted Lewis. Binary critic: Java Holy War ’98. Computer, 31(3):128, 126– 127, March 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://pdf.

computer.org/co/books/co1998/pdf/
r3128.pdf.

Anonymous:1998:LDC

[1627] Anonymous. Letter: Debunking cybermyths: The Stroustrup interview. Computer, 31(4):4, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4004.pdf. Notes

that a recent Computer ‘interview’ with C++ creator Bjarne Stroustrup is a hoax.

Heijink:1998:LIA

[1628] Rudolf Jan Heijink. Letter: ISDN available in The Netherlands. Computer, 31(4):4, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        dlib.computer.org/co/books/co1998/
                                        pdf/r4004.pdf.

Kubler:1998:LQM

[1629] Donald Kubler and Thomas Drake. Letter: Quality metrics don’t work. Computer, 31(4):4, 5, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r4004.pdf.

Steinhart:1998:LEE

[1630] Jon Steinhart. Letter: Electricalengineering vs. software design. Computer, 31(4):4, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        dlib.computer.org/co/books/co1998/
                                        pdf/r4004.pdf.

Parrish:1998:EME

[1631] Edward A. Parrish. Editor’s message; evolving a world-class membership magazine. Computer, 31(4):7, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r4007.pdf.

Lawton:1998:ITP

[1632] George Lawton. Industry trends: Paving the information superhighway’s last mile. Computer, 31(4):10–12, 14, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/r4010.pdf.

Clark:1998:TND

[1633] David Clark. Technology news:

Developments shake domain-name plan. Computer, 31(4):17, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4017.pdf.

Clark:1998:NBW

[1634] David Clark. News briefs; W3C approval gives XML a push; ISO adopts QuickTime for MPEG-4; group proposes antipiracy plan; delay for Gbit Ethernet specs. Computer, 31(4):18–19, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4018.pdf.

Anonymous:1998:ASd

[1635] Anonymous. Article summaries.

Computer, 31(4):20–21, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4020.pdf.

Wilkes:1998:RAE

[1636] Maurice V. Wilkes. A revisionist account of early language development. Computer, 31(4):22–25, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4022.pdf;
http://www.computer.org/computer/
co1998/r4022abs.htm.

Flynn:1998:CEY

[1637] Michael J. Flynn. Computer engineering 30 years after the IBM Model 91. Computer, 31(4):27–31, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4027.pdf;
http://www.computer.org/computer/
co1998/r4027abs.htm.

Yen:1998:TIM

[1638] I-Ling Yen, Raymond Paul, and Kinji Mori. Toward integrated methods for high-assurance systems. Computer, 31(4):32–34, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4032.pdf.

Yen:1998:KAH

[1639] I-Ling Yen and Ray Paul. Key applications for high-assurance systems. Computer, 31(4):35, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4035.pdf;
http://www.computer.org/computer/
co1998/r4035abs.htm.

Alkalai:1998:LLD

[1640] Leon Alkalai and Ann T. Tai. Long-life deep-space applications. Computer, 31(4):35–36, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4037.pdf.

Winter:1998:PSH

[1641] Victor L. Winter, John M. Covan, and Larry J. Dalton. Passive safety in high-consequence systems. Computer, 31(4):35–37, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r4035b.
pdf.

Harper:1998:CCA

[1642] Rick Harper and Barry Flahive. Challenges for continuously available systems. Computer, 31(4):39–40, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4039.pdf.

Tsai:1998:CSC

[1643] Wei-Tek Tsai, Ramin Mojdehbakhsh, and Sanjai Rayadurgam. Capturing safety-critical medical requirements. Computer, 31(4):40–42, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4040.pdf.

Mori:1998:ARC

[1644] Kinji Mori. Applications in rapidly changing environments. Computer, 31 (4):42–44, April 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1998/pdf/
r4042.pdf.

Lowry:1998:CBR

[1645] Michael R. Lowry. Component-based reconfigurable systems. Computer, 31 (4):44–45, April 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1998/pdf/
r4044.pdf.

Heimdahl:1998:SAI

[1646] Mats P. E. Heimdahl, Jeffrey M. Thompson, and Barbara J. Czerny. Specification and analysis of intercomponent communication. Computer, 31(4):47–54, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4047.pdf;
http://www.computer.org/computer/
co1998/r4047abs.htm.

Kang:1998:DAS

[1647] Myong H. Kang, Andrew P. Moore, and Ira S. Moskowitz. Design and assurance strategy for the NRL pump. Computer, 31(4):56–64, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4056.pdf;
http://www.computer.org/computer/
co1998/r4056abs.htm.

Khoshgoftaar:1998:UPH

[1648] Taghi M. Khoshgoftaar, Edward B. Allen, Robert Halstead, Gary P. Trio, and Ronald M. Flass. Using process history to predict software quality. Computer, 31(4):66–72, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4066.pdf;
http://www.computer.org/computer/
co1998/r4066abs.htm.

Anonymous:1998:CUD

[1649] Anonymous. CS update; draft software engineering accreditation criteria; Board of Governors approves reading of bylaws amendment on staff titles; Board of Governors approves reading of bylaws amendment on Membership Activities Board. Computer, 31(4):73–75, 77, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4073.pdf.

Westcott:1998:TAF

[1650] Douglas W. Westcott. Technical Activities Forum; designs for the computers of tomorrow. Computer, 31(4):76–77, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r4076.pdf.

Anonymous:1998:CCd

[1651] Anonymous. Call and calendar. Computer, 31(4):78, 80–81, 83, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4078.pdf.

Lutz:1998:NBB

[1652] Michael J. Lutz. New books:

Building databases for all media. Computer, 31(4):93, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4093.pdf.

Kroeker:1998:NTSa

[1653] Kirk L. Kroeker. New tools: Software development, hardware development, net development, component technology, systems analysis. Computer, 31(4):94–98, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r4094.pdf.

Gilbert:1998:IEI

[1654] John R. Gilbert. Integrated engineering: Integrating CAD tools for MEMS design. Computer, 31(4):99–101, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r4099.pdf.

Gray:1998:SRG

[1655] Lewis Gray. Software realities: Gray rebuts Bach: No cowboy programmers! Computer, 31(4):102–103, 105, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r4102.pdf.

Tredennick:1998:OCC

[1656] Nick Tredennick. Open channel: This cyberfarce stars Elvis and the Nerds. Computer, 31(4):104–105, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r4104.pdf.

Borchers:1998:IWG

[1657] Al Borchers, Jon Herlocker, Joseph Konstan, and John Riedl. Internet watch: Ganging up on information overload. Computer, 31(4):106–108, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/r4106.pdf.

Myers:1998:BCW

[1658] Ware Myers. Binary critic: Why software developers refuse to improve.

Computer, 31(4):112, 110–111, April 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r4112.pdf.

Brodine:1998:LTM

[1659] Dirk Brodine. Letters: Too much fun with system languages. Computer, 31(5):4, May 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1998/pdf/
r5004.pdf.

Ortiz:1998:ITH

[1660] Sixto Ortiz Jr. Industry trends:

Hardware-based networking widens the pipes. Computer, 31(5):8–9, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5008.pdf.

Clark:1998:TNA

[1661] David Clark. Technology news: Are ATM and Gigabit Ethernet ready for prime time? Computer, 31(5):11–13, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r5011.pdf.

Garber:1998:NBA

[1662] Lee Garber. News briefs: Auction sets stage for new broadband wireless service chip; makers delay move to 300mm wafer; are Europeans ready for the Euro?; W3C pushes net access for the disabled. Computer, 31(5):14–19, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5014.pdf.

Anonymous:1998:ASe

[1663] Anonymous. Article summaries. Computer, 31(5):18–19, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5018.pdf.

Zelkowitz:1998:CPE

[1664] Marvin V. Zelkowitz and Dolores Wallace. Computing practices: Experimental models for validating technology. Computer, 31(5):23–31, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5023.pdf;
http://www.computer.org/computer/
co1998/r5023abs.htm.

Tichy:1998:CSC

[1665] Walter F. Tichy. Cybersquare: Should computer scientists experiment more? Computer, 31(5):32–40, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5032.pdf;
http://www.computer.org/computer/
co1998/r5032abs.htm.

Bose:1998:CFP

[1666] Pradip Bose and Thomas M. Conte. Performance analysis and its impact on design. Computer, 31(5):41–49, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5041.pdf;
http://www.computer.org/computer/
co1998/r5041abs.htm.

Reilly:1998:CFP

[1667] Matt Reilly and John Edmondson. Performance simulation of an Alpha microprocessor. Computer, 31(5):50–58, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5050.pdf;
http://www.computer.org/computer/
co1998/r5050abs.htm.

Black:1998:CFC

[1668] Bryan Black and John Paul Shen. Calibration of microprocessor performance models. Computer, 31(5):59–65, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5059.pdf;
http://www.computer.org/computer/
co1998/r5059abs.htm.

Maly:1998:BHC

[1669] Kurt Maly, Ajay Gupta, and Satish Mynam. BTU: A host communication benchmark. Computer, 31(5):66–74, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r5066.pdf;
 http://www.computer.org/computer/
co1998/r5066abs.htm.

Ng:1998:NAD

[1670] Spencer W. Ng. News: Advances in disk technology: Performance issues. Computer, 31(5):75–81, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5075.pdf;
http://www.computer.org/computer/
co1998/r5075abs.htm.

Anonymous:1998:CUF

[1671] Anonymous. CS update; free service links members who share interests; initiative to support US undergraduate women; 1997 IEEE CS Student Award winners. Computer, 31(5):82–83, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r5082.pdf.

Anonymous:1998:CCe

[1672] Anonymous. Call and calendar. Computer, 31(5):84–86, 89–94, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r5084.pdf;
                                        http://www.computer.org/computer/
                                        co1998/issues.

Lutz:1998:NBS

[1673] Michael J. Lutz. New books: Software architecture defined. Computer, 31(5):95, May 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1998/pdf/
                                        r5095.pdf.

Kroeker:1998:NTSb

[1674] Kirk L. Kroeker. New tools: Software development, hardware development. Computer, 31(5):96–98, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r5096.pdf.

Severance:1998:SGC

[1675] Charles Severance. Standards: Grab the chance to work on the leading edge. Computer, 31(5):99, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5099.pdf.

McConnell:1998:MPP

[1676] Steve McConnell. Management:

The power of process. Computer, 31(5):100–102, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5100.pdf.

Glass:1998:OCS

[1677] Robert L. Glass. Open channel: Success? failure? or both? Computer, 31(5):103, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5103.pdf.

Meyer:1998:OTP

[1678] Bertrand Meyer, Christine Mingins, and Heinz Schmidt. Object technology: Providing trusted components to the industry. Computer, 31(5):104–105, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5104.pdf.

Turau:1998:IWW

[1679] Volker Turau. Internet watch: What practices are being adopted on the Web? Computer, 31(5):106–108, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5106.pdf.

Lewis:1998:BCWb

[1680] Ted Lewis. Binary critic: Why the economy is so good. Computer, 31(5):112, 110–111, May 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r5112.pdf.

Rosenwald:1998:LCP

[1681] Jeffrey A. Rosenwald, George Heilmeier, Bill Sneed, Bruce Walker, Igore Markov, Tom Hildebrandt, Brian O’Hearn, Alex B. Cannara, Akhatar Jameel, Matthias Stuempfle, Daniel Jiang, Axel Fuchs, Frank Preston, and Don Greenberg. Letters: Continuing the productivity paradox debat; post-2008 economic driver; “computer science” an oxymoron?; underestimating the cost of experiments; improving C++ runtime; safe Internet cars; dangerous concept; not that many transistors. Computer, 31(6):4–7, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6004.pdf.

Garber:1998:ITP

[1682] Lee Garber and David Sims. Industry trends: In pursuit of hardware-software codesign. Computer, 31(6):12–14, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6012.pdf.

Lawton:1998:TNI

[1683] George Lawton. Technology news: The Internet’s challenge to privacy. Computer, 31(6):16–18, June 1998.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6016.pdf.

Garber:1998:NBHa

[1684] Lee Garber. News briefs: HP enters the Java fray: Toward a more intelligent Web: Mobile IP on the move: Dispute over reported IT worker shortage. Computer, 31(6):19–21, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6019.pdf.

Anonymous:1998:ASf

[1685] Anonymous. Article summaries.

Computer, 31(6):22–23, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6022.pdf.

Wirth:1998:HCT

[1686] Niklaus Wirth. Hardware compilation: Translating programs into circuits. Computer, 31(6):25–31, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://computer.org/
computer/r6025abs.htm;   http://
dlib.computer.org/co/books/co1998/
pdf/r6025.pdf.

Jacob:1998:VMI

[1687] Bruce Jacob and Trevor Mudge.

Virtual memory: Issues of implementation. Computer, 31(6):33– 43, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

computer.org/computer/r6033abs.

                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r6033.pdf.

Voas:1998:CUC

[1688] Jeffrey M. Voas. The challenges of using COTS software in componentbased development. Computer, 31 (6):44–45, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        computer.org/computer/r6044abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r6044.pdf.

Talbert:1998:IJM

[1689] Nancy Talbert. Interview with John McDermid: The cost of COTS. Computer, 31(6):46–52, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /computer.org/computer/r6046abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r6046.pdf.

Voas:1998:CSS

[1690] Jeffrey M. Voas. Certifying off-the-shelf software components. Computer, 31(6):53–59, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1998/pdf/
                                        r6053.pdf.

Lindqvist:1998:MSR

[1691] Ulf Lindqvist and Erland Jonsson. A map of security risks associated with using COTS. Computer, 31 (6):60–66, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        computer.org/computer/r6060abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r6060.pdf.

Zhong:1998:SCC

[1692] Qun Zhong and Nigel Edwards.

Security control for COTS components. Computer, 31(6):67– 73, June 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

computer.org/computer/r6067abs.
htm;   http://dlib.computer.org/
co/books/co1998/pdf/r6067.pdf.

Anonymous:1998:CUCb

[1693] Anonymous. CS update: Computer Society recognizes pioneers in Central and Eastern Europe. Computer, 31(6):79–84, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6079.pdf.

Anonymous:1998:CCf

[1694] Anonymous. Call and calendar.

Computer, 31(6):84–87, 90–94, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6084.pdf.

Hurd:1998:SWD

[1695] John Hurd. Standards: Why does digital participate in standards? Computer, 31(6):95–96, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6095.pdf.

Lutz:1998:NBM

[1696] Michael J. Lutz. New books:

Making wavelets in the real world; protecting the OB; Diffie on privacy; digital design explained; Ecommerce du jour; seeing software; structures for XML; Smalltalk patterns. Computer, 31(6):97, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6097.pdf.

Anonymous:1998:NTSa

[1697] Anonymous. New tools: Software development: Uniscape’s internationalization library; Global Technologies’ Unix-to-NT solution; KAI’s multithreaded Java debugging tool; Price Systems’ parametric forecasting tool. Computer, 31(6):98, 102, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6098.pdf.

Anonymous:1998:NTH

[1698] Anonymous. New tools: Hardware development: JTAG’s flash memory tool; Accelerated Technology’s embedded development environment; Motorola’s reconfigurable embedded system technology. Computer, 31(6):99, 103, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6098.pdf.

Anonymous:1998:NTN

[1699] Anonymous. New tools: Net development: Lotus’ tools for Intranet/ Extranet development; Novera Software’s Java application server. Computer, 31(6):100, 103, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6098.pdf.

Anonymous:1998:NTC

[1700] Anonymous. New tools: Component technology: Rational’s componentsavvy visual tool; ICL’s Javabased ORBs. systems analysis: Network Communication’s network probe; Seagate’s server health monitor. Computer, 31(6):101, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6098.pdf.

Anonymous:1998:NTSb

[1701] Anonymous. New tools: Systems analysis: Network Communication’s network probe; Seagate’s server health monitor. Computer, 31(6):102, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6098.pdf.

Bollig:1998:IET

[1702] Scott Bollig and Dan Xiao. Integrated engineering: Throwing off the shackles of a legacy system. Computer, 31(6):104–106, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r6104.pdf.

Bach:1998:SRP

[1703] James Bach and Dave W. Smith. Software realities: Plans, lies, and videotape. Computer, 31(6):107–109, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r6107.pdf.

Stroustrup:1998:OCR

[1704] Bjarne Stroustrup and Scott Hamilton. Open channel: The real Stroustrup interview. Computer, 31(6):110–114, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/r6110.pdf.

Bork:1998:IWW

[1705] Alfred Bork and David R. Britton Jr. Internet watch: The Web is not yet suitable for learning. Computer, 31(6):115–116, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r6115.pdf.

Rodgers:1998:BCB

[1706] T. J. Rodgers. Binary critic: Building the US workforce, one engineer at a time. Computer, 31(6):120, 117, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r6120.pdf.

Bass:1998:LSD

[1707] Len Bass, Robert L. Glass, and Gary A. Hill. Letters: Success differs in research and academia. Computer, 31(7):4, 6, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r7004.pdf.

Phillips:1998:ITS

[1708] Barry Phillips. Industry trends:

Have storage area networks come of age? Computer, 31(7):10–12, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7010.pdf.

Lawton:1998:TNM

[1709] George Lawton. Technology news:

Multicasting: Will it transform the Internet? Computer, 31(7):13–15, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7013.pdf.

Garber:1998:NBM

[1710] Lee Garber. News briefs; moving 3D beyond VRML; US, Microsoft will square off on September 8; satellite processor failure cuts net, pager service; US seeks encryption standard; vendors unveil Bluetooth wireless technology; the end of MS-DOS. Computer, 31(7):18–21, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7018.pdf.

Anonymous:1998:ASg

[1711] Anonymous. Article summaries.

Computer, 31(7):22–23, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7022.pdf.

Dulong:1998:IAW

[1712] Carole Dulong. The IA-64 architecture at work. Computer, 31(7):24–32, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7024.pdf;
http://www.computer.org/computer/
co1998/r7024abs.htm.

Boehm:1998:UWS

[1713] Barry Boehm, Alexander Egyed, Julie Kwan, Dan Port, Archita Shah, and Ray Madachy. Using the Win-Win spiral model: A case study. Computer, 31(7):33–44, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7033.pdf;
http://www.computer.org/computer/
co1998/r7033abs.htm.

Lengyel:1998:CGV

[1714] Jed Lengyel. The convergence of graphics and vision. Computer, 31(7):46–53, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r7046.pdf;
 http://www.computer.org/computer/
co1998/r7046abs.htm.

McKee:1998:SMI

[1715] Sally A. McKee, Robert H. Klenke, Kenneth L. Wright, William A. Wulf, Maximo H. Salinas, James H. Aylor, and Alan P. Batson. Smarter memory: Improving bandwidth for streamed references. Computer, 31(7):54–63, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7054.pdf;
http://www.computer.org/computer/
co1998/r7054abs.htm.

Agrawal:1998:MST

[1716] Hira Agrawal, James L. Alberi, Joseph R. Horgan, J. Jenny Li, Saul London, W. Eric Wong, Sudipto Ghosh, and Norman Wilde. Mining system tests to aid software maintenance. Computer, 31(7):64–73, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7064.pdf;
http://www.computer.org/computer/
co1998/r7064abs.htm.

Anonymous:1998:COb

[1717] Anonymous. Career opportunities.

Computer, 31(7):74–76, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reed:1998:TAF

[1718] Karl Reed. Technical activities forum:

Why the CS should help chart the future of IT. Computer, 31(7):77–78, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7077.pdf.

Israel:1998:CUG

[1719] Michel Israel. CS update: Growth and stability: 1997 financial results. Computer, 31(7):79–80, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7079.pdf.

Coopers-Lybrand:1998:CUR

[1720] Coopers & Lybrand. CS update:

Report of independent accountants. Computer, 31(7):80–84, July 1998.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7079.pdf.

Burgess:1998:CUB

[1721] Angela Burgess and Carl K. Chang. CS update: Board of Governors approves slate for 1998 election. Computer, 31(7):85, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7079.pdf.

Anonymous:1998:CCg

[1722] Anonymous. Call and calendar. Computer, 31(7):85–87, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7085.pdf.

Lutz:1998:NBN

[1723] Michael J. Lutz. New books: Network management with scripts; Ethernet backbone; Web technologies; total cache knowledge; putting on the constraints; Java GUI details; doing business on the net; programming Netscape. Computer, 31(7):89, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7089.pdf.

Kroeker:1998:NTSc

[1724] Kirk L. Kroeker. New tools: Software development: TakeFive Software’s source code tool; RST’s test tools; Blue Sky’s help file authoring solution. hardware development: Green Hills Software’s single-source solution for tools and RTOS; Tanner ERA’s interconnect design tools; Xynetix Design Systems’ autorouter. net development: Aonix’s Web-to-legacy tool; Inktomi’s commercial network cache. component technology: Visual Edge’s enterprise application integration products; Rogue Wave Software’s suite of JavaBeans components. Computer, 31(7):90–93, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7090.pdf.

Severance:1998:SSR

[1725] Charles Severance. Standards: Standardizing real-time streaming protocols: Interview with Rob Lanphier, RealNetworks. Computer, 31(7):94–95, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r7094.pdf.

Meyer:1998:OTT

[1726] Bertrand Meyer. Object technology:

Tell less and say more: The power of implicitness. Computer, 31(7):97–98, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r7097.pdf.

Bassett:1998:MMF

[1727] Paul Bassett. Management: Managing for flexible software manufacturing. Computer, 31(7):100–102, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r7100.pdf.

Korpela:1998:IWL

[1728] Jukka Korpela. Internet watch: Lurching toward Babel: HTML and CSS and XML. Computer, 31(7):103–104, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r7103.pdf.

Holmes:1998:OCMa

[1729] Neville Holmes. Open channel: Mitigating Microsoft with virtual consoles. Computer, 31(7):105–106, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r7105.pdf.

Lewis:1998:BCJb

[1730] Ted Lewis. Binary critic: Joe Sixpack and Larry Lemming and Ralph Nader. Computer, 31(7):107–109, July 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r7107.pdf.

Black:1998:BCB

[1731] Dominique S. Black. Binary critic: Behind the hype about the IT labor and “shortage”. Computer, 31(7):112, 110–111, July 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1998/pdf/
                                        r7112.pdf.

Parrish:1998:RMMb

[1732] Ed Parrish, Tom Clune, Phil Armour, Jeffrey A. Rosenwald, Joachim Feise, Terry Vaughn, Frank Preson, R. G. (Ben) Bennetts, Jeffrey Van Myers, Robert A. Rivers, Ramesh Narayanaswamy, John Podlasek, Stig Nilsson, and Giovanni Jaramillo. Report to members; members see shades of truth in technology worker debate. Computer, 31(8):7–12, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8007.pdf.

Stein:1998:OCJ

[1733] Richard Marlon Stein. Open channel: T. J. Rodgers, H-1B visas, and me. Computer, 31(8):8, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8008.pdf.

Lawton:1998:ITB

[1734] George Lawton. Industry trends:

Biometrics: A new era in security. Computer, 31(8):16–18, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8016.pdf.

Ortiz:1998:TNA

[1735] Sixto Ortiz Jr. Technology news:

Active networks: The programmable pipeline. Computer, 31(8):19–21, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r8019.pdf.

Anonymous:1998:NBP

[1736] Anonymous. News briefs: Private domain-name system. FTC’s Intel case. Microsoft moves toward object interoperability. Computer, 31(8):22–25, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r8022.pdf.

Anonymous:1998:ASC

[1737] Anonymous. Article summaries: Cybersquare. Computer, 31(8):26–27, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/r8026.pdf.

Hill:1998:CPM

[1738] Mark D. Hill. Computing practices: Multiprocessors should support simple memory-consistency models. Computer, 31(8):28–34, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r8028.pdf;
                                        http://www.computer.org/computer/
                                        co1998/r8028abs.htm.

Long:1998:CFI

[1739] Earl Long, Amit Misra, and Janos Sztipanovits. Cover feature: Increasing productivity at Saturn. Computer, 31(8):35–43, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r8035.pdf;
                                        http://www.computer.org/computer/
                                        co1998/r8035abs.htm.

Schlett:1998:TEM

[1740] Manfred Schlett. Trends in embeddedmicroprocessor design. Computer, 31 (8):44–49, August 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1998/pdf/
r8044.pdf;   http://www.computer.
org/computer/co1998/r8044abs.htm.

Eyre:1998:DPH

[1741] Jennifer Eyre and Jeff Bier. DSP processors hit the mainstream. Computer, 31(8):51–59, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8051.pdf;
http://www.computer.org/computer/
co1998/r8051abs.htm.

Schulz:1998:RFM

[1742] Stephan Schulz, Jerzy W. Rozenblit, Michael Mrva, and Klaus Buchenrieder. Research features: Model-based codesign. Computer, 31(8):60–67, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r8060.pdf;
http://www.computer.org/computer/
co1998/r8060abs.htm.

Savor:1998:TAD

[1743] Tony Savor and Rudolph E. Seviora. Toward automatic detection of software failures. Computer, 31(8):68–74, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8068.pdf;
http://www.computer.org/computer/
co1998/r8068abs.htm.

Chen:1998:NIV

[1744] Hsinchun Chen, Jay Nunamaker, Jr., Richard Orwig, and Olga Titkova.

News: Information visualization for collaborative computing. Computer, 31(8):75–82, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8075.pdf;
http://www.computer.org/computer/
co1998/r8075abs.htm.

Anonymous:1998:CUA

[1745] Anonymous. CS update: Awards: Agerwala receives 1998 McDowell Award. Agrawal receives 1998 Goode Memorial Award. Wah, Ichikawa, and Bourbakis earn Technical Achievement Awards. TAB election results. science and engineering fair winners. Computer, 31(8):83–86, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8083.pdf.

Anonymous:1998:CCh

[1746] Anonymous. Call and calendar. Computer, 31(8):86–95, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8086.pdf.

Lutz:1998:NBL

[1747] Michael J. Lutz. New books: The legalities of inventing software. Computer, 31(8):102, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8102.pdf.

Lutz:1998:NTM

[1748] Michael J. Lutz. New tools; membership news. Computer, 31(8):103–107, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/r8103.pdf.

Holmes:1998:STD

[1749] Neville Holmes. Standards: Toward decent text encoding. Computer, 31(8):108–109, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8108.pdf.

Smith:1998:IES

[1750] Mark T. Smith. Integrated engineering: Smart cards: Integrating for portable complexity. Computer, 31(8):110–112, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8110.pdf.

Bach:1998:SRH

[1751] James Bach. Software realities:

The highs and lows of change control. Computer, 31(8):113–115, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8113.pdf.

Wilson:1998:IWI

[1752] Janet Wilson. Internet watch: The IETF: Laying the Net’s asphalt. Computer, 31(8):116–117, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r8116.pdf.

Napper:1998:BCE

[1753] Simon Napper. Binary critic: Embedded-system design plays catchup. Computer, 31(8):120, 118–119, August 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/r8120.pdf.

Covington:1998:L

[1754] Michael A. Covington, Ted Lewis, Hernan Astudillo, Jukka Korpela, Bruce Wilner, Nigel Edwards, and Qun Zhong. Letters. Computer, 31(9):4, 6, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r9004.pdf.

Parrish:1998:RMM

[1755] Edward A. Parrish. Report to members: Members react to privacy and encryption survey. Computer, 31(9):12–15, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r9012.pdf.

Goldberg:1998:TNA

[1756] Lee Goldberg. Technology news: The advent of ‘green’ computer design. Computer, 31(9):16–19, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/r9016.pdf.

Garber:1998:NBG

[1757] Lee Garber. News briefs: Global Internet summit: Home networking plan: Hyperlink Web-searching engine. Computer, 31(9):20–21, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9020.pdf.

Anonymous:1998:ASh

[1758] Anonymous. Article summaries. Computer, 31(9):22–23, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9022.pdf.

Dowd:1998:NST

[1759] Patrick W. Dowd and John T.

McHenry. Network security: It’s time to take it seriously. Computer, 31 (9):24–28, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/computer.org/computer/r9024abs.
htm;   http://dlib.computer.org/
co/books/co1998/pdf/r9024.pdf.

Schneier:1998:CDV

[1760] Bruce Schneier. Cryptographic design vulnerabilities. Computer, 31 (9):29–33, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

/computer.org/computer/r9029abs.
htm;   http://dlib.computer.org/
co/books/co1998/pdf/r9029.pdf;
http://www.counterpane.com/design-
vulnerabilities.pdf.

Rubin:1998:SWS

[1761] Aviel D. Rubin and Daniel E. Geer Jr.

A survey of Web security. Computer, 31(9):34–42, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /computer.org/computer/r9034abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r9034.pdf.

Oppliger:1998:SIL

[1762] Rolf Oppliger. Security at the Internet layer. Computer, 31(9):43– 47, September 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

                                        computer.org/computer/r9043abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r9043.pdf.

Arbaugh:1998:SVP

[1763] William A. Arbaugh, James R. Davin, David J. Farber, and Jonathan M. Smith. Security for virtual private intranets. Computer, 31(9): 48–56, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /computer.org/computer/r9048abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r9048.pdf.

Tarman:1998:AAE

[1764] Thomas D. Tarman, Robert L. Hutchinson, Lyndon G. Pierson, Peter E. Sholander, and Edward L. Witzke. Algorithm-agile encryption in ATM networks. Computer, 31 (9):57–64, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:/

                                        /computer.org/computer/r9057abs.
                                        htm;   http://dlib.computer.org/
                                        co/books/co1998/pdf/r9057.pdf.

Petitpierre:1998:SCL

[1765] Claude Petitpierre. Synchronous C++: A language for interactive applications. Computer, 31(9):65– 72, September 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

computer.org/computer/r9065abs.
htm;   http://dlib.computer.org/
co/books/co1998/pdf/r9065.pdf.

Bendz:1998:TAF

[1766] Diana J. Bendz. Technical activities forum: Electronics and the environment. Computer, 31(9):73, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9073.pdf.

Anonymous:1998:CCi

[1767] Anonymous. Call and calendar. Computer, 31(9):74–81, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9074.pdf.

Anonymous:1998:CEC

[1768] Anonymous. 1998 CS election: Candidate statements for office. Computer, 31(9):82–89, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9082.pdf.

Anonymous:1998:COc

[1769] Anonymous. Career opportunities.

Computer, 31(9):90–100, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9101.pdf.

Lutz:1998:NBPa

[1770] Michael J. Lutz. New books: Peering into the digital crystal ball. Computer, 31(9):101, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9101.pdf.

Kroeker:1998:NTSd

[1771] Kirk L. Kroeker. New tools: Software development: Rational Software’s Ada development environment; Pragmatix Software’s Meta-CASE tool; Wibu Systems’ copy protection tools. hardware development: Diab Data’s suite of analysis tools for embedded applications; Mustang Software’s message center tool; RSI’s AS/400 clientserver tool; SES’s client-server simulation tool; CACI Products’ network analysis tools. Computer, 31 (9):102–104, 106, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9102.pdf.

Resmer:1998:SIA

[1772] Mark Resmer. Standards: Internet architectures for learning. Computer, 31(9):105–106, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9105.pdf.

Cravener:1998:IWE

[1773] Patricia Cravener. Internet watch: Education on the Web: A rejoinder. Computer, 31(9):107–111, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9107.pdf.

Lewis:1998:BCWc

[1774] Ted Lewis. Binary critic: What to do about Microsoft. Computer, 31(9):112, 109–111, September 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/r9112.pdf.

Alfred:1998:LLS

[1775] Charlie Alfred, Ted Lewis, John O.

Goyo, Ted Lewis, Garth Klatt, and George M. Samaras. Letters: Lemmings and software quality; US Navy C2 rating; the imperfection of engineering; success and failure revisited. Computer, 31(10):4, 6, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/rx004.

Edwards:1998:ITCb

[1776] John Edwards. Industry trends: The changing face of freeware. Computer, 31(10):11–13, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx011.pdf.

Phillips:1998:TND

[1777] Barry Phillips. Technology news:

Designers: The browser war casualties. Computer, 31(10):14–16, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx014.pdf.

Lawton:1998:NBS

[1778] George Lawton. News briefs: Sun joins tetherless network fray; private doorbells: Solution to encryption controversy; Web extensions promise distributed computing; Wall Street conducts major Y2K test; Sun prepares to submit first Java standard; Netscape loses more ground in browser war. Computer, 31(10):17–19, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx017.pdf.

Phillips:1998:ASC

[1779] Barry Phillips. Article summaries: Computing practices. Computer, 31 (10):20–21, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co1998/
pdf/rx020.pdf.

McGhan:1998:CPP

[1780] Harlan McGhan and Mike O’Connor. Computing practices: PicoJava: A direct execution engine for Java bytecode. Computer, 31(10):22–30, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/rx022.pdf;
 http://www.computer.org/computer/
co1998/rx022abs.htm.

Goldin:1998:CBI

[1781] Daniel S. Goldin, Samuel L. Venneri, and Ahmed K. Noor. Cybersquare: Beyond incremental change. Computer, 31(10):31–39, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx031.pdf;
http://www.computer.org/computer/
co1998/rx031abs.htm.

Solomon:1998:CFW

[1782] David A. Solomon. Cover feature:

The Windows NT kernel architecture. Computer, 31(10):40–47, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx040.pdf;
http://www.computer.org/computer/
co1998/rx040abs.htm.

Cabrera:1998:CFA

[1783] Luis Felipe Cabrera, Brian Andrew, Kyle Peltonen, and Norbert Kusters. Cover feature: Advances in Windows NT storage management. Computer, 31(10):48–54, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx048.pdf;
http://www.computer.org/computer/
co1998/rx048abs.htm.

Gamache:1998:CFW

[1784] Rod Gamache, Rob Short, and Mike Massa. Cover feature: Windows NT clustering service. Computer, 31(10):55–62, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx055.pdf;
http://www.computer.org/computer/
co1998/rx055abs.htm.

Eick:1998:RFV

[1785] Stephen G. Eick. Research feature:

A visualization tool for Y2K. Computer, 31(10):63–69, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx063.pdf;
http://www.computer.org/computer/
co1998/rx063abs.htm.

Mukherjee:1998:RFM

[1786] Shubhendu S. Mukherjee and Mark D. Hill. Research feature: Making network interfaces less peripheral. Computer, 31(10):70–76, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx070.pdf;
http://www.computer.org/computer/
co1998/rx070abs.htm.

Bagrodia:1998:RFP

[1787] Rajive Bagrodia, Richard Meyer, Mineo Takai, Yu an Chen, Xiang Zeng, Jay Martin, and Ha Yoon Song News. Research feature: Parsec: A parallel simulation environment for complex systems. Computer, 31(10):77–85, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx077.pdf;
http://www.computer.org/computer/
co1998/rx077abs.htm.

Ranganathan:1998:TAF

[1788] N. Ranganathan. Technical activities forum: A forum for VLSI practitioners. Computer, 31(10):86, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx086.pdf.

Anonymous:1998:CCj

[1789] Anonymous. Call and calendar. Computer, 31(10):87–108, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx087.pdf.

Lutz:1998:NBPb

[1790] Michael J. Lutz. New books: Programming for high-end JavaBeans; Intel secrets; Ada and concurrency; datamodeling primer; technology and economics; dissecting a microprocessor; real-world ATM; when computers run amok. Computer, 31(10):109, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx109.pdf.

Kroeker:1998:NTSe

[1791] Kirk L. Kroeker. New tools: Software development: Rational Software’s development tools for Ada, Artlandia’s Mathematica graphics tool, ISE’s EiffelBase library available for free, Annasoft System’s Windows CE development tool. hardware development: White Mountain’s DSP tools, Analogy’s mixed-signal IC simulator, Dynamic Soft Analysis’ thermal analysis tool; net development: Framework Technologies’ project management system, CAI’s message broker, Gordian’s network redirector software; component technology: Black & White’s usage control software for multi-ORB CORBA implementations, IONA Technologies’ middleware development tools, GreenTree Technologies’ ActiveX controls, KL Group’s JavaBeans with automatic data binding.

Computer, 31(10):110–114, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rx110.pdf.

Zelnick:1998:SNT

[1792] Nate Zelnick. Standards: Nifty technology and nonconformance: The Web in crisis. Computer, 31(10):115–116, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/rx115.pdf.

Lawson:1998:IEB

[1793] Harold W. Lawson. Integrated engineering: From busyware to stableware. Computer, 31(10):117–119, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rx117.pdf.

Bosak:1998:IWM

[1794] Jon Bosak. Internet watch: Mediaindependent publishing: Four myths about XML. Computer, 31(10):120– 122, October 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1998/pdf/
                                        rx120.pdf.

Notkin:1998:OCL

[1795] David Notkin. Open channel: The lifeblood of our field. Computer, 31(10):123, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rx123.pdf.

Bach:1998:SRF

[1796] James Bach. Software realities: A framework for good enough testing. Computer, 31(10):124–127, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx124.pdf.

Departments:1998:BCS

[1797] Robert N. Britcher Departments. Binary critic: Software reliability: Numbers aren’t the whole story. Computer, 31(10):128, 127, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rx128.pdf.

Clinton:1998:LBM

[1798] Terry Clinton, Tom Parsons, Capers Jones, William Adams, Garth Klatt, Eric Haines, Ted Lewis, Philip Machanik, Stig Nilsson, Karl Reed, Howard R. Stearns, Neville Holmes, and John Brownie. Letters: The benefits of model-based integration; documentation is not green; picking on the overdog; buggy, slow windoze; there’s no such thing as free software; Linus’ law of open source development; bugfree development? no way; governmental IT planning and the Computer Society; text encoding questions; encoding the world’s languages. Computer, 31(11):4, 5–7, 11, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/ry004.pdf. Two

letters discuss Unicode and Multicode

Callahan:1998:ITM

[1799] John Callahan. Industry trends: Moving toward effortless networking. Computer, 31(11):12–15, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry012.pdf.

Koprowski:1998:TNE

[1800] Gene Koprowski. Technology news: Emerging uncertainty over IPv6. Computer, 31(11):16–17, 20, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry016.pdf.

Garber:1998:NBHb

[1801] Lee Garber. News briefs: HTTP nextgeneration; faster PCI bus; smartcard technology standard; new chip technologies. Computer, 31(11):18– 23, November 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1998/pdf/
                                        ry018.pdf.

Kozyrakis:1998:CND

[1802] Christoforos E. Kozyrakis and David A. Patterson. Cybersquare: A new direction for computer architecture research. Computer, 31(11):24–32, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry024.pdf;
                                        http://www.computer.org/computer/
                                        co1998/ry024abs.htm.

Barry:1998:CPS

[1803] Douglas Barry and Torsten Stanienda.

Computing practices: Solving the Java object storage problem. Computer, 31(11):33–41, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/ry033.pdf;
http://www.computer.org/computer/
co1998/ry033abs.htm.

Chien:1998:CFD

[1804] Andrew A. Chien, Mark D. Hill, and Shubhendu S. Mukherjee. Cover feature: Design challenges for highperformance network interfaces. Computer, 31(11):42–45, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/ry042.pdf;
http://www.computer.org/computer/
co1998/ry042abs.htm.

Walton:1998:HSD

[1805] Simon Walton, Anne Hutton, and Joe Touch. High-speed data paths in hostbased routers. Computer, 31(11):46– 52, November 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1998/pdf/
ry046.pdf;   http://www.computer.
org/computer/co1998/ry046abs.htm.

Bhoedjang:1998:ULN

[1806] Raoul A. F. Bhoedjang, Tim Rühl, and Henri E. Bal. User-level network interface protocols. Computer, 31(11):53–60, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry053.pdf;
                                        http://www.computer.org/computer/
                                        co1998/ry053abs.htm.

vonEicken:1998:EVI

[1807] Thorsten von Eicken and Werner Vogels. Evolution of the Virtual Interface Architecture. Computer, 31(11):61–68, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry061.pdf;
                                        http://www.computer.org/computer/
                                        co1998/ry061abs.htm.

Lee:1998:EPM

[1808] Whay Sing Lee, William J. Dally, Stephen W. Keckler, Nicholas P. Carter, and Andrew Chang. An efficient, protected message interface. Computer, 31(11):69–75, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry069.pdf;
                                        http://www.computer.org/computer/
                                        co1998/ry069abs.htm.

Douglis:1998:TAF

[1809] Fred Douglis, Steve Chapin, and Jim Isaak. Technical activities forum: Internet research on Internet time. Computer, 31(11):76–77, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry076.pdf.

Anonymous:1998:CCk

[1810] Anonymous. Call and calendar. Computer, 31(11):78–111, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/ry078.pdf.

Lutz:1998:NBD

[1811] Michael J. Lutz. New books: Doing business with use cases. Computer, 31(11):112, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/ry112.pdf.

Kroeker:1998:NTSf

[1812] Kirk L. Kroeker. New tools: Software development: Galorath’s parametric estimation tool; Uniscape’s multilingual API and runtime library; Object International’s UML modeler; ObjecTime Limited’s real-time code generation tool. hardware development: Diab Industry’s Java compiler for embedded systems; Verilog’s embedded systems tool; Intusoft’s simulationbased test synthesis tool. net development: Rational Software’s Web-based load-testing tool; IONA Technologies’ Internet security tool; Sonic Systems’ ICSA-certified low-cost firewall. Computer, 31(11):113–116, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/ry113.pdf.

Severance:1998:SFF

[1813] Charles Severance. Standards:

Firewire finally comes home. Computer, 31(11):117–118, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry117.pdf.

Turau:1998:IWF

[1814] Volker Turau. Internet watch: The Fortune 500 Web. Computer, 31(11):119–120, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry119.pdf.

Holmes:1998:OCMb

[1815] Neville Holmes. Open channel: The myth of the computer revolution. Computer, 31(11):121–122, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry121.pdf.

Meyer:1998:OTR

[1816] Bertrand Meyer. Object technology: The role of object-oriented metrics. Computer, 31(11):123–125, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry123.pdf.

Lewis:1998:BCL

[1817] Ted Lewis. Binary critic: The legacy maturity model. Computer, 31(11):128, 125–127, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/ry128.pdf.

Phoha:1998:LLI

[1818] Vir V. Phoha, Norman S. Clerman, William R. Duncan, Daniele Farnedi, and Bryan Kelly. Letters: Limitations of IPv4 ensure future of IPv6; proprietary product coverage; what Brooks’ Law says; upgrading with COTS; NT storage management. Computer, 31(12):4, 6, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz004.pdf. See

correction [1870] about the author of the Brooks Law letter.

Anonymous:1998:RMMb

[1819] Anonymous. Report to members: Members elect officers and board members for 1999. Computer, 31(12):10–11, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz010.pdf.

Garber:1998:ITU

[1820] Lee Garber. Industry trends: US will grant more high-tech visas; Internet pioneer Jon Postel dies. Computer, 31(12):15–16, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz015.pdf.

Clark:1998:TNH

[1821] David Clark. Technology news: Heavy traffic drives networks to IP over Sonet. Computer, 31(12):17–20, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz017.pdf.

Lawgon:1998:NBA

[1822] George Lawgon. News briefs: Alliance wants to give voice to the net; Intel pushes for Unix standards; groups seek to end browser wars; initiative hopes to drive basic chip research; HTML+TIME promises better multimedia; keyword browsing promises simple Web access. Computer, 31(12):21–23, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz021.pdf.

Irvine:1998:ISC

[1823] Cynthia E. Irvine, Shiu-Kai Chin, and Deborah Frincke. Integrating security into the curriculum. Computer, 31(12):25–30, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz025.pdf;
                                        http://www.computer.org/computer/
                                        co1998/rz025abs.htm.

Falchuk:1998:VMA

[1824] Benjamin Falchuk and Ahmed Karmouch. Visual modeling for agentbased applications. Computer, 31(12): 31–38, December 1998. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1998/pdf/
                                        rz031.pdf;   http://www.computer.
                                        org/computer/co1998/rz031abs.htm.

Singh:1998:ACL

[1825] Munindar P. Singh. Agent communication languages: Rethinking the principles. Computer, 31(12):40–47, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/rz040.pdf;
 http://www.computer.org/computer/
co1998/rz040abs.htm.

Brereton:1998:HNM

[1826] Pearl Brereton, David Budgen, and Geoff Hamilton. Hypertext: The next maintenance mountain. Computer, 31(12):49–55, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz049.pdf;
http://www.computer.org/computer/
co1998/rz049abs.htm.

Anonymous:1998:AIV

[1827] Anonymous. Annual index: Volume 31, 1998. Computer, 31(12):56–67, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1998/pdf/rz056.pdf.

Sztipanovits:1998:TAF

[1828] Janos Sztipanovits. Technical activities forum: Integrated engineering of computer-based systems. Computer, 31(12):68–69, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz068.pdf.

Piner:1998:CUW

[1829] Mary-Louise G. Piner. CS update; WW II cryptologist [Jack Good] receives Pioneer Award; Colossus builder Tommy Flowers (1905–1998); DCI founder [George Schussel] receives Entrepreneur Award; 1998 Bell Prize recognizes advances in parallel processing.

Computer, 31(12):70–73, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz070.pdf.

Parrish:1998:CR

[1830] Edward A. Parrish. Computer reviewers, 1998. Computer, 31(12):73–76, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/rz077.pdf.

Anonymous:1998:CCl

[1831] Anonymous. Call and calendar. Computer, 31(12):77–79, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz077.pdf.

Anonymous:1998:COd

[1832] Anonymous. Career opportunities. Computer, 31(12):80–116, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz080.pdf.

Lutz:1998:NBPc

[1833] Michael J. Lutz. New books: Placing user needs first, technology last. Computer, 31(12):117, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz117.pdf.

Kroeker:1998:NTSg

[1834] Kirk L. Kroeker. New tools: Software development: Genitor’s C/C++ tools; Distributive Data Systems’ metric tool for enterprise-wide measurement. Computer, 31(12):118, 121, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz118.pdf.

Kroeker:1998:NTH

[1835] Kirk L. Kroeker. New tools: Hardware development: Green Hills’ development environment for ARM processors; Metrowerks’ embedded GUI framework. Computer, 31(12):119, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz118.pdf.

Kroeker:1998:NTN

[1836] Kirk L. Kroeker. New tools:

Net development: IONA Technologies’ middleware messaging system; Banyan Software’s directory-enabled user management system. Computer, 31(12):120, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz118.pdf.

Kroeker:1998:NTC

[1837] Kirk L. Kroeker. New tools: Component technology: 4Developers’ tool for managing ActiveX and COM files; SoftWired’s Java-based object delivery API. Computer, 31(12):121, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz118.pdf.

Carpenter:1998:IEA

[1838] Todd Carpenter. Integrated engineering: Avionics integration for CNS… ATM. Computer, 31(12):124–126, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/rz124.pdf.

Rothman:1998:SRC

[1839] Johanna Rothman. Software realities: Of crazy numbers and release criteria. Computer, 31(12):127–128, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz127.pdf.

Isaak:1998:SRG

[1840] Jim Isaak. Standards: The role of government in IT standards. Computer, 31(12):129, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz129.pdf.

Monticino:1998:IWW

[1841] Michael Monticino. Internet watch: Web analysis: Stripping away the hype. Computer, 31(12):130–132, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1998/pdf/rz130.pdf.

Tracz:1998:OCC

[1842] Will Tracz. Open channel: Changing the channel. Computer, 31(12):133, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1998/pdf/rz133.pdf.

Halla:1998:BCH

[1843] Brian Halla. Binary critic: How the PC will disappear. Computer, 31(12):136, 134–135, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1998/pdf/rz136.pdf.

Aylor:1999:EMC

[1844] James H. Aylor. Editor’s message: Computer for the 21st Century. Computer, 32(1):4, 6, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1004.pdf.

Tripp:1999:PML

[1845] Leonard L. Tripp. President’s message: Lengthening our stride: Our 1999 plan. Computer, 32(1):7–8, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1007.pdf.

Li:1999:LWS

[1846] Ambrose Li, Justus Pendelton, Norman Hines, Wayne Ding, Phil Crooker, Jim Aylor, Dilip K. Banerji, Lexis H. Rivera-Rios, Steve Masticola, and Ted Lewis. Letters: When is software free and when is it freeware? Computer, 32(1):9–11, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1009.pdf.

Charles:1999:ITN

[1847] John Charles. Industry trends: Neural interfaces link the mind and the machine. Computer, 32(1):16–18, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1016.pdf.

Edwards:1999:TNE

[1848] John Edwards. Technology news: Employment 1999: Opportunities amid challenges. Computer, 32(1):19–22, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r1019.pdf.

Garber:1999:NBA

[1849] Lee Garber. News briefs: Apple engineers a major turnaround; Java battles brew on several fronts; IBM and SGI unveil ultrafast supercomputers; Y2K briefs. Computer, 32(1):23–25, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1023.pdf.

Goth:1999:YBC

[1850] Greg Goth. Y2K in brief: Concern rising about Y2KA (the Y2K Aftermath). Computer, 32(1):26, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1023.pdf.

Anonymous:1999:ASa

[1851] Anonymous. Article summaries. Computer, 32(1):28–29, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1028.pdf.

Kavi:1999:CSR

[1852] Krishna Kavi, James Browne, and Anand Tripathi. Computer systems research: The pressure is on. Computer, 32(1):30–39, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1030.pdf.

Smith:1999:PNS

[1853] Jonathan M. Smith. Programmable networks: Selected challenges in computer networking. Computer, 32(1): 40–42, January 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r1040.pdf;   http://www.computer.
org/computer/co1999/r1040abs.htm.

Hamilton:1999:SRC

[1854] Scott Hamilton. Semiconductor research corporation; taking Moore’s Law into the next century. Computer, 32(1):43–50, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1043.pdf;
http://www.computer.org/computer/
co1999/r1043abs.htm.

Rubin:1999:DYC

[1855] Howard Rubin. Diary of a Y2K consultant: Bracing for the millennium. Computer, 32(1):51–56, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1051.pdf;
http://www.computer.org/computer/
co1999/r1051abs.htm.

Billinghurst:1999:WDN

[1856] Mark Billinghurst and Thad Starner. Wearable devices: New ways to manage information. Computer, 32(1): 57–64, January 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1999/pdf/
                                        r1057.pdf;   http://www.computer.
                                        org/computer/co1999/r1057abs.htm.

Schilit:1999:NWM

[1857] Bill N. Schilit, Morgan N. Price, Gene Golovchinsky, Kei Tanaka, and Catherine C. Marshall. News: As we may read: The reading appliance revolution. Computer, 32(1):65–73, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r1065.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r1065abs.htm.

Anonymous:1999:IEC

[1858] Anonymous. IEEE engineering of computer based systems symposium and workstations (ECBS99): Advance tutorial and technical program. Computer, 32(1):74–75, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r1076.pdf.

Voas:1999:TAF

[1859] Jeffrey M. Voas. Technical activities forum: The future of computer assurance. Computer, 32(1):76–77, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r1076.pdf.

Piner:1999:CSCa

[1860] Mary-Louise G. Piner. Computer Society connection; activities and opportunities: 1999 outlook; volunteer opportunities; impact event; bylaws amendment; Parrish receives IEEE service award. Computer, 32(1):78–83, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1078.pdf.

Anonymous:1999:CCa

[1861] Anonymous. Call and calendar. Computer, 32(1):84–89, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1084.pdf.

Lutz:1999:NBM

[1862] Michael J. Lutz. New books:

Minding your database manners. Computer, 32(1):124, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1124.pdf.

Kroeker:1999:NTT

[1863] Kirk L. Kroeker. New tools: Today Comdex, tomorrow the world. Computer, 32(1):125–130, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1125.pdf.

Severance:1999:SPM

[1864] Charles Severance. Standards; Posix:

A model for future computing. Computer, 32(1):131–132, January 1999.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1131.pdf.

Choi:1999:IWF

[1865] Soon-Tong Choi and Andrew B. Whinston. Internet watch: The future of E-commerce: Integrate and customize. Computer, 32(1):133–134, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1133.pdf.

Boehm:1999:MCI

[1866] Barry Boehm and Chris Abts. Management: COTS integration: Plug and pray? Computer, 32(1):135–138, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r1135.pdf.

Meyer:1999:COTa

[1867] Bertrand Meyer. Component and object technology: On to components. Computer, 32(1):139–140, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1139.pdf.

Lewis:1999:BCFa

[1868] Ted Lewis and Benjamin C. Fuller. Binary critic: Fast-lane browsers put the Web on wheels. Computer, 32(1):144, 141–143, January 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r1144.pdf.

Montgomery:1999:LYL

[1869] Warren Montgomery, Hugh R. Kress, Hannes Lubich, Michael Scheffler, and Norman Lorrain. Letters: Y2K lessons for developers; Y2K is not the millennium; what will drive IPv6 adoption?; friendly freeware service; layers of common sense. Computer, 32(2):4, 6, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2004.pdf.

Anonymous:1999:Ca

[1870] Anonymous. Correction. Computer, 32(2):6, February 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/

r2004.pdf. See [1818].

Ortiz:1999:ITN

[1871] Sixto Ortiz Jr. Industry trends:

New chips move networking onto silicon. Computer, 32(2):7–9, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2007.pdf.

Lawton:1999:TNV

[1872] George Lawton. Technology news:

Vendors battle over mobile OS market. Computer, 32(2):13–15, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2013.pdf.

Garber:1999:NBN

[1873] Lee Garber. News briefs: Net access becomes key factor in selling PCs; universal networking language; new standard for expanding IP telephony; Y2K report. Computer, 32(2):16–18, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2016.pdf.

Goth:1999:YBT

[1874] Greg Goth. Y2K briefs: UN takes Y2K action. Computer, 32(2):19, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2019.pdf.

Rubin:1999:RYDb

[1875] Howard Rubin and James Sanders. Rubin’s Y2K diary: Mantras, blacklists, and precedents. Computer, 32 (2):20, February 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r2020.pdf.

Anonymous:1999:ASb

[1876] Anonymous. Article summaries. Computer, 32(2):22–23, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2022.pdf.

Srinivasan:1999:DPO

[1877] Savitha Srinivasan. Design patterns in object-oriented frameworks. Computer, 32(2):24–32, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2024.pdf;

http://www.computer.org/computer/
co1999/r2024abs.htm.

Hughes:1999:WIC

[1878] Thomas P. Hughes and Jerry R.

Sheehan. What has influenced computing innovation? Computer, 32(2):33–43, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2033.pdf;
http://www.computer.org/computer/
co1999/r2033abs.htm.

Schatz:1999:DLT

[1879] Bruce Schatz and Hsinchun Chen.

Digital libraries: Technological advances and social impacts. Computer, 32(2):45–50, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2045.pdf.

Schatz:1999:FSS

[1880] Bruce Schatz, William Mischo, Timothy Cole, Ann Bishop, Susan Harum, Eric Johnson, Laura Neumann, Hsinchun Chen, and Dorbin Ng. Federated search of scientific literature. Computer, 32(2):51–59, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2051.pdf;
http://www.computer.org/computer/
co1999/r2051abs.htm.

Thomas:1999:TCJ

[1881] Spencer W. Thomas, Ken Alexander, and Kevin Guthrie. Technology choices for the JSTOR online archive.

Computer, 32(2):60–65, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2060.pdf;
http://www.computer.org/computer/
co1999/r2060abs.htm.

Wactlar:1999:LLB

[1882] Howard D. Wactlar, Michael G. Christel, Yihong Gong, and Alexander G. Hauptmann. Lessons learned from building a terabyte digital video library. Computer, 32(2):66–73, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r2066.pdf;
 http://www.computer.org/computer/
co1999/r2066abs.htm.

Witten:1999:MCD

[1883] Ian H. Witten, Rodger J. McNab, Steve Jones, Mark Apperley, David Bainbridge, and Sally Jo Cunningham. Managing complexity in a distributed digital library. Computer, 32(2):74–79, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2074.pdf;
http://www.computer.org/computer/
co1999/r2074abs.htm.

Paepcke:1999:UDO

[1884] Andreas Paepcke, Michelle Q. Wang Baldonado, Chen-Chuan K. Chang, Steve Cousins, and Hector Garcia-Molina. Using distributed objects to build the Stanford Digital Library Infobus. Computer, 32(2):80–87, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r2080.pdf;
 http://www.computer.org/computer/
co1999/r2080abs.htm.

Piner:1999:CSCb

[1885] Mary-Louise G. Piner. Computer Society connection: International design competition; 1999 fellows; officer appointments. Computer, 32(2):88–93, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2088.pdf.

Anonymous:1999:CCb

[1886] Anonymous. Call and calendar. Computer, 32(2):94–97, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2094.pdf.

Anonymous:1999:COb

[1887] Anonymous. Career opportunities.

Computer, 32(2):98–109, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2098.pdf.

Lutz:1999:NBL

[1888] Michael J. Lutz. New books: Learning from our mistakes. Computer, 32(2):110, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2110.pdf.

Kroeker:1999:NTSb

[1889] Kirk L. Kroeker. New tools: Software development; net development; hardware development. Computer, 32(2):111–113, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r2111.pdf.

Kessler:1999:IWW

[1890] Gary C. Kessler, Ken Rosenblad, and Steven D. Shepard. Internet watch: The Web can be suitable for learning. Computer, 32(2):114–115, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r2114.pdf.

Fleischmann:1999:IEP

[1891] Josef Fleischmann and Klaus Buchenrieder. Integrated engineering: Prototyping networked embedded systems. Computer, 32(2):116–119, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r2116.pdf.

Bach:1999:SRRa

[1892] James Bach. Software realities: Reframing requirements analysis. Computer, 32(2):120–122, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r2120.pdf.

Williams:1999:OCI

[1893] T. W. Williams. Open channel: IEEE-USA and the issue of member choice.

Computer, 32(2):123–124, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2123.pdf.

Lewis:1999:BCO

[1894] Ted Lewis. Binary critic: The open source acid test. Computer, 32(2):128, 125–127, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r2128.pdf. See

continuation in [1942].

Kephart:1999:LLF

[1895] Gary Kephart, David J. Pfaltzgraff, William B. Adams, Stephen Blanchette Jr., and Barry Boehm. Letters: Looking forward can hurt your brain; alone on the road less traveled; what’s new about COTS headaches? Computer, 32(3):4, 6, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3004.pdf.

Lawton:1999:ITN

[1896] George Lawton. Industry trends:

New technologies take the network home. Computer, 32(3):11–14, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3011.pdf.

Leavitt:1999:TNW

[1897] Neal Leavitt. Technology news:

Will 1999 be the year of IP telephony? Computer, 32(3):15–17, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3015.pdf.

Goth:1999:NBG

[1898] Greg Goth. News briefs: Groups duel over new I/O standards; Oracle, Sun team up to dump the OS; bringing the net to mobile appliances; Irish girl invents new E-mail encryption; group releases draft biometric-API specs; Toshiba announces smallest memory chip. Computer, 32(3):18–20, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3018.pdf.

Goth:1999:YBUa

[1899] Greg Goth. Y2K in brief: US corporations’ Y2K efforts go global. Computer, 32(3):21, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3021.pdf.

Rubin:1999:RYDc

[1900] Howard Rubin and James Sanders. Rubin’s Y2K diary: Doing the Y2K spin: Is image everything? Computer, 32(3):22–23, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3022.pdf.

Anonymous:1999:ASc

[1901] Anonymous. Article summaries. Computer, 32(3):24–26, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3024.pdf.

Krol:1999:CWW

[1902] Marina Krol. Cybersquare: Have we witnessed a real-life Turing test? Computer, 32(3):27–30, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3027.pdf;
http://www.computer.org/computer/
co1999/r3027abs.htm.

Bhargava:1999:CPB

[1903] Hemant K. Bhargava, Suresh Sridhar, and Craig Herrick. Computing practices: Beyond spreadsheets: Tools for building decision support systems. Computer, 32(3):31–39, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3031.pdf;
http://www.computer.org/computer/
co1999/r3031abs.htm.

Pathak:1999:GEI

[1904] Girish Pathak and Shri K. Goyal.

Guest Editors’ introduction: Transforming business through information technology. Computer, 32(3):40–41, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r3040.pdf.

Srinivasan:1999:CRI

[1905] Krishnamurthy Srinivasan and Sundaresan Jayaraman. The changing role of information technology in manufacturing. Computer, 32(3):42–49, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r3042.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r3042abs.htm.

Brandau:1999:RGI

[1906] Richard Brandau, Tony Confrey, Alin D’Silva, Christopher Matheus, and Robert Weihmayer. Reinventing GTE with information technology. Computer, 32(3):50–58, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r3050.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r3050abs.htm.

Norris:1999:BII

[1907] Mark Norris, Paul Muschamp, and Steve Sim. The BT intranet: Information by design. Computer, 32(3):59–66, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r3059.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r3059abs.htm.

Jutla:1999:MBS

[1908] Dawn Jutla, Peter Bodorik, Catherine Hajnal, and Charles Davis. Making business sense of electronic commerce. Computer, 32(3):67–75, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r3067.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r3067abs.htm.

Piner:1999:CSCc

[1909] Mary-Louise G. Piner. Computer society connection; society activities attract new members; two presented technical achievement awards; Michael C. Mulder tribute. Computer, 32(3):76–79, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3076.pdf.

Kaner:1999:YCW

[1910] Cem Kaner, Robert Colwell, Richard Stern, and Kevin W. Bowyer. Your call: When a product doesn’t measure up to its claims. Computer, 32(3):80–82, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3080.pdf.

Anonymous:1999:CCc

[1911] Anonymous. Call and calendar. Computer, 32(3):83–84, 86–89, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3083.pdf.

Anonymous:1999:Cb

[1912] Anonymous.  Careers.  Com-

puter, 32(3):92–102, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3092.pdf.

Lutz:1999:NBY

[1913] Michael J. Lutz. New books:

Are you listening, Dave?; cultural team chemistry; Java image processing; Y2K preparation; making distributed objects work for COM-CORBA; fuzzy system use. Computer, 32(3):103, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3103.pdf.

Kroeker:1999:NTSc

[1914] Kirk L. Kroeker. New tools: Software development; net development; hardware development. Computer, 32(3):104–107, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3104.pdf.

Schneier:1999:IWC

[1915] Bruce Schneier. Internet watch: Cryptography: The importance of not being different. Computer, 32(3):108–109, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r3108.pdf.

Lawson:1999:SDS

[1916] Harold W. Lawson. Standards: Defining stakeholder relationships. Computer, 32(3):110–112, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3110.pdf.

Boehm:1999:MMR

[1917] Barry Boehm. Management: Making RAD work for your project. Computer, 32(3):113–114, March 1999.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3113.pdf.

Dubois:1999:COT

[1918] Paul F. Dubois. Component and object technology: Scientific components are coming. Computer, 32(3):115–117, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r3115.pdf.

Lewis:1999:BCI

[1919] Ted Lewis. Binary critic: Innovation in the small. Computer, 32(3):120, 118–119, March 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r3120.pdf.

Slimick:1999:LCI

[1920] John Slimick and T. W. Williams. Letters: In crediting innovation, don’t forget entrepreneurs. Computer, 32(4):4, 6, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r4004.pdf.

Garber:1999:ITS

[1921] Lee Garber. Industry trends: Steve Deering on IP Next Generation. Computer, 32(4):11–13, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4011.pdf.

Greenberg:1999:TNF

[1922] Ilan Greenberg. Technology news: Facing up to new interfaces. Computer, 32 (4):14–16, April 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r4014.pdf.

Goth:1999:NBX

[1923] Greg Goth and George Lawton. News briefs: XML specs duel for E-commerce; Pentium III privacy flap continues; new chips promise speed, adaptability; researchers work on interactive agents; groups work on standards for storage networks. Computer, 32(4):17–19, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4017.pdf.

Goth:1999:YRY

[1924] Greg Goth. Y2K report: Y2K in brief; groups help Third World fight Y2K problem; lawmakers seek Y2K immunity; Russian Y2K effort hampered by low funding. Computer, 32(4):20, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4020.pdf.

Rubin:1999:RYDd

[1925] Howard Rubin and James Sanders. Rubin’s Y2K diary: Getting down and dirty. Computer, 32(4):21, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4021.pdf.

Anonymous:1999:ASd

[1926] Anonymous. Article summaries.

Computer, 32(4):22–23, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4022.pdf.

Nichols:1999:ASP

[1927] Kenneth Nichols. The age of software patents. Computer, 32(4):25–31, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4025.pdf;
http://www.computer.org/computer/
co1999/r4025abs.htm.

Smith:1999:ANP

[1928] Jonathan M. Smith, Kenneth L.

Calvert, Sandra L. Murphy, Hilarie K. Orman, and Larry L. Peterson. Activating networks: A progress report. Computer, 32(4):32–41, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4032.pdf;
http://www.computer.org/computer/
co1999/r4032abs.htm.

Wetherall:1999:ANS

[1929] David Wetherall, John Guttag, and David Tennenhouse. ANTS: Network services without the red tape. Computer, 32(4):42–48, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4042.pdf;
http://www.computer.org/computer/
co1999/r4042abs.htm.

Hartman:1999:JPL

[1930] John J. Hartman, Peter A. Bigot, Patrick Bridges, Brady Montz, Rob Piltz, Oliver Spatscheck, Todd A. Proebsting, Larry L. Peterson, and Andy Bavier. Joust: A platform for liquid software. Computer, 32(4):50–56, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/r4050.pdf;
                                         http://www.computer.org/computer/
                                        co1999/r4050abs.htm.

Jajodia:1999:SIW

[1931] Sushil Jajodia, Paul Ammann, and Catherine D. McCollum. Surviving information warfare attacks. Computer, 32(4):57–63, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r4057.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r4057abs.htm.

Fan:1999:WBF

[1932] Ming Fan, Jan Stallaert, and Andrew B. Whinston NEWS. A Web-based financial trading system. Computer, 32(4):64–70, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r4064.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r4064abs.htm.

Piner:1999:CSCd

[1933] Mary-Louise G. Piner. Computer Society connection: Computer Society tool a finalist for Codie award; program brings expert tutorials to CS chapters; any member can start a chapter; read the Raymond interview online; Computer Society honors seven outstanding students; international science and engineering fair. Computer, 32(4):72–75, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4072.pdf.

Anonymous:1999:CCd

[1934] Anonymous. Call and calendar. Computer, 32(4):76–78, 80, 82–84, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4076.pdf.

Anonymous:1999:COd

[1935] Anonymous. Career opportunities.

Computer, 32(4):86–92, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4086.pdf.

Lutz:1999:NBSa

[1936] Michael J. Lutz. New books:

Software engineering success stories. Computer, 32(4):93, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4093.pdf.

Kroeker:1999:NTC

[1937] Kirk L. Kroeker. New tools: Component technology: Harlequin’s OO environment for enterprise development; Tom Sawyer Software’s Graph Editor Toolkit for ActiveX. net development: GoAhead Software’s remote upgrade software; Blue Lobster’s legacy-to-Web application. software development: Mozart Consortium’s development platform; InstallShield’s Windows CE installation tool. Computer, 32(4):94–96, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r4094.pdf.

Burch:1999:IWM

[1938] Hal Burch and Bill Cheswick. Internet watch: Mapping the Internet. Computer, 32(4):97–98, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r4097.pdf.

Czapiewski:1999:IEI

[1939] Piotr Czapiewski, Rajaraman Ganesan, Dean Volkert, and Bruce Weber. Integrated engineering: Integrated, computer-aided order fulfillment. Computer, 32(4):99–102, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r4099.pdf.

McGraw:1999:SRS

[1940] Gary McGraw. Software realities: Software assurance for security. Computer, 32(4):103–105, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r4103.pdf.

Kostek:1999:OCI

[1941] Paul J. Kostek. Open channel: IEEE-USA and technical professionalism. Computer, 32(4):106–107, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4106.pdf.

Lewis:1999:BCA

[1942] Ted Lewis. Binary critic: Asbestos pajamas: An open source dialogue. Computer, 32(4):112, 108–111, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r4112.pdf. Con-

tinues the debate about open source software development, notably in GNU/Linux [1894].

Gimlan:1999:LRE

[1943] Gideon Gimlan, Rex Page, Bruce E.

Hayden, and David A. E. Wall. Letters: Relax, not everyone is out to sue you; patent education is in order; the claims hold the key; the state of software reuse. Computer, 32(5):4–10, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5004.pdf.

Piner:1999:RMT

[1944] Mary-Louise G. Piner. Report to members: With two new awards, we honor Unix, RISC innovators. Computer, 32(5):11–13, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5011.pdf.

Edwards:1999:ITC

[1945] John Edwards. Industry trends: Convergence reshapes the networking industry. Computer, 32(5):14–16, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r5014.pdf.

Charles:1999:TNM

[1946] John Charles. Technology news: Middleware moves to the forefront. Computer, 32(5):17–19, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r5017.pdf.

Garber:1999:NBL

[1947] Lee Garber. News briefs: Linux support ranges from GUI to Big Blue; new XML standards; partnership for advanced computational infrastructure. Computer, 32(5):20–22, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r5020.pdf.

Goth:1999:YBUb

[1948] Greg Goth. Y2K in brief: Utilities say they are ready for Y2K. Computer, 32(5):23, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r5023.pdf.

Rubin:1999:RYDe

[1949] Howard Rubin. Rubin’s Y2K diary: Just say YES Corps. Computer, 32(5):24, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1999/pdf/
                                        r5024.pdf.

Anonymous:1999:ASe

[1950] Anonymous. Article summaries.

Computer, 32(5):26–27, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5026.pdf.

Grimshaw:1999:WAC

[1951] Andrew Grimshaw, Adam Ferrari, Frederick Knabe, and Marty Humphrey. Wide-area computing: Resource sharing on a large scale. Computer, 32 (5):29–37, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r5029.pdf;   http://www.computer.
org/computer/co1999/r5029abs.htm.

Gao:1999:EIG

[1952] Jerry Z. Gao, Cris Chen, Yasufumi Toyoshima, and David K. Leung. Engineering on the Internet for global software production. Computer, 32(5):38–47, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5038.pdf;
http://www.computer.org/computer/
co1999/r5038abs.htm.

Wall:1999:SRR

[1953] Larry Wall, David A. Taylor, Chris Horn, Paul Bassett, John K. Ousterhout, Martin L. Griss, Richard Mark Soley, Jim Waldo, and Charles Simonyi. Software [R]evolution: A roundtable. Computer, 32(5):48–57, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.

                                        org/co/books/co1999/pdf/r5048.pdf;
                                         http://www.computer.org/computer/
                                        co1999/r5048abs.htm.

Cooke:1999:UBI

[1954] Daniel Cooke, Joseph Urban, and Scott Hamilton. Unix and beyond: An interview with Ken Thompson. Computer, 32(5):58–64, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r5058.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r5058abs.htm.

Boyle:1999:DYT

[1955] James M. Boyle, R. Daniel Resler, and Victor L. Winter. Do you trust your compiler? Computer, 32(5):65–73, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/r5065.pdf;
                                         http://www.computer.org/computer/
                                        co1999/r5065abs.htm.

Ram:1999:SMS

[1956] Sudha Ram, Jinsoo Park, and George L. Ball. Semantic-model support for Geographic Information Systems. Computer, 32(5):74–81, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r5074.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r5074abs.htm.

Scherrer:1999:TAF

[1957] Deborah Scherrer. Technical activities forum: Wearable computing comes of age. Computer, 32(5):82–83, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5082.pdf.

Piner:1999:CSCe

[1958] Mary-Louise G. Piner. Computer Society connection: Advancing software engineering as a profession. Computer, 32(5):84–87, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5084.pdf.

Anonymous:1999:CCe

[1959] Anonymous. Call and calendar.

Computer, 32(5):88–93, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5088.pdf.

Anonymous:1999:COe

[1960] Anonymous. Career opportunities.

Computer, 32(5):94–101, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5094.pdf.

Lutz:1999:NBN

[1961] Michael J. Lutz. New books:

No character: How Apple lots its way. Computer, 32(5):102, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5102.pdf.

Kroeker:1999:NTNa

[1962] Kirk L. Kroeker. New tools: Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s processoriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5):103–107, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5103.pdf.

Throop:1999:SOS

[1963] Joe Throop. Standards: OpenMP: Shared-memory parallelism from the ashes. Computer, 32(5):108–109, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5108.pdf.

Myers:1999:MEC

[1964] Ware Myers. Management: Early communication key to software project success. Computer, 32(5):110–111, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5110.pdf.

Vetter:1999:IWW

[1965] Ron Vetter. Internet watch: Webbased enterprise computing. Computer, 32(5):112–113, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5112.pdf.

Siegel:1999:COT

[1966] Jon Siegel. Component and object technology: A preview of CORBA 3. Computer, 32(5):114–116, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5114.pdf.

Dalcher:1999:OCS

[1967] Darren Dalcher. Open channel: Simplicity and power: When more means less. Computer, 32(5):117, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r5117.pdf.

Lewis:1999:BCS

[1968] Ted Lewis. Binary critic: Something for nothing. Computer, 32(5):120, 118–119, May 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r5120.pdf.

Tripp:1999:PM

[1969] Leonard L. Tripp and Thomas J.

Plevyak. Presidents’ message. Computer, 32(6):4, 9, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co1999/
pdf/r6004.pdf.

Frankston:1999:LHN

[1970] Bob Frankston, Alec Effrat, Fred L.

Schneider, Ed Viguerie, William D. Jackson, Jim Geringer, Kenneth Nichols, Alexander Fronk, Pearl Brereton, David Budgen, and Geoff Hamilton. Letters: The home network will be an IP network; pining for the perfect job; what the dickens make Thompson fault Linux?; UNIX Systems Laboratory; presto! A true test of AI; are software patents really different?; support for hypertext maintenance. Computer, 32(6):6–9, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6006.pdf.

Ortiz:1999:ITB

[1971] Sixto Ortiz Jr. Industry trends: The battle over real-time Java. Computer, 32(6):13–15, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6013.pdf.

Garber:1999:TNM

[1972] Lee Garber. Technology news: Melissa virus creates a new type of threat. Computer, 32(6):16–19, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6016.pdf.

Lawton:1999:NBB

[1973] George Lawton. News briefs: Battle begins for device network supremacy; VXML Forum is created; Advanced Internet Center opens. Computer, 32(6):20–21, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6020.pdf. Dis-

cusses Sun/Sony/Philips Jini/HAVi vs. Microsoft Universal Plug and Play (UPnP) vs. Lucent’s InfernoSpaces.

Lear:1999:YRYa

[1974] Anne Lear. Y2K report: Y2K in brief:

Latin America faces Y2K difficulties; federal agencies redefine “mission critical”; Y2K market adapting for small business. Computer, 32(6):22, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6022.pdf.

Rubin:1999:RYDf

[1975] Howard Rubin. Rubin’s Y2K diary: Lashing down the world. Computer, 32(6):24–25, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6024.pdf.

Anonymous:1999:ASf

[1976] Anonymous. Article summaries.

Computer, 32(6):26–27, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6026.pdf.

Stankovic:1999:MAR

[1977] John A. Stankovic, Sang Hyuk Son, and Jorgen Hansson. Misconceptions about real-time databases. Computer, 32(6):29–36, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6029.pdf;
http://www.computer.org/computer/
co1999/r6029abs.htm.

Anonymous:1999:VLH

[1978] Anonymous. Vying for the lead in high-performance processors: An interview with Gary Lauterbach. Computer, 32(6):38–41, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r6038.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r6038abs.htm. Lauterbach is

the Chief Architect of the Sun Ultra-SPARC III chip.

Birnbaum:1999:HVA

[1979] Mark Birnbaum and Howard Sachs. How VSIA answers the SOC dilemma. Computer, 32(6):42–50, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r6042.pdf;
                                        http://www.computer.org/computer/
                                        co1999/R6042abs.htm.

Zorian:1999:TEC

[1980] Yervant Zorian, Erik Jan Marinissen, and Sujit Dey. Testing embedded-core-based system chips. Computer, 32(6):52–60, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r6052.pdf;
                                        http://www.computer.org/computer/
                                        co1999/R6052abs.htm.

Silcott:1999:SDN

[1981] Gary Silcott, Janet Wilson, Neil Peterson, William Peisel, and Kirk L. Kroeker. SOCs drive new product development: SOC challenges for wireless handsets; thin clients benefit form SOC; networking the office with SOCs; MEMS technology emerges. Computer, 32(6):61–66, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6061.pdf;
http://www.computer.org/computer/
co1999/R6061abs.htm.

Lawrence:1999:DLA

[1982] Steve Lawrence, C. Lee Giles, and Kurt Bollacker. Digital libraries and autonomous citation indexing. Computer, 32(6):67–71, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6067.pdf;
http://www.computer.org/computer/
co1999/R6067abs.htm.

Dahlgren:1999:COM

[1983] Fredrik Dahlgren and Josep Torrellas. Cache-only memory architectures. Computer, 32(6):72–79, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6072.pdf;
http://www.computer.org/computer/
co1999/R6072abs.htm.

Weiss:1999:EPP

[1984] Shlomo Weiss and Ehud Finkelstein.

Extending PCI performance beyond the desktop. Computer, 32(6):80–87, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r6080.pdf;
 http://www.computer.org/computer/
co1999/R6080abs.htm.

Piner:1999:CSCf

[1985] Mary-Louise G. Piner. Computer Society connection: Fall conference preview; computer architect prize awarded to Smith; Past Computer Society President receives service award. Computer, 32(6):88–92, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r6088.pdf.

Anonymous:1999:CCf

[1986] Anonymous. Call and calendar. Computer, 32(6):93–105, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r6093.pdf.

Lutz:1999:NBTa

[1987] Michael J. Lutz. New books: Towrad software you can depend on. Computer, 32(6):106, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r6106.pdf.

Kroeker:1999:NTNb

[1988] Kirk L. Kroeker. New tools: Net development: Entrust’s Open Toolkits; IBM’s Parallel Enterprise Server; Talarian’s message-queueing product; SGI’s open source SAN management software. software development: ParaSoft’s automatic error-detection tool; Aonix’s GUI management system for Solaris 7. component technology: IONA Technologies’ EJB Toolkit; ProtoView’s diagramming tool. Computer, 32(6):107–110, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6107.pdf.

Carpenter:1999:IWW

[1989] Brian E. Carpenter. Internet watch: What’s all the fuss about differentiated services? Computer, 32(6):111–112, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6111.pdf.

Bach:1999:SRRb

[1990] James Bach. Software realities:

Risk and requirements-based testing. Computer, 32(6):113–114, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6113.pdf.

Figueiredo:1999:IEE

[1991] Marco A. Figueiredo, Patrick H.

Stakem, Thomas P. Flatley, and Tonjua M. Hines. Integrated engineering: Extending NASA’s data processing to spacecraft. Computer, 32(6):115–118, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6115.pdf.

Lewis:1999:BCJ

[1992] Ted Lewis. Binary critic: Jumpstarting the global vending machine. Computer, 32(6):120, 118–119, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r6120.pdf. See

letter and reply [2044].

Leavitt:1999:TNO

[1993] Neal Leavitt. Technology news: Online 3D: Still waiting after all these years. Computer, 32(7):4, 6–7, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7004.pdf.

Lawton:1999:NBS

[1994] George Lawton. News briefs: Standard to bring order to IP traffic; European Parliament wants surveillance of net traffic; ATM forum seeks to make ATM more IP-friendly; chatterbots: The Web gets help from AI. Computer, 32(7):11–13, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7011.pdf.

Lear:1999:YRYb

[1995] Anne C. Lear. Y2K report: Y2K in brief: Y2K command centers focus on containment. Computer, 32(7):14, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7014.pdf.

Rubin:1999:RYDg

[1996] Howard Rubin. Rubin’s Y2K diary: The Declaration of Interdependence. Computer, 32(7):15, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7015.pdf.

Anonymous:1999:ASg

[1997] Anonymous. Article summaries.

Computer, 32(7):16–17, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7016.pdf.

Sipper:1999:ECC

[1998] Moshe Sipper. The emergence of cellular computing. Computer, 32(7):18–26, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r7018.pdf;
http://www.computer.org/computer/
co1999/r7018abs.htm.

Reilly:1999:DAM

[1999] Matt Reilly. Designing an Alpha microprocessor. Computer, 32(7):27–34, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7027.pdf;
http://www.computer.org/computer/
co1999/r7027abs.htm.

Meyer:1999:CBD

[2000] Bertrand Meyer and Christine Mingins. Component-based development: From buzz to spark. Computer, 32(7):35–37, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7035.pdf.

Beugnard:1999:MCC

[2001] Antoine Beugnard, Jean-Marc Jézéquel, Noël Plouzeau, and Damien Watkins. Making components contract aware.

Computer, 32(7):38–45, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7038.pdf;
http://www.computer.org/computer/
co1999/r7038abs.htm.

Cicalese:1999:BSD

[2002] Cynthia Della Torre Cicalese and Shmuel Rotenstreich. Behavioral specification of distributed software component interfaces. Computer, 32(7):46– 53, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r7046.pdf;
 http://www.computer.org/computer/
co1999/r7046abs.htm.

Stets:1999:CBA

[2003] Robert J. Stets, Galen C. Hunt, and Michael L. Scott. Component-based APIs for versioning and distributed applications. Computer, 32(7):54–61, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r7054.pdf;
 http://www.computer.org/computer/
co1999/r7054abs.htm.

Narasimhan:1999:UIE

[2004] Priya Narasimhan, Louise E. Moser, and P. M. Melliar-Smith. Using interceptors to enhance CORBA. Computer, 32(7):62–68, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7062.pdf;
http://www.computer.org/computer/
co1999/r7062abs.htm.

Anonymous:1999:CPIa

[2005] Anonymous. Call for participation:

1999 international conference on parallel processing (ICPP-99), the University of Aizu, Aizu, Fukushima, Japan, September 21–24, 1999. Computer, 32(7):69, July 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1999:CPIb

[2006] Anonymous. Call for papers: International symposium on information technology: Coding and computing, March 27–29, 2000, Las Vegas, NV, USA. Computer, 32(7):70, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:1999:CAO

[2007] Guijun Wang, Liz Ungar, and Dan Klawitter. Component assembly for OO distributed systems. Computer, 32(7):71–78, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7071.pdf;
http://www.computer.org/computer/
co1999/r7071abs.htm.

Baker:1999:TAF

[2008] Mark Baker, Rajkumar Buyya, and Dan Hyde. Technical activities forum: Cluster computing: A highperformance contender. Computer, 32 (7):79–80, July 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r7079.pdf.

Piner:1999:CSCg

[2009] Mary-Louise G. Piner. Computer Society connection: Teacher and innovator receives award; board approves ballot for Computer Society offices. Computer, 32(7):81–83, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r7081.pdf.

Israel:1999:FRC

[2010] Michel Israel. Financial report: Computer Society financially strong: 1998 financial results. Computer, 32(7):84– 89, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/r7084.pdf.

Anonymous:1999:COf

[2011] Anonymous. Career opportunities. Computer, 32(7):90–93, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r7094.pdf.

Anonymous:1999:CCg

[2012] Anonymous. Call and calendar. Computer, 32(7):94–98, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r7094.pdf.

Lutz:1999:NBCa

[2013] Michael J. Lutz. New books: Case-driven object modelling with UML; virtual-human integration; genetic programming; Java image processing; WAP programming. Computer, 32(7):99, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7099.pdf.

Kroeker:1999:NTNc

[2014] Kirk L. Kroeker. New tools: Net development: SoftQuad Software’s XML tool; Rebol’s Internet messaging language. component technology: Dart Communications’ terminal emulation component; Artisoft’s telephony components; Digital Design’s controls for creating OLAP apps. hardware development: Hewlett– Packard’s circuit synthesis tool; SGI’s MIPS compiler; Applied Microsystems’ tools for Motorola processors. Computer, 32(7):100–102, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7100.pdf.

Knight:1999:IWR

[2015] Stefan Knight. Internet watch:

The role of ADSL in Internet access. Computer, 32(7):103, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7103.pdf.

Eddon:1999:COT

[2016] Guy Eddon. Component and object technology: COM+: The evolution of component services. Computer, 32(7):104–106, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r7104.pdf.

Carrier:1999:MML

[2017] Lynn Carrier. Management: Managing at light speed. Computer, 32 (7):107–109, July 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1999/pdf/
                                        r7107.pdf.

Macedonia:1999:BCS

[2018] Michael Macedonia and Ted Lewis. Binary critic: Sony versus Wintel: Mortal combat. Computer, 32(7):112, 109–111, July 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/r7112.pdf.

Greenberg:1999:ITS

[2019] Ilan Greenberg and Lee Garber. Industry trends: Searching for new search technologies. Computer, 32(8):4, 6–7, 11, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8004.pdf.

Clark:1999:TNP

[2020] David Clark. Technology news: Preparing for a new generation of wireless data. Computer, 32(8):8–11, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/r8008.pdf.

Anonymous:1999:ICM

[2021] Anonymous. IEEE CS membership application. Computer, 32(8):12–14, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r8012.pdf.

Lear:1999:NBE

[2022] Anne C. Lear. News briefs: Explorer worm targets networks; H-1B visa cap reached; higher limit proposed; encryption industry grows outside US; software design flaw spurs lawsuits. Computer, 32(8):15–17, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8015.pdf.

Lear:1999:YRYc

[2023] Anne C. Lear. Y2K report: Y2K in brief: Global banking simulation is successful. Computer, 32(8):18, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8018.pdf.

Rubin:1999:RYDh

[2024] Howard Rubin. Rubin’s Y2K diary:

Paying the price for ubiquitous connectivity. Computer, 32(8):19, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8019.pdf.

Anonymous:1999:ASh

[2025] Anonymous. Article summaries.

Computer, 32(8):20–21, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8020.pdf.

Silverman:1999:EMM

[2026] Robert D. Silverman. Exposing the mythical MIPS year. Computer, 32(8):22–26, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8022.pdf;
http://www.computer.org/computer/
co1999/r8022abs.htm.

Hennessy:1999:FSR

[2027] John Hennessy. The future of systems research. Computer, 32(8):27–33, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/r8027.pdf;
 http://www.computer.org/computer/
co1999/r8027abs.htm.

Ramakrishnan:1999:GEI

[2028] Naren Ramakrishnan and Ananth Y. Grama. Guest Editors’ introduction: Data mining: From serendipity to science. Computer, 32(8):34–37, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8034.pdf.

Ganti:1999:MVL

[2029] Venkatesh Ganti, Johannes Gehrke, and Raghu Ramakrishnan. Mining very large databases. Computer, 32(8):38–45, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8038.pdf;
http://www.computer.org/computer/
co1999/r8038abs.htm.

Han:1999:CBM

[2030] Jiawei Han, Laks V. S. Lakshmanan, and Raymond T. Ng. Constraintbased, multidimensional data mining. Computer, 32(8):46–50, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8046.pdf;
http://www.computer.org/computer/
co1999/r8046abs.htm.

Hellerstein:1999:IDA

[2031] Joseph M. Hellerstein, Ron Avnur, Andy Chou, Christian Hidber, Chris Olston, Vijayshankar Raman, Tali Roth, and Peter J. Haas. Interactive data analysis: The control project. Computer, 32(8):51–59, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8051.pdf;
http://www.computer.org/computer/
co1999/r8051abs.htm.

Chakrabarti:1999:MWL

[2032] Soumen Chakrabarti, Byron E. Dom, S. Ravi Kumar, Prabhakar Raghavan, Sridhar Rajagopalan, Andrew Tomkins, David Gibson, and Jon Kleinberg. Mining the Web’s link structure. Computer, 32(8):60–67, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8060.pdf;
http://www.computer.org/computer/
co1999/r8060abs.htm.

Karypis:1999:CHC

[2033] George Karypis, Eui-Hong (Sam)

Han, and Vipin Kumar NEWS.

Chameleon: Hierarchical clustering using dynamic modeling. Computer, 32(8):68–75, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8068.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r8068abs.htm.

Piner:1999:CSCh

[2034] Mary-Louise G. Piner. Computer Society connection: Computer Society recognizes Intel ISEF high school achievers; Goode award presented to Purdue scholar; lecture series shares members’ expertise; Board approves reading of bylaws amendment on Press Activities Board. Computer, 32(8):76–80, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8076.pdf.

Anonymous:1999:CCh

[2035] Anonymous. Call and calendar. Computer, 32(8):81–85, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8081.pdf.

Anonymous:1999:COh

[2036] Anonymous. Career opportunities. Computer, 32(8):86–88, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8086.pdf.

Anonymous:1999:API

[2037] Anonymous. Advertiser/product index. Computer, 32(8):89, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8086.pdf.

Lutz:1999:NBSb

[2038] Michael J. Lutz. New books: Software development from the outside in. Computer, 32(8):90, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8090.pdf.

Kroeker:1999:NTSd

[2039] Kirk L. Kroeker. New tools: Software development: Premia’s editing system; Diab Data’s compiler suites for Coldfire; Borland’s visual development tools for Java 2; Sun outlines Java 2 at JavaOne. net development: InfoAccess’ HTML production tool; GoAhead Software’s free embedded Web server; Planetweb and Paradise Innovations produce next-gen set-top box. Computer, 32(8):91–93, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8091.pdf.

Ritter:1999:IWC

[2040] Terry Ritter. Internet watch: Cryptography: Is staying with the herd really best? Computer, 32(8):94–95, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r8094.pdf.

Levia:1999:IEP

[2041] Oz Levia. Integrated engineering: Programming system architectures with Java. Computer, 32(8):96–98, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8096.pdf.

Bach:1999:SRP

[2042] Jonathan Bach. Software realities: Playing the expert game. Computer, 32(8):99–103, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8099.pdf.

Lewis:1999:BCM

[2043] Ted Lewis. Binary critic: Mainframes are dead, long live mainframes. Computer, 32(8):104, 102–103, August 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r8104.pdf.

Maurer:1999:LCR

[2044] Peter M. Maurer, Mark K. Gardner, Darren Dalcher, Mordechai Ben-Menachem, Ted Lewis, A. B. Cannara, and Ken Nichols. Letters: Components are the revolutionary legacy of VBX; focus on answers, not features; startups can pop up anywhere; genetics and software patents. Computer, 32(9):4, 6, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9004.pdf. See

[1992].

Ortiz:1999:ITA

[2045] Sixto Ortiz Jr. Industry trends: Appliances challenge the server market. Computer, 32(9):11–13, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9011.pdf.

Charles:1999:TNI

[2046] John Charles. Technology news: Interplanetary network aims for the stars. Computer, 32(9):16–18, 21, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9016.pdf.

Lear:1999:NBW

[2047] Anne C. Lear. News briefs: Will XML plans compete or cooperate?; speech technology advances demonstrated; networking processors drive chip market; HP pitches net service. Computer, 32(9):19–20, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9019.pdf.

Lear:1999:YRYd

[2048] Anne C. Lear. Y2K report: Y2K in brief: UN conference highlights concerns, progress. Computer, 32(9):22, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9022.pdf.

Rubin:1999:RYDi

[2049] Howard Rubin. Rubin’s Y2K diary: Are the media crying “sheep”? Computer, 32(9):23, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9023.pdf.

Anonymous:1999:ASi

[2050] Anonymous. Article summaries. Computer, 32(9):24–25, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9024.pdf.

Bellini:1999:CPM

[2051] Pierfrancesco Bellini, Fabrizio Fioravanti, and Paolo Nesi Cybersquare. Computing practices: Managing music in orchestras. Computer, 32(9):26–35, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9026.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r9026abs.htm.

Holmes:1999:MEC

[2052] W. Neville Holmes. The myth of the educational computer. Computer, 32(9):36–42, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9036.pdf;
                                        http://www.computer.org/computer/
                                        co1999/r9036abs.htm.

Harasim:1999:FOL

[2053] Linda Harasim. A framework for online learning: The Virtual-U. Computer, 32(9):44–49, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9044.pdf;

http://www.computer.org/computer/
co1999/r9044abs.htm.

Roschelle:1999:DES

[2054] Jeremy Roschelle, Chris DiGiano, Manolis Koutlis, Alexander Repenning, Jonathan Phillips, Nicholas Jackiw, and Dan Suthers. Developing educational software components. Computer, 32(9):50–58, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9050.pdf;
http://www.computer.org/computer/
co1999/r9050abs.htm.

Suzuki:1999:LDS

[2055] Junichi Suzuki and Yoshikazu Yamamoto. Leveraging distributed software development. Computer, 32(9): 59–65, September 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
r9059.pdf;   http://www.computer.
org/computer/co1999/r9059abs.htm.

Nunamaker:1999:CCN

[2056] Jay F. Nunamaker. Collaborative computing: The next millennium. Computer, 32(9):66–71, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9066.pdf;
http://www.computer.org/computer/
co1999/r9066abs.htm.

Anonymous:1999:CSE

[2057] Anonymous. Computer society election. Computer, 32(9):72–79, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/r9072.pdf.

Piner:1999:CSCi

[2058] Mary-Louise G. Piner. Computer Society connection: IEEE Presidentelect candidates answer key computer society questions; Curriculum 2001: Bringing the future to the classroom. Computer, 32(9):80–85, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9080.pdf.

Lollock:1999:CSB

[2059] Guylaine M. Lollock. Computer Society Board endorses Parrish for IEEE President-elect. Computer, 32(9): 84, September 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1999/pdf/
                                        r9080.pdf.

Chang:1999:CBF

[2060] Carl K. Chang, Gerald Engel, Willis King, Eric Roberts, Russ Shackelford, Robert H. Sloan, and Pradip K. Srimani. Curriculum 2001: Bringing the future to the classroom. Computer, 32(9):85–88, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9080.pdf.

Anonymous:1999:CCi

[2061] Anonymous. Call and calendar. Computer, 32(9):89–94, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9089.pdf.

Anonymous:1999:COi

[2062] Anonymous. Career opportunities.

Computer, 32(9):96–104, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9096.pdf.

Lutz:1999:NBTb

[2063] Michael J. Lutz. New books: Take A stroll down IT memory lane; computers in society; software development; function in programming; Verilog design; taking the Web to the workplace. Computer, 32(9):105, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9105.pdf.

Kroeker:1999:NTSe

[2064] Kirk L. Kroeker. New tools:

Software development: SoftQuest Systems’ project management software; Metrowerks’ Java development tool for Mac OS; CAD-UL’s embedded-centric C/C++ toolkit; KL Group’s Java components. Computer, 32(9):106–107, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/r9106.pdf.

Snow:1999:IWI

[2065] Andrew P. Snow. Internet watch: Internet implications of telephone access. Computer, 32(9):108–110, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9108.pdf. See

correction [2069].

Boehm:1999:MMS

[2066] Barry Boehm. Management: Managing software productivity and reuse. Computer, 32(9):111–113, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9111.pdf.

Spinellis:1999:COT

[2067] Diomidis Spinellis. Component and object technology: Explore, excogitate, exploit: Component mining. Computer, 32(9):114–116, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/r9114.pdf.

Lewis:1999:BCFb

[2068] Ted Lewis. Binary critic: Fast, expensive, and horribly complex. Computer, 32(9):120, 118–119, September 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://developer.intel.com/
                                        design/ia64/index.htm;   http://
                                        dlib.computer.org/co/books/co1999/
                                        pdf/r9120.pdf.

Anonymous:1999:CIA

[2069] Anonymous. Correction: Internet applications of telephone access. Computer, 32(10):4, 6–7, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx004.pdf. See

Rodgers:1999:LLP

[2070] R. P. C. Rodgers, W. Lowell Morgan, Ware Myers, Bob Schuchman, Todd Martin, Charles C. Philipp, Ted Lewis, Jean-François Colonna, Rich Babcock, and James Boyle. Letters: Little place for computers in a rigorous curriculum; A welcome reality check on copying with overload; how to cope when you are a few neurons short of infinity; to be an expert: Just listen; how a good index can make an expert; technologists care about privacy; errors in program transformation table; problems with theorems. Computer, 32(10):4, 6–7, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx004.pdf.

Leavitt:1999:ITD

[2071] Neal Leavitt. Industry trends: Domain name solution causes new problems. Computer, 32(10):11–15, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx011.pdf.

Lear:1999:NBA

[2072] Anne C. Lear. News briefs: Agreement reached on new I/O architecture; CORBA component model approved; student develops smallest Web server; membership news. Computer, 32(10):16–18, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx016.pdf;
http://www-ccs.cs.umass.edu/~shri/
iPic.html.

Lear:1999:YBC

[2073] Anne C. Lear. Y2K in brief: Contingency plans should extend into 2000. Computer, 32(10):19, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx019.pdf.

Rubin:1999:RYDj

[2074] Howard Rubin. Rubin’s Y2K diary: Uncovering weak links in the readiness chain. Computer, 32(10):20, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx020.pdf.

Anonymous:1999:ASCa

[2075] Anonymous. Article summaries. Computer, 32(10):22–23, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx022.pdf.

Markus:1999:MSF

[2076] Karen W. Markus and Kaigham J. Gabriel Cybersquare. MEMS: The systems function revolution. Computer, 32(10):25–31, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx025.pdf;
                                        http://www.computer.org/computer/
                                        co1999/rx025abs.htm.

Pfleeger:1999:AEE

[2077] Shari Lawrence Pfleeger. Albert Einstein and empirical software engineering. Computer, 32(10):32–37, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx032.pdf;
http://www.computer.org/computer/
co1999/rx032abs.htm.

Moitra:1999:SES

[2078] Deependra Moitra. Software engineering in the small: Practical software engineering and management. Computer, 32(10):39–40, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx039.pdf.

Kelly:1999:PIS

[2079] Declan P. Kelly and Bill Culleton. Process improvement for small organizations. Computer, 32(10):41–47, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/rx041.pdf;
http://www.computer.org/computer/
co1999/rx041abs.htm.

Woodward:1999:EPM

[2080] Stuart Woodward. Evolutionary project management. Computer, 32 (10):49–57, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://

dlib.computer.org/co/books/co1999/
pdf/rx049.pdf;   http://www.computer.
org/computer/co1999/rx049abs.htm.

Cusumano:1999:SDI

[2081] Michael A. Cusumano and David B.

Yoffie. Software development on Internet time. Computer, 32(10):60–69, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/rx060.pdf;
 http://www.computer.org/computer/
co1999/rx060abs.htm.

Feature:1999:ECE

[2082] Kent Beck. Embracing change with extreme programming. Computer, 32(10):70–77, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx070.pdf;
http://www.computer.org/computer/
co1999/rx070abs.htm.

Arthur:1999:EEI

[2083] James D. Arthur, Markus K. Gröner, Kelly J. Hayhurst, and C. Michael Holloway. Evaluating the effectiveness of independent verification and validation. Computer, 32(10):79–83, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx079.pdf;
http://www.computer.org/computer/
co1999/rx079abs.htm.

Piner:1999:CSCj

[2084] Mary-Louise G. Piner. Computer society connection: Computer Society and ACM approve software; engineering code of ethics; opportunities available for professional and student chapters; student awards programs seek entrants by 31 October. Computer, 32(10):84–90, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx084.pdf.

Anonymous:1999:CCj

[2085] Anonymous. Call and calendar. Computer, 32(10):91–93, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx091.pdf.

Anonymous:1999:C

[2086] Anonymous.  Career.  Com-

puter, 32(10):94–111, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx094.pdf.

Kroeker:1999:NTN

[2087] Kirk L. Kroeker. New tools: Net development: Macromedia’s Web application and E-commerce development family; Anark Interactive Media’s 3D Web tool. embedded development: Force Computer’s singleboard computer for embedded applications; CAD-UL’s toolkit for embedded processors; AMC’s embedded software visibility tools. Computer, 32(10):113–115, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rx113.pdf.

Lutz:1999:NBCb

[2088] Michael J. Lutz. New books: Creating a virtual window to the world; optical networks in layered multiwavelength; automated software testing; software improvement through refactoring; getting a feel for Java. Computer, 32(10):116, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx116.pdf.

NMMDT:1999:IWN

[2089] Network Maturity Model Development Team. Internet watch: The network maturity model for Internet development. Computer, 32(10):117–118, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/rx117.pdf.

Bullock:1999:SRI

[2090] James Bullock. Software realities: Improving the development system model. Computer, 32(10):119–120, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx119.pdf.

Do:1999:IES

[2091] Sung-Hee Do and Nam P. Suh. Integrated engineering: Systematic OO programming with axiomatic design. Computer, 32(10):121–124, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rx121.pdf.

Lewis:1999:BCU

[2092] Ted Lewis. Binary critic: UbiNet: The ubiquitous Internet will be wireless. Computer, 32(10):128, 126–127, October 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/rx128.pdf.

Pelley:1999:LAI

[2093] Mike Pelley and Ted Lewis. Letters:

Ask Intel: Price and cost aren’t related. Computer, 32(11):9, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry009.pdf.

Lawton:1999:ITS

[2094] George Lawton. Industry trends: Storage technology takes center stage. Computer, 32(11):10–12, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry010.pdf.

Goldberg:1999:TNV

[2095] Lee H. Goldberg. Technology news: Vendors are counting on appliance-on-chip technology. Computer, 32(11):13–16, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry013.pdf.

Lear:1999:NBT

[2096] Anne C. Lear. News briefs:

Tentative domain-name agreement reached; chip makers stake out highperformance ground; lip-reading for better speech recognition. Computer, 32(11):17–18, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry017.pdf.

Lear:1999:YBU

[2097] Anne C. Lear. Y2K in brief: US issues Y2K travel forecast. Computer, 32(11):19, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry017.pdf.

Rubin:1999:RYDk

[2098] Howard Rubin. Rubin’s Y2K diary: A crisis of arrogance. Computer, 32(11):20–21, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry020.pdf.

Anonymous:1999:AS

[2099] Anonymous. Article summaries. Computer, 32(11):22–23, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry022.pdf.

Sylvester:1999:RDS

[2100] Dennis Sylvester and Kurt Keutzer. Cybersquare: Rethinking deep-submicron circuit design. Computer, 32(11):25– 33, November 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
ry025.pdf;   http://www.computer.
org/computer/co1999/ry025abs.htm.

Shriver:1999:JCI

[2101] Bruce Shriver and Peter Capek. Just curious: An interview with John Cocke. Computer, 32(11):34–41, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry034.pdf;

http://www.computer.org/computer/
co1999/ry034abs.htm.

Shriver:1999:SCC

[2102] Bruce Shriver and Peter Capek.

The Seymour Cray computer science and engineering award. Computer, 32(11):35, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry034.pdf;
http://www.computer.org/computer/
co1999/ry034abs.htm.

Buchholz:1999:PS

[2103] Werner Buchholz. Project Stretch.

Computer, 32(11):36, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry034.pdf;
http://www.computer.org/computer/
co1999/ry034abs.htm.

Sussenguth:1999:IAM

[2104] Ed Sussenguth. IBM’s ACS-1 machine. Computer, 32(11):38, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry034.pdf;
http://www.computer.org/computer/
co1999/ry034abs.htm.

Kapur:1999:GEI

[2105] Rohit Kapur and Thomas W. Williams.

Guest editors’ introduction: Tough challenges as design and test go nanometer. Computer, 32(11):42– 45, November 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1999/pdf/
                                        ry042.pdf.

Aitken:1999:NTE

[2106] Robert C. Aitken. Nanometer technology effects on fault models for IC testing. Computer, 32(11):46–51, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/ry046.pdf;
                                         http://www.computer.org/computer/
                                        co1999/ry046abs.htm.

Needham:1999:NTC

[2107] Wayne M. Needham. Nanometer technology challenges for test and test equipment. Computer, 32(11):52– 57, November 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

                                        computer.org/co/books/co1999/pdf/
                                        ry052.pdf;   http://www.computer.
                                        org/computer/co1999/ry052abs.htm.

Cheng:1999:CDA

[2108] Kwang-Ting Cheng and Angela Krstic. Current directions in automatic test-pattern generation. Computer, 32(11):58–64, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/ry058.pdf;
                                        http://www.computer.org/computer/
                                        co1999/ry058abs.htm.

Wagner:1999:RSB

[2109] Kenneth D. Wagner. Robust scanbased logic test in VDSM technologies. Computer, 32(11):66–74, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry066.pdf;
http://www.computer.org/computer/
co1999/ry066abs.htm.

Bowyer:1999:YCB

[2110] Kevin Bowyer. Your call: ‘Big Brother’ — or proper vigilence? Computer, 32(11):75, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry075.pdf.

Piner:1999:CSCk

[2111] Mary-Louise G. Piner. Computer society connection: Society competition offers US$25,000 for top student designs; innovator in consumer computing receives entrepreneur award. Computer, 32(11):76–77, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry076.pdf.

Anonymous:1999:CSN

[2112] Anonymous. Computer Society notes: Computer Editor-in-Chief named Rader Professor. Computer, 32(11):78, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry076.pdf.

Rosenthal:1999:TAF

[2113] Charles Rosenthal and Joseph Damore.

Technical activities forum: Hot topics in electronic design automation. Computer, 32(11):79, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/ry079.pdf.

Anonymous:1999:CCk

[2114] Anonymous. Call and calendar. Computer, 32(11):80–82, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/ry080.pdf.

Anonymous:1999:NBb

[2115] Anonymous. Advertiser/product index. Computer, 32(11):83, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1999:COj

[2116] Anonymous. Career opportunities. Computer, 32(11):84–118, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lutz:1999:NBF

[2117] Michael J. Lutz. New books: Folding all things virtual into urban life; performance modeling in network design; practical automated software testing; Java patterns; testing software step by step; working with UNIX shells. Computer, 32(11):119, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/ry119.pdf.

Kroeker:1999:NTSf

[2118] Kirk L. Kroeker. New tools: Software development: Monash University’s free Java development environment; ParaSoft’s error prevention tool; Technosoft’s C++ development kit; VideoSoft’s spell-checking and thesaurus controls. network development: Network Security’s security management tool; Fluke’s network inspector; MetaCreation’s Web-oriented painting tool. Computer, 32(11):120–122, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry120.pdf. See

correction [2150].

Marden:1999:IWT

[2119] Philip M. Marden, Jr. and Ethan V.

Munson. Internet watch: Today’s style sheet standards: The great vision blinded. Computer, 32(11):123– 125, November 1999. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://dlib.

computer.org/co/books/co1999/pdf/
ry123.pdf.

Severance:1999:SIW

[2120] Charles Severance. Standards: IEEE 802.11: Wireless is coming home. Computer, 32(11):126–127, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry126.pdf.

Reifer:1999:MTT

[2121] Donald J. Reifer. Management:

A tale of three developers. Computer, 32(11):128–130, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/ry128.pdf.

Meyer:1999:COTb

[2122] Bertrand Meyer. Component and object technology: Every little bit counts: Toward more reliable software. Computer, 32(11):131–133, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/ry131.pdf.

Lewis:1999:BCW

[2123] Ted Lewis. Binary critic: Where the smart money is? Computer, 32(11):136, 134–135, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/ry019.pdf;
                                        http://dlib.computer.org/co/books/
                                        co1999/pdf/ry136.pdf.

Gustafson:1999:LOF

[2124] John Gustafson, Rajat Todi, Robert Silverman, Collin Park, Ted Lewis, Ken Zuhr, and Ted Lewis. Letters: Operations are free: Data motion isn’t; Déjà vu processor complexities. Computer, 32(12):4, 6, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz004.pdf.

Piner:1999:RMN

[2125] Mary-Louise G. Piner. Report to members: New Society leadership for 2000. Computer, 32(12):8–9, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz008.pdf.

Hellman:1999:ITS

[2126] Reed Hellman. Industry trends: A semantic approach adds meaning to the Web. Computer, 32(12):13–16, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/rz013.pdf.

Clark:1999:TNT

[2127] David Clark. Technology news: Toward a new generation of simpler PCs. Computer, 32(12):17–19, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz017.pdf.

Lear:1999:NBI

[2128] Anne C. Lear. News briefs: IETF rejects net wiretap proposal; new plans ease net congestion; standard API to simplify network chip development; group releases realtime Java specification. Computer, 32(12):20–21, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz020.pdf.

Lear:1999:YBY

[2129] Anne C. Lear. Y2K in brief: Y2K failures could occur throughout 2000. Computer, 32(12):22, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz022.pdf.

Rubin:1999:RYDl

[2130] Howard Rubin. Rubin’s Y2K diary:

Will fast food accelerate Y2K collapse? Computer, 32(12):23, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz023.pdf.

Anonymous:1999:ASCc

[2131] Anonymous. Article summaries. Computer, 32(12):24–25, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz024.pdf.

Thakkar:1999:ISS

[2132] Shreekant (Ticky) Thakkar and Tom Huff. Internet streaming SIMD extensions. Computer, 32(12):26–34, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://dlib.computer.
org/co/books/co1999/pdf/rz026.pdf;
 http://www.computer.org/computer/
co1999/rz026abs.htm.

Foster:1999:DVD

[2133] Ian Foster, Joseph Insley, Gregor von Laszewski, Carl Kesselman, and Marcus Thiebaux. Distance visualization: Data exploration on the grid. Computer, 32(12):36–43, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz036.pdf;
http://www.computer.org/computer/
co1999/rz036abs.htm.

Shaffer:1999:VPV

[2134] Eric Shaffer, Daniel A. Reed, Shannon Whitmore, and Benjamin Schaeffer. Virtue: Performance visualization of parallel and distributed applications.

Computer, 32(12):44–51, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz044.pdf;
http://www.computer.org/computer/
co1999/rz044abs.htm.

Allen:1999:SEE

[2135] Gabrielle Allen, Tom Goodale, Gerd Lanfermann, Thomas Radke, Edward Seidel, Werner Benger, Hans-Christian Hege, Andre Merzky, Joan Massó, and John Shalf. Solving Einstein’s equations on supercomputers. Computer, 32(12):52–58, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz052.pdf;
http://www.computer.org/computer/
co1999/rz052abs.htm.

Johnson:1999:ISV

[2136] Christopher Johnson, Steven G.

Parker, Charles Hansen, Gordon L. Kindlmann, and Yarden Livnat. Interactive simulation and visualization. Computer, 32(12):59–65, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz059.pdf;
http://www.computer.org/computer/
co1999/rz059abs.htm.

Leigh:1999:VTE

[2137] Jason Leigh, Andrew E. Johnson, Maxine Brown, Daniel J. Sandin, and Thomas A. DeFanti. Visualization in teleimmersive environments. Computer, 32(12):66–73, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz066.pdf;
                                        http://www.computer.org/computer/
                                        co1999/rz066abs.htm.

Anonymous:1999:AIV

[2138] Anonymous. Annual index: Volume 32, 1999. Computer, 32(12):74–85, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/rz074.pdf.

Piner:1999:CSN

[2139] Mary-Louise G. Piner. Computer Society notes: David Huffman (1925– 1999). Computer, 32(12):86, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz086.pdf.

Piner:1999:CSCl

[2140] Mary-Louise G. Piner. Computer society connection: Inaugural Cray award presented in a November ceremony; meritorious service and outstanding contribution awards show; appreciation of volunteer efforts; Computer recognizes reviewers’ expertise. Computer, 32(12):86–90, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz086.pdf.

Anonymous:1999:CCl

[2141] Anonymous. Call and calendar. Computer, 32(12):91–94, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz091.pdf.

Anonymous:1999:APIb

[2142] Anonymous. Advertiser/product index. Computer, 32(12):95, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz091.pdf.

Anonymous:1999:COk

[2143] Anonymous. Career opportunities.

Computer, 32(12):96–136, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz096.pdf.

Lutz:1999:NBW

[2144] Michael J. Lutz. New books: Where teamwork and software meet; mapping out software with logic; time traveling; putting COM and CORBA together; using OO for the network; history of artificial intelligence. Computer, 32(12):137, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz137.pdf.

Kroeker:1999:NTH

[2145] Kirk L. Kroeker. New tools: Hardware development: Transim’s Internetbased IC simulation tool; 3DSP’s development tools for system-on-chip designs; International Test Technologies’ tool for PowerPC testing. software development: Indigo Rose’s Setup Factory 5.0: The next-generation installation builder; SRC’s tool for converting 4GL code to COBOL; Sorokin’s tools to manage DLLs and OLE links. network development: Network Security’s security management tool; Microtest’s fiber-optic tester; US Software’s distributed operating environment. Computer, 32(12):138–140, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz138.pdf.

Chadwick:1999:IWS

[2146] David Chadwick. Internet watch: Smart cards aren’t always the smart choice. Computer, 32(12):142–143, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://dlib.computer.
                                        org/co/books/co1999/pdf/rz142.pdf.

Meyer:1999:COTc

[2147] Bertrand Meyer. Component and object technology: A really good idea. Computer, 32(12):144–147, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz144.pdf. See

correction [2151].

Bach:1999:SRW

[2148] James Bach. Software realities: What software reality is really about. Computer, 32(12):148–149, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.org/
                                        co/books/co1999/pdf/rz148.pdf.

Lewis:1999:BCD

[2149] Ted Lewis. Binary critic: Directories don’t get no respect. Computer, 32(12):152, 150–151, December 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co1999/pdf/rz152.pdf.

Anonymous:2000:CNT

[2150] Anonymous. Correction: New tools.

Computer, 33(2):7, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2004.pdf. See

[2118].

Anonymous:2000:CRG

[2151] Anonymous. Correction: A really good idea. Computer, 33(2):7, February 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.org/
co/books/co2000/pdf/r2004.pdf. See